A Complete Bibliography of Publications in the Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches

sur les pêcheries du Canada (1960–1969)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

29 August 2024 Version 1.05

Title word cross-reference

−0.8 [1528]. −29 [1538]. −4 [830]. −5 [1528]. ^{14} [69, 1184, 282, 1212]. ^{65} [1233, 1258]. _{1} [433]. _{24} [412]. α [834, 1440, 668, 818]. β [558, 655, 645, 842, 1042]. M_{2} [1548, 1512].

-Hydroxycorticosterone [1440, 668, 818]. -labelled [69, 282]. -Propiothetin [558, 655, 645, 842, 1042]. -Tocopherol [834].

1-Phosphate [107]. 145^{o} [983]. 16-Docosapentaenoic [335]. 17-Dehydrocholesterol [44]. 170^{o} [883]. 18-Hydroxycorticosterone [1056].

2-Phenoxyethanol [326]. 222 [1022]. 29-Isofucosterol [44].

3- [509]. 316 [781].

4-Dinitrophenylhydrazones [479]. 4-Inch [509]. 42^{o} [1112].

5 [944]. 5’-Monophosphate [944]. 50^{o} [983].

6-Phosphate [1576]. 66 [1217].

= [1369].

Âge [1187].

Abdominal [532]. Ability [295, 1422]. Abnormal [1236]. Abnormalities [97, 499]. Abnormally [1338]. above [1040]. Absence [416]. Absorbing [1381]. Absorption [1237, 1063]. Abundance [1524, 1511, 648, 3, 398]. Aburatsubo [1079]. Acanthocephala [1362]. Acanthocephalan [1368, 938]. Acantholiparis [699, 1384]. Acanthosentis [1368]. acatenella [730]. Accelerated [453, 297]. Accessory [1409, 472, 798]. Accidents [1017]. Acclimated [616, 724, 761, 1282]. Acclimation [1265, 602, 927]. Account [11]. Accumulation [1093, 1184]. Accuracy [1192]. Acetate [186]. Acetylated [902, 1022]. Acetylcholinesterase [1078]. Acid [218, 335, 628, 707, 1009, 844, 1107, 1465, 1202, 1098, 1159, 1178, 456, 935, 679, 762, 501, 231, 119, 1458, 236, 783, 856, 10, 395, 1486, 896, 107, 82, 1160, 391, 1028, 1283, 276, 182]. Acid-Fast [501]. Acid-Pyrimidine [231]. Acid-soluble [182]. Acidosis [1021]. Acids [187, 218, 271, 367, 362, 378, 405, 412, 359, 858, 1164, 1204, 1415, 1538, 619, 1105, 488, 618, 1228, 576]. Acipenser [1248]. Acoustically [1224]. Acrocirridae [1527]. Across [973, 1440]. ACTH [1441]. Action [246]. Actions [68]. Activities [93, 146, 109]. Activity [546, 23, 634, 1443, 34, 172, 687, 1576, 455, 1129, 1473, 1516, 1001, 656, 818, 1008, 1587, 1282, 486, 559, 349]. aculeatus [1313]. Acute [793]. acutorostrata [330]. acutus [1591]. Adaptation [112, 469]. Adapter [1032]. Adaptive [275]. Addendum [1073]. Addition [605]. Additional [147, 1335, 249]. Adenosine [944, 212, 416]. Adipose [820, 224]. Adjacent [1541, 525, 1304, 1481, 873, 225]. Adjustment [1148, 765, 510]. Adrenal [668]. Adult [812, 1566, 156, 328, 1101, 237, 810, 1379, 1441, 1209, 1004, 1211, 1431, 729, 1471, 874, 1011, 948, 912, 911, 1515, 1372]. adulte [1372]. Advanced [237]. Advent [660]. Adverse [1181]. aeglefinus [756, 1574, 4]. aeneus [1430]. aequispina [171]. Aerial [106, 683]. Aeromonas [1155, 1499, 1330, 754]. aestuarina [1567]. Affected [1451]. Affecting [816, 1391, 1320, 536, 153]. affinis [1505, 1225]. Affinity [760]. Africa [243]. After [695, 811, 1060, 155, 753, 868, 712, 1086, 710, 1538, 1265, 1379, 1571].

Against [539, 1275]. agassizi [1322]. Age [1388, 1331, 1382, 136, 65, 520, 1574, 963, 228, 1564, 84, 1130, 1336, 1334, 79, 1081, 1491, 1543, 895, 1153, 1248, 899, 372, 906, 1501, 1187]. Age-Groups [1564, 963]. Aged [1202, 1580]. Ageing [1192]. Agent [539, 1399]. Agents [1486]. Ages [910]. aggregata [1501]. Aggregations [1426, 637]. Aggression [1226]. Aggressive [217]. Aging [1538, 1108]. Agonidae [1411]. Agonistic [1020]. Agricultural [1337]. Aid [86]. Ainslie [717]. Air [200, 236, 413]. Air-Lift [413]. Aklavik [87]. Alaska [742, 820, 380, 1247, 1061, 417, 424, 880, 625, 1029, 1313, 18, 926]. alaskensis [428]. alata [384, 548]. albacares [861]. Albacore [144]. Albatross [167]. Alberta [1307, 540, 1452]. Alcohol [929, 1090, 881]. Alcohols [135]. Aldosterone [1056]. Aldrin [1184]. Alectridium [1019, 804]. Alepocephalidae [579]. Aleutian [1456, 585]. aleuticus [73]. Alewife [907, 913]. Alexander [607]. Alga [68]. Algae [807, 734, 601]. Algal [1463, 1544, 1201]. Algonquin [1359, 38, 567, 889]. Aliasing [953]. Alimentary [507]. Aliphatic [479]. Alkaline [626]. Allocation [1540]. Allometric [1285]. Along [1345, 1305, 376, 528, 601, 646]. alosoides [20, 1249]. alpenae [214]. Alpha [1574]. Alpha-Glycerophosphate [1574]. alpinus [496, 129, 562, 763]. Alsea [885]. Alterations [1465, 34]. Alternative [379]. altivelis [1238]. America [1121, 129, 689, 591, 426, 1106, 1335, 1348, 489, 646]. American [1051, 1052, 1300, 1469, 1252, 771, 1096, 319, 113, 194, 1522, 1170, 233, 1209, 1163, 312, 387, 671, 895, 1389, 526, 1383, 1143, 1370, 595, 1268, 466]. americanus [1549, 1469, 69, 98, 768, 1218, 746, 1083, 269, 513, 1170, 565, 398, 574, 1257, 1055, 313, 1213, 473, 738, 992, 1181, 1275, 1514, 1401, 1268, 1132]. Amino [935, 619, 1105, 231, 119, 10, 391, 276]. Aminohydrolase [944]. Aminotransferases [1127]. Ammocoete [911]. Ammocoetes [21]. Ammodytes [591, 1182]. Ammonia [921]. Amoebophrya [1216]. Among [1234, 538, 1281, 726, 1193, 1198, 1226]. Amount [779]. Amounts [292]. Ampeliscid [833]. Amphibians [1350]. Amphipod [833, 823]. Amphipoda [1106, 1497, 574, 1335]. Amphipods [862, 1310, 1390, 1166, 225]. Anadromous [36, 469, 56]. Anaesthetic [326]. Anal [1092, 716]. Analog [743]. Analogy [7]. Analyses [858, 1522, 405]. Analysing [223]. Analysis [412, 1168, 1142, 661, 683, 1066, 1085, 1579]. Analytical [1462]. Anarhichas [819]. Anatomical [454]. Anatomy [1454]. Anchor [336, 42]. Anchored [1122]. Anchovy [1327]. Ancient [530]. Anemia [1283]. Anesthetic [933]. Angler [1119]. Angling [392, 52, 258]. Anguilla [952, 1332, 1209, 1163, 466]. Animal [1074, 301]. Animals [162, 69, 98, 178, 407, 896, 42, 70]. Anisakis [857]. Annual [122, 617, 812, 985, 794, 206]. annulata [960]. Annulus [742, 1207, 680]. Anodonta [860, 1258]. Anomalies [495]. Anomalodesmata [1484]. Anomalous [937, 5, 363, 218]. Anomura [1058, 366]. Anonyx [574, 1106, 1335]. Anoplogaster [638]. Anoplogasteridae [638]. Anoplopoma [1095, 228]. Anoxia [951]. Anteiso [858]. Antibiotics [188, 487, 1013, 875]. Antibody [795]. Anticholinesterase [1478]. Antigens [754]. antillarum [1533]. Antimora [100]. Aorta [1508]. Aortic [1442]. Aperture [898, 393]. Aphanopus [253]. Aporocotyle [947]. Aporocotylidae [947]. Apostle [1141]. Apparatus [241, 259, 909, 799, 368]. Apparent [121, 416]. Application [1496, 1295, 490, 1555, 283, 167, 223, 493, 471]. Applications [481]. Applied [1526]. Appraisal [1458]. Appreciation [1344]. Approach [393]. Approaches [1462, 151]. Appropriate [657]. Apron [152]. Aquaria [1020]. Aquarium [1277, 280, 24, 53, 731, 301]. Aquatic [69, 98, 178, 407, 1526, 872, 870, 996, 867, 1380, 731, 276]. Aqueous [186]. Aquiferous [1064]. Arctic [175, 496, 1150, 1493, 122, 1396, 127, 170, 525, 319, 1305, 958, 1247, 1221, 1476, 1447, 1106, 1335, 464, 7]. Arctic-Subarctic [127]. arcticus [128, 368]. Arcto [815]. Arcto-Norwegian [815]. Area [1345, 65, 19, 957, 606, 895, 374, 174, 204, 499, 770, 466, 363]. Areas [1385, 51, 387, 671, 505]. arenaria [22]. Argentina [722, 1360]. Argentine [722, 1360]. Argentinid [147]. Argis [377, 483]. Ariomma [1546]. Arlis [1305]. Arm [384]. Arrangement [308]. arrniger [1591]. artedii [1452]. Artedius [331]. arthrosis [1370]. Articles [337, 356, 375, 397, 420, 452, 475, 503, 533, 557, 575, 592, 614, 630, 641, 653, 669, 686, 702, 718, 735, 752, 769, 786, 802, 824, 845, 866, 882, 903, 923, 940, 956, 970, 984, 1002, 1038, 1057, 1077, 1097, 1115, 1134, 1156, 1180, 1199, 1219, 1243, 1298, 1324, 1342, 1404, 1432, 1460, 1495, 1523, 1547, 1570]. Artificial [272, 1277, 1207, 89, 1301, 1257, 946, 543]. Artificially [897, 349]. Ascanius [722]. Ascarid [64]. Ascending [360]. Ascidians [464]. ASDIC [570]. Ash [734]. Aspects [1407, 255, 988, 580, 1454, 1354, 1299, 1390, 993]. asper [1196]. Assemblages [1142]. Assessed [647]. Assessment [33, 458, 712, 427]. Assisted [390]. Associated [1234, 484, 955, 403, 163, 874, 514, 1338]. Association [374]. Associations [1255]. Asteroidea [773]. asteroides [587]. Astyanax [460, 386]. Asymmetrical [371]. Asymptotic [1133]. Atlantic [1034, 376, 528, 957, 655, 707, 1009, 834, 94, 568, 1, 281, 299, 333, 352, 643, 624, 1098, 901, 1262, 760, 724, 527, 1467, 990, 1120, 1376, 103, 102, 1279, 556, 1448, 693, 1372, 1423, 722, 137, 708, 800, 210, 1222, 1535, 322, 1131, 1314, 727, 247, 927, 1018, 888, 1189, 1405, 705, 190, 871, 84, 345, 809, 593, 1112, 925, 1416, 1568, 839, 1522, 586, 117, 118, 930, 269, 306, 419, 418, 822, 1377, 728, 1016, 1446, 1251, 778, 1211, 1431, 765, 522, 1282, 1468, 317, 512, 1572, 1521, 1362, 312, 895, 86, 819, 282, 332, 288, 1485]. Atlantic [32, 260, 342, 510, 529, 611, 780, 1011, 806, 1337, 286, 587, 1103, 1110, 1360, 330, 519, 341, 366, 477, 1106, 1335, 738, 1226, 1444, 2, 253, 723, 676, 1102, 181, 307, 321, 1582, 365, 1047, 1584, 1230, 1273, 1580, 1254]. atlanticus [332]. Atlantique [1372, 593]. Atoll [1366]. atratulus [1561]. Atresia [291]. atrum [638]. Attack [829]. Attempt [52]. Attempted [44]. Attraction [1144]. Audiogram [990]. Auditory [1376]. August [104]. aurantiacum [1019]. auranticum [804]. auratus [719, 602]. aurora [1238]. Australian [1349, 1375]. Automatically [308]. Autoradiography [1233]. Autumn [1446]. Autumn-Spawning [1446]. Autumnal [1026]. autumnalis [625]. Auxiliary [598]. Availability [1473]. Average [1564, 1540, 1190]. Avoidance [388, 918]. Axial [1294]. Ayres [467, 468]. Azeotrope [1449].

B [667, 694]. Babine [835, 972, 1543]. Background [1120]. Backscatter [1074]. Bacteria [733, 1526, 501, 173, 1577, 1573, 391]. Bacterial [817, 455, 1544, 276, 1514, 1284]. Bacterial-Toxin [817]. Bacterially [43]. Bactericidal [902]. Bacteriological [211]. Bacteriophages [1330]. Bacteriovorous [1304]. Baffin [1287, 193, 890, 891, 1413, 1414, 1010, 823, 1326]. Bag [995]. baie [593]. bairdii [451]. Bait [906]. Baja [1378]. Baker [500]. Balaenoptera [330]. Balance [572, 25]. balani [1121]. balanoides [1252]. Balanus [1252]. balteatus [1136, 91, 298]. Band [1376]. Bank [1053, 691, 329, 387, 671, 1389, 595]. Bankia [1099]. Banks [411, 1387, 357, 289, 1182]. Barbados [861]. Barnacle [1262]. Barriers [1548, 1275, 1512]. Barrow [1452]. Based [444, 674, 238, 623]. Bases [107]. Basin [344, 170, 1305]. Basins [95, 1445]. Bass [812, 1024, 701, 1277, 1474, 213, 846]. Batfish [332]. Bathymaster [585]. Bathymetric [773, 600]. Bathypterois [676]. Bathythermograph [898, 393]. Bay [55, 1185, 1507, 976, 987, 1140, 1293, 193, 256, 1266, 636, 1006, 527, 15, 80, 1524, 63, 1396, 635, 633, 648, 59, 1555, 1548, 1477, 1125, 813, 14, 92, 1389, 296, 1116, 1213, 483, 959, 1447, 318, 1355, 823, 61, 74, 652, 29, 205, 207, 1512]. Beaches [1061]. Beagle [35]. Beaked [1003]. Bean [1053]. Bear [704]. Beaufort [1447]. Become [453]. Bedrock [987]. Beds [457]. Before [155]. Behavior [217, 811, 1020, 1095, 280, 571, 974, 1084, 388, 661, 386, 1061, 1583, 512, 1198, 737]. Behavioral [491]. Behaviour [54, 1328, 1208, 835, 51, 1466, 717, 161, 1310, 1497, 1337, 1348, 1069]. Bella [606, 1353, 991, 1087]. Belle [1273]. Belly [1381, 715]. belonging [21]. Belwood [1126]. Benedict [171]. Benthic [1303, 1185, 1255, 807, 1172, 1280, 1304, 1123, 892]. benzoquinone [1562]. Bering [1111]. Bertalanffy [631, 1496, 1563, 1291]. Betaine [69, 98, 407]. Between [1136, 299, 103, 847, 820, 691, 210, 91, 1416, 194, 289, 213, 1027, 692, 650, 1011, 1079, 899, 1293, 561, 1202, 1252, 1427, 839, 38, 1291, 167]. Beverton [126]. Bias [1325]. bicuspidatus [1418]. Bideford [551, 1268, 1338]. Big [414, 251, 1565]. Bigmouth [324]. bilineata [467, 468]. Binding [695, 905, 1167]. binoculata [414]. Binomen [1371]. Bioassay [677]. Biochemical [40, 1552, 125, 8, 10, 24, 105, 493, 623, 1227, 26]. Biological [688, 818, 775, 647, 890, 1362, 1360, 308, 263, 29]. Biology [581, 1369, 722, 1043, 443, 262, 1354, 598, 833, 78, 1010, 1572]. Bioluminescence [862]. Bioluminescenee [850]. Biomass [1305, 1502, 565]. Biomicroscopic [542]. Biosynthesis [69, 98, 178, 407, 1056]. Biota [1571]. bison [1091]. Bivalve [1544].

Bivalves [1296]. Bivalvia [1484]. Black [1581, 1569, 1043, 253]. Blackberry [374]. Blacknose [1561]. Blackspotted [1319, 1466, 1261]. Blackstripe [739]. Blacktail [305]. Bladder [1440, 1357]. Bleached [1562, 1451]. Bleachery [690]. Blended [1080]. Blocks [348]. Blood [1159, 687, 616, 724, 761, 768, 865, 836, 1316, 1060, 285, 668, 1245, 189, 560, 705, 586, 1402, 402, 744, 980, 493, 543, 911, 70]. blooded [347]. Bloom [660, 730]. Blowing [200, 236]. Blubber [247]. Blue [417, 419, 893, 204]. Bluefin [966]. Bluegill [794, 1078, 1048, 1439]. Bluespot [837]. Board [75, 76, 148, 229, 338, 337, 474, 476, 356, 375, 397, 420, 452, 475, 503, 533, 557, 575, 592, 614, 630, 641, 653, 669, 686, 702, 718, 735, 752, 769, 786, 802, 824, 845, 866, 882, 903, 923, 940, 956, 970, 984, 1002, 1179, 1038, 1057, 1077, 1097, 1115, 1134, 1156, 1180, 1199, 1219, 1243, 1494, 1542, 11]. Bocaccio [801]. Bodies [299, 333, 560]. Body [352, 1504, 322, 790, 692, 726, 1439, 1525, 167]. Bonavista [520]. bondi [1546]. Bonito [552]. Bonnaterre [306]. Border [789]. borealis [844, 1075]. Borer [1099, 829]. Boring [1409]. Borneo [1350]. both [1415]. Bothrimonus [1369]. Bothrocara [639]. Bottle [63, 511]. Bottlenose [247]. Bottom [627, 1336, 957, 47]. botulinum [1176, 954, 854, 1065]. Bound [242]. Boundary [1263, 883]. Bounty [1273]. Brachydanio [710, 910]. Bradyidius [689]. Brain [347, 1078, 1154]. Braincon [781]. Branched [858]. Branding [1545]. Bras [64]. Breeding [1461, 49, 1014, 466]. Breton [357]. brevis [1427]. brevispinis [415]. Brief [892]. Brines [343, 173]. Brining [942]. Brisaster [916]. Bristol [1125]. Britain [281, 299, 333, 352]. britannicus [723]. British [1037, 700, 1353, 989, 776, 160, 1023, 149, 434, 1434, 151, 294, 438, 1175, 1506, 536, 552, 384, 1418, 972, 981, 1543, 1392, 161, 447, 435, 729, 1278, 1436, 826, 431, 740, 1087, 309, 201, 1248, 62, 448, 791, 975, 1059, 176, 445, 547, 144, 1500, 147, 287, 588, 1296, 1581, 1519, 171, 459, 373, 801, 807, 104, 305, 1297, 192, 239, 206, 381, 91, 1045, 1560, 100, 1567, 244, 608, 82, 583, 423, 606, 640, 732, 138, 9, 168, 440, 331, 1196, 372, 1072, 197, 843, 977, 415, 639, 1238, 396, 922, 663]. Brito [1225]. Brittlestars [600]. Broad [194, 1491]. Brood [431, 606]. Brood-Year [431, 606]. Brook [1256, 1410, 616, 1054, 1513, 720, 708, 800, 291, 290, 1442, 1412, 190, 270, 1433, 1231, 1183, 38, 1253, 924, 1583, 1509, 938, 32, 510, 1337, 258, 946, 1039, 1067, 670]. Brooks [380]. Brosme [1315]. Brown [1129, 1473, 1092, 280, 1590, 1244, 115, 349]. Brown-water [115]. Browning [541]. brunneum [639]. Brunswick [67, 869, 872, 870, 267, 190, 871, 868, 867, 987, 1374, 12, 153, 806, 1302, 52]. Bryozoa [1221, 960]. Bryozoan [968]. bubalus [1554]. Bubble [876]. Budget [123]. Buffalo [324, 1091]. Buffalofishes [1554]. Buffer [473]. Buffering [285]. Buoyancy [1467, 1148, 765, 1468, 510]. buoyant [20]. Buoyed [1122]. Buoys [537]. Burbot [262, 1242]. burchami [588]. Burke [699]. Burrard [1506]. Burrowing [1310, 1390, 42]. Burrows [1268]. Butter [1175, 1111]. byrkelange [204].

C [412, 1538, 817, 69, 1184, 830, 1528, 282, 28, 1212]. C-Aldrin [1184].

C-Dieldrin [1184]. C-Labelled [1212]. C. [451, 1027, 1452, 1075]. côte [1372, 593]. Caeca [889, 789]. Caecum [1099]. caerulea [202, 400]. Caesium [157]. Caged [868]. Calanoid [851, 1289, 826]. Calanoida [1391, 689]. calanoides [1524]. Calanus [80, 127, 193, 959]. Calcium [461]. Calculating [167]. Calculation [1530]. Calculations [1000]. Calibration [314, 781]. California [973, 354, 937, 1003, 1394, 1378]. californianus [354, 937]. californicus [1394]. californiensis [860, 1258]. Caligidae [423, 1191]. Caligoida [1405, 423]. Caligus [423, 1191, 1371]. callarias [659]. Callorhinus [408]. Calocaris [477]. Caloric [1185, 1458, 1502]. Calorific [1239]. Calycopsis [549]. Cambodian [89]. Cambridge [1447]. Cameron [1344]. cameroni [1368]. camtschatica [171, 1014, 1539, 485]. Canada [75, 337, 474, 356, 375, 397, 420, 452, 475, 503, 533, 557, 575, 592, 614, 630, 641, 653, 669, 686, 702, 718, 735, 752, 769, 786, 802, 824, 845, 866, 882, 903, 923, 940, 956, 970, 984, 1002, 1179, 1038, 1057, 1077, 1097, 1115, 1134, 1156, 1180, 1199, 1219, 1243, 1494, 1493, 691, 1359, 221, 500, 1247, 392, 252, 1587, 60, 839, 953, 451, 625, 1402, 826, 1221, 64, 738, 5, 663, 1408, 76, 148, 229, 338, 476]. canadense [1197]. Canadian [175, 1009, 568, 901, 1150, 122, 1034, 137, 210, 127, 170, 525, 1411, 1535, 376, 528, 1405, 56, 120, 957, 1585, 117, 118, 419, 418, 203, 530, 1377, 332, 355, 529, 587, 1476, 253, 321, 464, 39]. canadien [1150]. Canary [1089]. Cancer [49, 136, 1088, 1311, 737, 1075]. Candling [555]. Canned [695, 905, 936, 1339]. Canning [566]. Canthaxanthin [1309]. Capacity [285, 491, 457]. Cape [357, 1310]. Capelin [271, 1465, 1464, 716, 1492, 652]. Capello [1225]. Capilano [62]. capillata [1160]. Captive [54]. Captivity [209, 766, 544]. Caranx [1546]. Carassius [719, 602]. carbo [253]. Carbohydrate [887, 943, 1073, 654]. Carbon [1237, 687, 1024, 572, 736, 1502, 1424]. Carbon-14 [1237]. Carbonate [461]. Carbonyl [479, 480]. Carcharhinus [307]. carcharias [168]. Carcharodon [168]. Cardiovascular [948]. Cards [898, 393]. Careproctus [305]. Caridea [1461]. Caridean [226]. carlottensis [1152]. Carp [1508, 388, 808, 1264, 599]. carpio [1508, 808]. Carrier [1410, 1049]. Cartilaginous [668, 818]. Cases [414]. Catalyzed [453]. cataractae [1561, 1148]. Catch [520, 492, 232]. Catchabilities [235]. Catches [509, 13, 713, 3, 1039, 1059, 904]. Catecholamine [943, 1073]. catenella [1216]. Catfish [1021, 1078]. Catheterization [1442]. Cations [1314]. Catostomid [1358, 776]. Catostomidae [1554, 1027, 832]. Catostomus [1388, 154, 495, 1027]. Caught [103, 102, 484, 1018, 1194, 590, 458, 859, 1086, 1431, 756]. Caulolatilus [1394]. caurinum [1136]. caurinus [747]. Causative [1399]. Cause [679, 217]. Caused [762]. Causes [264, 251]. cavirostris [1003]. Cavities [1358, 1560]. Cedar [577]. celestis [1370]. Cell [350]. Cells [664, 1320, 1397, 436]. celui [593]. Census [683]. Central [1353, 1087, 700, 580, 1569, 438, 740]. Centrarchidae [772]. Centrifuged [767]. Centrolophidae [723]. Centrolophus [723]. centroscyllii [399]. Cephalopoda [728, 288]. ceratii [1216]. Certain [1009, 1500, 501, 567, 1457, 899]. cervinum [822]. Cessation [1236]. Cestoda [1369, 1359]. Cestodes [1366, 1372]. Cetacean [1246]. Chaetognatha [1435]. Chain [218, 412, 1443]. Chaleurs [593]. Chalkiness [535, 672, 697]. challengeri [1033]. Chamber [1017]. Change [158, 678, 1282]. Changes [1469, 172, 687, 360, 1556, 1105, 985, 1508, 1552, 484, 955, 206, 856, 855, 1118, 859, 1379, 924, 548, 361, 523, 729, 1436, 1528, 740, 1468, 1383, 201, 510, 514, 1293, 1028, 1571, 1147, 142, 1355, 263, 105, 1227, 236]. Changing [1084, 765]. Channel [1021, 1078, 1590, 445]. Channels [491]. Char [496, 919, 562]. Characterisation [281]. Characteristics [879, 700, 133, 6, 688, 1247, 1489, 891, 435, 1531, 309, 926, 463, 1075, 1399, 27, 150, 646]. Characterization [347, 1330]. Characters [444, 651, 451]. Charlotte [96, 975, 67, 13]. Chatham [1481]. Checks [603]. Chemical [162, 133, 771, 1557, 1247, 1391, 458, 647, 891, 1533, 883, 740, 139, 150]. Chemically [1065]. Chemotherapeutic [501]. chez [1197, 1329, 1367]. Chignik [1029, 1313]. Chill [778]. Chill-Stored [778]. Chilled [1259]. Chilling [1518]. Chimaerae [137]. Chinook [1234, 820, 1145, 1192, 1060, 1101, 16, 201, 1198]. Chionoecetes [1170, 1171]. Chlamydomonas [68]. Chloride [564]. Chlorination [1562]. Chlorine [45]. Chlorophthalmidae [1053, 1322]. Chlorophthalmus [1322]. Chlorophyll [1063]. Chlorophylls [1420]. Chlortetracycline [1178]. Choice [1101]. Cholesterol [8, 70, 109]. Choline [178]. Chondracanthidae [1045, 1560]. Chondracanthus [1560]. Chondrostei [137]. Chorioidal [437]. Chromatographic [858]. Chromatography [367, 362, 378, 558, 929, 955, 236, 471, 479]. Chromosome [1114]. Chromosomes [978, 682]. Chronic [1510]. Chrosomus [1341]. Chrysomonad [817]. chrysops [1474]. Chub [1136, 87]. Chum [603, 742, 1108, 1382, 112, 1334, 1421, 425, 637, 1072, 396]. Cichlidae [243]. Ciguatoxin [1478]. Ciliate [1177]. ciliatus [1238]. cincinnatus [848, 1363]. Circulating [343]. Circulation [743, 15, 1026, 704]. Circuli [680]. Cisco [214]. Clam [1165]. Clams [22, 1232, 203]. clarki [1023, 1023, 753, 1385, 971]. Class [678, 521]. Classes [582, 893]. Claws [108]. Clearance [1312]. clemensi [423, 426]. climat [593]. climate [593]. Clines [1317]. Clinical [1515]. Clinocardium [1165]. clipping [1011]. Close [1092]. Closely [366]. Clostridium [1176, 954, 854, 1065]. Clotting [586]. Club [436]. Clupea [994, 55, 58, 57, 191, 1449, 1448, 209, 210, 78, 197, 74, 300, 1056]. clupeaformis [1517, 48, 1205, 763, 254]. coarctatus [1075]. Coast [144, 1241, 1121, 691, 192, 688, 294, 839, 1280, 689, 9, 1572, 591, 426, 1059, 5, 601, 646, 593, 773, 600, 849, 1372, 1034, 376, 957, 269, 530, 364, 138, 66, 873]. Coastal [1345, 281, 299, 333, 352, 985, 414, 1250, 1522, 423, 1039, 25]. Coastrange [73]. Coasts [104, 953, 1106, 1335]. Coating [902]. Cockle [1165]. Cod [187, 218, 367, 362, 378, 655, 628, 842, 1204, 1202, 1538, 1098, 764, 1007, 88, 709, 1328, 788, 990, 1120, 1376, 659, 133, 103, 102, 741, 1104, 1498, 545, 684, 131, 132, 292, 401, 1105, 693, 944, 385, 346, 65, 520, 340, 642, 125, 484, 696, 955, 231, 815, 1577, 119, 727, 200, 236, 618, 783, 856, 855, 376, 705, 100, 120, 434, 345, 1387, 925, 382, 458, 403, 348, 516, 712, 814, 859, 1086, 586, 357, 371, 930, 1251, 478, 1157, 1528, 750, 789, 1066, 1044, 1572, 86, 216, 260, 374, 1147, 174, 176, 303, 278]. Cod [365, 220, 1273, 1580, 159, 1589, 479, 480, 534, 1310]. Codfish [1]. Codroy [1211]. Codworm [35, 1273]. Codworms [86]. Coefficient [486, 559]. Coefficients [1063]. cognatus [451]. Coho [268, 711, 1094, 217, 820, 1024, 1566, 632, 571, 1381, 24, 715, 491, 1281, 1419, 201, 1028, 1283, 111, 1227, 900, 922]. Cold [347, 344, 1486, 1545]. Cold-blooded [347]. Cold-Branding [1545]. Cold-Stored [1486]. Coliform [173]. Collagen [71]. Collected [600, 873, 1367]. Collecting [1429, 42]. Collections [959]. Cololabis [335]. Color [1485]. Colour [936, 801, 1336]. Colouration [982]. Coloured [325]. Columbia [826, 138, 663, 144, 1500, 147, 287, 588, 1037, 1296, 700, 1581, 1519, 1353, 171, 459, 373, 801, 807, 989, 104, 305, 1297, 192, 239, 776, 206, 381, 160, 1023, 149, 91, 1045, 1560, 100, 434, 1434, 151, 294, 438, 1175, 1506, 1567, 536, 552, 244, 608, 82, 583, 384, 1418, 972, 981, 1543, 1392, 161, 447, 435, 729, 1278, 1436, 431, 423, 606, 740, 1087, 640, 732, 309, 9, 168, 201, 440, 331, 1248, 1196, 62, 372, 448, 791, 1072, 197, 843, 975, 1059, 176, 445, 547, 977, 415, 639, 1238, 396, 922]. Columns [858]. Combined [749]. Comments [1406, 1335, 873]. Commercial [707, 1107, 520, 516, 647, 678, 893]. Commercially [1577]. commersoni [495]. commersonii [1027]. Common [1371, 1005, 302]. Communication [818]. Communities [1380, 1406]. Community [1463, 1406]. Comparaison [593]. Comparative [1107, 951, 507, 636, 1137, 1178, 1450, 1020, 239, 128, 1310, 791, 493, 763, 832, 772, 911, 109]. Compared [377, 1444]. Comparison [1410, 1108, 509, 1516, 160, 113, 1373, 351, 72, 184, 1291, 543, 108, 593, 126]. Comparisons [1387, 919, 723]. Compensation [794]. Competition [91, 273, 1281]. Complex [231, 293, 129, 625]. Component [271, 367, 362, 378, 412, 1562, 488]. Components [1314, 432, 1245, 1333]. Composition [218, 707, 1009, 844, 907, 1107, 1465, 1504, 952, 935, 771, 1557, 311, 137, 119, 236, 783, 856, 158, 896, 1464, 117, 118, 244, 913, 1533, 774, 740, 139, 140, 1472, 1160, 391, 276, 150]. Compounds [69, 98, 178, 407, 1498, 40, 830, 1157, 1203, 1339, 462, 182, 508, 1076, 479, 480]. Computations [222]. Computer [1172, 1000, 828]. Computers [1295]. Computing [1260]. Comté [1186]. Concentrate [929, 1090, 1449, 1448, 881, 216]. Concentrates [463]. Concentration [836, 68, 846, 212, 853]. Concentrations [1098, 1060, 1183, 943, 1073, 341]. Concept [674, 1406]. Concerning [182]. Conchoderma [1262]. Conchoecia [384, 548]. Condition [695, 573, 855, 1086, 364]. Conditioning [621]. Conditions [162, 1541, 1105, 691, 1207, 725, 1457, 1420, 1181, 90, 363]. Conductivity [41]. connue [1372]. Conrad [354]. Consecutive [1423]. Consequences [1442]. Consideration [1479]. Considerations [215].

Constancy [494]. Constant [1563, 53]. Constituents [134, 705]. Constitution [1511]. Constructing [379]. Consumed [609]. Consumption [604, 1551, 757, 755, 1170, 1282]. containing [28]. Contamination [590]. Content [1185, 762, 1556, 1458, 855, 196, 71, 27, 142, 827, 26]. Contents [1279, 1151, 417, 622, 1481, 1004]. Contiguous [1229]. Continental [93, 355, 1329, 1329, 1255, 957]. Continua [1406]. Continuous [884]. Contributed [15]. Contribution [450, 78, 1235]. Contributions [893, 1302]. Control [679, 1012, 1290, 1200]. Controlled [1068]. Controlled-Environment [1068]. Controlling [308]. Convenient [389]. Conversion [187]. Cooked [825, 1485]. Cooking [1, 385]. Coola [1353, 991, 606, 1087]. Cooling [297]. Copepod [1296, 633, 648, 423, 1371, 658, 1326]. Copepoda [1391, 399, 848, 1045, 1560, 1306, 1363, 1559, 1405, 689, 1191, 1370]. Copepods [851, 1524, 127, 1289, 826, 594]. Copper [1549, 453, 1510, 341, 518]. Copper-Catalyzed [453]. Coppermine [1491]. Core [1174]. Coregonidae [961]. Coregonids [1207]. Coregonine [1168]. Coregonus [1517, 48, 194, 625, 1491, 1452, 1205, 763, 254]. corinus [553]. Cornea [1100]. Corneal [542]. cornuta [638]. Corpuscles [1332, 727]. Correct [1019]. Corrections [75, 474, 1179, 1494]. Correlated [386]. Correlation [94, 1173, 1197]. Correlations [1031]. Corticosterone [1056]. Cortisol [1025, 1437]. Coryphaena [861, 181]. Coryphopterus [837]. costatum [405]. cotio [1368]. Cottid [1351, 331]. Cottidae [1323]. Cotton [219]. Cottus [1270, 451, 73, 1196]. coulteri [583]. Counting [1237, 1094]. Counts [106, 1577, 101]. County [67, 1186, 13, 1536]. Cover [706]. cowani [1015]. Crab [49, 136, 171, 1088, 1170, 1171, 1539, 1311, 737]. Crabs [1014, 678, 485]. Cranchiidae [728]. Cranial [915, 1315]. Crappie [1043]. Crassostrea [978, 573, 664, 839, 793]. crassus [1493, 555, 880]. Crayfish [771]. Creatine [40, 1263, 1264]. Creek [1412, 72, 315, 404, 431]. Cribrilina [960]. Cribrilinidae [960]. Cribrilinidae-Bryozoa [960]. cristata [1228]. Cristivomer [919]. Critical [1376, 1528, 16]. croissance [1187]. Crop [1185, 1280]. Crouch [852]. crumenopthalmus [1546]. Crustacea [178, 193, 959, 1046, 1447, 1109, 1391, 1310, 1497, 1390, 574, 377, 483, 477, 1335]. Crustacean [736, 833, 1472]. Crustaceans [1233, 567, 1135, 261]. Cryo [1403]. Cryo-Preserved [1403]. Cryogenic [1251]. Cryoprotective [1486]. Cryptobia [796]. crysos [1546]. Culaea [1509]. Cultivated [1320]. Cultivation [664]. Culture [1289, 1397, 1311, 274]. Cultured [121]. Cultures [1420]. Current [497, 781, 59, 973, 1305]. curtus [1191]. Curve [1563, 524]. Curves [631, 616, 724, 761, 781, 379, 1291]. Cusk [822]. Cusk-eel [822]. Cuspidaria [1015]. cuspidata [1001]. Cutthroat [1023, 753, 1385, 971]. Cutting [241]. Cuvier [1003]. Cyanea [1160]. cyanellus [597]. Cycle [122, 408, 1364, 794, 290, 1490, 1390, 25]. Cycles [1469, 1524, 386, 349, 770]. Cyclic [206]. Cyclohexanetetrol [1556]. Cyclopecten [1152]. Cyclopid [594]. Cyclopoida [1370]. Cyclops [1418].

Cyclopterid [1586]. Cyema [638]. Cyemidae [638]. cylindraceum [1081]. Cymatogaster [1501]. cynoglossus [610]. cyprinellus [1554, 324]. Cyprinid [1550]. Cyprinodontids [1265]. Cyprinodontiform [1316]. Cyprinus [1508, 808]. Cystic [1580]. Cystidicola [1357]. Cystophora [1228]. Cysts [555]. Cytological [1379]. Cytotaxonomic [1168].

Dace [1561, 239, 1148, 1341]. Daily [1118]. Damage [1431]. Dana [49, 136, 737]. Dangeard [68]. dans [1150, 1187]. Daphnia [1237, 361]. Dart [1215]. Dartmoor [1473, 629]. Dasyatis [1110]. Data [631, 1172, 734, 524, 393, 167]. d’autres [593]. Day [1118]. DDT [1256, 1576, 1279, 1034, 869, 1222, 870, 871, 1231, 1183, 522, 282]. deani [1581]. Death [544]. Decapod [844, 193, 959, 1046, 1447]. Decapoda [477, 1461, 373, 1058, 721, 226, 377, 366, 483]. Decay [99]. decipiens [1364, 231, 64, 278]. Deck [713]. Decomposition [618, 1424]. Deep [1519, 1428, 1267]. Deep-Sea [1519, 1428]. Defense [1083]. Deferment [52]. Deficient [1028]. Defined [1065]. Degradation [1279, 787, 1417, 955, 1040, 1222, 1161, 609, 1158]. Dehydrocholesterol [44]. Dehydrogenase [1137, 1576, 1561, 1516, 1574, 1276]. Delay [1218]. Delayed [1539]. Delphinapterus [1525]. Delta [1153]. DeLury [1308]. Demersal [1462, 624, 1142, 467, 303]. Denaturation [535]. Denatured [242]. Densities [53]. Density [704]. dentata [377, 483]. Deoxynucleosides [107]. Deoxyribonucleic [107]. Deoxyribose [107]. Dependence [617]. Depensatory [674]. d’éperlans [1186]. Deposition [655, 1333, 1286]. Deposits [551]. Depot [1228]. Deprived [1226]. Depth [721]. Depths [1000, 896, 912]. Derivation [1415, 1564]. Derivatives [1063]. Derived [524, 576]. Dermocystidium [1234, 874]. Desalting [132]. Described [126]. Describing [1190]. Description [20, 1369, 579, 667, 1058, 419, 1191, 591, 21, 418]. Descriptions [521]. Desiccation [821]. Design [745, 1068]. Detect [1566, 1397]. Detecting [1410]. Detection [555, 130, 86]. Determination [558, 929, 1331, 1382, 136, 1025, 755, 1334, 79, 677]. Determined [367, 362, 378, 1233, 1474, 236]. Determining [607, 920, 142]. Detoxification [282]. Developed [667]. Developing [202]. Development [1062, 1080, 1498, 1096, 340, 468, 642, 720, 889, 237, 808, 540, 716, 611, 562, 1348, 827, 368]. Developmental [20, 758, 1391, 265, 227]. Device [1566]. Devices [1455]. Diagnoses [1484]. Diagnostic [1127]. diaphanus [1265]. Diatom [405]. Dibranchus [332]. Dichromate [1424]. Diclidophorinean [1373]. Diel [656, 1321, 729, 1196]. Dieldrin [1453, 1184]. Dielectric [814]. Diet [169, 1028]. Dietary [1118, 81, 109]. Diets [1438]. Differences [897, 1252, 989, 137, 1416, 839, 561, 238]. Different [1564, 1421, 117, 1163, 793, 910, 143]. Differentiated [400]. Diffuser [445]. Digenea [1351, 947]. Digenetic [1356]. Digestion [788, 213, 220]. Digestive [764, 1438]. Digitize [898]. Dimensional [743]. Dimensions [1292]. Dimethyl [558, 655, 645, 842, 1042, 1232, 374]. Dimethyl- [558, 655, 645, 842, 1042]. Diminutive [384]. Dimorphism [538, 1586, 1092]. Dinitrophenylhydrazones [479]. Dinoflagellate [934, 1216]. Dioxide [687, 1024]. Diphyllobothriid [1366]. Diphyllobothrium [1345, 1359]. Diplocotyle [1369]. Dipped [1147]. Direct [1224]. Disc [295]. Discards [267]. Discrimination [1470]. Discussion [1530]. Disease [121, 876, 1521, 539, 1399, 1514, 595]. Diseases [1035]. Disguised [1133]. Disinfectants [1013]. Disinfection [1499]. Dispersal [1444]. Dispersion [445]. Displacement [753]. Disposal [547]. Dissociation [616, 724, 761]. Dissolution [1210]. Dissolved [268, 1024, 1513, 461, 27, 142, 1067]. Distinctions [194]. Distinctive [1407]. Distinguishing [820, 451]. Distribution [1204, 711, 773, 144, 162, 600, 281, 333, 352, 581, 1127, 277, 784, 865, 55, 456, 459, 40, 573, 1121, 1233, 292, 1511, 1400, 579, 1384, 1535, 1307, 656, 439, 376, 528, 1061, 1244, 1433, 424, 567, 1301, 548, 939, 1377, 18, 1051, 1509, 447, 1472, 721, 440, 549, 398, 1055, 1182, 873, 197, 1190, 223, 182, 2, 823, 601, 663]. Distributional [1408]. Distributions [1524, 807, 800, 1023, 1271, 841, 723, 977]. Disturbance [943, 1073]. Diurnal [55, 548, 1067]. Diving [146, 1017, 300, 190]. Division [350]. DNA [1168]. Docosapentaenoic [335]. Dogfish [849, 164, 66]. Dollar [1459]. dolomieui [1277]. Dolphin [861, 9, 181]. Domestic [612]. dont [1372]. d’Or [64]. d’Orbigny [288]. Dorsal [365]. Dose [1184, 665]. Doses [825]. Double [712, 814, 1086]. Double-Frozen [814]. Downstream [629, 270, 51]. Drag [1208]. Draggers [267]. Drags [509]. Drainage [796, 1408]. Dredge [42]. Dried [541, 1589]. Drift [63, 1305, 511, 957]. Drift-Bottle [63]. Drifting [873]. Drills [1377]. Drip [482, 545, 1096, 382, 458, 403, 672]. Driven [743]. Dry [734, 1114]. Drying [89, 920, 925]. dubius [1182, 676]. Due [315, 678, 1548]. Dugesia [1300]. Dulse [694]. d’une [1186]. Dungeness [1311, 737]. Duration [932]. During [655, 1465, 1469, 172, 687, 1353, 849, 40, 1541, 1364, 1096, 1552, 125, 570, 280, 991, 8, 888, 10, 196, 51, 740, 184, 174, 28, 310, 1227, 1212, 26, 949, 994, 1250, 1189, 1573]. Dusky [588, 307]. dussumieri [288]. Dwarf [73]. Dwarfism [960]. Dyes [1425]. Dylox [1475]. Dynamics [635, 815, 958, 251, 1553, 1140, 948].

E. [433]. Early [1062, 20, 1353, 991, 1391, 491, 740, 227, 365, 1089]. Earthworms [1037, 922]. East [243, 1121, 1305, 530, 1004, 591, 873]. East-Coast [530]. Eastern [369, 162, 281, 299, 333, 352, 1484, 985, 1319, 699, 1384, 291, 290, 38, 451, 1402, 660, 1033, 595, 551, 1378, 1529, 1445]. Eastward [1447]. Eaten [922]. eaux [1150]. Ecdysis [1181]. Echinarachnius [1459]. Echinoderms [873]. Echinoidea [917]. Echinorhynchus [938]. Echo [570]. Echo-Sounder [570]. Ecological [979, 1479]. Ecology [1407, 581, 1126, 239, 571, 324, 790, 120, 491, 835, 1413, 526, 1553, 1310, 1390]. Economic [67]. Economics [48]. Edible [49, 136]. Editor [421]. EDTA [1531, 1203]. Edward [1241, 32, 1293, 551, 1268, 256, 1341]. Eel [1332, 1209, 1163, 355, 466, 822]. Eelpout [639]. Eels [952, 717, 779]. Effect [1256, 1, 1159, 749, 760, 887, 709, 188, 482, 1576, 952, 455, 619, 1498, 573, 295, 35, 6, 1149, 902, 1070, 1146, 1516, 605, 1312, 621, 602, 134, 1482, 704, 1486, 705, 1217, 953, 1475, 1558, 404, 522, 1531, 1044, 317, 512, 825, 32, 195, 1039, 654, 562, 183, 1518, 547, 1386, 70, 1512, 564, 382, 516]. Effective [1499]. Effectiveness [1178, 316, 1275]. Effects [951, 268, 1062, 470, 1165, 272, 976, 1437, 1513, 869, 642, 720, 872, 406, 1050, 870, 871, 56, 867, 814, 13, 810, 1534, 275, 1539, 203, 506, 478, 540, 327, 1135, 1311, 251, 1325, 1011, 1439, 1293, 1537, 90, 1284, 1067, 39, 1273, 1259, 692]. Efficiencies [767]. Efficiency [1208, 35, 202, 726, 219]. Effluent [1562, 1451, 445, 977]. Effort [183, 599]. Egg [414, 101, 1333, 1499, 457, 227, 1286, 710]. Eggs [1068, 20, 490, 467, 468, 1403, 1131, 1175, 1499, 562, 303, 284, 248, 819]. Egmont [256]. Eigenmann [287]. Eight [276]. Ekman [853]. Elasmobranch [1440, 1245]. Elasmobranchs [1352]. Electric [816]. Electrical [743, 41]. Electroencephalographic [1471]. Electrolyte [952]. Electromagnetic [205]. Electron [1549]. Electronic [403]. Electropherograms [493, 494, 763, 832, 772, 911]. Electrophoresis [698, 1235, 111]. Electrophoretic [1137, 1263, 1154, 358, 980, 543]. elegans [548, 1370]. Elevating [413]. Elevations [1482]. Eleven [1304, 139, 140]. Elizabeth [1046]. Ellerslie [32]. Ellesmere [851]. ellipticum [732]. elongatus [180, 1560, 1371, 46, 747]. Elvers [1163]. Embiotocid [979]. Embryo [1501]. Embryonic [340, 642]. Embryos [720, 202, 710]. Emergence [1429, 491, 1281, 1286]. Emigration [217]. Emphasis [1127, 1123]. Employing [107]. Endemic [1345]. Endless [360]. Endogenous [1439]. Endoparasitic [1216]. Endotoxins [1284]. Endrin [836]. Endurance [643]. Energy [241, 1555, 1157, 123, 770]. England [914, 1374, 1550, 852]. English [1062, 1480]. Engraulis [1327]. Enophrys [1091]. Enrichment [1378]. Ensilage [82]. entomelas [608]. entomon [175, 1029]. Entosphenus [21]. entre [1197]. Entry [536]. Enumeration [173]. Environment [1068, 1483, 1014, 1123, 1571, 663]. Environmental [807, 491, 1497, 1537]. Environments [339, 1301]. Enzymatic [1078]. Enzyme [1162, 107]. Enzymes [659, 1516, 82]. Enzymic [220]. eos [1341]. eperlanus [1186]. Epibenthic [1185, 1463]. Epidemics [121]. Epipelagic [18]. Epipleurals [1294]. Epithelial [789]. Epizoic [965]. Epizootic [263]. Eptatretus [1581]. Equation [1496, 283, 492]. Equatorial [1378]. Equilibrium [684, 828, 1383]. Equivalents [1502]. Ergasilus [1370]. Erie [636, 1511, 97, 101, 582, 893, 214, 1445]. erinacea [561]. Erratum [1019]. Erroneous [762]. Error [163]. Erxleben [519, 681]. Escapements [1333]. Eschrichtius [199]. esculenta [243]. Esocidae [772, 208, 584]. Esox [757, 755, 50, 472, 596, 782, 798, 72, 315, 404, 433, 1187, 1119]. espèces [1372]. Especially [649]. Essex [852]. Esterases [1154, 1520]. Esters [187, 135, 673]. Estimate [1125, 315, 446, 1132]. Estimates [325, 1288, 1325, 565]. Estimating [759, 481, 158, 1041, 1299, 283, 3, 1286, 1272]. Estimation [1295, 921, 1224, 696, 243, 1458, 1130, 1540, 1333, 126, 235, 431, 606, 829, 1424]. Estimations [192, 185]. Estimator [1308]. Estradiol [134]. Estrogen [1437]. Estrus [408]. Estuarine [1316, 780, 1338, 7, 770]. Estuary [1223, 703, 931, 29, 852]. état [1372]. États [1367, 1372]. Ethylenediaminetetraacetic [1178]. Etiologic [539]. Études [1367]. Eucalia [270]. Euchaeta [1391]. Eulachon [450]. Eumicrotremus [1586]. Euphausiacea [1109]. Euphausid [844]. Euphausiid [736]. Euphausiids [908]. Euphotic [973]. European [1300, 1252]. Euryhaline [1265]. Eurythenes [823]. Eutrophic [1026, 767]. Evaluating [453]. Evaluation [1462, 1176, 458, 1565, 244]. Evaluations [1589]. Evidence [695, 1162, 394]. Evolution [1575]. Examination [1382]. Example [1133]. Examples [251]. Exchange [547]. Excreted [921]. Excretion [1022, 1439]. Exercise [1098, 749, 1060, 654]. Exophthalmia [693]. Exotic [922]. Expedition [1350]. Expeditions [80, 193]. Experimental [360, 1435, 1026, 1590, 755, 863, 1253, 524, 163, 1457]. Experiments [655, 63, 572, 1209, 308, 301, 1378]. Exploitation [1532, 635, 1555, 1130, 273, 644, 315, 1005, 1140, 1082]. Exploited [759, 1295, 1383]. Explorations [1462, 61]. Exploratory [1462]. Exposed [1184, 706]. Expressing [1291]. Extender [1584]. Extending [410]. Extension [1456, 1319, 797, 1567, 87, 1036, 747, 1111]. Extensions [1394, 226, 498, 751]. Extent [291]. Extract [1001]. Extractability [1202, 346, 382]. Extraction [929, 944, 1157, 1420]. Extractives [292, 401]. Extracts [186, 877, 165, 473, 954]. Eyelid [224].

F [526, 433]. F. [1265]. Fabr [387]. Fabricius [312, 671]. Facilities [36]. Facility [1290]. Factor [864]. Factors [807, 816, 800, 1336, 1391, 1320, 536, 316, 153, 1497, 142, 601]. Failure [163]. Fall [612, 382, 1198]. fam [1527]. fam.nov [1405]. Families [772]. Family [584, 198, 763, 832]. Fast [501]. Fat [788, 292, 401, 432, 1228, 159, 8]. Fath [1267]. Fathead [1510]. Fatigue [756, 945, 169]. Fatty [187, 218, 271, 367, 362, 378, 405, 412, 359, 628, 707, 1009, 844, 858, 1107, 1164, 1465, 1204, 1415, 1202, 1538, 1159, 456, 135, 236, 488, 618, 783, 856, 1228, 1486, 896, 1160, 576, 395]. Fauna [1353, 221, 790, 47, 1479, 331, 892, 1355, 1338]. Fawn [822]. Features [1107, 141, 469, 883, 138, 577]. Feces [1410]. Fecundity [1461, 885, 999, 406, 329, 1488, 1550, 930, 387, 502, 195, 176, 1132, 1445, 670]. Fed [1231, 1183, 1028, 349]. Feed [35]. Feeding [655, 1042, 1328, 1576, 1473, 209, 1364, 701, 1020, 1246, 1435, 1095, 1050, 1590, 790, 1118, 1043, 1124, 766, 736, 1421, 1327, 237, 810, 1479, 199, 201, 799, 1103, 1072, 852]. Female [408, 1437, 406, 372, 1501]. ferruginea [1071]. Fertility [605, 932]. Fertilization [1403, 1131, 1200, 1571]. Fertilized [467, 1175, 303]. Fibers [241]. Fibrous [1580]. Field [816, 1419]. Fifth [1423]. Figularia [968]. Fillets [1, 741, 545, 6, 292, 401, 385, 1096, 1577, 712, 1531, 86, 410, 1203, 159].

Fillipi [460, 386]. Filter [1042, 1327]. Filter-Feeding [1042]. Filters [1212]. fimbria [1095, 228]. Fin [711, 1382, 820, 1092, 472, 495, 798, 716, 1011, 365]. Fin-clipping [1011]. Fin-Marked [711]. Final [1507]. Finclipping [1419, 1386]. Fine [1409]. Fingerling [1504, 295, 1214, 1412]. Fingerlings [1193, 876]. finmarchicus [127]. finned [100]. Fiord [890, 891, 1413, 1414, 1010]. Fiords [1326]. First [287, 588, 1266, 354, 1396, 1270, 1428, 537, 1546, 728, 732, 168, 288, 1072, 1102, 639, 1238, 1341]. Fischer [1357]. Fish [929, 907, 834, 1090, 745, 1068, 1035, 1462, 1351, 1586, 147, 347, 624, 1032, 36, 277, 188, 343, 482, 733, 821, 1178, 1176, 607, 215, 241, 234, 103, 102, 454, 455, 564, 679, 762, 619, 1080, 626, 921, 1224, 292, 401, 1449, 1448, 589, 490, 186, 1034, 986, 1030, 173, 902, 1428, 886, 1142, 1246, 137, 1530, 1263, 1162, 570, 1368, 881, 1577, 627, 285, 1161, 388, 8, 668, 818, 267, 395, 1210, 1486, 790, 113, 257, 868, 867, 1154, 1374, 1395, 81, 325, 383, 1041, 89, 942, 1194, 1373, 316, 453, 909, 117, 118, 1393, 647, 969, 1534]. Fish [1292, 540, 1138, 1240, 227, 1033, 731, 1215, 216, 463, 211, 1144, 251, 413, 331, 110, 157, 1290, 1114, 954, 349, 182, 541, 1158, 253, 1102, 105, 212, 274, 389, 993, 962, 109, 297]. Fisheries [1496, 75, 76, 148, 229, 338, 474, 476, 1179, 1494, 11, 84, 1302, 337, 356, 375, 397, 420, 452, 475, 503, 533, 557, 575, 592, 614, 630, 641, 653, 669, 686, 702, 718, 735, 752, 769, 786, 802, 824, 845, 866, 882, 903, 923, 940, 956, 970, 984, 1002, 1038, 1057, 1077, 1097, 1115, 1134, 1156, 1180, 1199, 1219, 1243]. Fishery [568, 272, 67, 520, 48, 612, 12, 283, 163, 893, 1005, 1572, 565, 1082]. Fishes [1407, 979, 1500, 643, 1168, 865, 360, 1255, 1358, 1450, 1316, 1400, 776, 1095, 1535, 1307, 1167, 1564, 560, 1045, 1306, 1405, 1118, 1124, 56, 1357, 1587, 60, 1550, 1349, 567, 850, 965, 682, 1350, 18, 540, 128, 1520, 524, 692, 726, 327, 423, 364, 640, 732, 1370, 961, 1291, 224, 1346, 1348, 975, 823, 1284, 1173, 1408, 1367]. Fishing [1462, 1519, 1105, 1130, 983, 863, 145, 404, 678, 52, 183, 697]. Fishways [360]. Fitting [631]. Five [973, 444, 1420, 1497]. Fixation [1206]. Fjord [851, 206, 381]. Flag [1456]. Flatfish [1107]. Flatfishes [523]. Flathead [87]. flavescens [656, 1422, 782, 1354, 1355]. flavidus [1479, 751]. Flesh [367, 1204, 1557, 311, 137, 856, 117, 1309, 672, 827]. Floor [725]. Flora [1536, 1544]. Flounder [1107, 1549, 1498, 1454, 1577, 269, 1004, 1071, 565, 1132, 246]. Flour [81]. Flow [205]. Fluctuating [339]. Fluctuations [1513, 582, 1537, 1067]. Fluid [322, 1275]. Fluorescence [1393]. Fluorescent [1214, 1425, 1419]. Fluorescing [1381]. Fluoride [473]. Fluorimetric [1025]. Foetal [666]. Folic [1028, 1283]. Following [34, 172, 687, 237, 263]. fontinalis [1256, 616, 1054, 259, 708, 291, 290, 1442, 1433, 1183, 1253, 924, 938, 763]. Food [834, 1473, 406, 91, 757, 755, 1091, 1118, 871, 809, 1387, 19, 596, 30, 861, 1113, 729, 524, 692, 726, 799, 195, 1103, 1573, 1248, 1226, 254]. Foods [165, 1544]. Foot [860]. Forecasts [1462]. Forest [872, 870, 871, 868, 867]. Forests [869]. Foreword [421, 230, 1343]. Form [790]. Formalin [327]. Formation [628, 98, 742, 1207, 618, 1486, 808, 1347, 680, 258, 1039, 69]. formed [46]. Formerly [1371]. Forms [1162, 25]. Formula [126]. Formulation [1563]. Forster [154]. Found [1241, 414]. Four [23, 1207, 1306, 583, 261, 498, 1187]. Foxe [344, 170]. FP [1217]. FP-66 [1217]. Fractionation [673, 186]. Fractions [784, 28]. franciscana [1533]. Fraser [239, 796, 1248]. fraye [1186]. FRB [1298, 1324, 1342, 1404, 1432, 1460, 1495, 1523, 1547, 1570]. Free [187, 218, 628, 1415, 1202, 1538, 1032, 1159, 762, 1105, 633, 648, 1022, 618, 1486, 45, 402, 182, 300, 46, 672, 1049, 26]. Free-Diving [300]. Free-Swimming [1032, 633, 648, 402]. Freeze [925, 920, 541, 1589]. Freeze-Dried [541, 1589]. Freeze-Drying [920, 925]. Freezing [516, 712, 1086, 1528, 1044, 675, 566, 1584]. French [1329, 1150, 1367, 1372, 593, 1186, 1187]. Frequency [1144, 223]. Fresh [1415, 1300, 684, 816, 1040, 1074, 1574, 704, 1247, 316, 805, 1072]. Freshly [242]. Freshwater [907, 952, 921, 986, 1030, 1450, 1368, 51, 734, 1520, 860, 1258, 211, 961, 261, 493]. Frobisher [193]. Frozen [1465, 1538, 1, 695, 88, 709, 482, 821, 849, 741, 1104, 902, 385, 346, 787, 1096, 1417, 1577, 1131, 196, 382, 458, 348, 516, 814, 1157, 1528, 242, 310, 159]. Fry [897, 1094, 811, 629, 974, 1381, 1281, 835, 51, 1301, 1286, 1012, 637, 1072]. Fuca [141]. Function [191, 1146, 1530, 350, 351, 1292]. Fundulus [1265, 739]. Fundy [1266, 15, 63, 59, 1555, 1548, 92, 296, 1116, 318, 61, 74, 652, 207, 1512]. Fungus [1234, 874]. Further [914, 305, 1578, 339, 394, 1579, 449, 74, 102]. Furunculosis [1253, 211]. Furunculosis-Infected [1253]. Fury [497].

G. [667]. Gadidae [915, 1315]. Gadus [1098, 764, 1007, 1328, 990, 1120, 1376, 659, 292, 401, 1105, 693, 65, 340, 642, 727, 783, 376, 705, 120, 434, 345, 586, 1251, 750, 789, 1010, 1572, 286, 176, 303, 365, 1580]. Gaffkemia [1399, 1514]. Gaffkya [1083, 1088, 738, 1013, 1399, 1275]. gairdneri [1159, 1443, 34, 172, 470, 749, 687, 887, 885, 295, 988, 1146, 621, 571, 1023, 112, 1031, 1022, 927, 91, 1336, 1578, 943, 1073, 161, 1278, 754, 795, 539, 195, 62, 654, 980, 416, 1438, 646]. gairdnerii [784, 634, 189]. Gambusia [1505]. Gamma [665, 805, 410, 409, 825]. Ganglia [1173]. Garlic [1565]. Gas [367, 362, 378, 558, 929, 858, 236, 876, 471]. Gas-Bubble [876]. Gas-Liquid [378]. Gaseous [693]. Gasterosteus [1319, 941, 1466, 1575, 1313, 1261]. Gastric [213, 1275]. Gastropoda [1064, 1409]. Gatineau [1186]. géants [1186]. Gel [1263, 698]. Gelatin [164]. Gempylid [901]. Genera [1349, 276, 894]. General [855, 326]. Generalized [828]. Genetic [1012]. Genital [840]. Genus [1351, 444, 1384, 1559, 1357, 766, 424, 598, 1361, 1191, 874, 1335, 1102, 623]. Geographic [233, 1509]. Georgia [552, 826, 657, 994, 730, 1573]. Georgian [1006, 1524, 635]. Germination [954, 854]. germo [144]. giant [1186]. gibbosus [515, 847, 1361, 199]. gigas [978, 573, 793]. Gilbert [147, 305, 699]. Gill [277, 1373, 889, 1292, 219, 649]. Gilled [553]. Gillnet [1105]. Gillnets [1398, 1534, 1431]. Gillraker [967, 1052]. Gills [180, 1453]. Girard [414, 857, 202, 552]. Given [1130]. Gizzard [836]. Glacial [1223, 127]. gladius [787, 1103, 1582]. Gland [1379, 1441, 949]. Glands [389]. Glass [1461]. glauca [417, 419]. Glendale [700]. Globicephala [998]. Glucose [34, 1576, 621, 1099, 10]. Glutamic [1127]. Glutamic-Oxalacetic [1127]. Glyceraldehyde [1516]. Glyceraldehyde-3-Phosphate [1516]. Glycerin [79]. Glycerin-Stored [79]. Glycerophosphate [1574]. Glycerylphosphorylcholine [830]. Glycogen [1098, 34, 172, 470, 749, 360, 696, 85, 155, 169, 1157, 1158, 212]. Glycogenolytic [659]. Glycolysis [310]. Glycolytic [484, 1528]. Glyptocephalus [610]. Gmelin [578, 313, 723]. Goby [837]. Goldeye [20, 1249]. Goldfish [1149, 719, 1184, 602, 1169]. Gompertz [1291]. Gonad [237]. Gonadectomized [1437]. Gonadectomy [1379, 1558]. Gonads [8]. Gonyaulax [730, 934, 1216]. Goode [1053]. goodei [498]. gorbuscha [121, 603, 700, 1353, 991, 438, 680, 1195, 932]. Grab [334]. Grades [103]. Gradients [512]. Grading [936, 103]. Graham [243]. Grand [595, 329, 411, 387, 671, 1389]. Grande [593]. Grande-Rivière [593]. grandis [1428]. Granular [436]. Gray [199, 159]. Grayling [128, 368]. Great [168, 1123, 313, 95, 507, 879, 1318, 1168, 154, 704, 1124, 113, 257, 392, 177, 1155, 1082]. Greater [1444]. Green [597, 68]. Greenland [1305, 873, 1009, 1266, 1287, 439, 1109, 1306, 529, 1166, 639, 302]. Greiner [221]. Greysole [610]. groenlandica [1228]. groenlandicus [519, 681]. Gross [542]. Groundfish [864, 667, 590, 13, 675]. Grounds [167]. Group [1142]. Groups [1564, 963]. Growth [631, 1496, 1388, 1504, 136, 459, 1165, 617, 812, 1556, 35, 209, 1513, 1070, 65, 794, 210, 597, 154, 1482, 1236, 755, 252, 1217, 84, 434, 1133, 1563, 4, 345, 766, 736, 266, 438, 1421, 1231, 1081, 1361, 583, 68, 808, 1301, 275, 506, 72, 1491, 1347, 524, 692, 726, 860, 895, 934, 1205, 1085, 1579, 1311, 184, 425, 1325, 1153, 1011, 1182, 1248, 1291, 1200, 166, 372, 906, 846, 1201, 1065, 993, 1067, 256, 1187]. gryllus [823]. Guanidine [1076]. Guanine [1189]. Guanotrophy [914]. Guinea [1142]. Gulf [264, 1250, 511, 725, 1446, 1066, 1461, 742, 743, 1229, 1541, 1142, 1387, 417, 1546, 424, 1177, 289, 92, 1323, 18, 61, 74, 1273]. Gulls [1375]. Gunnel [426]. Gunnerus [127]. Günther [638, 100, 723]. Guppy [406]. Gutted [103, 102, 1574]. Gutting [1492]. Gymnocanthus [1269]. Gyrocotyle [1225].

habitat [1197]. Habits [1246, 1103]. Haddock [756, 103, 102, 1574, 329, 4, 1387, 14, 289, 410, 1203]. Haematological [719]. Haematozoa [1374]. Haemoglobins [1254]. Hagfish [1581]. Hake [556, 1588]. Halargyreus [1102]. Halibut [1009, 1266, 1516, 439, 411, 809, 117, 1394, 167, 166, 535, 672, 697, 1503, 639]. Halifax [1536]. Hand [1492]. Hansen [399]. Harbour [1317, 700, 153, 838, 841, 977, 1273]. Hardiness [394]. harengus [1056, 55, 58, 57, 191, 1449, 1448, 209, 210, 78, 74, 300, 1056]. Harp [1228, 1402, 519, 681]. Harpoon [1032]. Harvest [1119]. Hatched [918]. Hatchery [811, 1020, 169, 72, 569, 831, 1444]. Hatchery-Reared [1020, 569, 831, 1444]. Hatchery-stocked [72]. Hatchetfish [587]. Hatching [210, 248]. Haustoriidae [1310, 1497, 1390]. Hawaiian [1366]. Head [1565, 153]. Heading [1492]. Heart [916]. Heat [733, 602]. Heavily [297, 480]. Heavy [1080]. Hecate [590, 314, 434, 176]. Hecla [497]. Held [1040]. helvomaculatus [498]. Hematin [1498]. Hematological [1554, 556, 988, 1031, 620, 1028]. Hemilepidotus [1351]. Heming [262, 383, 495, 596, 782, 1041, 1242, 1354, 1587, 254]. Hemioniscus [1121]. Hemiurid [1349, 1361]. Hemocytes [992]. Hemoglobin [687, 760, 494]. Hemoglobins [493, 1503]. Hemolymph [746, 1401]. Hemolymphs [1075]. Hemolysis [760]. Heney [1186, 1186]. Hepatic [1576]. Hepatopancreas [992]. Hermaphrodite [1459]. Hermaphroditism [1517]. Herring [707, 1009, 1415, 994, 55, 58, 57, 54, 191, 1467, 1449, 1448, 209, 67, 210, 1122, 1236, 688, 30, 83, 244, 71, 82, 12, 92, 353, 1446, 78, 446, 413, 197, 1069, 61, 74, 263, 300]. heteroclitus [1265]. Heterogeneity [857]. Heterotrophs [1573]. Hexachlorocyclohexane [99]. Hexanchus [553]. Hierarchy [1226]. High [344, 1157, 261]. High-Energy [1157]. High-Salinity [344]. Highly [767]. Hilaire [1380]. Hiodon [20, 1249]. Hippoglossoides [312, 387, 671, 895, 526, 1383, 1266, 439, 639]. Hippoglossus [117, 1503]. hippurus [861, 181]. Histidine [26]. Histogram [781]. Histological [1549]. Histology [1409, 408]. Histopathology [876]. Histories [776, 1310]. History [450, 48, 324, 298, 38, 526, 64, 1230, 646, 1369]. Holarctic [894]. Holding [834, 745]. Holocephali [1225]. Holophryxus [428]. Holt [126]. homari [1083, 1088, 738, 1013, 1399, 1275]. Homarine [508]. Homarus [1469, 69, 98, 768, 1218, 746, 1083, 513, 1170, 398, 574, 1257, 1055, 473, 738, 992, 1181, 1275, 1514, 1401, 1268]. Home [328]. Homing [1101, 380, 1385, 971, 1471]. Homogenates [1176, 1210]. Homogeneity [359]. Hooded [1228]. Hook [431]. hopliticus [287]. Horizontal [1472, 512]. Hormone [1001]. Host [180, 1363]. Hosts [1373]. Hour [1118]. Howard [1367, 1367]. Hudson [80, 959]. hudsonius [650]. Humber [527]. Humpback [279]. Hunger [674]. Huron [37, 1507, 976]. Hyas [1075]. Hybrid [1561, 38, 433]. Hybridization [584, 1027]. Hybrids [1136, 543]. Hydroelectric [540]. Hydrographic [153]. Hydrography [931, 1305]. Hydroid [965]. Hydrolagus [1225]. Hydrolysis [88, 346, 382, 82, 159]. hydrophila [754]. Hydrostatic [191]. Hydroxycorticosterone [1440, 668, 818, 1056]. Hyperprosopon [732]. Hypertrophy [1558]. Hypocholesterolemic [109]. Hypomesus [857, 1036]. Hypophysectomized [952]. Hypophysectomy [1149]. Hypotheses [1454]. Hypothetical [906]. Hypoxanthine [986, 1093, 1189, 778]. Hypoxial [1281].

I. [1554]. Ice [1202, 1538, 695, 31, 1040, 1093, 656, 33, 873]. Ice-stored [33]. Iced [986, 787, 1574]. Iced-Storage [1574]. Icelinus [588]. Ichthyophonus [1071]. Ictiobus [1554, 324]. Identification [1294, 633, 1240, 471, 480, 347]. Igloolik [122]. II [299, 98, 122, 1367, 454, 1305, 236, 118, 169, 404, 692, 211, 66, 480]. III [333, 178, 455, 328, 137, 395, 726]. Ileum [789]. Iliamna [1061]. illecebrosus [906, 488, 906]. Illex [488, 906]. Illumination [460, 308]. Imagotaria [1188]. Immature [304, 561]. Immigrant [1300]. Immune [1515]. Immunization [539]. Immunohistochemical [1441]. Implications [1479]. Importance [1246, 1377]. Impoundment [32]. Impressions [714]. Improve [52]. Improved [216, 1584]. In-Vitro [278]. Inactivating [108]. Inch [509]. Incidence [857, 383, 738, 2, 697]. Incidental [717]. Including [384, 561]. Incomplete [130]. inconstans [270, 1509]. Increase [1200, 1226, 5]. Increased [514]. Incubator [1068]. Index [76, 148, 229, 338, 476, 615, 803, 1274, 1592, 1574]. Indian [384]. Indiana [794]. Indicator [352]. Indicators [1360]. Indies [861]. Indirect [1309]. Individual [1240]. Individuals [561]. Indophenol [921]. Induced [1556]. Induction [1076]. Industrial [1377]. Industry [39]. Inertial [1112]. Infauna [1280]. Infected [1253]. Infection [1054, 1155, 1071, 1480]. Infections [1346, 594]. Infectious [1410, 1489, 1585, 1049, 1515]. Infertile [670]. Infestation [2]. Infestations [1273]. Inflammatory [793]. Influence [1129, 1024, 597, 290, 1084, 112, 928, 1189, 1086, 1282, 637, 258, 846, 1514, 697]. Influencing [800, 316]. Information [339, 598, 449]. Infraoral [143]. Ingestion [35, 6]. ingolfianus [355]. Inhabiting [1526]. Inheritance [1136]. Inhibiting [455]. Inhibition [43]. Inhibitors [1181]. Iniomi [1053, 1322]. Injury [365]. Inlet [807, 776, 160, 1175, 1506, 536, 161, 435, 1079, 149]. Inlets [138, 309, 926]. Inner [636]. Inoculation [1515]. Inorganic [1206, 679, 1167]. Inosine [31]. Inositol [40]. Input [241]. Insect [1001, 1380]. Insecta [521, 1143]. Insecticide [1165, 1505]. Insecticides [868, 867, 900]. Insects [872, 1429, 870]. Inshore [855, 998]. Instability [1406]. Instance [914]. Instructional [997]. Insular [1400, 961]. Intake [35]. Integument [888, 1275]. Intended [566]. Intensities [706]. Intensity [1321, 536, 351, 68, 361]. Inter [623]. Inter-Species [623]. Interacting [85]. Interaction [1083, 571, 1193]. Interactions [68, 53]. Interference [696]. Interior [981]. Intermittent [654]. Internal [8]. International [11]. Interpopulation [1553]. Interpreting [1246]. Interrenal [1146, 1558, 1056]. Interrenalectomy [1245]. Interspecific [273]. Intertidal [1310, 601]. Intestinal [1279]. intestinalis [383, 1490]. Intestine [1368]. Intra [94]. Intra-ocular [94]. Intraperitoneal [507]. Intraseasonal [266]. Introduced [85, 440]. Introduction [867, 890, 448, 843]. Introductions [540]. Introductory [11]. Invasion [313]. Invasive [1365]. Invertebrates [1303, 1185, 1425, 996, 867, 1239]. Investigation [1532, 247, 300]. Investigations [337, 356, 375, 397, 420, 452, 475, 503, 533, 557, 575, 592, 614, 630, 641, 653, 669, 686, 702, 718, 735, 752, 769, 786, 802, 824, 845, 866, 882, 903, 923, 940, 956, 970, 984, 1002, 1038, 1057, 1077, 1097, 1115, 1134, 1156, 1180, 1199, 1219, 1243, 1298, 1324, 1342, 1404, 1432, 1460, 1495, 1523, 1547, 1570, 527, 1176, 1172, 11]. Involved [1157]. Involving [938]. Iodate [1458]. Iodide [1167]. Iodine [1009]. Ion [486, 559]. Ionic [602, 124]. Ions [1080]. IPN [1410]. Irminger [1018]. Iron [679, 762, 1223, 110, 663]. Irradiation [628, 343, 410]. irroratus [1088, 1075]. Island [1220, 1400, 149, 221, 1061, 151, 294, 1004, 435, 309, 1287, 1241, 851, 813, 890, 891, 1413, 1414, 678, 1010, 565, 32, 1293, 193, 873, 551, 256, 1326, 1341]. islandica [177]. Islandq [1268]. Islands [1141, 553, 1046, 975, 1456, 527, 725, 585]. Isle [1273]. Iso [858]. Isofucosterol [44]. Isolated [347, 733, 746, 1453, 1440, 818]. Isolating [941, 1027, 1261]. Isolation [847, 1330]. isolepis [1175, 1111]. Isopod [428, 1121, 175, 1029]. Isopropanol [1449, 1448]. Isopropyl [929, 1090, 881]. Isopsetta [1175, 1111]. Isosmotic [1265]. Isozymes [1561, 1276]. Isurus [321]. IV [352, 407, 564]. IX [8].

Jack [1050]. James [1477]. Japan [1079]. Japanese [1001]. japonica [1391]. Jaschnov [127]. Jellyfish [1160]. John [1267, 29]. jordani [639]. Journal [76, 148, 229, 338, 476]. Juan [553, 141]. Judged [857]. July [994]. Juvenile [1365, 217, 820, 1145, 280, 1023, 96, 1421, 1481, 491, 1540, 165, 553, 161, 317, 512, 1521, 703, 1198, 565, 1337, 518, 846, 1226, 1444, 1076].

Kalamalka [843]. Kamloops [34]. Kananaskis [540]. Karluk [380]. Karyotype [1168]. Keeping [66]. Kennebecasis [29]. keta [603, 1108, 1382, 112, 1334, 1421, 637]. Ketone [560]. Key [444, 633, 1434, 894]. Kidney [121, 810, 1521, 246]. Kill [938]. Killifish [531]. Kinase [1263, 1264]. King [171, 1014, 1539, 678, 485]. Kirtland [270]. kisutch [1095, 571, 1381, 24, 715, 111, 1227]. Known [1394, 1372]. Kodiak [678]. Kokanee [1117, 536, 1051, 1052, 729]. Kootenay [448]. Kouchibouguac [987]. Kraft [690, 1562, 1451, 445, 977]. Kraft-Mill [445]. Krøyer [2]. Kvichak [880].

L. [496, 55, 58, 57, 191, 990, 847, 292, 401, 209, 693, 719, 65, 210, 602, 727, 376, 705, 4, 345, 19, 417, 265, 808, 78, 317, 512, 819, 433, 318, 349, 2, 198, 249, 74, 181, 300]. Labelled [1212, 69, 282]. Laboratory [340, 1058, 1327, 1391, 1209, 1163, 1490, 1347, 1311, 611]. Laboratory-Reared [1058]. Labrador [842, 461, 374, 912, 1273]. lac [1186, 47, 1205, 114]. Lacépède [213, 330]. lacs [1187]. Lactate [34, 172, 470, 687, 1137, 1561, 360, 1060, 621, 1276]. Lactic [1098, 1021]. laevis [1362]. Lagenorhynchus [9]. Lagocephalus [198, 249]. LaHave [1053]. Lake [1388, 636, 1206, 1220, 1507, 1517, 976, 635, 1141, 989, 1359, 776, 160, 542, 221, 293, 500, 958, 790, 1118, 1061, 262, 495, 596, 782, 1041, 1242, 1354, 1587, 266, 279, 429, 536, 38, 889, 1418, 972, 1392, 161, 447, 729, 1452, 219, 569, 831, 919, 1205, 114, 649, 677, 1153, 1140, 1213, 52, 115, 779, 1200, 1571, 448, 1190, 1355, 150, 997, 254, 577, 1445, 1220, 272, 1511, 1126, 243, 1463, 154, 48, 753, 1385, 757, 704, 790, 113, 257, 392, 97, 101, 383, 582, 266, 1186, 298, 177, 1565, 971, 890, 891, 1413, 1414, 717, 774, 435, 893]. Lake [1472, 1010, 1123, 219, 569, 123, 214, 1239]. Lakes [1168, 580, 785, 886, 794, 1026, 380, 91, 252, 767, 981, 1029, 1436, 64, 1196, 892, 1119, 791, 843, 1187, 95, 507, 37, 879, 1318, 1124, 1155, 1123, 313, 1082]. Lakeward [1278, 1012]. Lamina [143]. Lamna [418]. lamottenii [21]. Lampetra [1583, 489]. Lamprey [951, 1006, 976, 1565, 265, 1583, 436, 157, 358, 911, 143, 489]. Lampreys [538, 507, 1493, 41, 1155, 21, 894]. Lampricide [951]. Lampris [306]. Lance [591]. Landed [102, 1518]. Landlocked [760, 761, 1416, 1568, 890, 891, 1413, 1414, 1010, 358, 143, 1326]. Lapse [570]. lar [377]. l’Arctique [1150]. Lardeau [1278]. Large [1032, 589, 293, 1194, 561]. Large-Immature [561]. Largemouth [1024, 213, 846]. Larger [926]. Largest [113]. Larvae [1218, 340, 1058, 1422, 1109, 202, 1475, 1544, 878, 1311, 398, 918, 1055, 1451, 483, 197, 1089]. Larval [994, 886, 1122, 857, 74, 1582]. Larvicides [1565]. Last [184]. Late [396]. Late-Spawning [396]. Lateral [1256, 784, 1443, 599]. lateralis [938]. latifrons [916]. Launce [1182]. Lawrence [1223, 743, 264, 1412, 1387, 511, 725, 1546, 289, 1446, 1066, 1273]. Layer [955, 479]. Layers [1059]. Lead [1513]. Lean [454, 455, 564, 619, 453]. Lebistes [406]. Lecithaster [1361]. Lecithinase [817, 784, 634]. Leech [581]. Leidy [22]. Lemon [657]. lemur [728]. Length [218, 412, 1427, 999, 84, 4, 1540, 967, 327, 1272, 1285, 223]. Lengths [1564, 167]. Lepidophyllum [1351]. Lepidopsetta [467, 468]. Lepomis [515, 847, 597, 748, 1048, 1439]. Lepophidium [822]. Leptoclinus [732]. Lernaeoceridae [848, 1363]. Lernaeopoda [399]. Lernaeopodidae [399, 1306, 1559]. Lesions [542, 532, 1580]. LeSueur [466, 307]. Lethal [1318, 836, 1265, 265, 261, 341, 518, 284]. leucas [1525]. Leucichthys [214]. Leuroglossus [147]. leurolepis [585]. Level [1159, 1120]. Levels [470, 687, 360, 1265, 1189, 1231, 750, 518, 70]. lewisi [1064, 1409]. Liberated [256]. License [392]. Lichtenstein [823]. Life [450, 1353, 1524, 776, 324, 298, 814, 1361, 38, 1490, 740, 731, 410, 825, 526, 1310, 1390, 64, 646]. Life-History [64]. Lift [413]. Light [1129, 131, 173, 708, 1425, 290, 1050, 350, 1321, 1578, 942, 351, 68, 706, 361, 506, 317, 512]. Light-Salted [942]. Ligula [383, 1490]. like [1401]. Limanda [1071]. Limited [1403]. Limits [657, 1394]. Limnetic [886, 1196]. Limnocalanus [1524]. Limnological [580, 115, 577]. Limnology [636, 1026, 1436, 791]. limosus [771]. lindbergi [1361]. Line [369, 1256, 784, 1443, 1519, 983]. Linear [858, 283, 126]. lineolata [552]. Ling [204]. Lingcod [180, 999, 196, 28, 46, 242, 310, 159]. Linnaeus [538, 861, 358]. Lion [937]. Lions [1345, 372]. Liparidae [1384]. Liparis [640]. Lipid [88, 346, 855, 382]. Lipids [359, 1465, 1204, 768, 709, 456, 488, 783, 856, 1228, 395, 1160, 576]. Lipolytic [1443]. liquefaciens [1499]. Liquefaction [82]. Liquid [367, 362, 378, 558, 929, 858, 1237, 236, 83, 1545]. List [1408]. Lithodes [171].

Lithodidae [1058]. Little [1430, 1211]. Live [834, 1515]. Live-Holding [834]. Liver [362, 85, 1161, 8, 200, 236, 856, 705, 81]. Livers [628, 618]. Living [607, 823]. Lobster [1469, 69, 98, 768, 1498, 746, 1170, 825, 398, 918, 1055, 1451, 473, 738, 992, 1013, 1181, 1399, 1275, 1514, 1401, 39, 108]. Lobsters [532, 513, 235, 574, 517, 1257, 1267, 1268, 256]. Localities [1582]. Localization [1263, 1441]. Located [1014]. Location [536]. Locations [973]. Lockeport [4]. Locomotor [1282]. Lolliguncula [1427]. Long [636, 1443, 100, 953, 1537]. Long-Chain [1443]. Long-finned [100]. Long-Term [953, 1537]. Longear [748]. longimanus [181]. Longjaw [214]. Longline [520, 3, 823]. Longnose [1388, 1561, 1148, 154]. Longwave [1425]. Loon [160, 161]. Loss [887, 482, 315, 212, 416, 1212]. Losses [1, 868]. Lost [779]. Lota [262, 1242]. Lotinae [1315]. Lovén [448, 843]. Low [1009, 865, 58, 665, 1124, 1170, 750, 1144, 918, 1290, 261, 1338, 124, 225]. Low-Dose [665]. Low-Frequency [1144]. Low-Oxygen [225]. Low-Temperature [1290]. Lowe [861, 253]. Lower [703, 313]. lucius [757, 755, 50, 472, 596, 782, 798, 433]. Lucknow [1368]. lugubris [1300]. Lumpenus [969]. lumpi [658, 2]. lupus [819]. lurida [978]. lutheri [1556]. Lycodopsis [1506]. Lymphocystis [595]. Lysianassidae [1106]. Lysolecithin [1210]. Lysolecithinase [534].

M [1344]. M.S. [1022]. Macdonaldia [1033]. Machine [222]. Mackenzie [1491]. Mackerel [1040, 1050, 1522, 118]. macrocephalus [340, 642, 120, 434, 176, 303]. macrocheilus [1027]. macrochirus [847, 1048, 1439]. Macroscopic [47]. macrurus [1524]. maculatus [732]. Magdalen [725, 526]. magellanicus [578, 1208, 264, 1139]. magister [49, 136, 1311, 737]. Magnesium [1486, 486]. Magnetized [1566]. Maine [931, 1461, 1229, 1541, 1250, 1177, 92, 1323, 61, 74]. Mainland [138, 528]. Maintaining [53]. Maintenance [755, 1490, 1347]. Major [636, 40, 1314, 8, 384]. Making [714]. Mako [321]. Malaysia [1350]. Male [678, 611]. Malic [1162]. Mallotus [652]. malma [763]. Malpeque [1293]. Mammalian [212]. Man [1345, 70]. Management [1496]. Manayunkia [1567]. Manganese [1223, 551]. Manitoba [580, 785, 262, 495, 596, 782, 1041, 1242, 1354, 919, 892, 254, 577]. Manitou [1220]. Manitoulin [1220, 813]. Manual [1000]. Margaree [1521]. Margaret [1185]. margolisi [947]. Maria [907, 913]. marin [593]. Marine [405, 359, 558, 645, 1009, 858, 1164, 1068, 546, 817, 1500, 450, 178, 407, 456, 1233, 807, 572, 44, 991, 1458, 135, 200, 673, 350, 1374, 609, 896, 1099, 1223, 351, 384, 1406, 1087, 139, 140, 732, 1288, 934, 440, 1420, 1478, 1553, 1573, 301, 274, 493, 829, 225, 576, 663, 70, 109, 1329, 1150, 1367, 593]. Marine-Borer [829]. marines [1150]. marins [1329, 1367]. marinus [538, 19, 41, 265, 2, 358, 21, 143]. Marion [989, 1463, 1418]. Maritime [1585, 60]. Mark [110, 157]. Marked [711, 1220, 256]. Marking [1214, 1041, 1138, 1240, 1419, 875]. Marlin [1194]. Mary [1389]. Masking [1376]. Maskinongé [1187]. masquinongy [72, 315, 404, 433, 1187, 1119].

Mass [490, 264, 222, 1337, 1338]. Massachusetts [1396]. Masses [227]. Mate [963]. Material [881]. Mathematical [993]. Mathematics [1133]. Mating [1539, 737]. Matschie [1366]. Matter [996, 1079, 884, 779]. Maturation [1454, 1361, 278, 670, 949]. Mature [1140, 561]. Maturing [291, 134, 24, 544, 1558]. Maturity [1388, 49, 65, 4, 809, 650, 671, 1182, 1445]. Maximum [704, 1321, 909, 16, 323, 1537]. Meal [81, 244]. Meals [1030, 647, 913]. Mean [1272]. Means [410, 675, 566]. meanyi [331]. Measured [897, 205]. Measurement [589]. Measurements [94, 1427, 59, 864, 304, 1540, 167]. Measuring [1237, 241, 1151, 285, 996, 909, 166]. Meat [771, 409, 825, 1518]. Meats [554]. Mechanism [693, 430]. Mechanisms [1083, 941, 1027, 1261]. Media [1065]. Medical [1350]. Medusa [549]. Medusae [1435]. Megalopa [483]. megalotis [748]. Meganyctiphanes [844, 1109]. melaena [998]. Melanogrammus [756, 1574, 4]. melanops [1569]. melanura [665]. melanurus [305]. Melosira [177]. Members [238]. Membrane [532, 710, 1212]. Mendota [1239]. mentella [912]. Mere [914]. Meristic [785, 1522]. Merluccius [556, 947, 1588]. Meromictic [1436]. Mesidotea [175, 1029]. Mesopelagic [1535, 1479]. Metabolic [1284]. Metabolism [178, 749, 887, 215, 441, 515, 1530, 621, 1099, 692, 1258, 654, 465, 599, 109]. Metabolites [1034, 1070]. Metacercariae [1348]. Metal [1080]. metamorphosis [1363]. Meteorological [153]. Meter [781]. Method [631, 1260, 607, 1094, 921, 589, 1025, 385, 1151, 1425, 158, 79, 453, 126, 1286, 216, 677, 167, 1114, 223, 561, 205]. Methods [759, 1462, 690, 1108, 481, 1474, 767, 647, 351, 1311, 1190, 697, 108]. Methyl [187, 673]. Metopirone [1146]. mexicanus [460, 386]. Mexico [1367]. Mexique [1367]. Michigan [37, 1123]. Micro [973, 698]. Micro- [973]. microcephalus [1287]. Micromesistius [286]. Microorganisms [1035, 746]. Micropterus [1277, 213]. Microscopic [542]. Microscopical [1549]. Microsporidial [1480]. Microturbellarian [1241]. Midwater [975, 1069, 1059]. Midway [1366]. Migrant [897]. Migrating [360, 1566, 328, 1135]. Migration [1365, 711, 23, 624, 37, 40, 632, 570, 991, 8, 293, 10, 1395, 51, 430, 275, 1278, 1471, 199, 779, 852, 26]. Migrations [156, 966, 289, 442, 1389, 1012, 519]. Migratory [974, 1441, 161]. Milk [1228]. Milky [364]. Mill [690, 1451, 225, 445, 977]. Milt [856]. Mineralization [1079]. Minke [330]. Minnow [1510]. minor [548]. Mio [1188]. Mio-Pliocene [1188]. Miramichi [806, 1302, 190, 84]. Miscellaneous [137]. Misidentified [1371]. Missouri [1408]. Mitchill [259, 291, 290, 433, 254]. Mixed [963, 518]. Mixing [1250]. Mixtures [1448]. Mobility [865]. Model [314, 743, 294, 443, 828, 1285, 1553, 993, 997]. Modes [1356]. Modification [1548, 502]. Modifications [1526]. Modified [607, 334, 1340, 42]. Moisture [684, 292]. Molecular [1162]. Mollusc [1015]. Mollusca [568, 440]. Molluscs [899]. Molt [1469, 1001]. Molt-Hormone [1001]. Molting [908]. Molva [204]. Monachus [1366]. Monitored [955]. Monk [1366]. Monkfish [528]. Monobenzoate [134]. Monocarbonyl [479, 480]. Monochrysis [1556]. Monogenea [1358, 1225]. Monogenean [1373]. Monogenetic [1358]. Monoglycerides [902]. Monophosphate [944]. Mont [1380]. montagui [852]. Moody [547]. mordax [1186, 1327]. morhua [1098, 764, 1007, 1328, 990, 1120, 1376, 292, 401, 1105, 693, 65, 727, 783, 376, 705, 345, 586, 1251, 750, 789, 1010, 1572, 365, 1580]. Mori [354]. Morid [1102]. Morphological [94, 1552, 1489, 919, 1227]. Morphology [95, 764, 1316, 848, 549, 224, 1363]. Morphometric [419, 418]. Morphometrics [1182]. Morphometries [1322]. Morphometry [583]. Mortalities [264, 283, 126, 1337]. Mortality [1234, 756, 481, 243, 1060, 1026, 189, 252, 434, 1130, 158, 1125, 315, 404, 185, 431, 606, 1087, 874, 1383, 184, 425, 1325, 1119, 562, 1338, 993]. Mortem [484, 212, 105, 46]. Mortis [234, 125, 416, 105]. Mosquitofish [1505]. Moult [1218, 838]. Moulting [1364]. Mountain [620, 791]. Mounted [1114]. Mouth [790]. Movement [257, 361, 32]. Movements [1303, 629, 570, 753, 1385, 411, 270, 298, 269, 92, 569, 342, 258, 946, 256]. Moving [1263]. Moving-Boundary [1263]. MS [470]. mucosus [640]. Mucous [887]. Mud [551]. Mueller [1370]. Müll. [177]. Müller [1191, 1362]. multidentata [1461]. Multiple [858, 1005, 1254]. Multiple-Branched [858]. murrayi [1323]. Muscle [1202, 1538, 1136, 784, 634, 1443, 1007, 88, 709, 659, 241, 234, 935, 454, 455, 564, 679, 619, 1080, 1104, 1498, 31, 292, 401, 1105, 186, 944, 787, 1417, 125, 484, 696, 955, 1040, 1093, 1263, 1162, 432, 196, 830, 1008, 1210, 1486, 1194, 453, 1393, 877, 1139, 778, 1431, 1157, 1528, 1264, 992, 1147, 182, 1158, 28, 46, 105, 212, 242, 310, 416, 535, 278, 124, 238, 493, 494, 543, 1529, 911, 534]. Muscles [599]. Muscular [756, 1159, 34, 172, 687, 246]. Muskellunge [1294, 433, 1187]. Mussel [354, 860, 1258]. Mussels [506]. Mutant [50]. Mya [22]. Myctophid [965]. Mylocheilus [1136]. Myogen [238, 494]. Myogens [493, 543, 911]. Myosin [662, 935]. Myotomal [1007]. Myoxocephalus [1270, 1430]. Mysis [37, 448, 843]. Mytilicola [1296, 354, 573, 1488]. Mytilus [354]. Myxosporidian [550].

n [1527, 1368, 1560, 1370, 947, 1112, 983]. N.B. [55, 13]. N.S. [4]. N.W.T [221, 500]. N.W.T. [87, 1491]. Nabisipi [703]. namaycush [1150, 542, 429, 889, 919, 763]. Name [366]. narium [1560]. Narragansett [633, 648]. Nasal [1358, 1560]. Nassarius [551]. nasus [194, 418, 1491]. Natantia [721]. Native [72]. Natural [1083, 1321, 145, 1301, 404, 1470, 211, 1116, 1200, 1571, 738, 594, 663]. Naturally [897]. Nature [231]. Nauplii [633, 648]. Naval [1350]. Nealotus [901]. Near [294, 883, 529]. Nearby [1046]. Necrosis [1410, 1489, 1585, 1049, 1515]. Need [285]. Neighbouring [505]. Nematoda [1365, 22, 857, 1357]. Nematode [1364, 278]. Nematodes [1375, 1350]. nematophora [549]. Neotype [377]. nerka [1108, 1382, 604, 945, 1504, 1437, 982, 328, 1146, 605, 1060, 963, 134, 1276, 293, 189, 395, 1398, 536, 237, 544, 810, 1379, 1558, 967, 1301, 1441, 1051, 1052, 729, 1012, 827, 1386, 932, 949].

Nessorhamphus [355]. Net [388, 767, 163, 1487]. n’était [1372]. Nets [277, 1122, 219, 649]. Neurofibroma [969]. Neurosine [347]. Neuston [1487]. Newfoundland [271, 1009, 1465, 700, 545, 1269, 1430, 65, 520, 1319, 688, 488, 1306, 19, 1416, 382, 458, 403, 516, 859, 1086, 1464, 937, 1016, 998, 1211, 17, 78, 1572, 387, 671, 895, 961, 906, 1267, 174, 204, 198, 505, 499, 1352]. Newly [918]. nicholsi [837]. Nicola [536, 729]. niger [723]. Night [1590]. nigromaculatus [1043]. nilotica [1551]. Ninespine [233]. Nipigon [1140]. Nitinat [435]. Nitrogen [1206, 10, 1545, 1439]. No [1350]. Nocardiae [1054]. nodulosus [1242, 1354]. Nogies [72, 315, 404]. Noise [1120]. Nomenclature [1051, 1509]. Nomographs [613]. Non [10]. Non-protein [10]. Nonparasitic [507, 489]. Nonsense [1133]. Normal [762]. North [162, 281, 299, 333, 352, 1121, 1569, 689, 1033, 591, 426, 1103, 330, 2, 676, 1102, 489, 646, 1300, 700, 1252, 222, 319, 113, 194, 1522, 129, 233, 1350, 1051, 1052, 323, 804, 549, 1143, 1370, 1106, 1335, 723, 427, 498, 751, 853]. Northeast [1113, 77, 968, 27, 142, 1166, 962]. Northeastern [1462, 1015, 1152, 1428, 579, 638, 1095, 622, 792, 1271, 226, 25]. Northern [1294, 797, 472, 1327, 1394, 498, 747, 751, 1341, 773, 600, 1287, 886, 794, 267, 252, 590, 1003, 981, 1059]. Northumberland [511, 398]. Northward [87]. Northwest [612, 568, 643, 624, 1262, 1493, 392, 1112, 306, 822, 728, 312, 895, 288, 806, 1302, 1110, 519, 366, 477, 181, 307]. Northwestern [1247, 625]. norvegica [844, 1109]. Norwegian [815]. Nose [431]. Notacanthid [1033]. notatus [739]. Note [22, 95, 428, 286, 62, 182, 249, 124, 284]. Notes [811, 1384, 1411, 1535, 1533, 1315, 1509]. Notropis [650]. nov [423]. Nova [1185, 1400, 1053, 1322, 353, 717, 1055, 1451, 64, 411, 1387, 357, 289, 1163, 819, 1071, 1182, 115]. Nuclei [1446]. Nucleotide [787, 1417, 696, 955, 1040, 1528]. Nucleotides [1339, 508, 242]. Number [758, 967, 1052, 446, 680, 1501]. Numbers [785, 312]. Numerical [924]. Nursery [51]. Nutrient [244, 68, 1378, 770]. Nutrients [1414]. Nutrition [1533, 1356]. Nutritional [1532, 6, 463]. Nutritive [771, 1557, 83, 647, 913]. nuttalli [1165]. Nylon [1431, 219].

O [177, 1562]. O. [603, 1108, 1382, 1060, 932]. Obituary [250, 422, 504, 685, 116, 179, 245, 320, 370, 777, 964]. Objective [188]. obliquidens [9]. obscurus [307]. Observation [1112]. Observations [271, 1549, 497, 1287, 54, 1507, 1208, 340, 1359, 280, 627, 414, 542, 293, 190, 868, 120, 1578, 511, 424, 38, 175, 353, 17, 1347, 364, 224, 1355, 105, 165]. Observed [631, 25]. Observing [146]. obsoletus [551]. Obtained [568]. Obtaining [682, 1114, 389]. Occa [1411]. occidental [1150]. Occurrence [335, 645, 1042, 994, 1266, 171, 1569, 648, 699, 1053, 1322, 96, 1109, 279, 1223, 1177, 306, 1323, 447, 1452, 214, 1110, 1196, 181, 307, 321, 652, 1341, 1150]. Occurrences [1269, 1430, 640]. Occurring [125]. Ocean [711, 144, 1367, 481, 1125, 1014, 442, 185, 1394, 1531, 184, 425, 5, 1367, 369, 1462, 222, 691, 525, 579, 439, 622, 983, 1304, 1113, 77, 175, 1271, 660, 1033, 549, 27, 142, 1378, 1529, 7, 427, 240, 498, 751, 962, 853]. Oceanic [985, 1569, 1095, 953, 721, 355, 25, 198, 249, 912]. Oceanic-Coastal [985]. Oceanographic [122, 93, 537, 691, 141, 138, 309, 926, 90, 427]. Oceanographical [29]. Oceanography [206, 149, 240]. Oceans [1128]. ocellata [561]. ocular [94]. Odobenus [1016]. Odor [1232]. Off [627, 1280, 773, 144, 600, 1519, 1241, 1255, 104, 691, 1297, 192, 1053, 1322, 1109, 60, 953, 357, 353, 1402, 721, 9, 1572, 819, 1071, 1055, 1451, 1267, 873, 975, 1059, 1378, 5, 1238]. Off-Bottom [627]. Offshore [509, 357, 289, 1182]. Ogac [890, 891, 1413, 1414, 1010]. Oil [187, 218, 335, 271, 362, 412, 695, 905, 200, 236, 247, 618, 109]. Oils [707, 1009, 907, 1107, 1415, 1030, 135, 200, 673, 109]. Old [393]. Olfaction [1101, 971]. Olfactory [255, 328, 165, 1471, 1470, 368]. Oligochaetes [1526, 1532]. Oligopeptide [347]. Once [458]. Once-Frozen [458]. Oncholaimidae [22]. oncorhynchi [1365]. Oncorhynchus [23, 121, 1127, 444, 603, 700, 1108, 1382, 1353, 604, 945, 1504, 1437, 982, 328, 1146, 605, 1192, 1095, 1060, 1101, 963, 571, 991, 134, 1276, 112, 293, 189, 395, 1334, 1398, 438, 1421, 1381, 536, 24, 424, 598, 1361, 715, 237, 544, 1379, 1558, 967, 1301, 1441, 1051, 1052, 729, 16, 185, 680, 1012, 637, 827, 124, 494, 623, 1195, 111, 1227, 1386, 932, 949]. Oncorhyncus [810]. one [1372, 137]. Onset [416]. Ontario [1303, 1220, 272, 1126, 748, 790, 270, 38, 567, 889, 145, 72, 315, 404, 774, 1472, 919, 211, 123, 1213, 1119, 1355]. Onychoteuthidae [288]. Oocyte [1454]. Opah [306]. Open [858, 1474, 753, 1385]. Open-Tubular [858]. Open-Water [1474, 753, 1385]. Opening [52]. Operations [316]. Opercular [259]. opercularis [699]. Ophidioid [1428]. Ophiodon [180, 1560, 46]. Ophiuroidea [600, 1434]. opilio [1170, 1171]. Opinicon [790]. Optical [224]. Optimal [1420]. Optimum [1321]. Oral [539]. Orange [554]. Orange-Red [554]. orbis [1586]. Orcas [1004]. Orconectes [771]. Order [373, 226, 143]. Ordinary [787, 416]. Oregon [773, 600, 144, 885, 1255, 414, 917, 721]. Organ [1409]. Organic [996, 1079, 779]. Organisms [1042, 609, 390, 1288]. Organogenesis [259]. Organoleptic [1040]. Organophosphate [1475]. Organophosphates [1161]. Organs [8]. orientalis [1296, 354, 573, 1488]. Orientation [1474, 991, 971, 430, 1209]. Origin [405, 558, 858, 344, 501, 182, 576]. Oscillation [1112]. Oscillations [1116]. Osmerus [1186]. Osmolarity [1149]. Osmoregulatory [469]. Osmotic [1314]. Osteology [915, 1315, 128]. Ostracods [384, 548]. Ostrea [978]. Other [337, 356, 375, 397, 420, 452, 475, 503, 533, 557, 575, 592, 614, 630, 641, 653, 669, 686, 702, 718, 735, 752, 769, 786, 802, 824, 845, 866, 882, 903, 923, 940, 956, 970, 984, 1002, 1038, 1057, 1077, 1097, 1115, 1134, 1156, 1180, 1199, 1219, 1243, 1298, 1324, 1342, 1404, 1432, 1460, 1495, 1523, 1547, 1570, 347, 1206, 482, 564, 1427, 1535, 388, 200, 194, 1194, 413, 1573, 868, 593]. Otolith [1108, 1236, 1446, 167]. Otoliths [1246, 79, 371, 166]. Ottawa [1270]. Otter [756, 1487]. Otters [792]. Ouest [466]. Outer [710, 1506]. Outlet [160, 161]. Ova [605, 932, 1259]. Ovaries [291, 1580]. Ovary [1454, 1320]. Over-Winter [1026]. Overhead [1578, 706]. Overwinter [1412].

Overwintering [1268]. Ovine [952, 1149]. Owikeno [1392]. Oxalacetic [1127]. Oxidation [1458, 395]. Oxidative [482, 741]. Oxide [69, 98, 407]. Oxygen [268, 687, 760, 616, 724, 761, 604, 1024, 1513, 1551, 720, 1170, 1282, 910, 261, 846, 27, 142, 225, 1067]. oxyrinchus [321]. Oyster [573, 1340, 664, 530, 1377, 793, 878, 1293, 465, 770]. Oysters [978, 1241, 839, 506, 462]. Ozernaya [651].

P [77, 27, 142, 369, 691, 983]. P.E.I. [1338]. Pacific [369, 1462, 222, 579, 622, 1113, 77, 1271, 1033, 549, 27, 142, 1378, 1529, 427, 498, 751, 962, 853, 162, 23, 1127, 1015, 1152, 1484, 444, 849, 1467, 255, 49, 136, 1367, 573, 31, 985, 1569, 1428, 537, 691, 340, 642, 481, 638, 699, 1095, 1384, 439, 120, 434, 443, 202, 1304, 424, 552, 1113, 1481, 598, 1361, 71, 544, 51, 792, 442, 446, 689, 185, 323, 660, 364, 793, 804, 9, 968, 1478, 226, 933, 66, 167, 25, 838, 841, 176, 303, 566, 124, 1503, 5, 7, 240, 1588, 400, 875, 646, 159]. pacifica [1506]. pacificum [1345]. pacificus [450]. Pacifique [1367]. Packed [858]. Pagophilus [1228, 519, 681]. Pagurus [366]. Pair [371, 833]. Palatability [1, 311, 385]. Pallas [228]. pallasi [197]. pallasii [994]. Pallor [1076]. palmata [694]. Paloheimo [283]. Paludicola [1300]. Pancreas [810]. Pancreatic [1410, 1489, 1585, 1049, 1515]. Pandalid [459]. Pandalus [844, 852]. Panel [458, 712]. Paper [358]. Parabassogigas [1428]. Paralichthys [1394]. Paralithodes [171, 1058, 1014, 1539, 485]. Paralytic [578, 296, 730]. Parapenaeus [1557]. Parasite [1353, 550, 573, 231, 1560, 1355, 1363]. Parasites [979, 1500, 1373, 598, 961, 1352]. Parasitic [507, 1296, 428, 1121, 1364, 1306, 1405, 423, 1371, 658, 278]. Parasitism [1006, 1216]. Parasudis [1053]. Parental [543]. Paricelinus [287]. Park [1359, 38, 567, 889]. parma [1459]. Parophrys [1062, 657, 1480]. Parr [1552, 888, 1189, 765, 1468, 1227]. Parry [1524]. Part [1369, 169]. Partial [347, 887, 132, 675, 566]. Partially [695]. Particular [412, 858, 834, 776, 1026, 298, 128, 1346]. Particularly [215]. Particulate [1327, 884]. Pasiphaea [1461]. Passage [36]. Passages [205]. Passamaquoddy [11, 36, 55, 59, 56, 14, 203, 90, 205, 39]. Passive [469]. Pasteurizing [825]. Pathogen [1083, 1013]. Pathogenicity [1088]. Pathology [1340]. Pattern [1473, 806]. Patterns [1137, 1505, 1263, 841, 358, 980, 238, 852]. paucispinis [801, 751]. Peamouth [1136]. Pectinidae [1152]. Pectoral [1382]. Pelage [666, 841]. Pelagic [862, 423, 799, 1110, 912, 1166]. Pellet [1438]. Pen [1427]. Penobscot [931]. Perca [656, 1422, 782, 1354, 1355]. Perception [328, 165, 739]. Perch [617, 785, 656, 1084, 1422, 782, 1354, 1426, 732, 1531, 313, 1213, 1355, 1501, 87]. Percidae [772]. Performance [897, 441, 945, 1024]. Perfused [1453]. Periodicals [337, 356, 375, 397, 420, 452, 475, 503, 533, 557, 575, 592, 614, 630, 641, 653, 669, 686, 702, 718, 735, 752, 769, 786, 802, 824, 845, 866, 882, 903, 923, 940, 956, 970, 984, 1002, 1038, 1057, 1077, 1097, 1115, 1134, 1156, 1180, 1199, 1219, 1243, 1298, 1324, 1342, 1404, 1432, 1460, 1495, 1523, 1547, 1570]. Periodicities [1118]. Periodicity [1321, 177]. Periods [1587]. Peritoneal [1410]. perlans [1186]. Permanence [1145]. Permanent [1114]. Peru [1345]. Pesticide [1030]. Pesticides [1337]. Peters [406, 332]. Petersen [295, 334]. Petrale [1531]. Petromyzon [538, 41, 265, 358, 21, 143]. Petromyzonidae [489, 894]. Petromyzontidae [772]. pH [1314, 189, 535, 672]. Phenotypic [879, 1313]. Phenoxyethanol [326]. Philometridae [1365]. Philonema [1365]. Phoca [1317, 666, 838, 841]. Phocanema [1364]. Pholis [640, 426]. Phosphamidon [872]. Phosphatase [546, 626]. Phosphatases [1161]. Phosphate [1576, 1516, 621, 107, 1271]. Phospholipase [817, 1008]. Phospholipid [196]. Phospholipids [1415]. Phosphorus [40, 1157, 25, 1147, 182, 28]. Phosphorus-containing [28]. Photograph [731]. Photographic [146]. Photographing [166]. Photography [570]. Photoperiod [597, 1189, 637]. Photoperiods [317, 512]. Photosynthesis [1321, 351]. Phrixocephalus [848, 1363]. Phylogeny [128]. Physical [162, 191, 206, 621, 1247, 891, 943, 1073, 150]. Physicochemical [859]. Physiographic [1293]. Physiological [1442, 1193, 1497]. Physiology [1332]. Physoclist [607]. Phytanic [858]. Phytoplankter [1164]. Phytoplankters [546, 350, 139, 140]. Phytoplankton [645, 122, 80, 1126, 989, 572, 958, 1321, 351, 1299, 774, 1420, 613, 1378, 1212]. Phytoplanktonic [817]. Pickle [131]. Pictou [1055, 1451]. Pigeon [1119]. Pigment [140, 1419, 613]. Pigmentation [820, 1309]. Pigments [1556, 1214, 982, 715]. Pike [1294, 757, 50, 394, 472, 596, 782, 893, 433, 755]. Pikes [208]. Pilchard [412]. Pilot [673, 998]. Pimephales [1510]. Pinaus [843]. Pink [121, 603, 700, 742, 1353, 811, 991, 438, 457, 939, 431, 606, 740, 1087, 680, 425, 1195, 852, 932, 1259]. pinniger [1089, 751]. Pinniped [1188]. Pisces [1150, 243, 638, 1384, 1560, 1546, 1323, 1027]. Piscicola [581]. Piscicolidae [581]. Piscine [501]. Pituitary [1149, 1146, 1379, 1441, 389, 949]. Placopecten [578, 1208, 264, 1139]. Plaice [1096, 312, 387, 671, 895, 1389, 526, 1383, 1085, 1579, 595]. planeri [1583]. Plankton [1174, 281, 299, 333, 352, 950, 456, 767, 30, 384, 1472, 878, 1455]. Planktonic [1524, 1304, 390, 548, 734]. Plant [673]. Planted [1507]. Planting [155]. Plantings [1253]. Plasma [1159, 1149, 1025, 1167, 10, 111]. Plastic [995]. Plastic-Bag [995]. Plasticity [967]. Plate [1411]. Plateau [981, 1329]. platessa [1085, 1579]. platessoides [312, 387, 671, 895, 526, 1383]. Platichthys [1297]. platypus [1058]. Plecoptera [521, 1143]. Pleistocene [792]. Plerocercoid [1347]. Plerocercoids [6]. Plerocereoids [880]. Pleuronectes [1085, 1579]. Pliocene [1188]. Poacher [1411]. Poey [181]. Point [636]. Poison [578]. Poisoning [1549]. Poisonous [531]. poissons [1367]. Polar [93]. Polinices [1064, 1409]. Pollachius [318]. Pollett [152]. Pollock [318]. Polluted [1337]. Pollution [1397]. Polychaeta [1527, 1567]. Polychaete [1591]. Polyipnus [587]. Polymodal [223]. Polymorphic [1520]. Polymorphism [1402, 1264, 1529]. Polyphosphate [545, 346, 382, 458]. Polyphosphate-Treated [545]. Polyphosphates [482]. Polyteny [1326]. Polyzoa [1221]. Pomoxis [1043]. Pomphorhynchus [1362]. Pond [258, 946, 1039]. Ponds [145]. Pontonema [22]. Population [1308, 976, 1224, 243, 815, 963, 688, 757, 443, 325, 495, 924, 1413, 53, 446, 235, 1066, 73, 251, 32, 514, 841, 1049, 1186, 404]. Populations [759, 1295, 1252, 85, 1321, 582, 1041, 1416, 839, 1522, 583, 361, 72, 678, 919, 1360, 738]. Porbeagle [311, 418]. Porgy [223]. Porpoise [302]. Porrocaecum [35, 231, 64]. Port [547]. Portable [714, 731, 1017]. Portlock [167]. Position [1315, 1173]. Possible [1415, 203, 1293, 391, 39, 1273]. Post [484, 544, 611, 105, 212, 46]. Post-Mortem [484]. Post-Smolt [611]. Post-Spawning [544]. Postlarvae [591]. Postmortem [955, 859, 1158]. Postpartum [666]. Postrigor [944, 877]. Postulated [1070]. Pot [863]. Potassium [827]. Potential [1532, 794, 1333]. poutassou [286]. Power [1530, 56, 13, 203, 90, 39]. PPI [570]. PPI-Scope [570]. Precipitin [1401]. Precipitin-like [1401]. Precise [1290]. Precocious [611]. Precursors [407]. Predaceous [1418]. Predation [897, 1575, 574, 823]. Predator [208, 644, 53, 1200]. Predicted [13]. Preference [23, 430]. Preliminary [978, 1500, 1484, 1532, 573, 1536, 156, 1131, 818, 247, 1488, 434, 424, 1330, 123, 517, 1580]. Premating [737]. Premises [379]. Preparation [662, 1149, 942, 107, 216]. Preparations [83, 1114]. Preparing [79]. Prerigor [944, 877, 1528, 1147]. Presence [679, 1150]. Preservation [1178, 881, 665, 1251, 805, 409, 1203, 487, 1047]. Preserved [1403]. Preserving [675, 566]. Prespawning [874]. Press [714]. Pressure [1314, 1135]. pretiosus [1036, 857, 1036]. Prevent [1499]. Prey [208, 644, 53]. Prickleback [804]. Primary [381, 1414, 972, 981, 1201]. Prince [32, 1293, 551, 1268, 256, 1241, 1341]. princeps [1394]. Principal [1228]. Prionace [417, 419]. Prior [1281]. Pristanic [858]. Probable [90]. Problem [490, 323, 393, 374]. Problems [1496, 429, 1157]. Procedure [1458, 661]. Procedures [1526]. Procercoids [594]. Process [988, 674]. Processed [733]. Processes [125, 814, 1044]. Produced [405, 1577, 1301, 43]. Producing [1216]. Product [733]. Production [1202, 1104, 1449, 1448, 381, 1463, 1299, 1413, 1414, 981, 795, 1325, 1079, 1293, 449, 1065]. Productivity [989, 775, 972, 1392, 276, 1201]. Products [70]. productus [947, 1588]. Profundal [892]. Proglottid [1347]. Program [828]. Programs [1172]. Progressive [1508, 33]. Project [36, 93, 56, 203, 90]. Prolactin [952, 1149, 1441, 1149]. Prolonged [402]. promelas [1510]. Propagated [897]. Propagation [272]. Properties [1127, 191, 662, 944, 1078, 1393, 224]. Propiothetin [558, 655, 645, 842, 1042]. Proposed [56, 453, 13, 366, 90]. Prosobranchiata [1064, 1409]. Prosopium [1081, 583, 620]. Protective [902]. Protein [929, 1090, 1202, 1538, 1469, 292, 401, 1449, 1448, 881, 8, 10, 81, 382, 83, 647, 877, 478, 1402, 216, 463, 535, 1235, 1529, 346]. Proteins [1136, 865, 186, 1167, 795, 992, 980, 124, 543, 911, 111]. Proteolytic [82].

Protozoa [1304]. Provided [145]. Provinces [60, 1585]. Proximate [137, 1464, 117, 118]. Pseudocalanus [1326]. Pseudophyllidea [1369]. Pseudopleuronectes [1549, 269, 565, 1132]. Pterolamiops [181]. pubescens [366]. Publications [803, 1274, 1592, 75, 474, 615, 1179, 1494]. Published [337, 356, 375, 397, 420, 452, 475, 503, 533, 557, 575, 592, 614, 630, 641, 653, 669, 686, 702, 718, 735, 752, 769, 786, 802, 824, 845, 866, 882, 903, 923, 940, 956, 970, 984, 1002, 1038, 1057, 1077, 1097, 1115, 1134, 1156, 1180, 1199, 1219, 1243, 1298, 1324, 1342, 1404, 1432, 1460, 1495, 1523, 1547, 1570]. Puffer [198, 249]. Puget [1280, 1304]. Pugheadedness [782, 1568]. pulex [1237]. Pulp [225]. Pulsed [1144]. Pump [413]. Pumping [950]. Pumpkinseed [515]. punctatus [1396]. Pungitius [233]. Puppies [6]. Pupping [1317]. Pups [35]. Pure [1120]. Purine [107]. Purines [1450, 888]. Pycnogonida [319]. Pygmy [797, 583]. Pyloric [889, 789]. Pyrgopsis [728]. Pyrimidine [231, 107]. Pyruvate [172, 687].

Quahaug [530]. Quahaugs [506]. Qualitative [30]. Qualities [385]. Quality [188, 849, 103, 102, 545, 787, 467, 1040, 81, 516, 1086, 647, 145, 478, 1044, 66, 303, 1518, 1589]. Quantitative [1410, 1172, 192, 30, 1157, 446, 182]. Quantities [1025]. Quantity [195]. quatre [1187]. quaylei [968]. Quebec [1270, 1380, 1187, 1186, 1187]. Queen [96, 1046, 975]. Quelques [1372, 1329]. Quinte [1213, 1355]. Quoddy [63, 30, 207]. Quotients [1169].

R. [1561]. Rabbit [1516]. radiata [1245]. Radiation [665, 805, 409, 825]. Radioactive [110, 157]. Radioactivity [1212]. Radioisotope [1258]. Rafinesque [20, 321]. Rainbow [1006, 879, 1318, 784, 634, 1159, 1443, 34, 172, 470, 749, 687, 887, 988, 1146, 1453, 720, 621, 160, 1031, 1022, 927, 1590, 91, 189, 830, 1008, 1169, 1336, 1578, 1320, 1475, 1193, 706, 169, 943, 1073, 161, 1278, 754, 795, 539, 195, 654, 980, 1438]. Raja [414, 1245, 499, 561]. Raker [889]. Rancid [453]. Rancidity [482, 454, 455, 564, 741, 619, 1080, 1498, 453]. Randomness [939]. Range [1456, 1319, 797, 1567, 87, 1394, 1036, 226, 498, 747, 751, 1111]. rapax [1371]. Rapid [589, 1093, 1545, 1240, 854]. Rapids [1483]. Rare [732, 331, 505]. Rate [604, 1504, 1437, 468, 758, 1482, 1563, 202, 438, 361, 1347, 1200, 562, 248]. Rates [1260, 1184, 243, 350, 1118, 158, 609, 516, 72, 1528, 185, 431, 606, 681]. Ratio [84]. Ration [1504, 1085, 1579]. Rations [6]. Ratios [538]. Raw [881, 1485]. Rays [1382]. Reaction [1538, 793, 1401]. Reactions [317]. Reactor [558]. Reading [1550]. Reagent [1424]. Reared [1020, 1058, 1207, 569, 831, 611, 1444]. Rearing [491, 1544]. Recaptured [529]. Recently [538, 1194]. récoltés [1367]. Recompression [1017]. Reconnaissance [115, 29]. Record [287, 588, 1517, 354, 1396, 305, 797, 100, 798, 728, 937, 168, 610, 587, 204, 198, 676, 823, 1341]. Recorded [355, 1111]. Recording [146, 570]. Records [568, 147, 1262, 1037, 373, 1270, 1428, 638, 1411, 1535, 796, 113, 60, 1546, 717, 1016, 732, 288, 521, 253, 1102, 181, 639, 747, 1238, 1230]. Recovery [365]. Recruitment [1260, 272, 828]. Rectum [789]. Recurrent [1142]. Red [554, 234, 787, 696, 1194, 1177, 730]. Redbelly [1341]. Redd [457]. Redds [939]. Redescription [428, 399, 804, 331]. Redfish [1107, 1093, 766, 19, 814, 658, 2, 912]. Redmouth [539]. Redside [1136, 91, 298]. Reducing [690]. Reduction [907, 1418, 183]. Reed [667]. reedi [962]. Reef [1257]. Refeeding [705]. Reference [412, 858, 834, 851, 1121, 776, 1530, 1026, 48, 392, 582, 30, 298, 971, 128, 276, 1346]. Reflectivity [437]. Refrigerated [343]. Refrozen [1044]. Regeneration [837]. Regime [813]. Regimes [1421]. Region [507, 63, 1141, 699, 1384, 4, 30, 567, 14, 18, 1140, 587, 738, 253, 207]. Regions [427]. regius [306]. Regression [1285]. Regulation [602, 322, 924]. Rehabilitation [1141]. reinhardi [68]. Reinhardtius [1266, 439, 639]. Reinterpretation [1090]. Related [771, 807, 1305, 830, 808, 1339, 366, 508]. Relation [844, 144, 1202, 333, 1098, 515, 604, 1120, 234, 103, 1473, 1551, 720, 1148, 1245, 376, 1398, 361, 213, 943, 1073, 765, 692, 730, 1079, 910, 391, 946, 105, 535, 827, 945, 1504]. Relations [1427, 726]. Relationship [162, 352, 812, 701, 210, 999, 4, 644, 650, 1085, 1579, 899, 992, 1059, 1363]. Relationships [299, 208, 623, 1501]. Relative [790, 609, 219]. Relatives [1092]. Relaxed [955]. Release [811]. Released [1444]. Relict [221]. relicta [37, 448, 843]. Remarks [147]. Removal [45, 86, 710]. Renal [1332]. Report [1500, 1488, 1072, 1580]. Reporting [130]. Representative [128]. Representatives [276, 772]. Reproducibly [308]. Reproduction [459, 408, 154, 379, 1183, 145, 1539, 485, 1205, 1213, 658]. Reproductive [1469, 847, 290, 1466, 1261, 681]. Reptiles [1350]. Required [565]. Requirements [1288]. rerio [710, 910]. Research [568, 75, 76, 148, 229, 338, 337, 474, 476, 356, 375, 397, 420, 452, 475, 503, 533, 557, 575, 592, 614, 630, 641, 653, 669, 686, 702, 718, 735, 752, 769, 786, 802, 824, 845, 866, 882, 903, 923, 940, 956, 970, 984, 1002, 1179, 1038, 1057, 1077, 1097, 1115, 1134, 1156, 1180, 1199, 1219, 1243, 1494, 667, 713, 1350, 232, 904]. Resemblances [561]. Resembling [384]. Reserves [85, 169]. Reservoir [1126, 1457]. Resident [85, 757]. Residual [957]. Residue [1030]. Residues [1034, 1184]. Resistance [268, 733, 1505, 1231, 1171, 169, 433, 710, 910, 261]. resorts [593]. Resources [1532]. Respect [601]. Respiration [490, 1463, 1249, 736, 908, 1135, 260, 993]. Respiratory [215, 441, 515, 1530, 1169]. Response [816, 1117, 1578, 165, 754, 1069, 1438, 1515]. Responses [255, 719, 708, 1277, 1471, 1497]. Responsible [1450]. Resultant [449]. Results [573, 1131, 434, 703]. Resumption [237]. Retained [996]. Retention [1042, 1090, 1070, 605]. reticulatus [406]. retina [1329]. Retinal [496]. rétine [1329]. rétinienne [1197]. Retinomotor [386, 1048, 523]. Return [806]. Returning [1423]. Returns [831]. Revealed [1252]. Reversal [1149]. Review [359, 1496, 1407, 151, 1406, 1191, 676, 1102, 212, 70]. Revision [1559, 1357].

Rhabdocoel [1533]. Rheotaxis [1163]. Rhinichthys [1561, 239, 1148]. Rhode [565]. Rhodymenia [694]. Rhynchobdellae [581]. Rhythms [1048]. Ribose [46]. ricei [1270]. richardi [838, 841]. Richardson [428]. richardsoni [489]. Richardsonius [1136, 91, 298]. Ricker [828]. Rigor [234, 125, 1044, 416, 105]. Ringed [683]. Rings [509]. Ripening [810]. Risso [286]. River [700, 1303, 1353, 885, 931, 190, 880, 1211, 1543, 1313, 540, 1278, 1087, 1153, 806, 1302, 1248, 62, 551, 1268, 29, 852, 663, 700, 527, 239, 991, 796, 1223, 1550, 1565, 835, 1521, 703]. Rivers [1491]. Rivière [593]. Robustness [1308]. Roccus [1474, 313, 1213]. Rock [467, 468]. Rockfish [1456, 1569, 797, 96, 608, 1479, 1235, 1089, 415, 498, 639, 747, 751, 1238]. Rocky [791]. Roe [378, 856]. Rohtee [1368]. Role [352, 571, 41, 98]. Roles [1101]. Ronge [47, 1205, 114]. Ronquil [585]. rosmarus [1016, 1016]. Rostherne [914]. rostrata [952, 1332, 100, 1209, 1163, 466]. Roughly [1563]. Round [1081]. Route [1365]. rubrivinctus [1456, 498]. Rudjakov [384]. Running [45]. Runs [160]. Rupture [710].

S [223]. S. [1197, 1591, 1023, 927, 1533, 763, 498, 747, 751, 1238]. Saanich [149]. Sabellidae [1567]. Sable [1400]. Sablefish [1009, 228, 1217, 1529]. sagitta [969]. Saint [29]. saira [335]. Sakinaw [447]. salar [760, 761, 94, 760, 724, 1279, 708, 1222, 1131, 927, 888, 1189, 1416, 1568, 778, 522, 1282, 1468, 317, 512, 1521, 1362, 282, 1485, 1254]. Sales [392]. Saline [580, 785]. Salinities [1338]. Salinity [268, 1062, 23, 281, 299, 333, 352, 58, 344, 1556, 1551, 104, 642, 1265, 500, 151, 430, 750, 1311, 918, 637, 27, 142, 207, 601]. Salinity-Induced [1556]. Salmincola [1559]. Salmo [94, 784, 634, 1159, 1443, 34, 172, 470, 749, 687, 760, 724, 761, 887, 885, 1129, 1473, 1279, 295, 988, 629, 1146, 708, 1222, 1092, 621, 571, 1023, 112, 1031, 1022, 753, 1385, 927, 91, 888, 1189, 189, 1336, 1578, 1416, 1568, 1244, 971, 778, 943, 1073, 161, 1278, 522, 1282, 1468, 317, 512, 1521, 1362, 754, 795, 282, 539, 1485, 195, 62, 654, 349, 980, 416, 1254, 1438, 646]. salmoides [213]. Salmon [1365, 268, 94, 711, 1234, 23, 897, 121, 1127, 690, 106, 695, 905, 444, 603, 700, 742, 1108, 1382, 760, 724, 761, 527, 936, 1353, 550, 215, 255, 441, 604, 945, 1504, 40, 217, 1279, 811, 31, 820, 1024, 1562, 1149, 1437, 1423, 1566, 1145, 146, 156, 632, 869, 982, 328, 1146, 1020, 605, 481, 708, 800, 1192, 1222, 1060, 1101, 963, 380, 571, 974, 1117, 991, 134, 1276, 112, 322, 1131, 1314, 870, 8, 927, 714, 1018, 293, 888, 1189, 10, 189, 395, 796, 190, 871, 1061, 84, 1334, 1398, 443, 438, 609, 622, 1421, 1381, 1416, 1568]. Salmon [24, 107, 424, 1113, 1481, 880, 1361, 715, 491, 1281, 1125, 651, 1540, 1333, 165, 237, 544, 810, 1379, 1558, 835, 51, 430, 1441, 457, 939, 1545, 778, 1211, 1431, 442, 765, 522, 1471, 1470, 16, 185, 323, 431, 606, 740, 1087, 876, 874, 680, 1282, 1419, 1468, 317, 512, 1521, 1362, 703, 201, 282, 1012, 1198, 184, 425, 413, 1485, 32, 342, 510, 529, 611, 780, 1011, 806, 1302, 1337, 517, 1309, 326, 563, 637, 402, 933, 948, 1028, 1283, 1072, 341, 518, 1226, 1444, 566, 827, 1047, 1584, 124, 1195, 111, 1227, 900, 875, 1386]. Salmon [1230, 396, 922, 1254, 932, 1259, 26]. salmonicida [1330]. Salmonid [581, 360, 1025, 128, 1286, 1284]. Salmonidae [1150, 961, 238, 543, 763]. Salmonids [1576, 1214, 758, 469, 928, 1585, 744, 1076]. Salmonoid [560]. Salmons [598]. salmositica [581, 796]. Salt [684, 816, 124]. Salted [133, 132, 173, 942, 297, 479, 480]. Salting [131, 478]. Saltno [1131]. Salts [482, 564, 679]. Saltwater [51]. Salvelinus [496, 1256, 616, 1150, 1054, 259, 708, 291, 290, 542, 1442, 429, 1433, 1183, 889, 1253, 924, 129, 938, 562, 763]. Sampler [995, 1174, 1542, 878, 1487]. Samples [589]. Sampling [950, 1526, 1122, 713, 232, 1325, 1455, 402, 904]. San [553]. Sand [1459, 1004, 591, 1310, 1390, 1182]. Sand-Burrowing [1310, 1390]. Sarcoplasm [1162]. Sarda [552]. Sardine [665, 202, 400]. Sardinella [665]. Sardinops [202, 400]. Saskatchewan [47, 1205, 114, 1153]. Saury [335]. Saux [1565]. Scabbard [253]. Scale [444, 1108, 84, 651, 837, 808]. Scales [603, 700, 742, 887, 626, 1294, 1192, 1207, 714, 1189, 1334, 680, 1230]. Scallop [509, 1208, 432, 877, 1139, 409, 1518, 471]. Scallops [554, 578, 1498, 264, 203]. Scatterers [17]. Scattering [1059]. schauinslandi [1366]. Schedules [1087]. Schistocephalus [1347]. Schmidt [355]. School [1395]. Schooling [1084, 661, 1050, 460]. Schools [1069]. Schultz [587]. schultzi [640]. Scintillation [1237]. Scoloplos [1591]. Scomber [1522]. scombrus [1522]. Scope [570]. Scorpaenid [227, 962]. Scorpaenidae [1235]. Scotia [1387, 357, 64, 1182, 1185, 1400, 1053, 1322, 353, 717, 1163, 819, 1071, 1055, 1451]. Scotian [289, 722, 411, 115]. Sculpin [287, 588, 1269, 1430, 1091, 1323]. Sculpins [1270, 451, 73]. Sea [951, 773, 538, 1345, 554, 578, 1519, 1353, 102, 976, 1428, 1577, 1289, 112, 469, 267, 725, 1125, 265, 937, 792, 447, 431, 606, 740, 486, 559, 731, 884, 372, 301, 473, 358, 143, 99, 904, 1018, 461, 1447, 912, 1273, 1111]. Sea-Floor [725]. Seal [1317, 408, 1228, 1366, 838, 841, 1273]. Seals [683, 1402, 519, 681]. Season [1317, 210, 809, 52]. Seasonal [55, 1524, 102, 454, 1473, 573, 1511, 1096, 648, 432, 856, 855, 376, 528, 71, 269, 145, 548, 283, 18, 729, 1436, 1468, 201, 806, 1355, 427, 401]. Seasons [1084]. Seaward [23]. Seawater [294, 1518]. sebago [760, 761]. Sebastes [766, 19, 658, 2, 912]. Sebastodes [1456, 801, 1569, 797, 608, 1479, 1089, 415, 498, 639, 747, 751, 962, 1238]. Sebastolobus [227, 1238]. Second [798]. Secretion [1437]. Sectional [292, 137]. Sections [117]. sedentaria [1527]. Sediment [1542]. Sedimentation [563]. Sediments [1526, 1425, 1463, 1458, 987, 1223]. Segregated [135]. Seine [152]. Selachians [437]. Selar [1546]. Selected [928]. Selecting [352]. Selection [963, 927, 1336, 522, 1049]. Selective [1325]. Selectivity [1398, 1534, 649]. Self [1237]. Self-Absorption [1237]. Semi [20, 1326]. Semi-buoyant [20]. Semi-Landlocked [1326]. Seminal [1314]. Senecella [1524]. Sensitivity [501]. Separated [111]. Separating [1425, 561]. Separation [1154, 124, 480]. Septibranch [1015, 1484]. Sera [1276]. Series [1172, 706]. Serological [839]. Serologically [400]. Serum [1469, 134, 1520, 795, 992, 358, 980, 1503]. Seston [767]. setacea [1099].

Several [1450]. Severe [749, 687]. Severely [242]. Sevin [1165]. Sex [538]. Sexual [538, 1586, 65, 1092, 671, 960, 611, 827, 949]. Sexually [24, 810, 1558]. Shad [836]. Shallows [526]. Shanny [1396, 732]. Shape [1534]. Shark [1287, 417, 419, 553, 168, 181, 302]. Sharks [307, 321]. Shears [1492]. Sheepshead [907, 913]. Shelf [93, 1255, 814, 1479, 1329, 722, 957]. Shell [1232, 899, 22]. Shell-Weight [899]. Shellfish [834, 578, 153, 296, 730, 920, 954]. Shells [530, 551]. Shelter [1277]. Shiner [1136, 650, 1501]. Shiners [91, 298, 1426]. Shipworms [203]. Shock [719, 602]. Shocking [41]. Shortbelly [639]. Shown [868]. Shrimp [844, 1461, 1557, 33, 863, 1339, 487, 508, 852]. Shrimps [459, 373, 721, 799, 226]. Shrinkage [864]. Sibling [833]. Side [667]. Side-Trawl [667]. Significance [1232, 1223, 491, 850]. Silicate [1271]. Silt [514]. silus [722, 1360]. Silver [394, 732]. Silvering [1450]. Similar [413]. Similarity [1446]. Simple [743, 1483, 682]. Simplification [1260]. Simplified [997]. Simulate [1457]. Simulated [1537]. Simulating [1105, 1483]. Simulation [443]. Simultaneous [268]. Single [1184, 712, 1086]. Sinistrality [1297]. Site [1101]. situ [187]. Six [553]. Six-Gilled [553]. Sixteenmile [776, 298]. Sixth [1423]. Size [657, 1462, 604, 1504, 573, 520, 112, 809, 1395, 404, 1292, 1543, 678, 726, 16, 650, 610, 1383, 1325, 413, 1525, 884, 449, 1501, 1326]. Size-Class [678]. Size-Selective [1325]. Sizes [235]. Skate [414]. Skates [505, 499]. Skeena [449]. Skeletal [1105, 1162, 212]. Skeletonema [405]. Skeletons [1294]. Skidegate [1175]. Skin [982, 1189, 190, 1381, 715, 436, 1076]. Skin-diving [190]. Skyes [1368]. Slave [154, 113, 257, 392, 177]. Slicing [86]. Slope [1255, 1166, 363]. Slopes [332]. Small [745, 1068, 950, 1094, 1025, 1026, 790, 270, 1041, 909, 1380, 390, 981, 251, 342, 946]. Smallmouth [812, 701, 1277]. Smelt [857, 1036]. Smoking [478]. Smolt [1552, 888, 1189, 1211, 1468, 611, 1011, 1227]. Smolting [322]. Smolts [974, 880]. Snailfish [305]. Snails [551]. Snowdrift [612]. Social [1198, 1226]. Sockeye [897, 742, 1108, 1382, 441, 604, 945, 1504, 40, 1437, 632, 982, 328, 1146, 605, 1060, 963, 380, 974, 1117, 134, 1276, 8, 293, 10, 189, 395, 796, 1061, 1398, 880, 1125, 651, 1540, 1333, 165, 237, 544, 810, 1379, 1558, 835, 1301, 1441, 323, 1012, 184, 326, 449, 948, 827, 1386, 932, 1259, 26, 949]. Sodium [564, 1440, 45, 559, 473, 1147, 827]. Soft [22, 1232, 1510]. Soft-Shell [1232, 22]. Sole [1062, 657, 467, 468, 1175, 1004, 1531, 1480, 1111]. Solids [920]. solidus [1347]. Soluble [124, 182]. Solubles [401, 292]. Solution [492]. Solutions [518, 1147, 124]. Solvent [929, 1449]. Some [707, 858, 834, 1042, 268, 1062, 759, 1035, 546, 979, 643, 215, 255, 1532, 501, 684, 1556, 944, 580, 807, 719, 1552, 137, 1435, 481, 800, 141, 48, 1078, 1442, 1245, 350, 1045, 1247, 705, 1118, 1124, 379, 1587, 609, 896, 1391, 516, 1393, 651, 1375, 1534, 1406, 1544, 734, 18, 435, 1436, 364, 1330, 926, 703, 1390, 142, 1537, 499, 561, 1352, 238, 993, 1227, 1284, 1239, 1356, 1329, 134, 928, 1372]. Somniosus [1287]. Sound [1524, 1004, 17, 1144, 1059, 96, 1280, 1304, 1481]. Sounder [570]. Sounding [1117]. Source [598, 163, 296, 46]. Sources [1206, 15, 81]. South [1507, 976, 688, 813, 1348]. Southeast [926]. Southern [699, 819, 1006, 635, 1126, 1307, 1434, 270, 725, 12, 1446, 1436, 1066, 201, 831, 791]. Southwest [1572, 556, 353]. Southwestern [1270, 264, 1322, 1023]. sp [1359, 1368, 1560, 1155, 423, 1370, 574, 947]. Span [1361]. Spatial [277, 989]. Spawning [106, 40, 1423, 632, 776, 210, 1101, 160, 8, 753, 10, 1061, 809, 536, 544, 1583, 353, 457, 1446, 1323, 446, 671, 780, 449, 300, 1230, 396, 26, 949, 1186]. Special [851, 1121, 1530, 392, 582, 30, 971, 1492, 1356]. Speciation [1373, 663]. Specie [352]. Species [978, 546, 23, 299, 352, 1152, 1484, 444, 1137, 862, 1255, 1358, 1142, 137, 239, 776, 579, 1384, 96, 1207, 388, 279, 833, 1520, 689, 826, 423, 1191, 139, 140, 1420, 1143, 675, 226, 1497, 1390, 1082, 366, 477, 561, 676, 238, 494, 543, 623, 21, 489, 284, 498, 747, 751, 1238, 922, 1225, 1372]. Specificity [979, 494]. Specimen [1459]. Specimens [868, 1033]. Speckled [145]. Spectrum [884]. Speed [54, 515, 604, 781, 1551, 1303]. Speeds [632, 909]. Spelling [1019]. Sperm [1403, 1131, 1251, 1047, 932, 1259]. Spermatozoa [605, 1584]. Sphyrion [658, 2]. Spider [1170, 1171]. Spinal [1173]. Spinner [1397]. Spiruroidea [1357]. Spoilage [102, 33, 1531]. Spontaneous [1282]. Spore [1065]. Spores [954, 854]. Sporozoa [1346]. Spottail [1426, 650]. spp [1557, 658, 594]. Spraying [869, 872, 870, 871]. Sprayings [868, 867]. Spring [280, 336, 1446, 660, 201]. Spring- [1446]. Spurious [679]. Squanga [279]. Squid [1427, 488, 906]. St [743, 264, 1387, 511, 725, 1546, 1380, 289, 1446, 1066, 1273]. St. [1185, 1223, 1389, 1267]. Stable [1138]. Stage [1044]. Stages [20, 304, 1391, 265, 227, 1011, 365]. Staghorn [1269]. Stainless [336]. Stains [1340, 1041]. Standard [8]. Standardization [1411]. Standardized [936]. Standing [1185, 1280, 1418]. Stannius [1332, 727]. Starch [1128, 1263, 1158, 698]. Starry [1004]. Stars [773]. Starvation [1159, 749, 928, 705, 827]. State [796, 857, 1434]. States [1367, 1372]. Station [1145, 691, 983, 77, 27, 142]. Stations [1304, 593]. Statistical [1085]. Statistics [315, 404, 1272]. Status [579, 916, 1370, 676]. Steaks [925, 1589]. Steel [336]. Steelhead [885, 295, 1145, 1552, 1403, 571, 1023, 112, 622, 1113, 1543, 62, 416, 646]. Steganodermatidae [1351]. stellatus [1297]. stenolepis [1503]. Sterile [28]. Sterilized [628]. Steroid [668, 727]. Sterol [432, 465]. Sterols [44, 471]. Stichaeus [1396]. Stickleback [1319, 1466, 1575, 1509, 1261]. Sticklebacks [941, 270, 233, 1313]. stilbius [147]. Stimulants [328]. Stimulation [1440]. Stimuli [317]. Stingray [1110]. Stizostedion [1197, 1151, 1422, 1140]. Stochastic [1553]. Stock [1462, 1253, 1383, 449, 1049]. stocked [72]. Stocks [1418, 323, 1005, 62, 263]. Stomach [1151, 417, 622, 1481, 1004, 302]. Stomachs [1246]. Storage [1465, 695, 709, 343, 849, 1126, 1096, 1574, 196, 1044, 410, 825, 1290, 28, 310, 1212, 1259]. Stored [1538, 741, 31, 1093, 830, 1486, 79, 516, 778, 1259, 33]. Stores [155]. Storm [1482]. Strait [434, 511, 176, 657, 497, 994, 141, 590, 552, 826, 730, 398, 1573, 1273].

Stranded [1003]. Stratification [1250]. Stream [1407, 1473, 629, 328, 1483, 776, 1312, 1277, 1429, 280, 155, 1590, 270, 536, 1380, 491, 1301, 342, 514, 1337, 563, 946, 1039, 1201]. Streams [1094, 1023, 870, 753, 867, 298, 1163, 161, 1444, 670]. Strength [1226, 124]. Stress [1245, 1231, 1281, 1193]. Striated [1007, 877, 1139]. Strigeatoid [1348]. Striped [9]. Strontium [1138]. Structural [359]. Structure [496, 1329, 1409, 294, 1534, 678, 1348, 427, 1197, 1329]. Structures [151, 13, 39]. Stuart [445]. Studies [1168, 1526, 102, 545, 40, 771, 156, 1316, 8, 324, 10, 867, 434, 516, 953, 77, 14, 890, 53, 126, 502, 211, 64, 26]. Study [978, 507, 215, 1554, 556, 67, 1450, 1314, 30, 1327, 24, 620, 129, 123, 74, 1367]. Sturgeon [1153, 1248, 1119]. Sub [1047, 7]. Sub-Arctic [7]. Sub-Zero [1047]. Subarctic [369, 1206, 985, 127, 699, 1384, 883, 660, 240]. Subcellular [784]. Subcutaneous [1215]. Subdivisions [1123]. Subfamily [1315]. Subjective [188]. Sublethal [1256, 719, 1184, 602, 1231, 1183]. Sublittoral [1536]. Submarine [445]. Subpopulations [400]. Subsampler [390]. Subsequent [1482]. Subspecies [384, 150]. Substance [268]. Substances [501, 1381]. Substrates [1176]. Subsurface [1229]. Subtidal [1390]. Subtropical [1348]. Success [582, 258]. Successful [448]. Succession [1082, 143]. Sucker [1388, 1331, 976, 154, 495]. Sueur [488]. Sugars [182, 508]. Suggested [1193]. Sulfide [1232]. Sulphate [1486, 710, 910]. Sulphide [842, 374]. Sulphite [225]. Sulphur [462]. Sulphuric [1458]. Summarising [1172]. Summary [867]. Summer [636, 616, 724, 761, 1250, 382, 254]. Sunfish [847, 794, 597, 748, 1439]. Superimposition [457]. Superior [1388, 1141, 266, 219, 569, 831, 1140]. Supplementary [603]. Supply [406]. Surf [857, 1036]. Surface [281, 1032, 15, 104, 774, 1271, 1487, 739]. Surfaces [936, 829]. Surge [1482]. Surgical [933]. Survey [1536, 892, 601]. Survival [1062, 58, 1220, 1507, 1165, 1412, 237, 1301, 513, 1211, 1286, 1311, 1011, 197, 1386, 256]. Susceptibilities [1013]. Sustained [604, 1422, 339, 1537]. Swiftsure [691]. Swim [1357]. Swimbladder [191, 607, 510]. Swimming [897, 643, 1032, 54, 441, 515, 604, 945, 295, 1024, 632, 633, 648, 1551, 1422, 1398, 909, 402, 948]. Swordfish [787, 1417, 118, 1103, 1582]. symmetricus [1095, 1050]. Sympatric [1532, 239, 279]. Syndesmis [1533]. Synonymy [1369, 428, 1591, 439]. Synopsis [1434]. Synthesis [44]. Synthetic [1289]. Systaltic [246]. System [950, 1064, 239, 991, 898, 293, 880, 972, 1313, 540, 1039]. Systematic [129]. Systematics [1369, 493, 623, 1235]. Systems [840, 107, 1503].

T [223]. T. [1344]. Taeniopteryx [1143]. Tag [1217, 336, 163, 1362, 1215, 517, 315]. Tagged [257, 256]. Tagging [1032, 295, 315, 703, 1011, 1476, 174]. Taggings [357]. Tags [295, 1566, 325, 130]. Taius [223]. Taken [417, 1350, 975]. Taking [219]. tamarensis [934]. Tank [745, 811, 1483, 1457]. Tanquary [851]. Tanypleuridae [1405]. Tapeta [437]. Tapeworm [1345]. Taste [712, 458].

Tattoo [1214]. Tattoo-Marking [1214]. Taxonomic [579, 1370, 1173]. Taxonomy [147, 262, 1533, 391]. Taxus [1001]. Technique [570, 682, 1240, 402, 389]. Techniques [146, 933, 300]. Teeth [894]. Teleost [1167, 386, 969, 682, 1114, 1056, 1173]. téléostéens [1329]. Teleostei [1560]. Teleosts [407, 1146, 137, 1329]. Television [54]. Temperature [268, 1062, 1, 281, 299, 333, 352, 57, 945, 1504, 617, 812, 1541, 1513, 1516, 104, 642, 125, 1040, 720, 1312, 290, 928, 927, 376, 350, 704, 1189, 151, 68, 275, 1170, 1171, 213, 522, 692, 1282, 1311, 1439, 1290, 637, 261, 1514, 1190, 562, 27, 142, 207, 5, 427, 284, 248]. Temperatures [144, 1318, 616, 724, 761, 1229, 1265, 1124, 1000, 1194, 403, 265, 1282]. Temporal [989, 192, 142]. Tendency [453]. Term [953, 1537]. Terminology [1411, 894]. Terns [1375]. Terranova [278]. Territories [1493, 392, 612]. Territory [279]. Test [1419]. Tests [897, 188, 1192]. Tetrachloro [1562]. Tetrachloro-o-benzoquinone [1562]. Tetracycline [188, 487, 875]. Tetraodontidae [198]. Tetrarhynchus [1372]. Tétrarhynques [1372]. Tetronychoteuthis [288]. Thalassinidea [477]. Thaleichthys [450]. Thames [1550]. Thaw [1096, 458, 403, 545]. Thaw-Drip [458, 403, 545]. Thawed [1040, 1074, 1577]. Thawing [403, 348, 814, 1044]. Their [144, 1353, 762, 1561, 1423, 633, 1246, 1373, 51, 184, 1360, 873, 142, 238, 493, 543, 894, 1192, 1535, 433]. Theoretical [379, 565]. Theory [490, 1555]. Thermal [733, 719, 602, 927, 1193, 813, 433, 1201]. Thermoacclimatory [988]. These [561]. Thin [955, 479]. Thin-Layer [955, 479]. Thiobarbituric [679, 762]. Thiosulphate [45]. Third [1195]. thomasi [1418]. thompsonii [1270]. Thornback [287]. Three [1137, 1532, 1316, 1095, 1026, 252, 1349, 1125, 145, 919, 1390, 744, 253, 1582, 751, 1238, 108]. Threespine [941, 1313]. Thresholds [1120]. Thrust [1398]. Thunnus [144, 861]. Thymallus [128, 368]. Thyroid [23, 810]. Tidal [314, 1555, 13, 601, 39, 1512]. Tide [1548, 1512]. Tidewaters [1122]. Tilapia [1551, 243]. Tilesius [171, 340, 485, 176, 303]. Time [1, 945, 1454, 1070, 570, 586, 680]. Time-Lapse [570]. Times [742]. Timing [536, 1364]. tip [181]. Tissue [558, 952, 1074, 1489, 1558, 274, 1056]. Tissues [1127, 1159, 762, 40, 1233, 1184, 155, 395, 71, 793, 26]. Titanium [1223]. Toad [1440]. Tocopherol [834]. Tolerance [57, 750, 473]. Tolerances [1497]. Tone [1120]. Topminnow [739]. Total [617, 812, 10, 920]. Toxaphene [677]. Toxic [268, 1562]. Toxicities [900]. Toxicity [690, 1510, 153, 730, 934]. Toxin [817, 296, 1216]. Toxin-Producing [1216]. Toxins [1478]. Trachurus [1095, 1050]. Track [1305]. Tracking [1474]. Tract [507, 764]. Transaminase [1127]. Transaminases [1127]. Transatlantic [966]. Transfer [490]. Transferrin [1503, 1588]. Transformation [1552, 888, 1189, 1227]. Transformations [727]. Transformed [538]. Transition [663]. Transitional [601]. Transmission [1359, 1499, 1275]. transmontanus [1248]. Transplantation [1252]. Transplanted [700]. Transponding [537]. Transport [369, 222, 1440, 853]. Trap [484, 1429, 382, 458, 516, 859, 1086, 1286]. Trap- [484]. Trap-Caught [458, 859, 1086]. Traversed [1471]. Trawl [756, 1542, 484, 667, 627, 975, 1069, 1059, 904]. Trawl-Board [1542]. Trawl-Caught [484]. Trawler [103, 102]. Trawler-Caught [103, 102]. Treated [545]. Treatment [1437, 346, 382, 458, 1531, 1017]. Trematoda [1358]. Trematode [1351, 1360]. Trematodes [1358, 1349, 1361, 1356, 1367, 1367]. Trench [791]. Trends [1229, 520]. Triaenophorus [1493, 555, 6, 304, 1242, 1354, 880, 284, 248, 594]. Trials [517]. Tributary [636, 380]. Trichodinids [180]. Tricladida [1300]. tricuspis [1269]. Triglops [1323]. Triglycerides [1415, 1443]. Trimethylamine [69, 98, 407, 103, 1104, 43]. Trimethylammonium [69, 98, 178, 407]. tripes [901]. Triphosphate [212, 416]. Tripolyphosphate [1147]. Trips [1423]. Tris [473]. Tropical [1035, 1194]. Tropomyosin [119, 220]. Trout [1256, 1006, 879, 1318, 784, 634, 1159, 1443, 1410, 34, 172, 470, 749, 687, 616, 887, 1137, 1220, 1507, 885, 662, 935, 1054, 1129, 1473, 295, 988, 1513, 1141, 1100, 1145, 629, 1146, 1552, 816, 1359, 1453, 720, 708, 800, 1403, 1092, 1026, 621, 160, 280, 571, 1023, 291, 290, 840, 85, 155, 542, 112, 1031, 1442, 1022, 1412, 753, 1385, 927, 1590, 91, 189, 830, 1008, 190, 1169, 1336, 1578, 266, 622, 1320, 1489, 429, 1244, 1433, 1231, 1183, 1113, 38, 889, 1475, 87, 1193, 971, 145, 706, 1253, 924, 1499, 1545, 169, 943, 1073, 1543]. Trout [161, 1278, 938, 754, 795, 219, 569, 831, 114, 539, 510, 514, 1337, 1309, 195, 52, 62, 258, 946, 1039, 1200, 654, 349, 150, 980, 416, 1067, 1438, 646, 1049, 670]. Trout-perch [87]. Trouts [1515]. truculentus [1053]. True [466]. trutta [1129, 1473, 629, 1092, 1244, 349]. tshawytscha [1192, 1060, 1101, 16]. Tubificid [1526, 1532]. Tubular [858]. Tubules [246]. Tubulovesicula [1361]. Tullibee [907, 913]. tumifrons [223]. Tumorlike [860]. Tuna [144, 861, 118, 966]. Turbellarian [1533]. Turbot [1009]. Turnover [1022]. Twice [382]. Twice-Frozen [382]. Two [978, 121, 743, 862, 1576, 1554, 1233, 1358, 1146, 239, 1162, 579, 1340, 96, 91, 1416, 279, 1550, 1361, 548, 981, 1520, 826, 793, 1033, 1282, 799, 184, 251, 226, 1455, 1196, 62, 1190, 561, 150, 900, 21, 284, 747, 1326, 1225, 776, 839]. Two-Dimensional [743]. Twoline [639]. Type [631, 817, 1176, 781, 1429, 644, 1191, 954, 854, 1065]. Types [143].

U.S. [1350]. Ultrasonic [1474]. Ultrasound [1074]. Ultrastructure [1007, 1100, 1139, 789]. Ultraviolet [343, 1425]. Unacclimated [57]. Uncommon [60]. Undergoing [1193]. Underwater [54, 1208, 627]. Underyearling [571]. Unexploited [252]. Unfrozen [859]. Ungava [1081, 483]. Unis [1367, 1372]. Unit [1350]. United [1367, 1372]. Unsaponifiable [432]. Unsaturated [359]. Untreated [545]. Unusual [982, 231, 783, 590, 975]. Unusually [396]. Upon [760, 602, 112, 1512]. Upper [1265, 284, 1278, 1123]. Upstream [1303]. Uptake [1453, 1118, 282, 910, 1147]. Urchin [916]. Urinary [840, 1022]. Urosalpinx [1377]. ursinus [408]. USA [1168]. Use [1127, 1237, 936, 1176, 921, 820, 1340, 325, 1397, 706, 390, 1362, 731, 409, 1203, 1360, 997, 894]. Used [1544]. Useful [1455]. Using [352, 300, 1425]. Utilization [1206, 395, 202, 736, 1211]. UV [1381]. UV-Absorbing [1381]. UV-Fluorescing [1381].

V [619, 1364]. vacillatum [22]. Vacuum [390]. Vacuum-Assisted [390]. Vaillant [676]. Valenciennes [638, 324, 197]. Validity [1371]. Valley [1270]. Value [1009, 771, 1557, 83, 913, 275, 463]. Values [103, 679, 762, 684, 1239]. Vancouver [294, 1037, 149, 151, 435, 309]. Variability [106, 691, 953, 358, 646]. Variance [1333]. Variant [801]. Variants [1276, 1588]. Variation [268, 1096, 439, 71, 233, 1313, 1052, 1509, 1326, 117]. Variations [102, 454, 292, 401, 192, 432, 345, 27, 103]. Various [1524, 385, 1265, 395, 896, 1528, 463, 675, 1302, 473, 1013]. Varying [308]. Velocity [369, 1148, 765]. Venous [687]. Ventral [1508, 850]. verrucosa [1411]. versicolor [1370]. Vertebral [758, 312]. Vertebrates [347, 493]. Vertical [369, 624, 37, 55, 1524, 807, 656, 424, 275, 512, 1455, 1190]. Vertically [1135]. Very [1093]. Vessel [667, 713, 232, 502]. Vessels [568, 102]. vetulus [1062, 657, 1480]. Viability [1259]. Victims [1017]. Victoria [243, 221]. VII [44]. villosus [652]. violacea [1110]. virens [318]. virgatum [1262]. virginica [664, 839]. Virus [1410, 1585, 1049, 1515]. Virus-Free [1049]. Visceral [81]. Visible [156]. Vision [1101, 460, 1336, 971]. Visual [255]. Vital [315, 404]. Vitamin [109]. Vitamins [694]. vitreum [1197, 1151, 1422, 1140, 1197, 1151, 1422, 1140]. Vitro [664, 1013, 278, 1279, 1320]. vitulina [838, 841, 1317, 666]. Vivo [1478]. Volatile [462, 480]. Volume [76, 148, 229, 338, 476, 734, 510]. Volumes [607, 744]. Voluntary [35]. vs [1000].

W [1344, 983, 883]. Walbaum [439, 269, 565, 204]. Walleye [701, 1422]. Walleyes [1151, 1140, 1445]. Walrus [1016]. Warty [1411]. was [1372]. Washes [1410]. Washing [316]. Washington [144, 104, 192, 796, 857, 1434, 1280, 553, 1004]. Waste [547]. Wastes [690]. Water [995, 1174, 299, 333, 352, 58, 15, 344, 292, 401, 1449, 1448, 816, 605, 1148, 1474, 1577, 1289, 112, 469, 8, 753, 1385, 376, 704, 1397, 348, 814, 1177, 41, 45, 513, 1163, 213, 1510, 765, 447, 486, 559, 730, 677, 1455, 1072, 301, 473, 27, 142, 827, 205, 547, 99, 663, 363, 115]. Water-Solubles [401, 292]. Water-Thawing [348]. Waters [144, 1300, 287, 588, 281, 299, 333, 352, 1229, 1541, 801, 985, 1269, 1430, 328, 305, 127, 414, 1250, 775, 855, 1247, 100, 60, 1481, 608, 419, 418, 998, 774, 17, 78, 1471, 1470, 423, 201, 440, 211, 355, 286, 1447, 873, 197, 25, 1166, 464, 225, 415, 639, 1238, 663, 1150, 525, 56, 120, 203]. Waves [739]. Weather [272, 77, 27]. Weight [1151, 999, 4, 1540, 734, 692, 327, 1272, 1285, 1439, 899, 1291, 992, 1147]. West [861, 580, 1109, 1306, 294]. West-Central [580]. Western [1036, 1445, 1150, 1388, 748, 1250, 175, 426, 286, 1103, 330, 2, 253, 676, 1102, 1582, 489]. Wet [1458, 297, 1151]. Wet-Oxidation [1458]. Whale [247, 1003, 199].

Whales [1477, 998, 330, 1525, 1476]. wheatlandi [1319, 1466, 1261]. which [1372]. Whirling [502]. White [1331, 1007, 234, 976, 696, 1474, 1477, 495, 1194, 168, 313, 1248, 1525, 1476, 1213, 181]. White-tip [181]. Whitecoat [666]. Whitefish [555, 1517, 272, 6, 635, 48, 252, 97, 101, 582, 194, 279, 1081, 583, 620, 1491, 805, 1394, 1205, 649, 254]. Whitefishes [194]. Whole [244]. Widow [608]. Wild [869, 1020, 1412, 1444]. williamsoni [620]. Wilson [848, 1363, 1559, 1370]. wilsoni [797]. Wind [743]. Wind-Driven [743]. Windermere [757, 562]. Winnipeg [1249, 48]. Winter [1549, 616, 724, 761, 1367, 1454, 1026, 280, 357, 269, 565, 1573, 1132, 1367]. Wintering [1477, 998]. Wire [1566]. Wisconsin [886, 1412, 1239]. Within [1023, 1315, 623, 1341]. Without [834, 605]. wolffish [819]. Wooden [829]. Woodland [1312]. Woods [473]. World [1244, 1433].

X [10]. XI [26]. XIII [40]. Xiphias [787, 1103, 1582].

Year [582, 431, 606, 893, 1195]. Year-Classes [582, 893]. Yearling [94, 1277]. Years [1125, 184]. Yellow [617, 785, 656, 1084, 1422, 1354, 1426, 1355]. Yellowfin [861]. Yellowstone [753, 1385, 971]. Yellowtail [1071]. Yew [1001]. Yield [1462, 382, 16, 323, 828, 1325, 1248, 1200, 183, 1537]. Yields [907, 339, 1383]. Yirus [1489]. Yolk [202]. Young [1365, 268, 690, 441, 209, 1562, 869, 708, 800, 228, 293, 871, 1585, 966, 808, 1278, 522, 342, 510, 1196, 341, 900, 21]. Yukon [279].

zacentrus [498]. Zalophus [937]. Zebrafish [710, 910]. zenithicus [1452]. Zero [1047]. Zinc [186, 1288, 710, 910, 341, 518]. Ziphius [1003]. Zn [1233, 1258]. Zoarcid [1506]. Zoega [1362]. Zone [973, 1528, 493, 543, 427, 911, 111]. Zones [166]. Zoogeography [429]. Zoology [1406]. Zooplankton [333, 973, 851, 1511, 192, 170, 525, 1305, 77, 1299, 275, 1413, 1418, 1502, 1573, 853].

References

Armstrong:1960:ETT

[1] Iva L. Armstrong, Edna W. Park, and Barbara A. McLaren. The effect of time and temperature of cooking on the palatability and cooking losses of frozen Atlantic codfish fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(1):1–7, January 1960. CODEN JFRBAK. ISSN 0015-296X. URL

      https://cdnsciencepub.com/doi/full/10.1139/f60-001.

Templeman:1960:IDI

[2] Wilfred Templeman and H. J. Squires. Incidence and distribution of infestation by Sphyrion lumpi (Krøyer) on the redfish, Sebastes marinus (L.), of the Western North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(1):9–31, January 1960. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f60-002.

Murphy:1960:EAL

[3] Garth I. Murphy. Estimating abundance from longline catches. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(1):33–40, January 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f60-003.

Kohler:1960:GLW

[4] A. C. Kohler. The growth, length–weight relationship, and maturity of haddock (Melanogrammus aeglefinus L.) from the region of lockeport, N.S. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(1):41–60, January 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f60-004.

Tully:1960:AIT

[5] J. P. Tully, A. J. Dodimead, and S. Tabata. An anomalous increase of temperature in the ocean off the Pacific Coast of Canada through 1957 and 1958. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(1): 61–80, January 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-005.

Crampton:1960:EIT

[6] E. W. Crampton. Effect of the ingestion of Triaenophorus plerocercoids on the nutritional characteristics of whitefish fillets in the rations of puppies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(1):81–90, January 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f60-006.

Tully:1960:EAS

[7] J. P. Tully and F. G. Barber. An estuarine analogy in the sub-Arctic Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(1):91– 112, January 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f60-007.

Idler:1960:BSS

[8] D. R. Idler and I. Bitners. Biochemical studies on sockeye salmon during spawning migration: IX. Fat, protein and water in the major internal organs and cholesterol in the liver and gonads of the standard fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(1):113–122, January 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f60-008.

Pike:1960:PSD

[9] Gordon C. Pike. Pacific striped dolphin, Lagenorhynchus obliquidens, off the coast of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(1):123–124, January 1960. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f60-009.

Jonas:1960:BSS

[10] R. E. E. Jonas and R. A. MacLeod. Biochemical studies on sockeye salmon during spawning migration: X. Glucose, total protein, nonprotein nitrogen and amino acid nitrogen in plasma. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(1):125–126, January 1960. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f60-010.

Hart:1960:IPF

[11] J. L. Hart and D. L. McKernan. International Passamaquoddy Fisheries Board Fisheries Investigations 1956–59 Introductory Account. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(2):127–131, February 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f60-011.

McKenzie:1960:HFS

[12] R. A. McKenzie and S. N. Tibbo. Herring fishery in Southern New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(2):133–168, February 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f60-012.

Martin:1960:PEP

[13] W. R. Martin. Predicted effects of proposed tidal power structures on groundfish catches in Charlotte County, N.B. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(2):169–173, February 1960. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f60-013.

McCracken:1960:SHP

[14] F. D. McCracken. Studies of haddock in the Passamaquoddy Bay region. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(2):175–180, February 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f60-014.

Bumpus:1960:SWC

[15] Dean F. Bumpus. Sources of water contributed to the Bay of Fundy by surface circulation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(2): 181–197, February 1960. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f60-015.

Parker:1960:CSM

[16] Robert R. Parker. Critical size and maximum yield for Chinook salmon (Oncorhynchus tshawytscha). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(2):199–210, February 1960. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f60-016.

Olsen:1960:OSS

[17] Steinar Olsen. Observations on sound scatterers in Newfoundland waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(2):211–219, February 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f60-017.

Neave:1960:SDS

[18] Ferris Neave and M. G. Hanavan. Seasonal distribution of some epipelagic fishes in the Gulf of Alaska region. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(2):221–233, February 1960. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f60-018.

Lambert:1960:FRS

[19] D. G. Lambert. The food of the redfish Sebastes marinus (L.) in the Newfoundland area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(2): 235–243, February 1960. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f60-019.

Battle:1960:DSB

[20] Helen I. Battle and William M. Sprules. A description of the semibuoyant eggs and early developmental stages of the goldeye, Hiodon alosoides (Rafinesque). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(2):245–266, February 1960. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f60-020.

Vladykov:1960:DYA

[21] Vadim D. Vladykov. Description of young ammocoetes belonging to two species of lampreys: Petromyzon marinus and Entosphenus lamottenii. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(2):267–290, February 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f60-021.

Anderson:1960:NPV

[22] Roy C. Anderson and Neil Bourne. Note on Pontonema vacillatum Leidy 1855 (Nematoda: Oncholaimidae) from soft-shell clams (Mya arenaria). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(2):291–293, February 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f60-022.

Baggerman:1960:SPT

[23] Bertha Baggerman. Salinity preference, thyroid activity and the seaward migration of four species of Pacific salmon (Oncorhynchus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(3):295–322, March 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f60-023.

MacLeod:1960:BSC

[24] Robert A. MacLeod, R. E. E. Jonas, and E. Onofrey. A biochemical study of Coho salmon (Oncorhynchus kisutch) maturing sexually in an aquarium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(3):323– 335, March 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f60-024.

Strickland:1960:FBC

[25] J. D. H. Strickland and K. H. Austin. On the forms, balance and cycle of phosphorus observed in the coastal and oceanic waters of the Northeastern Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(3): 337–345, March 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-025.

Wood:1960:BSS

[26] J. D. Wood, D. W. Duncan, and M. Jackson. Biochemical studies on sockeye salmon during spawning migration: XI. The free histidine content of the tissues. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17 (3):347–351, March 1960. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f60-026.

Tabata:1960:CWV

[27] Susumu Tabata. Characteristics of water and variations of salinity, temperature, and dissolved oxygen content of the water at Ocean Weather Station “P” in the Northeast Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(3):353–370, March 1960. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f60-027.

Tomlinson:1960:PCF

[28] Neil Tomlinson, Vera M. Creelman, and K. G. Reid. The phosphoruscontaining fractions of sterile lingcod muscle during storage at 0 ^{o}C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(3):371–376, March 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f60-028.

Trites:1960:OBR

[29] R. W. Trites. An oceanographical and biological reconnaissance of Kennebecasis Bay and the Saint John River Estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(3):377–408, March 1960. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f60-029.

Legare:1960:QQS

[30] J. E. Henri Legaré and Delphine C. Maclellan. A qualitative and quantitative study of the plankton of the quoddy region in 1957 and 1958 with special reference to the food of the herring. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(3):409–448, March 1960. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f60-030.

Creelman:1960:IMP

[31] Vera M. Creelman and Neil Tomlinson. Inosine in the muscle of Pacific salmon stored in ice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17 (3):449–451, March 1960. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f60-031.

Saunders:1960:EIP

[32] J. W. Saunders. The effect of impoundment on the population and movement of Atlantic salmon in Ellerslie Brook, Prince Edward Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4):453–473, April 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f60-032.

Iyengar:1960:APS

[33] J. R. Iyengar, K. Visweswariah, M. N. Moorjani, and D. S. Bhatia. Assessment of the progressive spoilage of ice-stored shrimp. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4):475–485, April 1960. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f60-033.

Black:1960:AGG

[34] Edgar C. Black, Anne C. Robertson, Arthur R. Hanslip, and Wing-Gay Chiu. Alterations in glycogen, glucose and lactate in rainbow and Kamloops trout, Salmo gairdneri, following muscular activity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4):487–500, April 1960. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f60-034.

Crampton:1960:EIP

[35] E. W. Crampton, E. Donefer, and Donna J. Schad. Effect of the ingestion of Porrocaecum (codworm) on growth, voluntary intake and feed efficiency of beagle pups. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4):501–505, April 1960. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f60-035.

Bell:1960:FAF

[36] M. C. Bell and C. H. Clay. Facilities for anadromous fish passage, Passamaquoddy Project. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4): 507–516, April 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-036.

Beeton:1960:VMM

[37] Alfred M. Beeton. The vertical migration of Mysis relicta in Lakes Huron and Michigan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4): 517–539, April 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-037.

Martin:1960:OLH

[38] N. V. Martin and N. S. Baldwin. Observations on the life history of the hybrid between Eastern brook trout and lake trout in Algonquin Park, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4): 541–551, April 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-038.

Wilder:1960:PEP

[39] D. G. Wilder. Possible effects of Passamaquoddy tidal power structures on the Canadian lobster industry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4):553–563, April 1960. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f60-039.

Chang:1960:BSS

[40] V. M. Chang, H. Tsuyuki, and D. R. Idler. Biochemical studies on sockeye salmon during spawning migration: XIII. The distribution of phosphorus compounds, creatine and inositol in the major tissues. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4):565–582, April 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f60-040.

McCauley:1960:REC

[41] R. W. McCauley. The role of electrical conductivity of water in shocking lampreys (Petromyzon marinus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4):583–589, April 1960. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f60-041.

Thomas:1960:MAD

[42] M. L. H. Thomas. A modified anchor dredge for collecting burrowing animals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4):591– 594, April 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f60-042.

Stewart:1960:IBP

[43] James E. Stewart and C. H. Castell. Inhibition of bacterially produced trimethylamine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4): 595–596, April 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-043.

Fagerlund:1960:MSV

[44] U. H. M. Fagerlund and D. R. Idler. Marine sterols. VII. The synthesis of 29-isofucosterol and the attempted synthesis of 17-dehydrocholesterol. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4):597–599, April 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f60-044.

McCauley:1960:RFC

[45] R. W. McCauley and D. P. Scott. Removal of free chlorine from running water by sodium thiosulphate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(4):601, April 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-045.

Tomlinson:1960:SFR

[46] Neil Tomlinson and Vera M. Creelman. On the source of free ribose formed post mortem in the muscle of lingcod (Ophiodon elongatus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5):603–606, May 1960. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f60-046.

Oliver:1960:MBF

[47] D. R. Oliver. The macroscopic bottom fauna of Lac la Ronge, Saskatchewan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5): 607–624, May 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-047.

Hewson:1960:HLW

[48] L. C. Hewson. A history of the Lake Winnipeg fishery for whitefish, Coregonus clupeaformis, with some reference to its economics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5):625–639, May 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f60-048.

Butler:1960:MBP

[49] T. H. Butler. Maturity and breeding of the Pacific edible crab, Cancer magister Dana. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5): 641–646, May 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-049.

Lawler:1960:MPE

[50] G. H. Lawler. A mutant pike, Esox lucius. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5):647–654, May 1960. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f60-050.

McDonald:1960:BPS

[51] J. McDonald. The behaviour of Pacific salmon fry during their downstream migration to freshwater and saltwater nursery areas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5):655–676, May 1960. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f60-051.

Smith:1960:AIT

[52] M. W. Smith. An attempt to improve trout fishing in a New Brunswick lake by deferment of opening of angling season. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5):677–685, May 1960. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f60-052.

Nicholls:1960:AMC

[53] C. F. Nicholls and E. A. Doxtator. An aquarium for maintaining constant prey population densities in studies of predator–prey interactions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5):687–688, May 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f60-053.

Brawn:1960:UTO

[54] Vivien M. Brawn. Underwater television observations of the swimming speed and behaviour of captive herring. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5):689–698, May 1960. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f60-054.

Brawn:1960:SDV

[55] Vivien M. Brawn. Seasonal and diurnal vertical distribution of herring (Clupea harengus L.) in Passamaquoddy, Bay, N.B. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5):699–711, May 1960. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f60-055.

Kerswill:1960:EPP

[56] C. J. Kerswill. Effects of proposed Passamaquoddy Power Project on anadromous fishes in Canadian Waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5):713–720, May 1960. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f60-056.

Brawn:1960:TTU

[57] Vivien M. Brawn. Temperature tolerance of unacclimated herring (Clupea harengus L.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5):

721–723, May 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-057.

Brawn:1960:SHC

[58] Vivien M. Brawn. Survival of herring (Clupea harengus L.) in water of low salinity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5): 725–726, May 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-058.

Forrester:1960:CMP

[59] W. D. Forrester. Current measurements in Passamaquoddy Bay and the Bay of Fundy 1957 and 1958. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5):727–729, May 1960. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f60-059.

Leim:1960:RUF

[60] A. H. Leim. Records of uncommon fishes from waters off the Maritime Provinces of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5): 731–733, May 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-060.

Tibbo:1960:EHB

[61] S. N. Tibbo and V. M. Brawn. Explorations for herring in the Bay of Fundy and Gulf of Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5):735–737, May 1960. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f60-061.

Smith:1960:NTS

[62] S. B. Smith. A note on two stocks of steelhead trout (Salmo gairdneri) in Capilano River, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(5):739–742, May 1960. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f60-062.

Chevrier:1960:DBE

[63] J. R. Chevrier and R. W. Trites. Drift-bottle experiments in the quoddy region, Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6):

743–762, June 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-063.

Scott:1960:SLH

[64] D. M. Scott and W. F. Black. Studies on the life-history of the ascarid Porrocaecum decipiens in the Bras d’Or Lakes, Nova Scotia, Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6):763–774, June 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f60-064.

Fleming:1960:AGS

[65] A. M. Fleming. Age, growth and sexual maturity of cod (Gadus morhua L.) in the Newfoundland area, 1947–1950. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6):775–809, June 1960. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f60-065.

Southcott:1960:KQP

[66] B. A. Southcott, R. H. Moyer, E. G. Baker, and H. L. A. Tarr. Keeping quality of Pacific Coast dogfish. II. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6):811–814, June 1960. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f60-066.

Doucet:1960:ESH

[67] W. F. Doucet. Economic study of the herring fishery of Charlotte County, New Brunswick, 1956–1957. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6):815–870, June 1960. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f60-067.

McCombie:1960:AIT

[68] Alen Milne McCombie. Actions and interactions of temperature, light intensity and nutrient concentration on the growth of the green alga, Chlamydomonas reinhardi Dangeard. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6):871–894, June 1960. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f60-068.

Bilinski:1960:BTC

[69] E. Bilinski. Biosynthesis of trimethylammonium compounds in aquatic animals: I. Formation of trimethylamine oxide and betaine from C^{14}- labelled compounds by lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6):895–902, June 1960. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f60-069.

Wood:1960:EMP

[70] J. D. Wood. The effect of marine products on the blood cholesterol levels in man and in animals. A review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6):903–912, June 1960. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f60-070.

McBride:1960:SVC

[71] J. R. McBride, R. A. MacLeod, and D. R. Idler. Seasonal variation in the collagen content of Pacific herring tissues. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6):913–918, June 1960. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f60-071.

Muir:1960:CGR

[72] Barry S. Muir. Comparison of growth rates for native and hatcherystocked populations of Esox masquinongy in Nogies Creek, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6):919–927, June 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f60-072.

Ricker:1960:PDC

[73] W. E. Ricker. A population of dwarf coastrange sculpins (Cottus aleuticus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6):929– 932, June 1960. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f60-073.

Tibbo:1960:FSL

[74] S. N. Tibbo and J. E. Henri Legaré. Further study of larval herring (Clupea harengus L.) in the Bay of Fundy and Gulf of Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6):933–942, June 1960. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f60-074.

Anonymous:1960:CRP

[75] Anonymous. Corrections for recent publications of The Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6):943–944, June 1960. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f60-075.

Anonymous:1960:IVJ

[76] Anonymous. Index to volume 17 of the Journal of the Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 17(6): 945–953, June 1960. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f60-076.

McAllister:1961:ZSO

[77] C. D. McAllister. Zooplankton studies at Ocean Weather Station “P” in the Northeast Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(1):1–29, January 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-001.

Olsen:1961:CBH

[78] Steinar Olsen. Contribution to the biology of herring (Clupea harengus L.) in Newfoundland waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(1):31–46, January 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-002.

Lawler:1961:MPG

[79] G. H. Lawler and G. P. McRae. A method for preparing glycerin-stored otoliths for age determination. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(1):47–50, January 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-003.

Bursa:1961:PCE

[80] Adam Bursa. Phytoplankton of the Calanus Expeditions in Hudson Bay, 1953 and 1954. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(1): 51–83, January 1961. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f61-004.

Larsen:1961:QFF

[81] B. A. Larsen and W. W. Hawkins. The quality of fish flour, liver meal, and visceral meal as sources of dietary protein. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(1):85–91, January 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-005.

McBride:1961:LBC

[82] J. R. McBride, D. R. Idler, and R. A. MacLeod. The liquefaction of British Columbia herring by ensilage, proteolytic enzymes and acid hydrolysis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(1):93–112, January 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f61-006.

March:1961:PNV

[83] B. E. March, J. Biely, J. R. McBride, D. R. Idler, and R. A. MacLeod. The protein nutritive value of “liquid herring” preparations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(1):113–116, January 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f61-007.

Kerr:1961:SLR

[84] Robert B. Kerr. Scale to length ratio, age and growth of Atlantic salmon in Miramichi fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(1): 117–124, January 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-008.

Hochachka:1961:LGR

[85] P. W. Hochachka. Liver glycogen reserves of interacting resident and introduced trout populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(1):125–135, January 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-009.

Power:1961:SFA

[86] H. E. Power. Slicing of fillets as an aid in detection and removal of codworms from Atlantic cod fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(1):137–140, January 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-010.

McAllister:1961:NRE

[87] D. E. McAllister. Northward range extension of the flathead chub and trout-perch to Aklavik, N.W.T. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(1):141, January 1961. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f61-011.

Bligh:1961:LHF

[88] E. G. Bligh. Lipid hydrolysis in frozen cod muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(1):143–145, January 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-012.

Legendre:1961:ADC

[89] R. Legendre. Artificial drying of cambodian fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(2):147–162, February 1961. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f61-013.

Trites:1961:PEP

[90] R. W. Trites. Probable effects of proposed Passamaquoddy Power Project on oceanographic conditions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(2):163–201, February 1961. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f61-014.

Johannes:1961:CFB

[91] R. E. Johannes and P. A. Larkin. Competition for food between redside shiners (Richardsonius balteatus) and rainbow trout (Salmo gairdneri) in two British Columbia lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(2):203–220, February 1961. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f61-015.

McKenzie:1961:HMB

[92] R. A. McKenzie and S. N. Tibbo. Herring movements in the Bay of Fundy and Gulf of Maine, 1957 and 1958. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(2):221–252, February 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-016.

Collin:1961:OAP

[93] A. E. Collin. Oceanographic activities of the Polar Continental Shelf Project. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(2):253–258, February 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f61-017.

Ali:1961:CMI

[94] M. A. Ali, P. Copes, and W. R. Stevenson. Correlation of morphological and intra-ocular measurements in the Atlantic salmon (Salmo salar) yearling. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(2):259–272, February 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f61-018.

Anderson:1961:NMB

[95] D. V. Anderson. A note on the morphology of the basins of the Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(2):273–277, February 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f61-019.

Hitz:1961:OTS

[96] C. R. Hitz. Occurrence of two species of juvenile rockfish in Queen Charlotte Sound. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(2): 279–281, February 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-020.

Lawler:1961:ALE

[97] G. H. Lawler. Abnormalities in Lake Erie whitefish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(2):283–284, February 1961. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f61-021.

Bilinski:1961:BTC

[98] E. Bilinski. Biosynthesis of trimethylammonium compounds in aquatic animals. II. Role of betaine in the formation of trimethylamine oxide by lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(2):285–286, February 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-022.

Werner:1961:DHS

[99] A. E. Werner and M. Waldichuk. Decay of hexachlorocyclohexane in sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(2):287–289, February 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f61-023.

Ketchen:1961:NRL

[100] K. S. Ketchen and R. M. Wilson. A new record of the long-finned cod, Antimora rostrata Günther, from British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(2):291, February 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-024.

Lawler:1961:ECL

[101] G. H. Lawler. Egg counts of Lake Erie whitefish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(2):293–294, February 1961. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f61-025.

Castell:1961:SFV

[102] C. H. Castell and Jacqueline G. Giles. Spoilage of fish in the vessels at sea 7. Further studies on seasonal variations in the landed quality of gutted, trawler-caught Atlantic cod and haddock. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(3):295–302, March 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-026.

Castell:1961:GFQ

[103] C. H. Castell, Maxine F. Elson, and Jacqueline G. Giles. Grading fish for quality 4. Variations in the relation between trimethylamine values and grades for gutted, trawler-caught Atlantic cod and haddock. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(3):303–310, March 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f61-027.

Favorite:1961:STS

[104] Felix Favorite. Surface temperature and salinity off the Washington and British Columbia coasts, August, 1958 and 1959. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(3):311–319, March 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-028.

Tomlinson:1961:OPM

[105] N. Tomlinson, E. S. Arnold, Eve Roberts, and S. E. Geiger. Observations on post mortem biochemical changes in fish muscle in relation to Rigor Mortis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18 (3):321–336, March 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-029.

Bevan:1961:VAC

[106] Donald E. Bevan. Variability in aerial counts of spawning salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(3):337–348, March 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f61-030.

Mannan:1961:PDP

[107] Abdul Mannan and H. L. A. Tarr. Preparation of deoxynucleosides, purine and pyrimidine bases and deoxyribose 1-phosphate from deoxyribonucleic acid employing salmon enzyme systems. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(3):349–366, March 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-031.

Wilder:1961:CTM

[108] D. G. Wilder and D. W. McLeese. A comparison of three methods of inactivating lobster claws. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(3):367–375, March 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-032.

Wood:1961:DMF

[109] J. D. Wood and J. Topliff. Dietary marine fish oils and cholesterol metabolism 3. The comparative hypocholesterolemic activities of fish oil and vitamin A. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(3): 377–382, March 1961. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f61-033.

Scott:1961:RIF

[110] D. P. Scott. Radioactive iron as a fish mark. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(3):383–391, March 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-034.

Vanstone:1961:PPC

[111] W. E. Vanstone and F. Chung-Wai Ho. Plasma proteins of Coho salmon, Oncorhynchus kisutch, as separated by zone electrophoresis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(3):393–399, March 1961. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f61-035.

Houston:1961:ISU

[112] Arthur H. Houston. Influence of size upon the adaptation of steelhead trout (Salmo gairdneri) and chum salmon (Oncorhynchus keta) to sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(3):401–415, March 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f61-036.

Keleher:1961:CLG

[113] J. J. Keleher. Comparison of largest Great Slave Lake fish with North American records. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18 (3):417–421, March 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-037.

Rawson:1961:LTL

[114] D. S. Rawson. The lake trout of Lac la Ronge, Saskatchewan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(3):423–462, March 1961.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f61-038.

Smith:1961:LRN

[115] M. W. Smith. A limnological reconnaissance of a Nova Scotian brownwater lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(3):463– 478, March 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f61-039.

Anonymous:1961:O

[116] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(4):479–482, April 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-040.

Mannan:1961:PCCa

[117] A. Mannan, D. I. Fraser, and W. J. Dyer. Proximate composition of Canadian Atlantic fish: I. Variation in composition of different sections of the flesh of Atlantic halibut (Hippoglossus hippoglossus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(4):483–493, April 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-041.

Mannan:1961:PCCb

[118] A. Mannan, D. I. Fraser, and W. J. Dyer. Proximate composition of Canadian Atlantic fish: II. Mackerel, tuna and swordfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(4):495–499, April 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-042.

Hoogland:1961:AAC

[119] P. L. Hoogland, H. C. Freeman, Beryl Truscott, and A. E. Waddell. The amino acid composition of cod tropomyosin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(4):501–512, April 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-043.

Ketchen:1961:OEP

[120] K. S. Ketchen. Observations on the ecology of the Pacific cod (Gadus macrocephalus) in Canadian Waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(4):513–558, April 1961. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f61-044.

Bell:1961:TEA

[121] Gordon R. Bell. Two epidemics of apparent kidney disease in cultured pink salmon (Oncorhynchus gorbuscha). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(4):559–562, April 1961. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f61-045.

Bursa:1961:AOC

[122] Adam S. Bursa. The annual oceanographic cycle at Igloolik in the Canadian Arctic: II. The phytoplankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(4):563–615, April 1961. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f61-046.

Rodgers:1961:PSE

[123] G. K. Rodgers and D. V. Anderson. A preliminary study of the energy budget of Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(4):617–636, April 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-047.

Tsuyuki:1961:MPP

[124] H. Tsuyuki and Eve Roberts. Muscle proteins of Pacific salmon (Oncorhynchus): I. A note on the separation of muscle proteins soluble in low ionic strength salt solutions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(4):637–640, April 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-048.

Fraser:1961:TBP

[125] Doris I. Fraser, Somsee Punjamapirom, and W. J. Dyer. Temperature and the biochemical processes occurring during rigor mortis in cod muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(4):641– 644, April 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f61-049.

Paloheimo:1961:SEM

[126] J. E. Paloheimo. Studies on estimation of mortalities: I. Comparison of a method described by Beverton and Holt and a new linear formula. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(5):645–662, May 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f61-050.

Grainger:1961:CCG

[127] E. H. Grainger. The copepods Calanus glacial is Jaschnov and Calanus finmarchicus (Gunnerus) in Canadian Arctic-subarctic waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(5):663–678, May 1961. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f61-051.

Norden:1961:COR

[128] Carroll R. Norden. Comparative osteology of representative salmonid fishes, with particular reference to the grayling (Thymallus arcticus) and its phylogeny. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(5): 679–791, May 1961. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f61-052.

McPhail:1961:SSS

[129] J. D. McPhail. A systematic study of the Salvelinus alpinus complex in North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(5): 793–816, May 1961. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f61-053.

Paulik:1961:DIR

[130] Gerald J. Paulik. Detection of incomplete reporting of tags. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(5):817–832, May 1961. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f61-054.

Crean:1961:LPS

[131] P. B. Crean. The light pickle salting of cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(5):833–844, May 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-055.

Crean:1961:PDS

[132] P. B. Crean. The partial desalting of salted cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(5):845–850, May 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-056.

Cardin:1961:CCS

[133] A. Cardin, E. Bilinski, F. Maltais, M. A. Bordeleau, and A. Laframboise. Chemical characteristics of salted cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(5):851–858, May 1961. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f61-057.

Ho:1961:EEM

[134] F. Chung-Wai Ho and W. E. Vanstone. Effect of estradiol monobenzoate on some serum constituents of maturing sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(5):859– 864, May 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f61-058.

Jangaard:1961:FAM

[135] P. M. Jangaard and R. G. Ackman. Fatty alcohols from marine oils and segregated esters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(5): 865–872, May 1961. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f61-059.

Butler:1961:GAD

[136] T. H. Butler. Growth and age determination of the Pacific edible crab Cancer magister Dana. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(5):873–891, May 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-060.

Fraser:1961:PCC

[137] D. I. Fraser, A. Mannan, and W. J. Dyer. Proximate composition of Canadian Atlantic fish. III. Sectional differences in the flesh of a species of Chondrostei, one of Chimaerae, and of some miscellaneous teleosts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(6):893–905, June 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f61-061.

Pickard:1961:OFI

[138] G. L. Pickard. Oceanographic features of inlets in the British Columbia Mainland Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(6): 907–999, June 1961. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f61-062.

Parsons:1961:CCE

[139] T. R. Parsons, K. Stephens, and J. D. H. Strickland. On the chemical composition of eleven species of marine phytoplankters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(6):1001–1016, June 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-063.

Parsons:1961:PCE

[140] T. R. Parsons. On the pigment composition of eleven species of marine phytoplankters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18 (6):1017–1025, June 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-064.

Herlinveaux:1961:SOF

[141] R. H. Herlinveaux and J. P. Tully. Some oceanographic features of Juan de Fuca Strait. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(6): 1027–1071, June 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-065.

Tabata:1961:TCS

[142] Susumu Tabata. Temporal changes of salinity, temperature, and dissolved oxygen content of the water at Station “P” in the Northeast Pacific Ocean, and some of their determining factors. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(6):1073–1124, June 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-066.

Vladykov:1961:OSD

[143] V. D. Vladykov and G. N. Mukerji. Order of succession of different types of infraoral lamina in landlocked sea lamprey (Petromyzon marinus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(6):1125–1143, June 1961. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f61-067.

Alverson:1961:OTT

[144] Dayton L. Alverson. Ocean temperatures and their relation to albacore tuna (Thunnus germo) distribution in waters off the coast of Oregon, Washington and British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(6):1145–1152, June 1961. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f61-068.

McCrimmon:1961:SFQ

[145] Hugh R. McCrimmon. Seasonal fishing quality provided by the natural reproduction of speckled trout in three Ontario ponds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18(6):1153–1158, June 1961. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f61-069.

Ellis:1961:DPT

[146] D. V. Ellis. Diving and photographic techniques for observing and recording salmon activities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18 (6):1159–1166, June 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-070.

Barraclough:1961:ARA

[147] W. E. Barraclough and T. H. Butler. Additional records of the argentinid fish, Leuroglossus stilbius Gilbert, from British Columbia, with remarks on its taxonomy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18 (6):1167–1169, June 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-071.

Anonymous:1961:IVJ

[148] Anonymous. Index to volume 18 of the Journal of the Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 18 (6):1171–1178, June 1961. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f61-072.

Herlinveaux:1962:OSI

[149] R. H. Herlinveaux. Oceanography of Saanich Inlet in Vancouver Island, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(1): 1–37, January 1962. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f62-001.

Thurston:1962:PCC

[150] Claude E. Thurston. Physical characteristics and chemical composition of two subspecies of lake trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(1):39–44, January 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-002.

Lane:1962:RTS

[151] R. K. Lane. A review of the temperature and salinity structures in the approaches to Vancouver Island, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(1):45–91, January 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-003.

Elson:1962:PAS

[152] P. F. Elson. The Pollett apron seine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(1):93–100, January 1962. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f62-004.

Prakash:1962:HMF

[153] A. Prakash and J. C. Medcof. Hydrographic and meteorological factors affecting shellfish toxicity at Head Harbour, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(1):101–112, January 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f62-005.

Harris:1962:GRL

[154] Roy H. D. Harris. Growth and reproduction of the longnose sucker, Catostomus catostomus (Forster), in Great Slave Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(1):113–126, January 1962. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f62-006.

Hochachka:1962:GST

[155] P. W. Hochachka and A. C. Sinclair. Glycogen stores in trout tissues before and after stream planting. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(1):127–136, January 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-007.

Ellis:1962:PSV

[156] D. V. Ellis. Preliminary studies on the visible migrations of adult salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(1):137–148, January 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f62-008.

Scott:1962:RCF

[157] D. P. Scott. Radioactive caesium as a fish and lamprey mark. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(1):149–157, January 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f62-009.

Lander:1962:MEM

[158] Robert H. Lander. A method of estimating mortality rates from change in composition. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(1): 159–168, January 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-010.

Wood:1962:FHF

[159] J. D. Wood and S. A. Haqq. Fat hydrolysis in frozen fillets of lingcod and Pacific Gray cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(1): 169–171, January 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-011.

Hartman:1962:CIO

[160] G. F. Hartman, T. G. Northcote, and C. C. Lindsey. Comparison of inlet and outlet spawning runs of rainbow trout in Loon Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(2):173–200, February 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f62-012.

Northcote:1962:MBJ

[161] T. G. Northcote. Migratory behaviour of juvenile rainbow trout, Salmo gairdneri, in outlet and inlet streams of Loon Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(2):201–270, February 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f62-013.

Aron:1962:DAE

[162] William Aron. The distribution of animals in the Eastern North Pacific and its relationship to physical and chemical conditions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(2):271–314, February 1962. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f62-014.

Parker:1962:TFA

[163] Robert R. Parker. Tag failure associated with a net fishery as a source of experimental error. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(2): 315–320, February 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-015.

Geiger:1962:DG

[164] Shirley E. Geiger, Eve Roberts, and N. Tomlinson. Dogfish gelatin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(2):321–326, February 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f62-016.

McBride:1962:OPJ

[165] J. R. McBride, D. R. Idler, R. E. E. Jonas, and N. Tomlinson. Olfactory perception in juvenile salmon: I. Observations on response of juvenile sockeye to extracts of foods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(2):327–334, February 1962. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f62-017.

Southward:1962:PHO

[166] G. Morris Southward. Photographing halibut otoliths for measuring growth zones. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(2): 335–338, February 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-018.

Southward:1962:MCB

[167] G. Morris Southward. A method of calculating body lengths from otolith measurements for Pacific halibut and its application to Portlock– Albatross grounds data between 1935 and 1957. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(2):339–362, February 1962. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f62-019.

Pike:1962:FRG

[168] Gordon C. Pike. First record of the great white shark (Carcharodon carcharias) from British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(2):363, February 1962. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f62-020.

Miller:1962:DGR

[169] Richard B. Miller and Frances Miller. Diet, glycogen reserves and resistance to fatigue in hatchery rainbow trout. Part II. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(3):365–375, March 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-021.

Grainger:1962:ZFB

[170] E. H. Grainger. Zooplankton of Foxe Basin in the Canadian Arctic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(3):377–400, March 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f62-022.

Butler:1962:OKC

[171] T. H. Butler and Josephine F. L. Hart. The occurrence of the king crab, Paralithodes camtschatica (Tilesius), and of Lithodes aequispina Benedict in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19 (3):401–408, March 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-023.

Black:1962:CGP

[172] Edgar C. Black, Anne Robertson Connor, Kwok-Cheung Lam, and Wing-Gay Chiu. Changes in glycogen, pyruvate and lactate in rainbow trout (Salmo gairdneri) during and following muscular activity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(3):409–436, March 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-024.

Dussault:1962:ECB

[173] H. P. Dussault. Enumeration of coliform bacteria in light salted fish brines. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(3):437–444, March 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f62-025.

Templeman:1962:CTN

[174] Wilfred Templeman and A. M. Fleming. Cod tagging in the Newfoundland area during 1947 and 1948. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(3):445–487, March 1962. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f62-026.

McCrimmon:1962:OIM

[175] Hugh McCrimmon and Joseph Bray. Observations on the isopod Mesidotea entomon in the Western Canadian Arctic Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(3):489–496, March 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-027.

Thomson:1962:FPC

[176] J. Arthur Thomson. On the fecundity of Pacific cod (Gadus macrocephalus Tilesius) from Hecate Strait, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(3):497–500, March 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-028.

Lund:1962:PMI

[177] J. W. G. Lund. The periodicity of Melosira islandica o. Müll. in Great Slave Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(3):501– 504, March 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f62-029.

Bilinski:1962:BTC

[178] E. Bilinski. Biosynthesis of trimethylammonium compounds in aquatic animals: III. Choline metabolism in marine crustacea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(3):505–510, March 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-030.

Anonymous:1962:O

[179] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(3):511–514, March 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-031.

Bell:1962:TGN

[180] Gordon Bell. Trichodinids from the gills of a new host, the lingcod Ophiodon elongatus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19 (3):515–516, March 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-032.

Tibbo:1962:NRO

[181] S. N. Tibbo. New records for occurrence of the white-tip shark, Pterolamiops longimanus (Poey), and the dolphin, Coryphaena hippurus L., in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(3):517–518, March 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-033.

Tarr:1962:NCO

[182] H. L. A. Tarr and M. Leroux. A note concerning the origin and quantitative distribution of acid-soluble phosphorus compounds and free sugars in fish muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19 (3):519–520, March 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-034.

Tanaka:1962:ERF

[183] Syoiti Tanaka. The effect of reduction of fishing effort on yield. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4):521–529, April 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f62-035.

Ricker:1962:COG

[184] W. E. Ricker. Comparison of ocean growth and mortality of sockeye salmon during their last two years. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4):531–560, April 1962. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f62-036.

Parker:1962:EOM

[185] Robert R. Parker. Estimations of ocean mortality rates for Pacific salmon (Oncorhynchus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4): 561–589, April 1962. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f62-037.

Dingle:1962:PFM

[186] J. R. Dingle, J. A. Hines, and J. M. Neelin. Proteins in fish muscle: 17. Fractionation of aqueous extracts with zinc acetate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4):591–604, April 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-038.

Ackman:1962:CFF

[187] R. G. Ackman, L. R. Gallay, P. M. Jangaard, and M. L. Hughes. Conversion of the free fatty acids of cod oil to methyl esters in situ. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4):605–614, April 1962. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f62-039.

Boyd:1962:ETA

[188] J. W. Boyd and B. A. Southcott. Effect of tetracycline antibiotics on objective and subjective fish quality tests. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4):615–618, April 1962. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f62-040.

Jonas:1962:BPM

[189] R. E. E. Jonas, Harcharan S. Sehdev, and N. Tomlinson. Blood pH and mortality in rainbow trout (Salmo gairdnerii) and sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4): 619–624, April 1962. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f62-041.

Keenleyside:1962:SDO

[190] Miles H. A. Keenleyside. Skin-diving observations of Atlantic salmon and brook trout in the Miramichi River, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4):625–634, April 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-042.

Brawn:1962:PPH

[191] Vivien M. Brawn. Physical properties and hydrostatic function of the swimbladder of herring (Clupea harengus L.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4):635–656, April 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-043.

Frolander:1962:QET

[192] Herbert F. Frolander. Quantitative estimations of temporal variations of zooplankton off the coast of Washington and British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4):657–675, April 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f62-044.

Squires:1962:DCC

[193] H. J. Squires. Decapod crustacea of the Calanus Expeditions in Frobisher Bay, Baffin Island, 1951. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4):677–686, April 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-045.

Lindsey:1962:DBB

[194] C. C. Lindsey. Distinctions between the broad whitefish, Coregonus nasus, and other North American whitefishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4):687–714, April 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-046.

Scott:1962:EFQ

[195] D. P. Scott. Effect of food quantity on fecundity of rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4):715– 731, April 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f62-047.

Jonas:1962:PCL

[196] R. E. E. Jonas and N. Tomlinson. The phospholipid content of lingcod muscle during frozen storage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(4):733–734, April 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-048.

Stevenson:1962:DSH

[197] J. C. Stevenson. Distribution and survival of herring larvae (Clupea pallasi Valenciennes) in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5):735–810, May 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-049.

Templeman:1962:ROP

[198] Wilfred Templeman. Record of the oceanic puffer, Lagocephalus lagocephalus (L.), family Tetraodontidae, from Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5):811–814, May 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-050.

Pike:1962:MFG

[199] Gordon C. Pike. Migration and feeding of the Gray whale (Eschrichtius gibbosus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5):815– 838, May 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f62-051.

Jangaard:1962:ABC

[200] P. M. Jangaard and R. G. Ackman. Air blowing of cod liver oil and other marine oils. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5): 839–849, May 1962. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f62-052.

Prakash:1962:SCF

[201] A. Prakash. Seasonal changes in feeding of Coho and Chinook (spring) salmon in Southern British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5):851–866, May 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-053.

Lasker:1962:ERY

[202] Reuben Lasker. Efficiency and rate of yolk utilization by developing embryos and larvae of the Pacific sardine Sardinops caerulea (Girard). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5):867–875, May 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f62-054.

Medcof:1962:PEP

[203] J. C. Medcof. Possible effects of Passamaquoddy Power Project on clams, scallops and shipworms in Canadian Waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5):877–889, May 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-055.

Templeman:1962:RBL

[204] Wilfred Templeman and H. J. Squires. Record of the blue ling, Molva byrkelange (Walbaum), from the Newfoundland area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5):891–894, May 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-056.

Trites:1962:FWP

[205] R. W. Trites and D. G. MacGregor. Flow of water in the passages of Passamaquoddy Bay measured by the electromagnetic method. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5):895–919, May 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f62-057.

Gilmartin:1962:ACC

[206] Malvern Gilmartin. Annual cyclic changes in the physical oceanography of a British Columbia fjord. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5):921–974, May 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-058.

Trites:1962:TSQ

[207] R. W. Trites. Temperature and salinity in the quoddy region of the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5): 975–978, May 1962. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f62-059.

Crossman:1962:PPR

[208] E. J. Crossman. Predator–prey relationships in pikes (Esocidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5):979–980, May 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f62-060.

Das:1962:FGY

[209] Naresh Das and S. N. Tibbo. On the feeding and growth of young herring (Clupea harengus L.) in captivity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5):981–983, May 1962. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f62-061.

Graham:1962:RBG

[210] T. R. Graham. A relationship between growth, hatching and spawning season in Canadian Atlantic herring (Clupea harengus L.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(5):985–987, May 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-062.

Rabb:1962:BSF

[211] L. Rabb and L. A. McDermott. Bacteriological studies of freshwater fish: II. Furunculosis in Ontario fish in natural waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6):989–995, June 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-063.

Tomlinson:1962:GCP

[212] N. Tomlinson and S. E. Geiger. Glycogen concentration and post mortem loss of adenosine triphosphate in fish and mammalian skeletal muscle. A review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6):997– 1003, June 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f62-064.

Molnar:1962:RBW

[213] Gyula Molnár and István Tölg. Relation between water temperature and gastric digestion of largemouth bass (Micropterus salmoides Lacépède). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6):1005–1012, June 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f62-065.

Scott:1962:OLC

[214] W. B. Scott and Stanford H. Smith. The occurrence of the longjaw cisco, Leucichthys alpenae, in Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6):1013–1023, June 1962. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f62-066.

Brett:1962:SCS

[215] J. R. Brett. Some considerations in the study of respiratory metabolism in fish, particularly salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6):1025–1038, June 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-067.

Power:1962:IMP

[216] H. E. Power. An improved method for the preparation of fish protein concentrate from cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19 (6):1039–1045, June 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-068.

Chapman:1962:ABJ

[217] D. W. Chapman. Aggressive behavior in juvenile Coho salmon as a cause of emigration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19 (6):1047–1080, June 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-069.

Ackman:1962:FFA

[218] R. G. Ackman, R. D. Burgher, and M. L. Hughes. The free fatty acids of cod oil: 1. Anomalous composition by fatty acid chain length. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6):1081–1084, June 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f62-070.

Pycha:1962:REN

[219] Richard L. Pycha. The relative efficiency of nylon and cotton gill nets for taking lake trout in Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6):1085–1094, June 1962. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f62-071.

Truscott:1962:EDC

[220] B. Truscott, P. L. Hoogland, P. H. Odense, and A. E. Waddell. The enzymic digestion of cod tropomyosin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6):1095–1104, June 1962. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f62-072.

Johnson:1962:RFG

[221] Lionel Johnson. The relict fauna of Greiner Lake, Victoria Island, N.W.T., Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19 (6):1105–1120, June 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-073.

Fofonoff:1962:MCM

[222] N. P. Fofonoff. Machine computations of mass transport in the North Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6): 1121–1141, June 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-074.

Tanaka:1962:MAP

[223] Syoiti Tanaka. A method of analysing a polymodal frequency distribution and its application to the length distribution of the porgy, Taius tumifrons (T. and S.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19 (6):1143–1159, June 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-075.

Stewart:1962:OMO

[224] Kenneth W. Stewart. Observations on the morphology and optical properties of the adipose eyelid of fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6):1161–1162, June 1962. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f62-076.

Waldichuk:1962:ALO

[225] M. Waldichuk and E. L. Bousfield. Amphipods in low-oxygen marine waters adjacent to a sulphite pulp mill. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6):1163–1165, June 1962. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f62-077.

Ronholt:1962:RET

[226] Lael L. Ronholt. Range extensions for two species of caridean shrimps (order Decapoda) from the Northeastern Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6):1167, June 1962. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f62-078.

Pearcy:1962:EME

[227] W. G. Pearcy. Egg masses and early developmental stages of the scorpaenid fish, Sebastolobus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6):1169–1173, June 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-079.

Heyamoto:1962:AYS

[228] H. Heyamoto. Age of young sablefish, Anoplopoma fimbria (Pallas) 1811. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19(6):1175–1177, June 1962. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f62-080.

Anonymous:1962:IVJ

[229] Anonymous. Index to volume 19 of the Journal of the Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 19 (6):1179–1185, June 1962. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f62-081.

Anonymous:1963:F

[230] Anonymous. FOREWORD. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1), January 1963. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f63-001.

Freeman:1963:NUA

[231] H. C. Freeman, P. L. Hoogland, and P. H. Odense. The nature of an unusual amino acid-pyrimidine complex from the cod parasite, Porrocaecum decipiens. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1): 1–11, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f63-002.

Paloheimo:1963:SCR

[232] J. E. Paloheimo and L. M. Dickie. Sampling the catch of a research vessel. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):13–25, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-003.

McPhail:1963:GVN

[233] J. D. McPhail. Geographic variation in North American ninespine sticklebacks, Pungitius pungitius. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):27–44, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-004.

Buttkus:1963:RWM

[234] Hans Buttkus. Red and white muscle of fish in relation to rigor mortis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):45–58, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-005.

Paloheimo:1963:ECP

[235] J. E. Paloheimo. Estimation of catchabilities and population sizes of lobsters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):59–88, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f63-006.

Jangaard:1963:ABC

[236] P. M. Jangaard, R. G. Ackman, R. D. Burgher, and M. L. Hughes. The air blowing of cod liver oil. II. Changes in fatty acid composition as determined by gas–liquid chromatography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):89–94, January 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-007.

McBride:1963:RFS

[237] J. R. McBride, U. H. M. Fagerlund, M. Smith, and N. Tomlinson. Resumption of feeding by and survival of adult sockeye salmon (Oncorhynchus nerka) following advanced gonad development. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):95–100, January 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-008.

Tsuyuki:1963:SDS

[238] H. Tsuyuki and E. Roberts. Species differences of some members of Salmonidae based on their muscle myogen patterns. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):101–104, January 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-009.

Gee:1963:CET

[239] J. H. Gee and T. G. Northcote. Comparative ecology of two sympatric species of dace (Rhinichthys) in the Fraser River system, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):105–118, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f63-010.

Uda:1963:OSP

[240] Michitaka Uda. Oceanography of the subarctic Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):119–179, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-011.

Buttkus:1963:AME

[241] Hans Buttkus. Apparatus for measuring the energy input in cutting fibers of fish muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20 (1):181–186, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-012.

Tomlinson:1963:BNF

[242] Neil Tomlinson and S. E. Geiger. The bound nucleotides of freshly frozen and severely denatured frozen lingcod muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):187–194, January 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-013.

Garrod:1963:EMR

[243] D. J. Garrod. An estimation of the mortality rates in a population of Tilapia esculenta Graham (Pisces, Cichlidae) in Lake Victoria, East Africa. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):195–227, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f63-014.

March:1963:NCE

[244] B. E. March, J. Biely, and H. L. A. Tarr. Nutrient composition and evaluation of British Columbia whole herring meal. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):229–238, January 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-015.

Anonymous:1963:Oa

[245] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):239–242, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-016.

Townsley:1963:SMA

[246] P. M. Townsley and M. A. Scott. Systaltic muscular action of the kidney tubules of flounder. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20 (1):243–244, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-017.

Jangaard:1963:PIB

[247] P. M. Jangaard, R. D. Burgher, and R. G. Ackman. A preliminary investigation of the blubber oil from the Atlantic bottlenose whale. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1):245–247, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-018.

Watson:1963:TRH

[248] N. H. F. Watson and G. H. Lawler. Temperature and rate of hatching of Triaenophorus eggs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20 (1):249–251, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-019.

Templeman:1963:ANO

[249] W. Templeman. Additional note on the oceanic puffer Lagocephalus lagocephalus (L.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(1): 253, January 1963. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f63-020.

Anonymous:1963:Ob

[250] Anonymous. Obituary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):255–256, February 1963. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f63-021.

Ricker:1963:BES

[251] W. E. Ricker. Big effects from small causes: Two examples from fish population dynamics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2): 257–264, February 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-022.

Kennedy:1963:GMW

[252] W. A. Kennedy. Growth and mortality of whitefish in three unexploited lakes in Northern Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):265–272, February 1963. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f63-023.

Templeman:1963:TRB

[253] Wilfred Templeman and H. J. Squires. Three records of the black scabbard fish, Aphanopus carbo Lowe, from the Canadian region of the Western Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):273–278, February 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f63-024.

Watson:1963:SFL

[254] N. H. F. Watson. Summer food of lake whitefish, Coregonus clupeaformis Mitchill, from Heming Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):279–286, February 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-025.

Brett:1963:SAO

[255] J. R. Brett and C. Groot. Some aspects of olfactory and visual responses in Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2): 287–303, February 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-026.

Wilder:1963:MGS

[256] D. G. Wilder. Movements, growth and survival of marked and tagged lobsters liberated in Egmont Bay, Prince Edward Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):305–318, February 1963. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f63-027.

Keleher:1963:MTG

[257] J. J. Keleher. The movement of tagged Great Slave Lake fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):319–326, February 1963.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-028.

Smith:1963:IPF

[258] M. W. Smith. Influence of pond formation on brook trout movements and angling success. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2): 327–345, February 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-029.

Day:1963:OOA

[259] L. R. Day. Organogenesis of the opercular apparatus in Salvelinus fontinalis (Mitchill). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2): 347–372, February 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-030.

Saunders:1963:RAC

[260] Richard L. Saunders. Respiration of the Atlantic cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):373–386, February 1963. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f63-031.

Sprague:1963:RFF

[261] J. B. Sprague. Resistance of four freshwater crustaceans to lethal high temperature and low oxygen. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):387–415, February 1963. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f63-032.

Lawler:1963:BTB

[262] G. H. Lawler. The biology and taxonomy of the burbot, Lota lota, in Heming Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):417–433, February 1963. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f63-033.

Tibbo:1963:BCH

[263] S. N. Tibbo and T. R. Graham. Biological changes in herring stocks following an epizootic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):

435–449, February 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-034.

Dickie:1963:CMM

[264] L. M. Dickie and J. C. Medcof. Causes of mass mortalities of scallops (Placopecten magellanicus) in the Southwestern Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):451–482, February 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f63-035.

McCauley:1963:LTD

[265] R. W. McCauley. Lethal temperatures of the developmental stages of the sea lamprey, Petromyzon marinus L. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):483–490, February 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-036.

Lawrie:1963:IGL

[266] A. H. Lawrie. Intraseasonal growth in Lake Superior lake trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):491–496, February 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-037.

Jean:1963:DFS

[267] Yves Jean. Discards of fish at sea by Northern New Brunswick draggers. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):497–524, February 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f63-038.

Alderdice:1963:SES

[268] D. F. Alderdice. Some effects of simultaneous variation in salinity, temperature and dissolved oxygen on the resistance of young Coho salmon to a toxic substance. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2): 525–550, February 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-039.

McCracken:1963:SMW

[269] F. D. McCracken. Seasonal movements of the winter flounder, Pseudopleuronectes americanus (Walbaum), on the Atlantic Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):551–586, February 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-040.

Lamsa:1963:DMB

[270] A. Lamsa. Downstream movements of brook sticklebacks, Eucalia inconstans (Kirtland), in a small Southern Ontario stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(2):587–589, February 1963. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f63-041.

Ackman:1963:OOC

[271] R. G. Ackman, P. M. Jangaard, R. D. Burgher, M. L. Hughes, and W. A. MacCallum. Observations on the oil and component fatty acids of the oil from Newfoundland capelin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):591–596, March 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-042.

Christie:1963:EAP

[272] W. J. Christie. Effects of artificial propagation and the weather on recruitment in the Lake Ontario whitefish fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):597–646, March 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-043.

Larkin:1963:ICE

[273] P. A. Larkin. Interspecific competition and exploitation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):647–678, March 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-044.

Townsley:1963:MFT

[274] P. M. Townsley, H. G. Wight, and M. A. Scott. Marine fish tissue culture. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):679–684, March 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-045.

McLaren:1963:ETG

[275] Ian A. McLaren. Effects of temperature on growth of zooplankton, and the adaptive value of vertical migration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):685–727, March 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-046.

Stewart:1963:AAC

[276] James E. Stewart, P. L. Hoogland, H. C. Freeman, and A. E. J. Waddell. Amino acid composition of representatives of eight bacterial genera with reference to aquatic productivity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):729–734, March 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-047.

Berst:1963:SDF

[277] A. H. Berst and A. M. McCombie. The spatial distribution of fish in gill nets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):735– 742, March 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f63-048.

Townsley:1963:VMP

[278] P. M. Townsley, H. G. Wight, M. A. Scott, and M. L. Hughes. The invitro maturation of the parasitic nematode, Terranova decipiens, from cod muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):743– 747, March 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f63-049.

Lindsey:1963:SOT

[279] C. C. Lindsey. Sympatric occurrence of two species of humpback whitefish in Squanga Lake, Yukon Territory. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):749–767, March 1963. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f63-050.

Hartman:1963:OBJ

[280] G. F. Hartman. Observations on behavior of juvenile brown trout in a stream aquarium during winter and spring. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):769–787, March 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-051.

Bary:1963:TSPa

[281] B. McK. Bary. Temperature, salinity and plankton in the Eastern North Atlantic and coastal waters of Britain, 1957: I. The characterisation and distribution of surface waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):789–826, March 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-052.

Premdas:1963:UDC

[282] F. H. Premdas and J. M. Anderson. The uptake and detoxification of C^{14}-labelled DDT in Atlantic salmon, Salmo salar. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):827–837, March 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-053.

Muir:1963:APL

[283] Barry S. Muir and Harold White. Application of the paloheimo linear equation for estimating mortalities to a seasonal fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):839–840, March 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-054.

Watson:1963:NUL

[284] N. H. F. Watson. A note on the upper lethal temperature of eggs of two species of Triaenophorus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):841–844, March 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-055.

Hochachka:1963:NMB

[285] Peter W. Hochachka. On the need for measuring the buffering capacity of fish blood. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3): 845–847, March 1963. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f63-056.

Scott:1963:NGM

[286] W. B. Scott. A note on Gadus (Micromesistius) poutassou (Risso) from Western Atlantic waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):849–850, March 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-057.

Barraclough:1963:FRT

[287] W. E. Barraclough and K. S. Ketchen. First record of the thornback sculpin, Paricelinus hopliticus, Eigenmann and Eigenmann, in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20 (3):851–852, March 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-058.

Rees:1963:FRT

[288] E. I. S. Rees and Malcolm R. Clarke. First records of Tetronychoteuthis dussumieri (d’Orbigny) (Cephalopoda; Onychoteuthidae) from the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):853–854, March 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-059.

McCracken:1963:MHB

[289] F. D. McCracken. Migrations of haddock between the Gulf of St. Lawrence and offshore Nova Scotian Banks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(3):855–857, March 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-060.

Henderson:1963:ILT

[290] Nancy E. Henderson. Influence of light and temperature on the reproductive cycle of the Eastern brook trout, Salvelinus fontinalis (Mitchill). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):859–897, April 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-061.

Henderson:1963:EAM

[291] Nancy E. Henderson. Extent of atresia in maturing ovaries of the Eastern brook trout, Salvelinus fontinalis (Mitchill). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):899–908, April 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-062.

Dambergs:1963:EFM

[292] N. Dambergs. Extractives of fish muscle. 3. Amounts, sectional distribution, and variations of fat, water-solubles, protein and moisture in cod (Gadus morhua L.) fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):909–918, April 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-063.

Johnson:1963:OMY

[293] W. E. Johnson and C. Groot. Observations on the migration of young sockeye salmon (Oncorhynchus nerka) through a large, complex lake system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):919– 938, April 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f63-064.

Lane:1963:MSS

[294] Robert K. Lane. A model of seawater structure near the West Coast of Vancouver Island, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):939–967, April 1963. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f63-065.

Clancy:1963:ETP

[295] Dan W. Clancy. The effect of tagging with Petersen disc tags on the swimming ability of fingerling steelhead trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):969–981, April 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-066.

Prakash:1963:SPS

[296] A. Prakash. Source of paralytic shellfish toxin in the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):983–996, April 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-067.

Wood:1963:ACW

[297] A. L. Wood. Accelerated cooling of wet, heavily salted fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):997–1000, April 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-068.

Lindsey:1963:LHR

[298] C. C. Lindsey and T. G. Northcote. Life history of redside shiners, Richardsonius balteatus, with particular reference to movements in and out of Sixteenmile Lake streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):1001–1030, April 1963. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f63-069.

Bary:1963:TSPb

[299] B. McK. Bary. Temperature, salinity and plankton in the Eastern North Atlantic and coastal waters of Britain, 1957: II. The relationships between species and water bodies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):1031–1065, April 1963. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f63-070.

Tibbo:1963:IHC

[300] S. N. Tibbo, D. J. Scarratt, and P. W. G. McMullon. An investigation of herring (Clupea harengus L.) spawning using free-diving techniques. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):1067–1079, April 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-071.

Stewart:1963:SWA

[301] James E. Stewart and H. E. Power. A sea water aquarium for marine animal experiments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20 (4):1081–1084, April 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-072.

Williamson:1963:CPS

[302] G. R. Williamson. Common porpoise from the stomach of a Greenland shark. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):1085– 1086, April 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f63-073.

Thomson:1963:DQF

[303] J. Arthur Thomson. On the demersal quality of the fertilized eggs of Pacific cod, Gadus macrocephalus Tilesius. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):1087–1088, April 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-074.

Lawler:1963:MIS

[304] G. H. Lawler and N. H. F. Watson. Measurements of immature stages of Triaenophorus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20 (4):1089–1093, April 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-075.

Forrester:1963:FRB

[305] C. R. Forrester and R. M. Wilson. A further record of the blacktail snailfish, Careproctus melanurus Gilbert, from British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):1095–1096, April 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-076.

McKenzie:1963:OOL

[306] R. A. McKenzie and S. N. Tibbo. An occurrence of opah, Lampris regius (Bonnaterre), in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):1097–1099, April 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-077.

Tibbo:1963:ODS

[307] S. N. Tibbo and R. A. McKenzie. An occurrence of dusky sharks, Carcharhinus obscurus, (Lesueur) 1818, in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):1101–1102, April 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-078.

Stewart:1963:AAR

[308] James E. Stewart. An arrangement for automatically and reproducibly controlling and varying illumination in biological experiments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(4):1103–1107, April 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-079.

Pickard:1963:OCI

[309] G. L. Pickard. Oceanographic characteristics of inlets of Vancouver Island, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(5):1109–1144, May 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-080.

Tomlinson:1963:GLM

[310] N. Tomlinson, R. E. E. Jonas, and S. E. Geiger. Glycolysis in lingcod muscle during frozen storage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(5):1145–1152, May 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-081.

Dyer:1963:CPP

[311] W. J. Dyer, D. I. Fraser, and S. N. Tibbo. Composition and palatability of porbeagle flesh. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20 (5):1153–1158, May 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-082.

Pitt:1963:VNA

[312] T. K. Pitt. Vertebral numbers of American plaice, Hippoglossoides platessoides (Fabricius), in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(5):1159–1187, May 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-083.

Scott:1963:ILG

[313] W. B. Scott and W. J. Christie. The invasion of the Lower Great Lakes by the white perch, Roccus americanus (Gmelin). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(5):1189–1195, May 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-084.

Bell:1963:TCH

[314] W. H. Bell and N. Boston. Tidal calibration of the Hecate model. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(5):1197–1212, May 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-085.

Muir:1963:VSE

[315] Barry S. Muir. Vital statistics of Esox masquinongy in Nogies Creek, Ontario: I. Tag loss, mortality due to tagging, and the estimate of exploitation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(5):1213– 1230, May 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f63-086.

MacCallum:1963:FIE

[316] W. A. MacCallum, M. W. Mullan, and Isabel N. Plaunt. Factors influencing the effectiveness of fresh fish washing operations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(5):1231–1244, May 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-087.

Pinhorn:1963:EPR

[317] A. T. Pinhorn and C. W. Andrews. Effect of photoperiods on the reactions of juvenile Atlantic salmon (Salmo salar L.) to light stimuli. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(5):1245–1266, May 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-088.

Steele:1963:PPV

[318] D. H. Steele. Pollock (Pollachius virens (L.)) in the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(5):1267–1314, May 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-089.

Hedgpeth:1963:PNA

[319] Joel W. Hedgpeth. Pycnogonida of the North American Arctic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(5):1315–1348, May 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-090.

Anonymous:1963:Oc

[320] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(5):1349–1352, May 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-091.

Tibbo:1963:OMS

[321] S. N. Tibbo, R. A. McKenzie, and W. B. Scott. An occurrence of mako sharks, Isurus oxyrinchus, Rafinesque 1810, in the Canadian Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(5):1353–1354, May 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-092.

Houston:1963:BFR

[322] Arthur H. Houston and Lawrence T. Threadgold. Body fluid regulation in smolting Atlantic salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1355–1369, June 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-093.

Parker:1963:PMY

[323] Robert R. Parker. On the problem of maximum yield from North Pacific sockeye salmon stocks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1371–1396, June 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-094.

Johnson:1963:SLH

[324] R. P. Johnson. Studies on the life history and ecology of the bigmouth buffalo, Ictiobus cyprinellus (Valenciennes). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1397–1429, June 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-095.

Lawler:1963:UCT

[325] G. H. Lawler and G. F. M. Smith. Use of coloured tags in fish population estimates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20 (6):1431–1434, June 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-096.

Sehdev:1963:PGA

[326] H. S. Sehdev, J. R. McBride, and U. H. M. Fagerlund. 2-phenoxyethanol as a general anaesthetic for sockeye salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1435–1440, June 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-097.

Parker:1963:EFL

[327] Robert R. Parker. Effects of formalin on length and weight of fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1441–1455, June 1963. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f63-098.

Fagerlund:1963:OPM

[328] U. H. M. Fagerlund, J. R. McBride, M. Smith, and N. Tomlinson. Olfactory perception in migrating salmon: III. Stimulants for adult sockeye salmon (Oncorhynchus nerka) in home stream waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1457–1463, June 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-099.

Hodder:1963:FGB

[329] V. M. Hodder. Fecundity of Grand Bank haddock. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1465–1487, June 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-100.

Sergeant:1963:MWB

[330] D. E. Sergeant. Minke whales, Balaenoptera acutorostrata Lacépède, of the Western North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1489–1504, June 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-101.

Rosenblatt:1963:RRC

[331] Richard H. Rosenblatt and Donald Wilkie. A redescription of the rare cottid fish, Artedius meanyi, new to the fauna of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1505–1511, June 1963. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f63-102.

Rees:1963:BDA

[332] E. I. S. Rees. The batfish, Dibranchus atlanticus Peters, on the Canadian Atlantic slopes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20 (6):1513–1517, June 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-103.

Bary:1963:TSPc

[333] B. McK. Bary. Temperature, salinity and plankton in the Eastern North Atlantic and coastal waters of Britain, 1957: III. The distribution of zooplankton in relation to water bodies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1519–1548, June 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-104.

Flury:1963:MPG

[334] J. Flury. A modified Petersen grab. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1549–1550, June 1963. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f63-105.

Ackman:1963:ODA

[335] R. G. Ackman and M. Jangaard. On the occurrence of 4,7,10,13,16docosapentaenoic acid in saury (Cololabis saira) oil. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1551–1552, June 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-106.

Lawler:1963:SSS

[336] G. H. Lawler. Spring stainless steel anchor tag. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1553, June 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-107.

Anonymous:1963:RAI

[337] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20(6):1555, June 1963. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f63-108.

Anonymous:1963:IVJ

[338] Anonymous. Index to volume 20 of the Journal of the Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 20 (6):1557–1566, June 1963. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f63-109.

Larkin:1964:FIS

[339] P. A. Larkin and W. E. Ricker. Further information on sustained yields from fluctuating environments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1):1–7, January 1964. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f64-001.

Forrester:1964:LOE

[340] C. R. Forrester. Laboratory observations on embryonic development and larvae of the Pacific cod (Gadus macrocephalus Tilesius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1):9–16, January 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-002.

Sprague:1964:LCC

[341] J. B. Sprague. Lethal concentrations of copper and zinc for young Atlantic salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1): 17–26, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f64-003.

Saunders:1964:MYA

[342] Richard L. Saunders and John H. Gee. Movements of young Atlantic salmon in a small stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1):27–36, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-004.

Boyd:1964:UIC

[343] J. W. Boyd and B. A. Southcott. Ultraviolet irradiation of circulating refrigerated fish storage brines. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1):37–43, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-005.

Campbell:1964:OCH

[344] N. J. Campbell. The origin of cold high-salinity water in Foxe Basin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1):45–55, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-006.

Kohler:1964:VGA

[345] A. C. Kohler. Variations in the growth of Atlantic cod (Gadus morhua L.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1):57–100, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f64-007.

Dyer:1964:PTF

[346] W. J. Dyer, H. Brockerhoff, R. J. Hoyle, and D. I. Fraser. Polyphosphate treatment of frozen cod. I. Protein extractability and lipid hydrolysis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1):101–106, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-008.

Baslow:1964:NNO

[347] Morris H. Baslow. Neurosine, a new oligopeptide isolated from the brain of fish and other cold-blooded vertebrates: I. Identification and partial characterization. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1): 107–113, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-009.

MacCallum:1964:WTF

[348] W. A. MacCallum and D. G. Ellis. Water-thawing of frozen cod blocks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1):115–131, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-010.

Swift:1964:ACB

[349] D. R. Swift. Activity cycles in the brown trout (Salmo trutta L.): 2. fish artificially fed. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21 (1):133–138, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-011.

Jitts:1964:CDR

[350] H. R. Jitts, C. D. McAllister, K. Stephens, and J. D. H. Strickland. The cell division rates of some marine phytoplankters as a function of light and temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21 (1):139–157, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-012.

McAllister:1964:MPP

[351] C. D. McAllister, N. Shah, and J. D. H. Strickland. Marine phytoplankton photosynthesis as a function of light intensity: A comparison of methods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1):159–181, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f64-013.

Bary:1964:TSP

[352] B. McK. Bary. Temperature, salinity and plankton in the Eastern North Atlantic and coastal waters of Britain, 1957. IV. The species’ relationship to the water body; its role in distribution and in selecting and using indicator species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21 (1):183–202, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-014.

McKenzie:1964:OHS

[353] R. A. McKenzie. Observations on herring spawning off Southwest Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1):203–204, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f64-015.

Chew:1964:FRM

[354] K. K. Chew, A. K. Sparks, and S. C. Katkansky. First record of Mytilicola orientalis Mori in the California mussel Mytilus californianus Conrad. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1):205–207, January 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-016.

Rees:1964:NIS

[355] E. I. S. Rees. Nessorhamphus ingolfianus Schmidt, an oceanic eel, recorded in Canadian continental waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1):209–211, January 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-017.

Anonymous:1964:RAIa

[356] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(1):213–214, January 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-018.

Martin:1964:WCT

[357] W. R. Martin and Yves Jean. Winter cod taggings off Cape Breton and on offshore Nova Scotia Banks, 1959–62. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):215–238, February 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-019.

Thomas:1964:VPE

[358] M. L. H. Thomas and H. R. McCrimmon. Variability in paper electrophoretic patterns of the serum of landlocked sea lamprey, Petromyzon marinus Linnaeus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2): 239–246, February 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-020.

Ackman:1964:SHU

[359] R. G. Ackman. Structural homogeneity in unsaturated fatty acids of marine lipids. A review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):247–254, February 1964. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f64-021.

Connor:1964:CGL

[360] Anne R. Connor, Carl H. Elling, Edgar C. Black, Gerald B. Collins, Joseph R. Gauley, and Edward Trevor-Smith. Changes in glycogen and lactate levels in migrating salmonid fishes ascending experimental “endless” fishways. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2): 255–290, February 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-022.

McNaught:1964:RMP

[361] Donald C. McNaught and Arthur D. Hasler. Rate of movement of populations of Daphnia in relation to changes in light intensity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):291–318, February 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-023.

Ackman:1964:CLO

[362] R. G. Ackman and R. D. Burgher. Cod liver oil: Component fatty acids as determined by gas–liquid chromatography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):319–326, February 1964. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f64-024.

Worthington:1964:ACS

[363] L. V. Worthington. Anomalous conditions in the slope water area in 1959. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):327–333, February 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f64-025.

Patashnik:1964:OMC

[364] Max Patashnik and Herman S. Groninger, Jr. Observations on the milky condition in some Pacific Coast fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):335–346, February 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-026.

Townsley:1964:ESR

[365] P. M. Townsley and M. L. Hughes. Early stages in the recovery to injury in the dorsal fin of the Atlantic cod (Gadus morhua). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):347–354, February 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-027.

Squires:1964:PPP

[366] H. J. Squires. Pagurus pubescens and a proposed new name for a closely related species in the Northwest Atlantic (Decapoda: Anomura). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):355–365, February 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-028.

Ackman:1964:CFC

[367] R. G. Ackman and R. D. Burgher. Cod flesh: Component fatty acids as determined by gas–liquid chromatography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):367–371, February 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-029.

Watling:1964:DOA

[368] Harold Watling and Howard H. Hillemann. The development of the olfactory apparatus of the grayling (Thymallus arcticus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):373–396, February 1964. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f64-030.

Acara:1964:VTV

[369] A. Acara. On the vertical transport velocity on line “P” in the Eastern Subarctic Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):397–407, February 1964. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f64-031.

Anonymous:1964:Oa

[370] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):409–412, February 1964. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f64-032.

May:1964:APC

[371] A. W. May. An asymmetrical pair of cod otoliths. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):413–414, February 1964. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f64-033.

Spalding:1964:AGF

[372] D. J. Spalding. Age and growth of female sea lions in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):415–417, February 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f64-034.

Butler:1964:RSO

[373] T. H. Butler. Records of shrimps (order Decapoda) from British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):419–421, February 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f64-035.

Sipos:1964:ADS

[374] J. C. Sipos and R. G. Ackman. Association of dimethyl sulphide with the “blackberry” problem in cod from the Labrador area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):423–425, February 1964. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f64-036.

Anonymous:1964:RAIb

[375] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(2):427, February 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-037.

Jean:1964:SDC

[376] Yves Jean. Seasonal distribution of cod (Gadus morhua L.) along the Canadian Atlantic Coast in relation to water temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):429–460, March 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-038.

Squires:1964:NAL

[377] H. J. Squires. Neotype of Argis lar compared with Argis dentata (Crustacea, Decapoda). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21 (3):461–467, March 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-039.

Ackman:1964:CRC

[378] R. G. Ackman and R. D. Burgher. Cod Roe: Component fatty acids as determined by gas-liquid chromatography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):469–476, March 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-040.

Larkin:1964:SAP

[379] P. A. Larkin, R. F. Raleigh, and N. J. Wilimovsky. Some alternative premises for constructing theoretical reproduction curves. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):477–484, March 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-041.

Hartman:1964:THS

[380] Wilbur L. Hartman and Robert F. Raleigh. Tributary homing of sockeye salmon at Brooks and Karluk Lakes, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):485–504, March 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-042.

Gilmartin:1964:PPB

[381] Malvern Gilmartin. The primary production of a British Columbia fjord. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):505–538, March 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-043.

MacCallum:1964:PTFa

[382] W. A. MacCallum, H. S. Shieh, Dorothy A. Chalker, W. J. Dyer, and D. R. Idler. Polyphosphate treatment of frozen cod. 2. Effect on drip, yield, lipid hydrolysis and protein extractability in twice-frozen Newfoundland summer trap and fall cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):539–548, March 1964. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f64-044.

Lawler:1964:ILI

[383] G. H. Lawler. Incidence of Ligula intestinalis in Heming Lake fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):549–554, March 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-045.

McHardy:1964:MOP

[384] R. A. McHardy. Marine ostracods from the plankton of Indian Arm, British Columbia, including a diminutive subspecies resembling Conchoecia alata major Rudjakov. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):555–576, March 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-046.

Dyer:1964:CMP

[385] W. J. Dyer, Doris I. Fraser, Ruth G. MacIntosh, and Margaret Myer. Cooking method and palatability of frozen cod fillets of various qualities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):577–589, March 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-047.

John:1964:RCC

[386] Kenneth R. John and Michael Haut. Retinomotor cycles and correlated behavior in the teleost Astyanax mexicanus (Fillipi). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):591–595, March 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-048.

Pitt:1964:FAP

[387] T. K. Pitt. Fecundity of the American plaice, Hippoglossoides platessoides (Fabr.) from Grand Bank and Newfoundland areas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):597–612, March 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-049.

Hunter:1964:NAB

[388] John R. Hunter and Warren J. Wisby. Net avoidance behavior of carp and other species of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21 (3):613–633, March 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-050.

Tsuyuki:1964:CTO

[389] H. Tsuyuki, P. J. Schmidt, and M. Smith. A convenient technique for obtaining pituitary glands from fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):635–637, March 1964. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f64-051.

McHardy:1964:VAS

[390] Robert A. McHardy. A vacuum-assisted subsampler for use with small planktonic organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3): 639–640, March 1964. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f64-052.

Smith:1964:AAC

[391] Michael Smith. Amino acid composition of bacteria and its possible relation to taxonomy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3): 641–642, March 1964. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f64-053.

Keleher:1964:ALS

[392] J. J. Keleher. Angling license sales from Northwest Territories, Canada, with special reference to Great Slave Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):643–645, March 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-054.

Sauer:1964:BDA

[393] Charles D. Sauer. Bathythermograph data on aperture cards: A new approach to an old problem. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):647–650, March 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-055.

Lawler:1964:FEH

[394] G. H. Lawler. Further evidence of hardiness of “silver” pike. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):651–652, March 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-056.

Jonas:1964:ULF

[395] R. E. E. Jonas and E. Bilinski. Utilization of lipids by fish: III. Fatty acid oxidation by various tissues from sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):653–656, March 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-057.

Wickett:1964:ULS

[396] W. Percy Wickett. An unusually late-spawning British Columbia chum salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):657, March 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f64-058.

Anonymous:1964:RAIc

[397] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(3):659–660, March 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-059.

Scarratt:1964:ADL

[398] D. J. Scarratt. Abundance and distribution of lobster larvae (Homarus americanus) in Northumberland Strait. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):661–680, April 1964. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f64-060.

Kabata:1964:RLC

[399] Z. Kabata. Redescription of Lernaeopoda centroscyllii Hansen, 1923 (Copepoda: Lernaeopodidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):681–689, April 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-061.

Vrooman:1964:SDS

[400] Andrew M. Vrooman. Serologically differentiated subpopulations of the Pacific sardine, Sardinops caerulea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):691–701, April 1964. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f64-062.

Dambergs:1964:EFM

[401] N. Dambergs. Extractives of fish muscle. 4. Seasonal variations of fat, water-solubles, protein, and water in cod (Gadus morhua L.) fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):703–709, April 1964. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f64-063.

Smith:1964:TPB

[402] Lynwood S. Smith and Gordon R. Bell. A technique for prolonged blood sampling in free-swimming salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):711–717, April 1964. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f64-064.

MacCallum:1964:TTD

[403] W. A. MacCallum. Temperatures and thaw-drip associated with electronic thawing of Newfoundland cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):719–726, April 1964. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f64-065.

Muir:1964:VSE

[404] Barry S. Muir. Vital statistics of Esox masquinongy in Nogies Creek, Ontario: II. Population size, natural mortality, and the effect of fishing.

Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):727–746, April 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-066.

Ackman:1964:OMF

[405] R. G. Ackman, P. M. Jangaard, R. J. Hoyle, and H. Brockerhoff. Origin of marine fatty acids. I. Analyses of the fatty acids produced by the diatom Skeletonema costatum. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):747–756, April 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-067.

Hester:1964:EFS

[406] Frank J. Hester. Effects of food supply on fecundity in the female guppy, Lebistes reticulatus (Peters). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):757–764, April 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-068.

Bilinski:1964:BTC

[407] E. Bilinski. Biosynthesis of trimethylammonium compounds in aquatic animals: IV. Precursors of trimethylamine oxide and betaine in marine teleosts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):765– 771, April 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f64-069.

Craig:1964:HRE

[408] A. M. Craig. Histology of reproduction and the estrus cycle in the female fur seal, Callorhinus ursinus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):773–811, April 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-070.

Power:1964:UGR

[409] H. E. Power, D. I. Fraser, W. Neal, W. J. Dyer, C. H. Castell, H. C. Freeman, and D. R. Idler. Use of gamma radiation for the preservation of scallop meat. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4): 813–826, April 1964. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f64-071.

Power:1964:GIM

[410] H. E. Power, D. I. Fraser, W. Neal, W. J. Dyer, and C. H. Castell. Gamma irradiation as a means of extending the storage life of haddock fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):827–835, April 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-072.

Kohler:1964:MHN

[411] A. C. Kohler. Movements of halibut on the Nova Scotian and Grand Banks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):837– 840, April 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f64-073.

Ackman:1964:POA

[412] R. G. Ackman and J. C. Sipos. Pilchard oil: An analysis for component fatty acids with particular reference to the C_{24} chain length. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):841–843, April 1964. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f64-074.

Roach:1964:ALP

[413] S. W. Roach, F. G. Claggett, and J. S. M. Harrison. An air-lift pump for elevating salmon, herring, and other fish of similar size. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):845–849, April 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-075.

Hitz:1964:OEC

[414] Charles R. Hitz. Observations on egg cases of the big skate (Raja binoculata Girard) found in Oregon coastal waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):851–854, April 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-076.

Westrheim:1964:RSB

[415] S. J. Westrheim. Rockfish (Sebastodes brevispinis) in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):855– 856, April 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f64-077.

Tomlinson:1964:AOR

[416] N. Tomlinson, S. E. Geiger, and W. W. Kay. Apparent onset of rigor mortis in steelhead trout (Salmo gairdneri) in the absence of loss of adenosine triphosphate from the ordinary muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):857–859, April 1964. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f64-078.

LeBrasseur:1964:SCB

[417] R. J. LeBrasseur. Stomach contents of blue shark (Prionace glauca L.) taken in the Gulf of Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):861–862, April 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-079.

McKenzie:1964:MDP

[418] R. A. McKenzie and S. N. Tibbo. A morphometric description of porbeagle (Lamna nasus) from Canadian Atlantic waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):863–864, April 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-080.

McKenzie:1964:MDB

[419] R. A. McKenzie and S. N. Tibbo. A morphometric description of blue shark (Prionace glauca) from Canadian Atlantic waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):865–866, April 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-081.

Anonymous:1964:RAId

[420] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(4):867–868, April 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-082.

Anonymous:1964:EF

[421] Anonymous. Editor’s foreword. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5), May 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-083.

Anonymous:1964:Ob

[422] Anonymous. Obituary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):869–872, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-084.

Parker:1964:NSP

[423] R. R. Parker and L. Margolis. A new species of parasitic copepod, Caligus clemensi sp. nov. (Caligoida: Caligidae), from pelagic fishes in the coastal waters of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):873–889, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-085.

Manzer:1964:POV

[424] J. I. Manzer. Preliminary observations on the vertical distribution of Pacific salmon (genus Oncorhynchus) in the Gulf of Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):891–903, May 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-086.

Ricker:1964:OGM

[425] W. E. Ricker. Ocean growth and mortality of pink and chum salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):905–931, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-087.

Rosenblatt:1964:NGP

[426] Richard H. Rosenblatt. A new gunnel, Pholis clemensi, from the coast of Western North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):933–939, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-088.

Tully:1964:ORA

[427] John P. Tully. Oceanographic regions and assessment of temperature structure in the seasonal zone of the North Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):941–970, May 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-089.

Butler:1964:RPI

[428] T. H. Butler. Redescription of the parasitic isopod Holophryxus alaskensis Richardson, and a note on its synonymy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):971–976, May 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-090.

Lindsey:1964:PZL

[429] C. C. Lindsey. Problems in zoogeography of the lake trout, Salvelinus namaycush. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5): 977–994, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f64-091.

McInerney:1964:SPO

[430] John E. McInerney. Salinity preference: an orientation mechanism in salmon migration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5): 995–1018, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f64-092.

Parker:1964:ESM

[431] Robert R. Parker. Estimation of sea mortality rates for the 1960 broodyear pink salmon of Hook Nose Creek, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1019–1034, May 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-093.

Idler:1964:SVS

[432] D. R. Idler, T. Tamura, and T. Wainai. Seasonal variations in the sterol, fat and unsaponifiable components of scallop muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1035–1042, May 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-094.

Scott:1964:TRP

[433] D. P. Scott. Thermal resistance of pike (Esox lucius L.), muskellunge (E. masquinongy Mitchill), and their F_{1} hybrid. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1043–1049, May 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-095.

Ketchen:1964:PRS

[434] K. S. Ketchen. Preliminary results of studies on growth and mortality of Pacific cod (Gadus macrocephalus) in Hecate Strait, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1051–1067, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-096.

Northcote:1964:SCN

[435] T. G. Northcote, Mildred S. Wilson, and D. R. Hurn. Some characteristics of Nitinat Lake, an inlet on Vancouver Island, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1069–1081, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-097.

Pfeiffer:1964:CCG

[436] Wolfgang Pfeiffer and T. F. Pletcher. Club cells and granular cells in the skin of lamprey. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21 (5):1083–1088, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-098.

Nicol:1964:RCT

[437] J. A. C. Nicol. Reflectivity of the chorioidal tapeta of selachians. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1089–1100, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-099.

LeBrasseur:1964:GRC

[438] R. J. LeBrasseur and R. R. Parker. Growth rate of Central British Columbia pink salmon (Oncorhynchus gorbuscha). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1101–1128, May 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-100.

Hubbs:1964:DSP

[439] Carl L. Hubbs and Norman J. Wilimovsky. Distribution and synonymy in the Pacific Ocean, and variation, of the Greenland halibut, Reinhardtius hippoglossoides (Walbaum). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1129–1154, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-101.

Quayle:1964:DIM

[440] D. B. Quayle. Distribution of introduced marine Mollusca in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21 (5):1155–1181, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-102.

Brett:1964:RMS

[441] J. R. Brett. The respiratory metabolism and swimming performance of young sockeye salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1183–1226, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-103.

Neave:1964:OMP

[442] Ferris Neave. Ocean migrations of Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1227–1244, May 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-104.

Larkin:1964:MSP

[443] P. A. Larkin and A. S. Hourston. A model for simulation of the population biology of Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1245–1265, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-105.

Bilton:1964:KFS

[444] H. T. Bilton, D. W. Jenkinson, and M. P. Shepard. A key to five species of Pacific salmon (genus Oncorhynchus) based on scale characters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1267–1288, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-106.

Waldichuk:1964:DKM

[445] Michael Waldichuk. Dispersion of kraft-mill effluent from a submarine diffuser in Stuart Channel, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1289–1316, May 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-107.

Outram:1964:QEN

[446] D. N. Outram and F. H. C. Taylor. A quantitative estimate of the number of Pacific herring in a spawning population. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1317–1320, May 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-108.

Northcote:1964:ODS

[447] T. G. Northcote and W. E. Johnson. Occurrence and distribution of sea water in Sakinaw Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1321–1324, May 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-109.

Sparrow:1964:SIM

[448] R. A. H. Sparrow, P. A. Larkin, and R. A. Rutherglen. Successful introduction of Mysis relicta Lovén into Kootenay Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1325–1327, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-110.

Shepard:1964:FIS

[449] M. P. Shepard, F. C. Withler, J. McDonald, and K. V. Aro. Further information on spawning stock size and resultant production for skeena sockeye. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1329– 1331, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f64-111.

Barraclough:1964:CML

[450] W. E. Barraclough. Contribution to the marine life history of the eulachon Thaleichthys pacificus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1333–1337, May 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-112.

McAllister:1964:DCS

[451] D. E. McAllister. Distinguishing characters for the sculpins Cottus bairdii and C. cognatus in Eastern Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1339–1342, May 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-113.

Anonymous:1964:RAIe

[452] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(5):1343–1344, May 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-114.

MacLean:1964:RLF

[453] Jill MacLean and C. H. Castell. Rancidity in lean fish muscle: I. A proposed accelerated copper-catalyzed method for evaluating the tendency of fish muscle to become rancid. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1345–1359, June 1964. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f64-115.

Castell:1964:RLFa

[454] C. H. Castell and Jill MacLean. Rancidity in lean fish muscle: II. Anatomical and seasonal variations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1361–1369, June 1964. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f64-116.

Castell:1964:RLFb

[455] C. H. Castell and Jill MacLean. Rancidity in lean fish muscle: III. The inhibiting effect of bacterial activity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1371–1377, June 1964. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f64-117.

Brockerhoff:1964:FAD

[456] H. Brockerhoff, M. Yurkowski, R. J. Hoyle, and R. G. Ackman. Fatty acid distribution in lipids of marine plankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1379–1384, June 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-118.

McNeil:1964:RSE

[457] William J. McNeil. Redd superimposition and egg capacity of pink salmon spawning beds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1385–1396, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-119.

MacCallum:1964:PTFb

[458] W. A. MacCallum, Dorothy A. Chalker, J. T. Lauder, P. H. Odense, and D. R. Idler. Polyphosphate treatment of frozen cod. 3. Taste panel evaluation, chemical assessment and thaw-drip in once-frozen Newfoundland trap-caught cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21 (6):1397–1402, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-120.

Butler:1964:GRD

[459] T. H. Butler. Growth, reproduction, and distribution of pandalid shrimps in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21 (6):1403–1452, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-121.

John:1964:IVS

[460] Kenneth R. John. Illumination, vision, and schooling of Astyanax mexicanus (Fillipi). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6): 1453–1473, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-122.

MacIntyre:1964:DCC

[461] W. G. MacIntyre and R. F. Platford. Dissolved calcium carbonate in the Labrador Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21 (6):1475–1480, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-123.

Ronald:1964:VSC

[462] A. P. Ronald and W. A. B. Thomson. The volatile sulphur compounds of oysters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1481– 1487, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f64-124.

Power:1964:CNV

[463] H. E. Power. Characteristics and nutritional value of various fish protein concentrates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21 (6):1489–1504, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-125.

Trason:1964:ACA

[464] Winona Bethune Trason. Ascidians of the Canadian Arctic waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1505–1517, June 1964. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f64-126.

Tamura:1964:SMO

[465] T. Tamura, B. Truscott, and D. R. Idler. Sterol metabolism in the oyster. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1519–1522, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-127.

Vladykov:1964:OTB

[466] Vadim D. Vladykov. Ouest for the true breeding area of the American eel (Anguilla rostrata LeSueur). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1523–1530, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-128.

Forrester:1964:DQF

[467] C. R. Forrester. Demersal quality of fertilized eggs of rock sole (Lepidopsetta bilineata Ayres). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1531–1532, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-129.

Forrester:1964:RDE

[468] C. R. Forrester. Rate of development of eggs of rock sole (Lepidopsetta bilineata Ayres). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21 (6):1533–1534, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-130.

Houston:1964:PFO

[469] Arthur Hillier Houston. On passive features in the osmoregulatory adaptation of anadromous salmonids to sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1535–1538, June 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-131.

Black:1964:EMG

[470] Edgar C. Black and Anne R. Connor. Effects of MS 222 on glycogen and lactate levels in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1539–1542, June 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-132.

Wainai:1964:AGC

[471] T. Wainai, T. Tamura, B. Truscott, and D. R. Idler. The application of gas chromatography to the identification of the sterols of scallop. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1543–1546, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f64-133.

Lawler:1964:NPE

[472] G. H. Lawler. A northern pike, Esox lucius, with an accessory fin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1547, June 1964. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f64-134.

Stewart:1964:LHA

[473] James E. Stewart and John W. Cornick. Lobster (Homarus americanus) tolerance for tris buffer, sodium fluoride, and sea water extracts of various Woods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1549– 1551, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f64-135.

Anonymous:1964:CRP

[474] Anonymous. Corrections for recent publications of The Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1553–1554, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-136.

Anonymous:1964:RAIf

[475] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21(6):1555–1556, June 1964. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f64-137.

Anonymous:1964:IVJ

[476] Anonymous. Index to volume 21 of the Journal of the Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 21 (6):1557–1569, June 1964. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f64-138.

Squires:1965:NSC

[477] H. J. Squires. A new species of Calocaris (Crustacea: Decapoda, Thalassinidea) from the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):1–11, January 1965. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f65-001.

Munro:1965:ESS

[478] I. C. Munro and A. B. Morrison. Effects of salting and smoking on protein quality of cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1): 13–16, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f65-002.

Yurkowski:1965:CCSa

[479] M. Yurkowski. Carbonyl compounds in salted cod: I. Thin-layer chromatography of aliphatic monocarbonyl 2,4-dinitrophenylhydrazones. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):17–25, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-003.

Yurkowski:1965:CCSb

[480] M. Yurkowski and M. A. Bordeleau. Carbonyl compounds in salted cod: II. Separation and identification of volatile monocarbonyl compounds from heavily salted cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):27–32, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-004.

Fredin:1965:SME

[481] R. A. Fredin. Some methods for estimating ocean mortality of Pacific salmon and applications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):33–51, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-005.

Boyd:1965:EPO

[482] J. W. Boyd and B. A. Southcott. Effect of polyphosphates and other salts on drip loss and oxidative rancidity of frozen fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):53–67, January 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-006.

Squires:1965:LMA

[483] H. J. Squires. Larvae and megalopa of Argis dentata (Crustacea: Decapoda) from Ungava Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):69–82, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-007.

Fraser:1965:PMG

[484] Doris I. Fraser, H. M. Weinstein, and W. J. Dyer. Post-mortem glycolytic and associated changes in the muscle of trap- and trawl-caught cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):83–100, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-008.

Powell:1965:RKC

[485] Guy C. Powell and Richard B. Nickerson. Reproduction of king crabs, Paralithodes camtschatica (Tilesius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):101–111, January 1965. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f65-009.

Platford:1965:ACM

[486] R. F. Platford. Activity coefficient of the magnesium ion in sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):113–116, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-010.

Southcott:1965:TAS

[487] B. A. Southcott and J. W. Boyd. Tetracycline antibiotics in shrimp preservation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1): 117–129, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-011.

Jangaard:1965:LCF

[488] Peter M. Jangaard and R. G. Ackman. Lipids and component fatty acids of the Newfoundland squid, Illex illecebrosus (Le Sueur). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):131–137, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-012.

Vladykov:1965:LRN

[489] Vadim D. Vladykov and W. I. Follett. Lampetra richardsoni, a new nonparasitic species of lamprey (Petromyzonidae) from Western North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):139–158, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-013.

Daykin:1965:AMT

[490] P. N. Daykin. Application of mass transfer theory to the problem of respiration of fish eggs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22 (1):159–171, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-014.

Mason:1965:SEE

[491] J. C. Mason and D. W. Chapman. Significance of early emergence, environmental rearing capacity, and behavioral ecology of juvenile Coho salmon in stream channels. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):173–190, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-015.

Murphy:1965:SCE

[492] Garth I. Murphy. A solution of the catch equation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):191–202, January 1965. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f65-016.

Tsuyuki:1965:CZE

[493] H. Tsuyuki, E. Roberts, and W. E. Vanstone. Comparative zone electropherograms of muscle myogens and blood hemoglobins of marine and freshwater vertebrates and their application to biochemical systematics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):203–213, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-017.

Tsuyuki:1965:SSC

[494] H. Tsuyuki, E. Roberts, W. E. Vanstone, and J. R. Markert. The species specificity and constancy of muscle myogen and hemoglobin electropherograms of Oncorhynchus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):215–217, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-018.

Lawler:1965:FAW

[495] G. H. Lawler. Fin anomalies in the white sucker, Catostomus commersoni, population in Heming Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):219–220, January 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-019.

Ali:1965:RSA

[496] M. A. Ali. Retinal structure in the Arctic char (Salvelinus alpinus L.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):221–223, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-020.

Barber:1965:COF

[497] F. G. Barber. Current observations in Fury and Hecla Strait. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):225–229, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-021.

Westrheim:1965:NRE

[498] S. J. Westrheim. Northern range extensions for four species of rockfish (Sebastodes goodei, S. helvomaculatus, S. rubrivinctus, and S. zacentrus) in the North Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):231–235, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-022.

Templeman:1965:SAS

[499] Wilfred Templeman. Some abnormalities in skates (Raja) of the Newfoundland area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1): 237–238, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-023.

Johnson:1965:SBL

[500] Lionel Johnson. The salinity of Baker Lake, N.W.T., Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):239–241, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-024.

Conroy:1965:SSA

[501] D. A. Conroy and E. B. Solarolo. Sensitivity of some acid-fast bacteria of piscine origin to certain chemotherapeutic substances. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):243–245, January 1965. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f65-025.

Pitt:1965:MWV

[502] T. K. Pitt. Modification of the whirling vessel for fecundity studies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):247–251, January 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-026.

Anonymous:1965:RAIa

[503] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(1):253–254, January 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-027.

Anonymous:1965:O

[504] Anonymous. Obituary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):255–258, February 1965. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f65-028.

Templeman:1965:RSN

[505] Wilfred Templeman. Rare skates of the Newfoundland and neighbouring areas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):259–279, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-029.

Medcof:1965:ELG

[506] J. C. Medcof and C. J. Kerswill. Effects of light on growth of oysters, mussels, and quahaugs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2): 281–288, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-030.

Battle:1965:CSI

[507] Helen I. Battle and Kaye Hayashida. Comparative study of the intraperitoneal alimentary tract of parasitic and nonparasitic lampreys from the Great Lakes region. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2): 289–306, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-031.

Tarr:1965:NRC

[508] H. L. A. Tarr and A. G. Comer. Nucleotides and related compounds, sugars, and homarine in shrimp. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):307–311, February 1965. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f65-032.

Bourne:1965:CCI

[509] Neil Bourne. A comparison of catches by 3- and 4-inch rings on offshore scallop drags. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2): 313–333, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-033.

Saunders:1965:ABY

[510] Richard L. Saunders. Adjustment of buoyancy in young Atlantic salmon and brook trout by changes in swimbladder volume. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):335–352, February 1965. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f65-034.

Lauzier:1965:DBO

[511] L. M. Lauzier. Drift bottle observations in Northumberland Strait, Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2): 353–368, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-035.

Pinhorn:1965:EPB

[512] A. T. Pinhorn and C. W. Andrews. Effect of photoperiods on the behavior of juvenile Atlantic salmon (Salmo salar L.) in vertical and horizontal light gradients. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):

369–383, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-036.

McLeese:1965:SLH

[513] D. W. McLeese. Survival of lobsters, Homarus americanus, out of water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):385–394, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-037.

Saunders:1965:CSP

[514] J. W. Saunders and M. W. Smith. Changes in a stream population of trout associated with increased silt. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):395–404, February 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-038.

Brett:1965:RMP

[515] J. R. Brett and D. B. Sutherland. Respiratory metabolism of pumpkinseed (Lepomis gibbosus) in relation to swimming speed. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):405–409, February 1965. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f65-039.

MacCallum:1965:SQN

[516] W. A. MacCallum, Dorothy A. Chalker, E. J. Laishley, W. J. Dyer, and D. R. Idler. Studies on the quality of Newfoundland cod. 10. Effect of some commercial freezing rates on the frozen and stored quality of trap cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):411–420, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-040.

Scarratt:1965:PTT

[517] D. J. Scarratt and P. F. Elson. Preliminary trials of a tag for salmon and lobsters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):421–423, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-041.

Sprague:1965:LLM

[518] J. B. Sprague and B. Ann Ramsay. Lethal levels of mixed copper–zinc solutions for juvenile salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):425–432, February 1965. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f65-042.

Sergeant:1965:MHS

[519] D. E. Sergeant. Migrations of harp seals Pagophilus groenlandicus (Erxleben) in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):433–464, February 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-043.

Fleming:1965:TCA

[520] A. M. Fleming. Trends in catch, age, and size of cod from the commercial longline fishery at Bonavista, Newfoundland, 1952–62. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):465–474, February 1965. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f65-044.

Ricker:1965:NRD

[521] W. E. Ricker. New records and descriptions of Plecoptera (class Insecta). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):475–501, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-045.

Ogilvie:1965:EDT

[522] D. M. Ogilvie and J. M. Anderson. Effect of DDT on temperature selection by young Atlantic salmon, Salmo salar. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):503–512, February 1965. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f65-046.

Nicol:1965:RCF

[523] J. A. C. Nicol. Retinomotor changes in flatfishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):513–520, February 1965. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f65-047.

Paloheimo:1965:FGF

[524] J. E. Paloheimo and L. M. Dickie. Food and growth of fishes: I. A growth curve derived from experimental data. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):521–542, February 1965. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f65-048.

Grainger:1965:ZAO

[525] E. H. Grainger. Zooplankton from the Arctic Ocean and adjacent Canadian Waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):543–564, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-049.

Powles:1965:LHE

[526] P. M. Powles. Life history and ecology of American plaice (Hippoglossoides platessoides F.) in the Magdalen Shallows. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):565–598, February 1965. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f65-050.

Blair:1965:BIH

[527] A. A. Blair. Bay of Islands and Humber River Atlantic salmon investigations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):599–620, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-051.

Jean:1965:SDM

[528] Yves Jean. Seasonal distribution of monkfish along the Canadian Atlantic Mainland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):621–624, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-052.

Saunders:1965:CAS

[529] R. L. Saunders, C. J. Kerswill, and P. F. Elson. Canadian Atlantic salmon recaptured near Greenland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):625–629, February 1965. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f65-053.

Medcof:1965:ACE

[530] J. C. Medcof, Arthur H. Clarke, Jr., and John S. Erskine. Ancient Canadian East-Coast oyster and quahaug shells. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):631–634, February 1965. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f65-054.

White:1965:KP

[531] H. C. White, J. C. Medcof, and L. R. Day. Are killifish poisonous? Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):635–638, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-055.

McLeese:1965:LAM

[532] D. W. McLeese. Lesions on the abdominal membrane of lobsters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):639–640, February 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-056.

Anonymous:1965:RAIb

[533] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(2):641–642, February 1965. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f65-057.

Yurkowski:1965:LCM

[534] M. Yurkowski and H. Brockerhoff. Lysolecithinase of cod muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):643–652, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-058.

Tomlinson:1965:CHR

[535] N. Tomlinson, S. E. Geiger, and E. Dollinger. Chalkiness in halibut in relation to muscle pH and protein denaturation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):653–663, March 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-059.

Lorz:1965:FAS

[536] H. W. Lorz and T. G. Northcote. Factors affecting stream location, and timing and intensity of entry by spawning Kokanee (Oncorhynchus nerka) into an inlet of Nicola Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):665–687, March 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-060.

Favorite:1965:FTO

[537] F. Favorite, D. Fisk, and W. J. Ingraham, Jr. First transponding oceanographic buoys in the Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):689–694, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-061.

Applegate:1965:SRS

[538] Vernon C. Applegate and M. L. H. Thomas. Sex ratios and sexual dimorphism among recently transformed sea lampreys, Petromyzon marinus Linnaeus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3): 695–711, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f65-062.

Ross:1965:OIR

[539] A. J. Ross and G. W. Klontz. Oral immunization of rainbow trout (Salmo gairdneri) against an etiologic agent of “redmouth disease”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):713–719, March 1965. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f65-063.

Nelson:1965:EFI

[540] J. S. Nelson. Effects of fish introductions and hydroelectric development on fishes in the Kananaskis River System, Alberta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):721–753, March 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-064.

Tarr:1965:BFD

[541] H. L. A. Tarr and R. E. A. Gadd. Browning of freeze-dried fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):755–760, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-065.

Hoffert:1965:BGM

[542] J. Russell Hoffert and Paul O. Fromm. Biomicroscopic, Gross, and microscopic observations of corneal lesions in the lake trout, Salvelinus namaycush. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):761–766, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-066.

Tsuyuki:1965:ZEC

[543] H. Tsuyuki and Eve Roberts. Zone electrophoretic comparison of muscle myogens and blood proteins of artificial hybrids of Salmonidae with their parental species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22 (3):767–773, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-067.

McBride:1965:PSD

[544] J. R. McBride, U. H. M. Fagerlund, M. Smith, and N. Tomlinson. Post-spawning death of Pacific salmon: Sockeye salmon (Oncorhynchus nerka) maturing and spawning in captivity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):775–782, March 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-068.

Chalker:1965:SQN

[545] Dorothy A. Chalker, W. A. MacCallum, and D. R. Idler. Studies on the quality of Newfoundland cod: 11. Thaw-drip in polyphosphate-treated and untreated fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22 (3):783–791, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-069.

Antia:1965:PAS

[546] N. J. Antia and A. Watt. Phosphatase activity in some species of marine phytoplankters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3): 793–799, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f65-070.

Waldichuk:1965:WEP

[547] Michael Waldichuk. Water exchange in Port Moody, British Columbia, and its effect on waste disposal. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):801–822, March 1965. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f65-071.

McHardy:1965:DSC

[548] R. A. McHardy and B. McK. Bary. Diurnal and seasonal changes in distribution of two planktonic ostracods, Conchoecia elegans and Conchoecia alata minor. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22 (3):823–840, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-072.

Renshaw:1965:DMM

[549] R. Ward Renshaw. Distribution and morphology of the Medusa, Calycopsis nematophora, from the North Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):841–847, March 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-073.

Boyd:1965:MPS

[550] J. W. Boyd and N. Tomlinson. Myxosporidian parasite from salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):849–850, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-074.

Thomas:1965:MDS

[551] M. L. H. Thomas. Manganese deposits on shells of Eastern mud snails, Nassarius obsoletus, in Bideford River, Prince Edward Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):851–852, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-075.

Manzer:1965:SLG

[552] J. I. Manzer. Sarda lineolata (Girard), a Pacific bonito, in the Strait of Georgia, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):853–855, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-076.

Miller:1965:JSG

[553] Bruce Miller and David W. Greenfield. A juvenile six-gilled shark (Hexanchus corinus) from the San Juan Islands, Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):857–859, March 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-077.

Bourne:1965:ORM

[554] Neil Bourne and E. G. Bligh. Orange-red meats in sea scallops. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):861–864, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-078.

Budde:1965:CDT

[555] W. Budde. Candling for the detection of Triaenophorus crassus cysts in whitefish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):865– 867, March 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-079.

Cokkoy:1965:HSH

[556] D. A. Cokkoy and J. L. Rodriguez. Hematological study of the hake (Merluccius merluccius) from the Southwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):869–871, March 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-080.

Anonymous:1965:RAIc

[557] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(3):873–874, March 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-081.

Ackman:1965:RDD

[558] R. G. Ackman and J. Dale. Reactor for determination of dimethyl-β-propiothetin in tissue of marine origin by gas–liquid chromatography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):875–883, April 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-082.

Platford:1965:ACS

[559] R. F. Platford. Activity coefficient of the sodium ion in sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):885–889, April 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-083.

Jonas:1965:KBB

[560] R. E. E. Jonas and E. Bilinski. Ketone bodies in the blood of salmonoid fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):891–898, April 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-084.

Templeman:1965:SRD

[561] Wilfred Templeman. Some resemblances and differences between Raja erinacea and Raja ocellata, including a method of separating mature and large-immature individuals of these two species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):899–912, April 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-085.

Swift:1965:ETM

[562] D. R. Swift. Effect of temperature on mortality and rate of development of the eggs of the Windermere char (Salvelinus alpinus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):913–917, April 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-086.

Shapley:1965:SSS

[563] S. Philip Shapley and Daniel M. Bishop. Sedimentation in a salmon stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):919– 928, April 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-087.

Castell:1965:RLF

[564] C. H. Castell, Jill MacLean, and Barbara Moore. Rancidity in lean fish muscle: IV. Effect of sodium chloride and other salts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):929–944, April 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-088.

Saila:1965:ETB

[565] Saul B. Saila, Donald B. Horton, and Richard J. Berry. Estimates of the theoretical biomass of juvenile winter flounder, Pseudopleuronectes americanus (Walbaum) required for a fishery in Rhode Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):945–954, April 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-089.

Tomlinson:1965:PFM

[566] N. Tomlinson, S. E. Geiger, W. W. Kay, J. Uthe, and S. W. Roach. Partial freezing as a means of preserving Pacific salmon intended for canning. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):955–968, April 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-090.

Martin:1965:DCC

[567] N. V. Martin and L. J. Chapman. Distribution of certain crustaceans and fishes in the region of Algonquin Park, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):969–976, April 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-091.

Allen:1965:RMN

[568] J. A. Allen. Records of Mollusca from the Northwest Atlantic obtained by Canadian fishery research vessels, 1946–61. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):977–997, April 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-092.

Pycha:1965:MHR

[569] Richard L. Pycha, William R. Dryer, and George R. King. Movements of hatchery-reared lake trout in Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):999–1024, April 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-093.

Groot:1965:TLP

[570] C. Groot and W. L. Wiley. Time-lapse photography of an ASDIC echosounder PPI-scope as a technique for recording fish movements during migration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):1025– 1034, April 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-094.

Hartman:1965:RBE

[571] G. F. Hartman. The role of behavior in the ecology and interaction of underyearling Coho salmon (Oncorhynchus kisutch) and steelhead trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22 (4):1035–1081, April 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-095.

Eppley:1965:CBE

[572] Richard W. Eppley and Phillip R. Sloan. Carbon balance experiments with marine phytoplankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):1083–1097, April 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-096.

Chew:1965:PRS

[573] K. K. Chew, A. K. Sparks, and S. C. Katkansky. Preliminary results on the seasonal size distribution of Mytilicola orientalis and the effect of this parasite on the condition of the Pacific oyster, Crassostrea gigas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):1099–1101, April 1965.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-097.

Scarratt:1965:PLH

[574] D. J. Scarratt. Predation on lobsters (Homarus americanus) by Anonyx sp. (Crustacea, Amphipoda). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):1103–1104, April 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-098.

Anonymous:1965:RAId

[575] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(4):1105–1106, April 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-099.

Williams:1965:FAD

[576] P. M. Williams. Fatty acids derived from lipids of marine origin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5):1107–1122, May 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-100.

Webb:1965:LFC

[577] D. W. Webb. Limnological features of Cedar Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5):1123–1136, May 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-101.

Bourne:1965:PSP

[578] N. Bourne. Paralytic shellfish poison in sea scallops (Placopecten magellanicus, Gmelin). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22 (5):1137–1149, May 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-102.

Grinols:1965:DDT

[579] Richard B. Grinols and Hiromu Heyamoto. Description, distribution, and taxonomic status of two species of Alepocephalidae from the Northeastern Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22 (5):1151–1164, May 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-103.

Driver:1965:LAS

[580] E. A. Driver. Limnological aspects of some saline lakes in West-Central Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5):1165– 1173, May 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-104.

Becker:1965:DEB

[581] Clarence D. Becker and Max Katz. Distribution, ecology, and biology of the salmonid leech, Piscicola salmositica (Rhynchobdellae: Piscicolidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5):1175–1195, May 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-105.

Lawler:1965:FSY

[582] G. H. Lawler. Fluctuations in the success of year-classes of whitefish populations with special reference to Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5):1197–1227, May 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-106.

McCart:1965:GMF

[583] P. McCart. Growth and morphometry of four British Columbia populations of pygmy whitefish (Prosopium coulteri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5):1229–1259, May 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-107.

Crossman:1965:HFE

[584] E. J. Crossman and Keen Buss. Hybridization in the family Esocidae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5):1261–1292, May 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-108.

McPhail:1965:NRB

[585] J. D. McPhail. A new ronquil, Bathymaster leurolepis, from the Aleutian Islands. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5):1293– 1297, May 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-109.

Macnab:1965:BCT

[586] H. C. Macnab and K. Ronald. Blood clotting time of the Atlantic cod, Gadus morhua. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5): 1299–1301, May 1965. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f65-110.

Scott:1965:RHP

[587] W. B. Scott. A record of the hatchetfish, Polyipnus asteroides Schultz, from the Canadian Atlantic region. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5):1303–1304, May 1965. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f65-111.

Barraclough:1965:FRD

[588] W. E. Barraclough and T. H. Butler. First record of the dusky sculpin (Icelinus burchami) in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5):1305–1307, May 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-112.

Davenport:1965:MRM

[589] D. Davenport and W. R. Hakling. Method of rapid measurement for large samples of fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22 (5):1309–1310, May 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-113.

Lippa:1965:UCG

[590] E. J. R. Lippa. Unusual contamination of groundfish caught in Northern Hecate Strait. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5): 1311–1312, May 1965. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f65-114.

Richards:1965:DPS

[591] Sarah W. Richards. Description of the postlarvae of the sand lance (Ammodytes) from the East Coast of North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5):1313–1317, May 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-115.

Anonymous:1965:RAIe

[592] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(5):1319–1320, May 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-116.

Lauzier:1965:CDC

[593] L. M. Lauzier and A. Marcotte. Comparaison du climat marin de Grande-Rivière (baie des Chaleurs) avec celui d’autres stations de la côte atlantique. (French) [Comparison of the marine climate of Grande-Rivière (Baie des Chaleurs) with that of other resorts on the Atlantic coast]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1321–1334, June 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-117.

Watson:1965:NIC

[594] N. H. F. Watson and G. H. Lawler. Natural infections of cyclopid copepods with procercoids of Triaenophorus spp. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1335–1343, June 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-118.

Templeman:1965:LDA

[595] Wilfred Templeman. Lymphocystis disease in American plaice of the Eastern Grand Bank. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22 (6):1345–1356, June 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-119.

Lawler:1965:FPE

[596] G. H. Lawler. The food of the pike, Esox lucius, in Heming Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1357– 1377, June 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-120.

Gross:1965:IPG

[597] W. L. Gross, E. W. Roelofs, and P. O. Fromm. Influence of photoperiod on growth of green sunfish, Lepomis cyanellus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1379–1386, June 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-121.

Margolis:1965:PAS

[598] L. Margolis. Parasites as an auxiliary source of information about the biology of Pacific salmons (genus Oncorhynchus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1387–1395, June 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-122.

Wittenberger:1965:EML

[599] C. Wittenberger and I. V. Diaciuc. Effort metabolism of lateral muscles in carp. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1397– 1406, June 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-123.

Astrahantseff:1965:BDB

[600] Sergei Astrahantseff and Miles S. Alton. Bathymetric distribution of brittlestars (Ophiuroidea) collected off the Northern Oregon Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1407–1424, June 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-124.

Widdowson:1965:SDI

[601] Thomas B. Widdowson. A survey of the distribution of intertidal algae along a coast transitional in respect to salinity and tidal factors. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1425–1454, June 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-125.

Heinicke:1965:ETA

[602] Elizabeth Anne Heinicke and A. H. Houston. Effect of thermal acclimation and sublethal heat shock upon ionic regulation in the goldfish, Carassius auratus L. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22 (6):1455–1476, June 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-126.

Bilton:1965:SCS

[603] H. T. Bilton and W. E. Ricker. Supplementary checks on the scales of pink salmon (Oncorhynchus gorbuscha) and chum salmon (O. keta). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1477–1489, June 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-127.

Brett:1965:RSR

[604] J. R. Brett. The relation of size to rate of oxygen consumption and sustained swimming speed of sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1491–1501, June 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-128.

Foerster:1965:ERS

[605] R. E. Foerster. Effect of retention of spermatozoa and ova of sockeye salmon, Oncorhynchus nerka, in water and without addition of water, on fertility. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1503– 1521, June 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f65-129.

Parker:1965:ESM

[606] Robert R. Parker. Estimation of sea mortality rates for the 1961 broodyear pink salmon of the bella coola area, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1523–1554, June 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-130.

Brawn:1965:AMD

[607] Vivien M. Brawn. Alexander’s method of determining swimbladder volumes modified for living physoclist fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1555–1558, June 1965. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f65-131.

McAllister:1965:WRS

[608] D. E. McAllister and S. J. Westrheim. Widow rockfish, Sebastodes entomelas, new to British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1559–1561, June 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-132.

LeBrasseur:1965:RRD

[609] R. J. LeBrasseur and K. Stephens. Relative rates of degradation of some organisms consumed by marine salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1563–1564, June 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-133.

Powles:1965:NSR

[610] P. M. Powles. New size record for greysole (Glyptocephalus cynoglossus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1565–1566, June 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-134.

Saunders:1965:PSD

[611] Richard L. Saunders and Eugene B. Henderson. Precocious sexual development in male post-smolt Atlantic salmon reared in the laboratory. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1567–1570, June 1965. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f65-135.

Keleher:1965:FDF

[612] J. J. Keleher and C. G. Haight. The fall domestic fishery at Snowdrift, Northwest Territories. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1571–1573, June 1965. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f65-136.

Stephens:1965:PPN

[613] K. Stephens. Phytoplankton pigment nomographs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1575–1578, June 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-137.

Anonymous:1965:RAIf

[614] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1579–1580, June 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-138.

Anonymous:1965:IP

[615] Anonymous. Index to publications for 1965. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 22(6):1581–1615, June 1965. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f65-139.

Black:1966:ODCa

[616] Edgar C. Black, Donald Kirkpatrick, and Harold H. Tucker. Oxygen dissociation curves of the blood of brook trout (Salvelinus fontinalis) acclimated to summer and winter temperatures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(1):1–13, January 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-001.

Coble:1966:DTA

[617] Daniel W. Coble. Dependence of total annual growth in yellow perch on temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23 (1):15–20, January 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-002.

Jangaard:1966:DCL

[618] P. M. Jangaard and H. E. Power. Decomposition of cod livers and the formation of free fatty acids in cod oil. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(1):21–25, January 1966. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f66-003.

Castell:1966:RLF

[619] C. H. Castell, Jill MacLean, Barbara Moore, and Wanda Neal. Rancidity in lean fish muscle. V. The effect of amino acids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(1):27–43, January 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-004.

McKnight:1966:HSM

[620] Ida M. McKnight. A hematological study on the mountain whitefish, Prosopium williamsoni. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(1):45–64, January 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-005.

Hammond:1966:EPC

[621] Brian R. Hammond and Cleveland P. Hickman, Jr. The effect of physical conditioning on the metabolism of lactate, phosphate, and glucose in rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(1):65–83, January 1966. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f66-006.

LeBrasseur:1966:SCS

[622] R. J. LeBrasseur. Stomach contents of salmon and steelhead trout in the Northeastern Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(1):85–100, January 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-007.

Tsuyuki:1966:ISR

[623] H. Tsuyuki and E. Roberts. Inter-species relationships within the genus Oncorhynchus based on biochemical systematics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(1):101–107, January 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-008.

Beamish:1966:VMD

[624] F. W. H. Beamish. Vertical migration by demersal fish in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(1):109–139, January 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-009.

McPhail:1966:CAC

[625] J. D. McPhail. The Coregonus autumnalis complex in Alaska and Northwestern Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(1): 141–148, January 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-010.

Coble:1966:APF

[626] Daniel W. Coble. Alkaline phosphatase in fish scales. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(1):149–152, January 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-011.

High:1966:UOF

[627] William L. High and Larry D. Lusz. Underwater observations on fish in an off-bottom trawl. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23 (1):153–154, January 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-012.

Ackman:1966:FFA

[628] R. G. Ackman, J. E. Stewart, and H. E. Power. Free fatty acid formation in cod livers sterilized by irradiation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(1):155–156, January 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-013.

Elliott:1966:DMT

[629] J. M. Elliott. Downstream movements of trout fry (Salmo trutta) in a Dartmoor stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23 (1):157–159, January 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-014.

Anonymous:1966:RAIa

[630] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(1):161, January 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-015.

Allen:1966:MFG

[631] K. Radway Allen. A method of fitting growth curves of the von Bertalanffy type to observed data. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(2):163–179, February 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-016.

Ellis:1966:SSS

[632] Derek V. Ellis. Swimming speeds of sockeye and Coho salmon on spawning migration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(2): 181–187, February 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-017.

Faber:1966:FSC

[633] Daniel J. Faber. Free-swimming copepod nauplii of Narragansett Bay with a key to their identification. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(2):189–205, February 1966. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f66-018.

Bilinski:1966:LAM

[634] E. Bilinski and R. E. E. Jonas. Lecithinase activity in the muscle of rainbow trout (Salmo gairdnerii). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(2):207–220, February 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-019.

Cucin:1966:DEL

[635] Daniel Cucin and Henry A. Regier. Dynamics and exploitation of lake whitefish in Southern Georgian Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(2):221–274, February 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-020.

Berst:1966:CSL

[636] A. H. Berst and H. R. McCrimmon. Comparative summer limnology of inner long point Bay, Lake Erie, and its major tributary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(2):275–291, February 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-021.

Shelbourn:1966:ITS

[637] J. E. Shelbourn. Influence of temperature, salinity, and photoperiod on the aggregations of chum salmon fry (Oncorhynchus keta). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(2):293–304, February 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-022.

Grinols:1966:NPR

[638] Richard B. Grinols. Northeastern Pacific records of Anoplogaster cornuta Valenciennes (Anoplogasteridae: Pisces) and Cyema atrum Günther (Cyemidae: Pisces). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(2): 305–307, February 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-023.

Westrheim:1966:FRT

[639] S. J. Westrheim and F. T. Pletcher. First records of the twoline eelpout, Bothrocara brunneum, Greenland halibut, Reinhardtius hippoglossoides, and shortbelly rockfish, Sebastodes jordani, in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(2):309–312, February 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-024.

Peden:1966:OFP

[640] Alex E. Peden. Occurrences of the fishes Pholis schultzi and Liparis mucosus in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(2):313–316, February 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-025.

Anonymous:1966:RAIb

[641] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(2):317, February 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-026.

Forrester:1966:EST

[642] C. R. Forrester and D. F. Alderdice. Effects of salinity and temperature on embryonic development of the Pacific cod (Gadus macrocephalus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(3):319–340, March 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-027.

Beamish:1966:SES

[643] F. W. H. Beamish. Swimming endurance of some Northwest Atlantic fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(3):341–347, March 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-028.

Larkin:1966:ETP

[644] P. A. Larkin. Exploitation in a type of predator–prey relationship. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(3):349–356, March 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-029.

Ackman:1966:ODP

[645] R. G. Ackman, C. S. Tocher, and J. McLachlan. Occurrence of dimethylβ-propiothetin in marine phytoplankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(3):357–364, March 1966. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f66-030.

Withler:1966:VLH

[646] I. L. Withler. Variability in life history characteristics of steelhead trout (Salmo gairdneri) along the Pacific Coast of North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(3):365–393, March 1966. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f66-031.

March:1966:PNQ

[647] B. E. March, Jacob Biely, C. Goudie, and H. L. A. Tarr. Protein nutritive quality of commercial fish meals as assessed by biological and chemical methods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23 (3):395–414, March 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-032.

Faber:1966:SOA

[648] Daniel J. Faber. Seasonal occurrence and abundance of free-swimming copepod nauplii in Narragansett Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(3):415–422, March 1966. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f66-033.

Regier:1966:SGN

[649] H. A. Regier and D. S. Robson. Selectivity of gill nets, especially to lake whitefish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(3): 423–454, March 1966. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f66-034.

Peer:1966:RBS

[650] D. L. Peer. Relationship between size and maturity in the spottail shiner, Notropis hudsonius. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23 (3):455–457, March 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-035.

Mathisen:1966:SSC

[651] Ole A. Mathisen. Some scale characters of the ozernaya sockeye salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(3):459–462, March 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-036.

Tibbo:1966:OCM

[652] S. N. Tibbo and R. D. Humphreys. An occurrence of capelin (Mallotus villosus) in the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(3):463–467, March 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-037.

Anonymous:1966:RAIc

[653] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(3):469, March 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-038.

Stevens:1966:EIE

[654] E. Don Stevens and Edgar C. Black. The effect of intermittent exercise on carbohydrate metabolism in rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):471–485, April 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-039.

Ackman:1966:DDP

[655] R. G. Ackman, J. Dale, and J. Hingley. Deposition of dimethyl-β-propiothetin in Atlantic cod during feeding experiments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):487–497, April 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-040.

Hergenrader:1966:DAV

[656] Gary L. Hergenrader and Arthur D. Hasler. Diel activity and vertical distribution of yellow perch (Perca flavescens) under the ice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):499–509, April 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-041.

Allen:1966:ASL

[657] K. Radway Allen and C. R. Forrester. Appropriate size limits for lemon sole (Parophrys vetulus) in the Strait of Georgia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):511–520, April 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-042.

Squires:1966:RSL

[658] H. J. Squires. Reproduction in Sphyrion lumpi, a copepod parasitic on redfish (Sebastes spp.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):521–526, April 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-043.

Burt:1966:GEC

[659] J. R. Burt. Glycogenolytic enzymes of cod (Gadus callarias) muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):527–538, April 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-044.

Parsons:1966:ASB

[660] T. R. Parsons, L. F. Giovando, and R. J. LeBrasseur. The advent of the spring Bloom in the Eastern subarctic Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):539–546, April 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-045.

Hunter:1966:PAS

[661] John R. Hunter. Procedure for analysis of schooling behavior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):547–562, April 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-046.

Buttkus:1966:PPT

[662] H. Buttkus. Preparation and properties of trout myosin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):563–573, April 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-047.

Williams:1966:DSI

[663] P. M. Williams and K. S. Chan. Distribution and speciation of iron in natural waters: Transition from river water to a marine environment, British Columbia, Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):575–593, April 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-048.

Li:1966:VCC

[664] M. F. Li, James E. Stewart, and R. E. Drinnan. In vitro cultivation of cells of the oyster, Crassostrea virginica. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):595–599, April 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-049.

Hashish:1966:PSS

[665] S. Hashish, H. M. Roushdy, M. F. Hussein, M. M. Abdel-Baki, and Y. M. Hassan. Preservation of the sardine Sardinella melanura by low-dose gamma radiation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4): 601–606, April 1966. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f66-050.

Stutz:1966:FPW

[666] Sherwood S. Stutz. Foetal and postpartum whitecoat pelage in Phoca vitulina. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):607– 609, April 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-051.

Harling:1966:DGT

[667] W. R. Harling and S. J. Westrheim. Description of groundfish trawl developed for the side-trawl research vessel G. B. Reed. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):611–613, April 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-052.

Idler:1966:HCF

[668] D. R. Idler and B. Truscott. 1α-hydroxycorticosterone from cartilaginous fish: a new adrenal steroid in blood. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):615–619, April 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-053.

Anonymous:1966:RAId

[669] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(4):621, April 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-054.

Wydoski:1966:MFB

[670] Richard S. Wydoski and Edwin L. Cooper. Maturation and fecundity of brook trout from infertile streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5):623–649, May 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-055.

Pitt:1966:SMS

[671] T. K. Pitt. Sexual maturity and spawning of the American plaice, Hippoglossoides platessoides (Fabricius), from Newfoundland and Grand Bank areas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5):651– 672, May 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-056.

Tomlinson:1966:FDF

[672] N. Tomlinson, S. E. Geiger, and E. Dollinger. Free drip, flesh pH, and chalkiness in halibut. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5): 673–680, May 1966. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f66-057.

Jangaard:1966:PPF

[673] Peter M. Jangaard. Pilot plant fractionation of marine oils methyl esters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5):681–687, May 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-058.

Junge:1966:DPB

[674] Charles O. Junge. Depensatory process based on the concept of hunger. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5):689–699, May 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-059.

Roach:1966:PFM

[675] S. W. Roach, N. Tomlinson, S. E. Geiger, and E. Dollinger. Partial freezing as a means of preserving various species of groundfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5):701–713, May 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-060.

Templeman:1966:RBD

[676] Wilfred Templeman. A record of Bathypterois dubius Vaillant from the Western North Atlantic, and review of status of the species. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5):715–722, May 1966. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f66-061.

Royer:1966:BMD

[677] L. M. Royer. Bioassay method for the determination of toxaphene in lake water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5):723– 727, May 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-062.

Nickerson:1966:CSC

[678] Richard B. Nickerson, Frank J. Ossiander, and Guy C. Powell. Change in size-class structure of populations of Kodiak Island commercial male king crabs due to fishing. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5): 729–736, May 1966. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f66-063.

Castell:1966:CCS

[679] C. H. Castell, Barbara Moore, and Wanda Neal. Cause and control of spurious thiobarbituric acid values from fish muscle in the presence of inorganic iron salts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5): 737–746, May 1966. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f66-064.

Pearson:1966:NCT

[680] Roger E. Pearson. Number of circuli and time of annulus formation on scales of pink salmon (Oncorhynchus gorbuscha). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5):747–756, May 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-065.

Sergeant:1966:RRH

[681] D. E. Sergeant. Reproductive rates of harp seals, Pagophilus groenlandicus (Erxleben). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5): 757–766, May 1966. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f66-066.

McPhail:1966:STO

[682] J. D. McPhail and R. L. Jones. A simple technique for obtaining chromosomes from teleost fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5):767–768, May 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-067.

McLaren:1966:AAC

[683] I. A. McLaren. Analysis of an aerial census of ringed seals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5):769–773, May 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-068.

Cooper:1966:SEM

[684] D. L. Cooper and T. C. Noel. Some equilibrium moisture values of fresh and salt cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5): 775–778, May 1966. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f66-069.

Anonymous:1966:Oa

[685] Anonymous. Obituary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5):779, May 1966. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f66-070.

Anonymous:1966:RAIe

[686] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(5):781, May 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-071.

Black:1966:CLH

[687] Edgar C. Black, Glenville T. Manning, and Koichiro Hayashi. Changes in levels of hemoglobin, oxygen, carbon dioxide, pyruvate, and lactate in venous blood of rainbow trout (Salmo gairdneri) during and following severe muscular activity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):783–795, June 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-072.

Humphreys:1966:BCH

[688] R. D. Humphreys. Biological characteristics of a herring population on the South Coast of Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):797–804, June 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-073.

Park:1966:NSB

[689] Tai Soo Park. A new species of Bradyidius (Copepoda: Calanoida) from the Pacific Coast of North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):805–811, June 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-074.

Betts:1966:NMR

[690] J. L. Betts and G. G. Wilson. New methods for reducing the toxicity of kraft mill bleachery wastes to young salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):813–824, June 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-075.

Fofonoff:1966:VOC

[691] N. P. Fofonoff and S. Tabata. Variability of oceanographic conditions between Ocean Station P and Swiftsure Bank off the Pacific Coast of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):825– 868, June 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-076.

Paloheimo:1966:FGFa

[692] J. E. Paloheimo and L. M. Dickie. Food and growth of fishes. II. Effects of food and temperature on the relation between metabolism and body weight. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6): 869–908, June 1966. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f66-077.

Dehadrai:1966:MGE

[693] Padmakar V. Dehadrai. Mechanism of gaseous exophthalmia in the Atlantic cod, Gadus morhua L. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):909–914, June 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-078.

Kanazawa:1966:VBD

[694] Akio Kanazawa, Akira Saito, and D. R. Idler. Vitamins B in dulse (Rhodymenia palmata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):915–916, June 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-079.

Bilinski:1966:EBO

[695] E. Bilinski, G. Lang, and N. Tomlinson. Evidence for the binding of oil in salmon canned after storage in ice or in a partially frozen condition. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):917–919, June 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-080.

Fraser:1966:NIG

[696] Doris I. Fraser, E. Lo, and W. J. Dyer. Nucleotide interference with glycogen estimation in white and red muscle of cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):921–924, June 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-081.

Tomlinson:1966:IFM

[697] N. Tomlinson, S. E. Geiger, and E. Dollinger. Influence of fishing methods on the incidence of chalkiness in halibut. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):925–928, June 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-082.

Tsuyuki:1966:MSG

[698] H. Tsuyuki, E. Roberts, R. H. Kerr, and A. P. Ronald. Micro starch gel electrophoresis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6): 929–933, June 1966. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f66-083.

Grinols:1966:SOA

[699] Richard B. Grinols. Southern occurrence of Acantholiparis opercularis Gilbert and Burke in the Eastern subarctic Pacific region. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):935–937, June 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-084.

Bilton:1966:CSP

[700] H. T. Bilton. Characteristics of scales from pink salmon (Oncorhynchus gorbuscha) transplanted from the Glendale River in Central British Columbia to the North Harbour River in Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):939–940, June 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-085.

Fedoruk:1966:FRW

[701] Alex N. Fedoruk. Feeding relationship of walleye and smallmouth bass. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):941–943, June 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-086.

Anonymous:1966:RAIf

[702] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(6):945, June 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-087.

Power:1966:JSE

[703] Geoffrey Power and Gilles Shooner. Juvenile salmon in the estuary and Lower Nabisipi River and some results of tagging. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):947–961, July 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-088.

Johnson:1966:TMD

[704] L. Johnson. Temperature of maximum density of fresh water and its effect on circulation in Great Bear Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):963–973, July 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-089.

Kamra:1966:ESR

[705] S. K. Kamra. Effect of starvation and refeeding on some liver and blood constituents of Atlantic cod (Gadus morhua L.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):975–982, July 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-090.

McCrimmon:1966:UOC

[706] Hugh McCrimmon and Wen-Hwa Kwain. Use of overhead cover by rainbow trout exposed to a series of light intensities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):983–990, July 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-091.

Ackman:1966:SCA

[707] R. G. Ackman and C. A. Eaton. Some commercial Atlantic herring oils; fatty acid composition. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):991–1006, July 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-092.

Gibson:1966:RLY

[708] R. John Gibson and Miles H. A. Keenleyside. Responses to light of young Atlantic salmon (Salmo salar) and brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):1007–1024, July 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-093.

Bligh:1966:LCM

[709] E. G. Bligh and Margaret A. Scott. Lipids of cod muscle and the effect of frozen storage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23 (7):1025–1036, July 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-094.

Skidmore:1966:RZS

[710] J. F. Skidmore. Resistance to zinc sulphate of zebrafish (Brachydanio rerio) embryos after removal or rupture of the outer egg membrane. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):1037–1041, July 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-095.

Allen:1966:OMD

[711] George H. Allen. Ocean migration and distribution of fin-marked Coho salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):1043– 1061, July 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-096.

MacCallum:1966:TPA

[712] W. A. MacCallum, E. J. Laishley, W. J. Dyer, and D. R. Idler. Taste panel assessment of cod fillets after single and double freezing. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):1063–1081, July 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-097.

May:1966:DSR

[713] A. W. May and V. M. Hodder. Deck sampling of research vessel catches. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):1083–1088, July 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-098.

Jenkinson:1966:PPM

[714] D. W. Jenkinson and K. Sutherland. A portable press for making impressions of salmon scales. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):1089–1093, July 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-099.

Markert:1966:PBS

[715] J. R. Markert and W. E. Vanstone. Pigments in the belly skin of Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):1095–1098, July 1966. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f66-100.

Pitt:1966:DAF

[716] T. K. Pitt. Development of the anal fin of the capelin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):1099–1100, July 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-101.

Medcof:1966:IRB

[717] J. C. Medcof. Incidental records on behaviour of eels in Lake Ainslie, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):1101– 1105, July 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-102.

Anonymous:1966:RAIg

[718] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(7):1107, July 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-103.

Falkner:1966:SHR

[719] Neil Ward Falkner and A. H. Houston. Some haematological responses to sublethal thermal shock in the goldfish, Carassius auratus L. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1109–1120, August 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-104.

Garside:1966:EOR

[720] E. T. Garside. Effects of oxygen in relation to temperature on the development of embryos of brook trout and rainbow trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1121–1134, August 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-105.

Pearcy:1966:DDO

[721] William G. Pearcy and Carl A. Forss. Depth distribution of oceanic shrimps (Decapoda; Natantia) off Oregon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1135–1143, August 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-106.

Emery:1966:BAA

[722] A. R. Emery and F. D. McCracken. Biology of the Atlantic Argentine (Argentina silus Ascanius) on the Scotian Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1145–1160, August 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-107.

Templeman:1966:DCC

[723] Wilfred Templeman and Richard L. Haedrich. Distributions and comparisons of Centrolophus niger (Gmelin) and Centrolophus britannicus Günther (Centrolophidae) from the North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1161–1185, August 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-108.

Black:1966:ODCb

[724] Edgar C. Black, Harold H. Tucker, and Donald Kirkpatrick. Oxygen dissociation curves of the blood of Atlantic salmon (Salmo salar) acclimated to summer and winter temperatures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1187–1195, August 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-109.

Loring:1966:SFC

[725] D. H. Loring and D. J. G. Nota. Sea-floor conditions around the Magdalen Islands in the Southern Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1197–1207, August 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-110.

Paloheimo:1966:FGFb

[726] J. E. Paloheimo and L. M. Dickie. Food and growth of fishes: III. Relations among food, body size, and growth efficiency. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1209–1248, August 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-111.

Idler:1966:STC

[727] D. R. Idler and H. C. Freeman. Steroid transformations by corpuscles of Stannius of the Atlantic cod (Gadus morhua L.). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1249–1255, August 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-112.

Mercer:1966:FRP

[728] M. C. Mercer. First record of Pyrgopsis lemur (Cephalopoda; Cranchiidae) from the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1257–1258, August 1966. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f66-113.

Northcote:1966:SDC

[729] T. G. Northcote and H. W. Lorz. Seasonal and diel changes in food of adult Kokanee (Oncorhynchus nerka) in Nicola Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1259–1263, August 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-114.

Prakash:1966:RWB

[730] A. Prakash and F. J. R. Taylor. A “red water” Bloom of Gonyaulax acatenella in the Strait of Georgia and its relation to paralytic shellfish toxicity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1265–1270, August 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-115.

Pletcher:1966:PAU

[731] T. F. Pletcher. A portable aquarium for use at sea to photograph fish and aquatic life. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8): 1271–1275, August 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-116.

Peden:1966:RMF

[732] Alex E. Peden. Rare marine fishes from British Columbia with first records of silver perch, Hyperprosopon ellipticum, and shanny, Leptoclinus maculatus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8): 1277–1280, August 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-117.

Boyd:1966:TRB

[733] J. W. Boyd and B. A. Southcott. Thermal resistance of bacteria isolated from a heat processed fish product. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1281–1283, August 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-118.

Nalewajko:1966:DWA

[734] C. Nalewajko. Dry weight, ash, and volume data for some freshwater planktonic algae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8): 1285–1288, August 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-119.

Anonymous:1966:RAIh

[735] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(8):1289, August 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-120.

Lasker:1966:FGR

[736] Reuben Lasker. Feeding, growth, respiration, and carbon utilization of a euphausiid crustacean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1291–1317, September 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-121.

Snow:1966:PMB

[737] C. Dale Snow and John R. Neilsen. Premating and mating behavior of the Dungeness crab (Cancer magister Dana). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1319–1323, September 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-122.

Stewart:1966:IGH

[738] James E. Stewart, John W. Cornick, Doris I. Spears, and D. W. McLeese. Incidence of Gaffkya homari in natural lobster (Homarus americanus) populations of the Atlantic region of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1325–1330, September 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-123.

Schwartz:1966:PSW

[739] Erich Schwartz and Arthur D. Hasler. Perception of surface waves by the blackstripe topminnow, Fundulus notatus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1331–1352, September 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-124.

Parker:1966:CCC

[740] Robert R. Parker and W. E. Vanstone. Changes in chemical composition of Central British Columbia pink salmon during early sea life. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1353–1384, September 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-125.

Castell:1966:ORF

[741] C. H. Castell, B. A. Moore, P. M. Jangaard, and Wanda E. Neal. Oxidative rancidity in frozen stored cod fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1385–1401, September 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-126.

Bilton:1966:TAF

[742] H. T. Bilton and S. A. M. Ludwig. Times of annulus formation on scales of sockeye, pink, and chum salmon in the Gulf of Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1403–1410, September 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-127.

Blackford:1966:STD

[743] B. L. Blackford. A simple two-dimensional electrical analog model for wind-driven circulation in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1411–1438, September 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-128.

Smith:1966:BVT

[744] Lynwood S. Smith. Blood volumes of three salmonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1439–1446, September 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-129.

Alderdice:1966:DSH

[745] D. F. Alderdice, J. R. Brett, and D. B. Sutherland. Design of small holding tank for fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9): 1447–1450, September 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-130.

Cornick:1966:MIH

[746] John W. Cornick and James E. Stewart. Microorganisms isolated from the hemolymph of the lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1451–1454, September 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-131.

Westrheim:1966:NREa

[747] S. J. Westrheim. Northern range extension records for two rockfish species (Sebastodes caurinus and S. elongatus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1455–1456, September 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-132.

Gruchy:1966:LML

[748] C. G. Gruchy and W. B. Scott. Lepomis megalotis, the longear sunfish, in Western Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9): 1457–1459, September 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-133.

Black:1966:CES

[749] Edgar C. Black, N. J. Bosomworth, and Georgina K. Docherty. Combined effect of starvation and severe exercise on glycogen metabolism of rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1461–1463, September 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-134.

Odense:1966:TLC

[750] Paul Odense, Andre Bordeleau, and Raymond Guilbault. Tolerance levels of cod (Gadus morhua) to low salinity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1465–1467, September 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-135.

Westrheim:1966:NREb

[751] S. J. Westrheim. Northern range extensions for three species of rockfish (Sebastodes flavidus, S. paucispinis, and S. pinniger) in the North Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1469–1471, September 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-136.

Anonymous:1966:RAIi

[752] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(9):1473, September 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-137.

Jahn:1966:OWM

[753] Lawrence A. Jahn. Open-water movements of the cutthroat trout (Salmo clarki) in Yellowstone Lake after displacement from spawning streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1475–1485, October 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-138.

Post:1966:RRT

[754] George Post. Response of rainbow trout (Salmo gairdneri) to antigens of Aeromonas hydrophila. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1487–1494, October 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-139.

Johnson:1966:EDF

[755] L. Johnson. Experimental determination of food consumption of pike, Esox lucius, for growth and maintenance. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1495–1505, October 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-140.

Beamish:1966:MFM

[756] F. W. H. Beamish. Muscular fatigue and mortality in haddock, Melanogrammus aeglefinus, caught by otter trawl. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1507–1521, October 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-141.

Johnson:1966:CFR

[757] L. Johnson. Consumption of food by the resident population of pike, Esox lucius, in Lake Windermere. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1523–1535, October 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-142.

Garside:1966:DRV

[758] E. T. Garside. Developmental rate and vertebral number in salmonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1537–1551, October 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-143.

Allen:1966:SME

[759] K. Radway Allen. Some methods for estimating exploited populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1553–1574, October 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-144.

Black:1966:EHU

[760] Edgar C. Black, Harold H. Tucker, and Donald Kirkpatrick. The effect of hemolysis upon the oxygen affinity of hemoglobin in the Atlantic (Salmo salar) and landlocked salmon (Salmo salar sebago). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1575–1580, October 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-145.

Black:1966:ODCc

[761] Edgar C. Black, Donald Kirkpatrick, and Harold H. Tucker. Oxygen dissociation curves of the blood of landlocked salmon (Salmo salar sebago) acclimated to summer and winter temperatures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1581–1586, October 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-146.

Castell:1966:ETA

[762] C. H. Castell and G. A. Boyce. Erroneous thiobarbituric acid values in fish tissues caused by their normal content of free iron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1587–1598, October 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-147.

Tsuyuki:1966:CEC

[763] H. Tsuyuki, J. F. Uthe, E. Roberts, and L. W. Clarke. Comparative electropherograms of Coregonus clupeaformis, Salvelinus namaycush, S. alpinus, S. malma, and S. fontinalis from the family Salmonidae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1599–1606, October 1966.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-148.

Bishop:1966:MDT

[764] C. Bishop and P. H. Odense. Morphology of the digestive tract of the cod, Gadus morhua. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10): 1607–1615, October 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-149.

Neave:1966:ABA

[765] Nancy M. Neave, Carole L. Dilworth, J. G. Eales, and R. L. Saunders. Adjustment of buoyancy in Atlantic salmon parr in relation to changing water velocity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10): 1617–1620, October 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-150.

Kohler:1966:FGR

[766] A. C. Kohler. Feeding and growth of redfish (genus Sebastes) in captivity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1621–1623, October 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-151.

Kerekes:1966:ENP

[767] Joseph Kerekes. Efficiencies of the net plankton and the centrifuged seston methods in highly eutrophic lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1625–1628, October 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-152.

Bligh:1966:BLL

[768] E. G. Bligh and Margaret A. Scott. Blood lipids of the lobster, Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1629– 1631, October 1966. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f66-153.

Anonymous:1966:RAIj

[769] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(10):1633–1634, October 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-154.

Uyeno:1966:NEC

[770] Fukuzo Uyeno. Nutrient and energy cycles in an estuarine oyster area. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1635–1652, November 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-155.

Dabrowski:1966:SCC

[771] Teofil Dabrowski, Edward Kolakowski, Helena Wawreszuk, and Czeslawa Choroszucha. Studies on chemical composition of American crayfish (Orconectes limosus) meat as related to its nutritive value. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1653–1662, November 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-156.

Uthe:1966:CER

[772] J. F. Uthe, E. Roberts, L. W. Clarke, and H. Tsuyuki. Comparative electropherograms of representatives of the families Petromyzontidae, Esocidae, Centrarchidae, and Percidae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1663–1671, November 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-157.

Alton:1966:BDS

[773] Miles S. Alton. Bathymetric distribution of sea stars (Asteroidea) off the Northern Oregon Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1673–1714, November 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-158.

Nalewajko:1966:CPS

[774] C. Nalewajko. Composition of phytoplankton in surface waters of Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1715–1725, November 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-159.

Ivlev:1966:BPW

[775] V. S. Ivlev. The biological productivity of waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1727–1759, November 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-160.

Geen:1966:LHT

[776] G. H. Geen, T. G. Northcote, G. F. Hartman, and C. C. Lindsey. Life histories of two species of catostomid fishes in Sixteenmile Lake, British Columbia, with particular reference to inlet stream spawning. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1761–1788, November 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-161.

Anonymous:1966:Ob

[777] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1789–1793, November 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-162.

Murray:1966:HMC

[778] J. Murray, N. R. Jones, and J. R. Burt. Hypoxanthine in the muscle of chill-stored Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1795–1797, November 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-163.

Smith:1966:AOM

[779] M. W. Smith. Amount of organic matter lost to a lake by migration of eels. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1799–1801, November 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-164.

Saunders:1966:ESA

[780] J. W. Saunders. Estuarine spawning of Atlantic salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1803–1804, November 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-165.

Carrothers:1966:NSC

[781] P. J. G. Carrothers. New speed calibration curves for the Braincon Type 316 histogram current meter. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1805–1806, November 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-166.

Lawler:1966:PPP

[782] G. H. Lawler. Pugheadedness in perch, Perca flavescens, and pike, Esox lucius, of Heming Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1807–1808, November 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-167.

Jangaard:1966:UFA

[783] P. M. Jangaard, R. G. Ackman, and J. C. Sipos. Unusual fatty acid composition of lipids from a cod, Gadus morhua. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1809–1810, November 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-168.

Bilinski:1966:DLS

[784] E. Bilinski and R. E. E. Jonas. Distribution of lecithinase in the subcellular fractions of rainbow trout (Salmo gairdnerii) lateral line muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1811–1813, November 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-169.

Driver:1966:MNY

[785] E. A. Driver and E. T. Garside. Meristic numbers of yellow perch in saline lakes in Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11): 1815–1817, November 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-170.

Anonymous:1966:RAIk

[786] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(11):1819–1820, November 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-171.

Dyer:1966:NDQ

[787] W. J. Dyer, Doris I. Fraser, and Dianne P. Lohnes. Nucleotide degradation and quality in ordinary and red muscle of iced and frozen swordfish (Xiphias gladius). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12): 1821–1833, December 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-172.

Brockerhoff:1966:DFC

[788] H. Brockerhoff. Digestion of fat by cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12):1835–1839, December 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-173.

Odense:1966:UEB

[789] P. H. Odense and C. M. Bishop. The ultrastructure of the epithelial border of the ileum, pyloric caeca, and rectum of the cod, Gadus morhua. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12):1841–1843, December 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-174.

Keast:1966:MBF

[790] Allen Keast and Deirdre Webb. Mouth and body form relative to feeding ecology in the fish fauna of a small lake, Lake Opinicon, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12):1845–1874, December 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-175.

Sparrow:1966:CLL

[791] R. A. H. Sparrow. Comparative limnology of lakes in the Southern Rocky Mountain Trench, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12):1875–1895, December 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-176.

Mitchell:1966:NPP

[792] Edward Mitchell. Northeastern Pacific Pleistocene sea otters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12):1897–1911, December 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-177.

Pauley:1966:AIR

[793] Gilbert B. Pauley and Albert K. Sparks. The acute inflammatory reaction in two different tissues of the Pacific oyster, Crassostrea gigas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12):1913–1921, December 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-178.

Gerking:1966:AGC

[794] Shelby D. Gerking. Annual growth cycle, growth potential, and growth compensation in the bluegill sunfish in Northern Indiana lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12):1923–1956, December 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-179.

Post:1966:SPA

[795] George Post. Serum proteins and antibody production in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12): 1957–1963, December 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-180.

Katz:1966:RCS

[796] Max Katz, James C. Woodey, Clarence D. Becker, P. T. K. Woo, and J. R. Adams. Records of Cryptobia salmositica from sockeye salmon from the Fraser River drainage and from the State of Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12):1965–1966, December 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f66-181.

Harling:1966:NRE

[797] W. R. Harling. Northern range extension record for the pygmy rockfish (Sebastodes wilsoni). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12): 1967–1968, December 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-182.

Lawler:1966:SRA

[798] G. H. Lawler. A second record of an accessory fin on Esox lucius. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12):1969, December 1966. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f66-183.

Renfro:1966:FFA

[799] William C. Renfro and William G. Pearcy. Food and feeding apparatus of two pelagic shrimps. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12): 1971–1975, December 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-184.

Gibson:1966:SFI

[800] R. J. Gibson. Some factors influencing the distributions of brook trout and young Atlantic salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12):1977–1980, December 1966. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f66-185.

Davenport:1966:CVB

[801] D. Davenport. Colour variant of bocaccio (Sebastodes paucispinis) in British Columbia waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12):1981, December 1966. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f66-186.

Anonymous:1966:RAIl

[802] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12):1983–1984, December 1966. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f66-187.

Anonymous:1966:IP

[803] Anonymous. Index to publications for 1966. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 23(12):1985–2039, December 1966. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f66-188.

Peden:1967:RNP

[804] Alex E. Peden. Redescription of a North Pacific prickleback, Alectridium auranticum. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1): 1–8, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f67-001.

Ostovar:1967:PFW

[805] K. Ostovar, M. Slusar, and M. Vaisey. Preservation of fresh whitefish with gamma radiation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1): 9–19, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f67-002.

Saunders:1967:SPR

[806] Richard L. Saunders. Seasonal pattern of return of Atlantic salmon in the Northwest Miramichi River, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):21–32, January 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-003.

Druehl:1967:VDS

[807] Louis D. Druehl. Vertical distributions of some benthic marine algae in a British Columbia inlet, as related to some environmental factors. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):33–46, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-004.

McCrimmon:1967:SFR

[808] Hugh R. McCrimmon and U. Boon Swee. Scale formation as related to growth and development of young carp, Cyprinus carpio L. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):47–51, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-005.

Kohler:1967:SMS

[809] A. C. Kohler. Size at maturity, spawning season, and food of Atlantic halibut. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):53–66, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-006.

McBride:1967:EFT

[810] J. R. McBride. Effects of feeding on the thyroid, kidney, and pancreas in sexually ripening adult sockeye salmon (Oncorhyncus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):67–76, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-007.

Coburn:1967:HRT

[811] A. Coburn and P. McCart. A hatchery release tank for pink salmon fry with notes on behavior of the fry in the tank and after release. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):77–85, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-008.

Coble:1967:RTT

[812] Daniel W. Coble. Relationship of temperature to total annual growth in adult smallmouth bass. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):87–99, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-009.

McCombie:1967:TRS

[813] A. M. McCombie. The thermal regime of South Bay, manitoulin Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):101–125, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-010.

MacCallum:1967:EWD

[814] W. A. MacCallum, Dorothy A. Chalker, W. J. Dyer, and D. R. Idler. Effects of water and dielectric thawing processes on shelf life of doublefrozen cod and redfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):127–144, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-011.

Garrod:1967:PDA

[815] D. J. Garrod. Population dynamics of the arcto-Norwegian cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):145–190, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-012.

Flux:1967:FAR

[816] John E. C. Flux. Factors affecting the response of trout to an electric field in fresh and salt water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):191–199, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-013.

Antia:1967:BTT

[817] N. J. Antia and E. Bilinski. A bacterial-toxin type of phospholipase (lecithinase C) in a marine phytoplanktonic chrysomonad. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):201–204, January 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-014.

Idler:1967:PCB

[818] D. R. Idler, H. C. Freeman, and B. Truscott. A preliminary communication on the biological activity of 1α-hydroxycorticosterone isolated from cartilaginous fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24 (1):205–206, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-015.

Powles:1967:AWA

[819] P. M. Powles. Atlantic wolffish (Anarhichas lupus L.) eggs off southern Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1): 207–208, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-016.

Dahlberg:1967:UAF

[820] Michael L. Dahlberg and Duane E. Phinney. The use of adipose fin pigmentation for distinguishing between juvenile Chinook and Coho salmon in Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):209–210, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-017.

Boyd:1967:DFF

[821] J. W. Boyd, B. A. Southcott, and G. F. Boothby. Desiccation of frozen fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):211–212, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-018.

McKenzie:1967:FCE

[822] R. A. McKenzie. Fawn cusk-eel, Lepophidium cervinum, in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):213–214, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-019.

Templeman:1967:PLF

[823] Wilfred Templeman. Predation on living fishes on longline in Baffin Bay by the amphipod Eurythenes gryllus (Lichtenstein), and a new distribution record. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1): 215–217, January 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-020.

Anonymous:1967:RAIa

[824] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(1):219, January 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-021.

Power:1967:EPD

[825] H. E. Power, D. I. Fraser, W. J. Dyer, W. Neal, and C. H. Castell. Effect of pasteurizing doses of gamma radiation on storage life of cooked lobster meat. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2): 221–230, February 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-022.

Park:1967:TNS

[826] Tai Soo Park. Two new species of calanoid copepods from the Strait of Georgia, British Columbia, Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):231–242, February 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-023.

Tomlinson:1967:SPW

[827] N. Tomlinson, J. R. McBride, and S. E. Geiger. The sodium, potassium, and water content of the flesh of sockeye salmon (Oncorhynchus nerka) in relation to sexual development and starvation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):243–248, February 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-024.

Paulik:1967:GCP

[828] G. J. Paulik and William H. Bayliff. A generalized computer program for the Ricker model of equilibrium yield per recruitment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):249–259, February 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-025.

Walden:1967:EMB

[829] C. C. Walden, I. V. F. Allen, and P. C. Trussell. Estimation of marine-borer attack on wooden surfaces. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):261–272, February 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-026.

Jonas:1967:GRC

[830] R. E. E. Jonas and E. Bilinski. Glycerylphosphorylcholine and related compounds in rainbow trout muscle stored at −4 C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):273–280, February 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-027.

Pycha:1967:RHR

[831] Richard L. Pycha and George R. King. Returns of hatchery-reared lake trout in Southern Lake Superior, 1955–62. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):281–298, February 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-028.

Tsuyuki:1967:CEF

[832] H. Tsuyuki, E. Roberts, R. H. Kerr, J. F. Uthe, and L. W. Clarke. Comparative electropherograms of the family Catostomidae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):299–304, February 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-029.

Mills:1967:BAA

[833] Eric L. Mills. The biology of an ampeliscid amphipod crustacean sibling species pair. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2): 305–355, February 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-030.

Ackman:1967:TSA

[834] R. G. Ackman and M. G. Cormier. α-tocopherol in some Atlantic fish and shellfish with particular reference to live-holding without food. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):357–373, February 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-031.

McCart:1967:BES

[835] P. McCart. Behaviour and ecology of sockeye salmon fry in the Babine River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):375–428, February 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-032.

Brungs:1967:LEC

[836] W. A. Brungs and D. I. Mount. Lethal endrin concentration in the blood of gizzard shad. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):

429–432, February 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-033.

McCart:1967:SRB

[837] P. McCart. Scale regeneration in the bluespot goby, Coryphopterus nicholsi. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):433–434, February 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-034.

Stutz:1967:MPH

[838] Sherwood S. Stutz. Moult in the Pacific harbour seal Phoca vitulina richardi. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):435–441, February 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-035.

Li:1967:SDB

[839] M. F. Li, Carol Flemming, and James E. Stewart. Serological differences between two populations of oysters (Crassostrea virginica) from the Atlantic Coast of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):443–446, February 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-036.

Henderson:1967:UGS

[840] Nancy E. Henderson. The urinary and genital systems of trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):447–449, February 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-037.

Stutz:1967:PPP

[841] Sherwood S. Stutz. Pelage patterns and population distributions in the Pacific harbour seal (Phoca vitulina richardi). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):451–455, February 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-038.

Ackman:1967:DPD

[842] R. G. Ackman, J. Hingley, and A. W. May. Dimethyl-β-propiothetin dimethyl sulphide in Labrador cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):457–461, February 1967. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f67-039.

Stringer:1967:IMR

[843] G. E. Stringer. Introduction of Mysis relicta Lovén into kalamalka and pinaus lakes, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):463–465, February 1967. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f67-040.

Ackman:1967:FAC

[844] R. G. Ackman and C. A. Eaton. Fatty acid composition of the decapod shrimp, Pandalus borealis, in relation to that of the euphausid, Meganyctiphanes norvegica. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2): 467–471, February 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-041.

Anonymous:1967:RAIb

[845] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(2):473, February 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-042.

Stewart:1967:IOC

[846] Nelson E. Stewart, Dean L. Shumway, and Peter Doudoroff. Influence of oxygen concentration on the growth of juvenile largemouth bass. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):475–494, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-043.

Clark:1967:RIB

[847] Francis W. Clark and Miles H. A. Keenleyside. Reproductive isolation between the sunfish Lepomis gibbosus and L. macrochirus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):495–514, March 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-044.

Kabata:1967:MPC

[848] Z. Kabata. Morphology of Phrixocephalus cincinnatus Wilson, 1908 (Copepoda: Lernaeoceridae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):515–526, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-045.

Boyd:1967:QPC

[849] J. W. Boyd, B. A. Southcott, and P. J. Schmidt. Quality of Pacific Coast dogfish during frozen storage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):527–535, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-046.

McAllister:1967:SVB

[850] D. E. McAllister. The significance of ventral bioluminescenee in fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):537–554, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-047.

Cairns:1967:ZTF

[851] Alan A. Cairns. The zooplankton of Tanquary Fjord, Ellesmere Island, with special reference to calanoid copepods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):555–568, March 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-048.

Warren:1967:FMP

[852] P. J. Warren and R. W. Sheldon. Feeding and migration patterns of the pink shrimp, Pandalus montagui, in the Estuary of the River Crouch, Essex, England. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24 (3):569–580, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-049.

Wickett:1967:ETZ

[853] W. Percy Wickett. Ekman transport and zooplankton concentration in the North Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):581–594, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-050.

Strasdine:1967:RGC

[854] G. A. Strasdine. Rapid germination of Clostridium botulinum type E spores. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):595–606, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-051.

Jangaard:1967:SCG

[855] P. M. Jangaard, H. Brockerhoff, R. D. Burgher, and R. J. Hoyle. Seasonal changes in general condition and lipid content of cod from inshore waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):607–612, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-052.

Jangaard:1967:SCF

[856] P. M. Jangaard, R. G. Ackman, and J. C. Sipos. Seasonal changes in fatty acid composition of cod liver, flesh, Roe, and milt lipids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):613–627, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-053.

Kilambi:1967:HSS

[857] Raj V. Kilambi and Allan C. DeLacy. Heterogeneity of surf smelt, Hypomesus pretiosus (Girard), in the State of Washington, as judged by incidence of larval Anisakis (Nematoda). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):629–633, March 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-054.

Ackman:1967:SLI

[858] R. G. Ackman, J. C. Sipos, and C. S. Tocher. Some linear, Iso, Anteiso, and multiple-branched fatty acids of marine origin: Analyses on packed and open-tubular gas–liquid chromatographic columns with particular reference to pristanic and phytanic acids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):635–650, March 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-055.

MacCallum:1967:PPC

[859] W. A. MacCallum, June I. Jaffray, D. N. Churchill, D. R. Idler, and P. H. Odense. Postmortem physicochemical changes in unfrozen Newfoundland trap-caught cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24 (3):651–678, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-056.

Pauley:1967:TGF

[860] Gilbert B. Pauley. A tumorlike growth on the foot of a freshwater mussel (Anodonta californiensis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):679–682, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-057.

Lewis:1967:FDC

[861] John B. Lewis and Fritz Axelsen. Food of the dolphin, Coryphaena hippurus Linnaeus, and of the yellowfin tuna, Thunnus albacares (Lowe), from Barbados, West Indies. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):683–686, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-058.

Bowman:1967:BTS

[862] Thomas E. Bowman. Bioluminescence in two species of pelagic amphipods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):687–688, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-059.

McBride:1967:SPE

[863] Roland McBride and Louis Barr. A shrimp pot for experimental fishing. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):689–690, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-060.

Forrester:1967:SFM

[864] C. R. Forrester. The shrinkage factor in measurements of groundfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):691–693, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-061.

Bouck:1967:DLM

[865] G. R. Bouck and R. C. Ball. Distribution of low mobility proteins in the blood of fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24 (3):695–697, March 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-062.

Anonymous:1967:RAIc

[866] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(3):699, March 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-063.

Kerswill:1967:SEF

[867] C. J. Kerswill. Studies on effects of forest sprayings with insecticides, 1952–63, on fish and aquatic invertebrates in New Brunswick streams: Introduction and summary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):701–708, April 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-064.

Kerswill:1967:FLA

[868] C. J. Kerswill and H. E. Edwards. Fish losses after forest sprayings with insecticides in New Brunswick, 1952–62, as shown by caged specimens and other observations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):709–729, April 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-065.

Elson:1967:EWY

[869] P. F. Elson. Effects on wild young salmon of spraying DDT over New Brunswick forests. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4): 731–767, April 1967. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f67-066.

Ide:1967:EFS

[870] F. P. Ide. Effects of forest spraying with DDT on aquatic insects of salmon streams in New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):769–805, April 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-067.

Keenleyside:1967:EFS

[871] Miles H. A. Keenleyside. Effects of forest spraying with DDT in New Brunswick on food of young Atlantic salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):807–822, April 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-068.

Grant:1967:EAI

[872] C. D. Grant. Effects on aquatic insects of forest spraying with phosphamidon in New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):823–832, April 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-069.

Stendell:1967:ECD

[873] Rey Stendell. Echinoderms collected from a drifting ice Island off the East Greenland Coast, with comments on their distribution in adjacent waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):833– 842, April 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-070.

Pauley:1967:PAS

[874] Gilbert B. Pauley. Prespawning adult salmon mortality associated with a fungus of the genus Dermocystidium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):843–848, April 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-071.

Weber:1967:MPS

[875] Douglas Weber and George J. Ridgway. Marking Pacific salmon with tetracycline antibiotics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):849–865, April 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-072.

Pauley:1967:HGB

[876] Gilbert B. Pauley and Roy E. Nakatani. Histopathology of “gas-bubble” disease in salmon fingerlings. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):867–871, April 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-073.

Matsumoto:1967:PEP

[877] J. J. Matsumoto, W. J. Dyer, J. R. Dingle, and D. G. Ellis. Protein in extracts of prerigor and postrigor scallop striated muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):873–882, April 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-074.

Quayle:1967:PSO

[878] D. B. Quayle and L. D. B. Terhune. A plankton sampler for oyster larvae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):883–885, April 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-075.

Bidgood:1967:PCR

[879] B. F. Bidgood and A. H. Berst. Phenotypic characteristics of rainbow trout in the Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):887–892, April 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-076.

Margolis:1967:TCP

[880] L. Margolis. Triaenophorus crassus plerocereoids in sockeye salmon smolts from the Kvichak River System, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):893–894, April 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-077.

Guttmann:1967:PRM

[881] A. Guttmann, N. Dambergs, and P. M. Jangaard. Preservation of raw material for fish protein concentrate with isopropyl alcohol. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):895–896, April 1967. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f67-078.

Anonymous:1967:RAId

[882] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(4):897, April 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-079.

Park:1967:CFS

[883] Kilho Park. Chemical features of the subarctic boundary near 170^{o}W. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):899–908, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-080.

Sheldon:1967:CSS

[884] R. W. Sheldon and T. R. Parsons. A continuous size spectrum for particulate matter in the sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):909–915, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-081.

Bulkley:1967:FST

[885] Ross V. Bulkley. Fecundity of steelhead trout, Salmo gairdneri, from Alsea River, Oregon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5): 917–926, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f67-082.

Faber:1967:LLF

[886] Daniel J. Faber. Limnetic larval fish in Northern Wisconsin lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):927–937, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-083.

Black:1967:EPL

[887] Edgar C. Black and S. J. Tredwell. Effect of a partial loss of scales and mucous on carbohydrate metabolism in rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):939–953, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-084.

Johnston:1967:PIA

[888] C. E. Johnston and J. G. Eales. Purines in the integument of the Atlantic salmon (Salmo salar) during parr–smolt transformation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):955–964, May 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-085.

Martin:1967:PCG

[889] N. V. Martin and F. K. Sandercock. Pyloric caeca and gill raker development in lake trout, Salvelinus namaycush, in Algonquin Park, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):965–974, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-086.

McLaren:1967:IBS

[890] Ian A. McLaren. Introduction to biological studies of Ogac Lake, a landlocked fiord on Baffin Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):975–980, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-087.

McLaren:1967:PCC

[891] Ian A. McLaren. Physical and chemical characteristics of Ogac Lake, a landlocked fiord on Baffin Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):981–1015, May 1967. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f67-088.

Slack:1967:BSP

[892] H. D. Slack. A brief survey of the profundal benthic fauna of lakes in Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5): 1017–1033, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f67-089.

Parsons:1967:CYC

[893] John W. Parsons. Contributions of year-classes of blue pike to the commercial fishery of Lake Erie, 1943–59. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):1035–1066, May 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-090.

Vladykov:1967:TLP

[894] Vadim D. Vladykov and W. I. Follett. The teeth of lampreys (Petromyzonidae): Their terminology and use in a key to the Holarctic genera. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):1067–1075, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-091.

Pitt:1967:AGA

[895] T. K. Pitt. Age and growth of American plaice (Hippoglossoides platessoides) in the Newfoundland area of the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):1077–1099, May 1967. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f67-092.

Lewis:1967:FAC

[896] R. W. Lewis. Fatty acid composition of some marine animals from various depths. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):1101– 1115, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-093.

Bams:1967:DPN

[897] R. A. Bams. Differences in performance of naturally and artificially propagated sockeye salmon migrant fry, as measured with swimming and predation tests. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5): 1117–1153, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f67-094.

Hope:1967:SDB

[898] E. L. V. Hope and C. D. Sauer. A system to digitize bathythermograph aperture cards. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5): 1155–1164, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f67-095.

Sheldon:1967:RBS

[899] R. W. Sheldon. Relationship between shell-weight and age in certain molluscs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):1165– 1171, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-096.

Velsen:1967:TTI

[900] F. P. J. Velsen and D. F. Alderdice. Toxicities of two insecticides to young Coho salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24 (5):1173–1175, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-097.

Beckett:1967:GNT

[901] J. S. Beckett and B. E. Barrett. The gempylid Nealotus tripes new to the Canadian Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24 (5):1177–1178, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-098.

Dussault:1967:BEA

[902] H. P. Dussault. Bactericidal effect of acetylated monoglycerides as protective coating on frozen fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):1179–1183, May 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-099.

Anonymous:1967:RAIe

[903] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(5):1185, May 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-100.

Westrheim:1967:SRT

[904] S. J. Westrheim. Sampling research trawl catches at sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1187–1202, June 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-101.

Bilinski:1967:BOC

[905] E. Bilinski and P. Clement. Binding of oil in canned salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1203–1207, June 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-102.

Squires:1967:GHA

[906] H. J. Squires. Growth and hypothetical age of the Newfoundland bait squid Illex illecebrosus illecebrosus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1209–1217, June 1967. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f67-103.

Ackman:1967:FFO

[907] R. G. Ackman, C. A. Eaton, E. G. Bligh, and A. W. Lantz. Freshwater fish oils: Yields and composition of oils from reduction of sheepshead, tullibee, Maria, and alewife. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1219–1227, June 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-104.

Paranjape:1967:MRE

[908] Madhu A. Paranjape. Molting and respiration of euphausiids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1229–1240, June 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-105.

MacLeod:1967:NAM

[909] J. C. MacLeod. A new apparatus for measuring maximum swimming speeds of small fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24 (6):1241–1252, June 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-106.

Skidmore:1967:OUZ

[910] J. F. Skidmore. Oxygen uptake by zebrafish (Brachydanio rerio) of different ages in relation to zinc sulphate resistance. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1253–1267, June 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-107.

Uthe:1967:CZE

[911] J. F. Uthe and H. Tsuyuki. Comparative zone electropherograms of muscle myogens and blood proteins of adult and ammocoete lamprey. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1269–1273, June 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-108.

Templeman:1967:ARS

[912] Wilfred Templeman. Adult redfish, Sebastes mentella, pelagic over oceanic depths in the Labrador Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1275–1290, June 1967. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f67-109.

March:1967:CNV

[913] B. E. March, J. Biely, E. G. Bligh, and A. W. Lantz. Composition and nutritive value of meals from alewife, sheepshead, Maria, and tullibee. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1291–1298, June 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-110.

Brinkhurst:1967:RME

[914] Ralph O. Brinkhurst and Brenda Walsh. Rostherne Mere, England, a further instance of guanotrophy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1299–1313, June 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-111.

Mujib:1967:COG

[915] Khwaja Abdul Mujib. The cranial osteology of the Gadidae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1315–1375, June 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-112.

McCauley:1967:SHU

[916] James E. McCauley. Status of the heart urchin, Brisaster latifrons. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1377–1384, June 1967. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f67-113.

McCauley:1967:EO

[917] James E. McCauley and Andrew G. Carey, Jr. Echinoidea of Oregon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1385–1401, June 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-114.

Scarratt:1967:ALS

[918] D. J. Scarratt and G. E. Raine. Avoidance of low salinity by newly hatched lobster larvae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1403–1406, June 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-115.

Qadri:1967:MCT

[919] S. U. Qadri. Morphological comparisons of three populations of the lake char, Cristivomer namaycush, from Ontario and Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1407–1411, June 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-116.

Shaw:1967:FDD

[920] William K. Shaw, Haskell S. Tubiash, and Allan M. Barker. Freezedrying for determining total solids in shellfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1413–1417, June 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-117.

Cocking:1967:UIM

[921] A. W. Cocking. Use of indophenol method for estimation of ammonia excreted by freshwater fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1419–1420, June 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-118.

Wickett:1967:ESE

[922] W. Percy Wickett. Exotic species of earthworms in British Columbia eaten by Coho salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1421–1422, June 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-119.

Anonymous:1967:RAIf

[923] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(6):1423, June 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-120.

McFadden:1967:NCP

[924] James T. McFadden, Gaylord R. Alexander, and David S. Shetter. Numerical changes and population regulation in brook trout Salvelinus fontinalis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1425– 1459, July 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-121.

Legendre:1967:FDA

[925] R. Legendre and A. L. Wood. Freeze-drying of Atlantic cod steaks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1461–1473, July 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-122.

Pickard:1967:SOC

[926] G. L. Pickard. Some oceanographic characteristics of the larger inlets of Southeast Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24 (7):1475–1506, July 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-123.

Javaid:1967:TAT

[927] M. Yaqub Javaid and John M. Anderson. Thermal acclimation and temperature selection in Atlantic salmon, Salmo salar, and rainbow trout, S. gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1507– 1513, July 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-124.

Javaid:1967:ISS

[928] M. Yaqub Javaid and John M. Anderson. Influence of starvation on selected temperature of some salmonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1515–1519, July 1967. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f67-125.

Ackman:1967:DIA

[929] R. G. Ackman, H. J. Hingley, and H. E. Power. Determination of isopropyl alcohol in fish protein concentrate by solvent extraction and gas–liquid chromatography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1521–1529, July 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-126.

May:1967:FAC

[930] A. W. May. Fecundity of Atlantic cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1531–1551, July 1967. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f67-127.

Haefner:1967:HPR

[931] Paul A. Haefner, Jr. Hydrography of the Penobscot River (Maine) Estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1553– 1571, July 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-128.

Withler:1967:DFO

[932] F. C. Withler and R. M. Humphreys. Duration of fertility of ova and sperm of sockeye (Oncorhynchus nerka) and pink (O. gorbuscha) salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1573–1578, July 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-129.

Smith:1967:AST

[933] L. S. Smith and G. R. Bell. Anesthetic and surgical techniques for Pacific salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1579– 1588, July 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-130.

Prakash:1967:GTM

[934] A. Prakash. Growth and toxicity of a marine dinoflagellate, Gonyaulax tamarensis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1589– 1606, July 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-131.

Buttkus:1967:AAC

[935] H. Buttkus. Amino acid composition of myosin from trout muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1607–1612, July 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-132.

Bolton:1967:USC

[936] R. S. Bolton, J. H. Mann, and W. Gushue. Use of standardized colour surfaces in the grading of canned salmon for colour. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1613–1622, July 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-133.

Mercer:1967:ARC

[937] M. C. Mercer. Anomalous record of a California sea lion, Zalophus californianus, from Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1623–1626, July 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-134.

Pippy:1967:KBT

[938] J. H. C. Pippy and I. M. Sandeman. A kill of brook trout (Salvelinus fontinalis) involving the Acanthocephalan Echinorhynchus lateralis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1627–1628, July 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-135.

McNeil:1967:RDP

[939] William J. McNeil. Randomness in distribution of pink salmon redds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1629–1634, July 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-136.

Anonymous:1967:RAIg

[940] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(7):1635, July 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-137.

Hagen:1967:IMT

[941] D. W. Hagen. Isolating mechanisms in threespine sticklebacks (Gasterosteus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1637–1692, August 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-138.

Legendre:1967:PLS

[942] R. Legendre. Preparation of light-salted fish by brining. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1693–1699, August 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-139.

Nakano:1967:CCC

[943] T. Nakano and N. Tomlinson. Catecholamine and carbohydrate concentrations in rainbow trout (Salmo gairdneri) in relation to physical disturbance. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1701–1715, August 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-140.

Dingle:1967:ESP

[944] J. R. Dingle and J. A. Hines. Extraction and some properties of adenosine 5’-monophosphate aminohydrolase from prerigor and postrigor muscle of cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1717–1730, August 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-141.

Brett:1967:SPS

[945] J. R. Brett. Swimming performance of sockeye salmon (Oncorhynchus nerka) in relation to fatigue time and temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1731–1741, August 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-142.

Smith:1967:MBT

[946] M. W. Smith and J. W. Saunders. Movements of brook trout in relation to an artificial pond on a small stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1743–1761, August 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-143.

Smith:1967:AMS

[947] John W. Smith. Aporocotyle margolisi n. sp. (Digenea: Aporocotylidae) from Merluccius productus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1763–1773, August 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-144.

Smith:1967:CDS

[948] L. S. Smith, J. R. Brett, and J. C. Davis. Cardiovascular dynamics in swimming adult sockeye salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1775–1790, August 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-145.

vanOverbeeke:1967:PGS

[949] A. P. van Overbeeke and J. R. McBride. The pituitary gland of the sockeye (Oncorhynchus nerka) during sexual maturation and spawning. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1791–1810, August 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-146.

Beers:1967:PSS

[950] John R. Beers, Gene L. Stewart, and John D. H. Strickland. A pumping system for sampling small plankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1811–1818, August 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-147.

Agris:1967:CEL

[951] Paul F. Agris. Comparative effects of a lampricide and of anoxia on the sea lamprey. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8): 1819–1822, August 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-148.

Butler:1967:EOP

[952] D. G. Butler. Effect of ovine prolactin on tissue electrolyte composition of hypophysectomized freshwater eels (Anguilla rostrata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1823–1826, August 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-149.

Mann:1967:EAS

[953] C. R. Mann and G. T. Needler. Effect of aliasing on studies of longterm oceanic variability off Canada’s coasts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1827–1831, August 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-150.

Strasdine:1967:GSC

[954] G. A. Strasdine and Joanne M. Kelly. Germination of spores of Clostridium botulinum type E in fish and shellfish extracts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1833–1835, August 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-151.

Fraser:1967:NDM

[955] Doris I. Fraser, J. R. Dingle, J. A. Hines, Sandra C. Nowlan, and W. J. Dyer. Nucleotide degradation, monitored by thin-layer chromatography, and associated postmortem changes in relaxed cod muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1837–1841, August 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-152.

Anonymous:1967:RAIh

[956] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(8):1843, August 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-153.

Lauzier:1967:BRD

[957] L. M. Lauzier. Bottom residual drift on the Continental Shelf area of the Canadian Atlantic Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9):1845–1859, September 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-154.

Kalff:1967:PDA

[958] Jacob Kalff. Phytoplankton dynamics in an Arctic lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9):1861–1871, September 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-155.

Squires:1967:DCC

[959] H. J. Squires. Decapod crustacea from Calanus collections in Hudson Bay in 1953, 1954, and 1958–61. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9):1873–1903, September 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-156.

Powell:1967:SDC

[960] N. A. Powell. Sexual dwarfism in Cribrilina annulata (Cribrilinidae-Bryozoa). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9):1905–1910, September 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-157.

Sandeman:1967:PFF

[961] I. M. Sandeman and J. H. C. Pippy. Parasites of freshwater fishes (Salmonidae and Coregonidae) of insular Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9):1911–1943, September 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-158.

Westrheim:1967:SRN

[962] S. J. Westrheim and H. Tsuyuki. Sebastodes reedi, a new scorpaenid fish in the Northeast Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9):1945–1954, September 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-159.

Hanson:1967:MSP

[963] Arthur J. Hanson and Howard D. Smith. Mate selection in a population of sockeye salmon (Oncorhynchus nerka) of mixed age-groups. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9):1955–1977, September 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-160.

Anonymous:1967:O

[964] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9):1979–1984, September 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-161.

McCormick:1967:HEM

[965] J. Michael McCormick, R. Michael Laurs, and James E. McCauley. A hydroid epizoic on myctophid fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9):1985–1989, September 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-162.

Mather:1967:TMY

[966] Frank J. Mather III, Martin R. Bartlett, and James S. Beckett. Transatlantic migrations of young bluefin tuna. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9):1991–1997, September 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-163.

McCart:1967:PGN

[967] P. McCart and B. Andersen. Plasticity of gillraker number and length in Oncorhynchus nerka. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9):

1999–2002, September 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-164.

Powell:1967:FQN

[968] N. A. Powell. Figularia quaylei: A new bryozoan from the Northeast Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9):2003–2005, September 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-165.

McArn:1967:NTF

[969] Glynn McArn and S. R. Wellings. Neurofibroma in a teleost fish, Lumpenus sagitta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9): 2007–2008, September 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-166.

Anonymous:1967:RAIi

[970] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(9):2009, September 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-167.

McCleave:1967:HOC

[971] James D. McCleave. Homing and orientation of cutthroat trout (Salmo clarki) in Yellowstone Lake, with special reference to olfaction and vision. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10):2011–2044, October 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-168.

Narver:1967:PPB

[972] David W. Narver. Primary productivity in the Babine Lake System, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10): 2045–2052, October 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-169.

Beers:1967:MZE

[973] J. R. Beers and G. L. Stewart. Micro- zooplankton in the euphotic zone at five locations across the California Current. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10):2053–2068, October 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-170.

Hartman:1967:MBS

[974] W. L. Hartman, W. R. Heard, and B. Drucker. Migratory behavior of sockeye salmon fry and smolts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10):2069–2099, October 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-171.

Taylor:1967:UFT

[975] F. H. C. Taylor. Unusual fishes taken by midwater trawl off the Queen Charlotte Islands, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10):2101–2115, October 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-172.

Coble:1967:WSP

[976] Daniel W. Coble. The white sucker population of South Bay, Lake Huron, and effects of the sea lamprey on it. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10):2117–2136, October 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-173.

Werner:1967:DKM

[977] A. E. Werner and W. F. Hyslop. Distributions of kraft mill effluent in a British Columbia harbour. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10):2137–2153, October 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-174.

Ahmed:1967:PSC

[978] Muzammil Ahmed and Albert K. Sparks. A preliminary study of chromosomes of two species of oysters (Ostrea lurida and Crassostrea gigas). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10):2155–2159, October 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-175.

Arai:1967:ESP

[979] Hisao P. Arai. Ecological specificity of parasites of some embiotocid fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10):2161–2168, October 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-176.

Thurston:1967:EPB

[980] Robert V. Thurston. Electrophoretic patterns of blood serum proteins from rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10):2169–2188, October 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-177.

Narver:1967:PPT

[981] David W. Narver. Primary production in two small lakes of the Northern Interior Plateau of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10):2189–2193, October 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-178.

Evelyn:1967:PSS

[982] T. P. T. Evelyn. Pigments from a sockeye salmon (Oncorhynchus nerka) with unusual skin colouration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10):2195–2199, October 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-179.

LeBrasseur:1967:LFO

[983] R. J. LeBrasseur. Line fishing at Ocean Station P, 50^{o}00’ N, 145^{o}00’ W. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10):2201–2203, October 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-180.

Anonymous:1967:RAIj

[984] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(10):2205–2206, October 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-181.

Dodimead:1967:ACO

[985] A. J. Dodimead and G. L. Pickard. Annual changes in the oceaniccoastal waters of the Eastern subarctic Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2207–2227, November 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-182.

Dugal:1967:HIF

[986] L. C. Dugal. Hypoxanthine in iced freshwater fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2229–2239, November 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-183.

Kranck:1967:BSK

[987] Kate Kranck. Bedrock and sediments of Kouchibouguac Bay, New Brunswick. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2241–2265, November 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-184.

DeWilde:1967:HAT

[988] Mary Anne DeWilde and A. H. Houston. Hematological aspects of the thermoacclimatory process in the rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2267–2281, November 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-185.

Efford:1967:TSD

[989] Ian E. Efford. Temporal and spatial differences in phytoplankton productivity in Marion Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2283–2307, November 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-186.

Buerkle:1967:AAC

[990] Udo Buerkle. An audiogram of the Atlantic cod, Gadus morhua L. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2309–2319, November 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-187.

Healey:1967:OPS

[991] M. C. Healey. Orientation of pink salmon (Oncorhynchus gorbuscha) during early marine migration from Bella Coola River system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2321–2338, November 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-188.

Stewart:1967:MWR

[992] James E. Stewart, John W. Cornick, Diane M. Foley, M. F. Li, and C. M. Bishop. Muscle weight relationship to serum proteins, hemocytes, and hepatopancreas in the lobster, Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2339–2354, November 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-189.

Ursin:1967:MMS

[993] Erik Ursin. A mathematical model of some aspects of fish growth, respiration, and mortality. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11): 2355–2453, November 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-190.

Barraclough:1967:OLH

[994] W. E. Barraclough. Occurrence of larval herring (Clupea pallasii) in the Strait of Georgia during July 1966. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2455–2460, November 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-191.

Anderson:1967:NPB

[995] R. Stewart Anderson. A new plastic-bag water sampler. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2461–2465, November 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-192.

Johannes:1967:MOM

[996] R. E. Johannes and Masako Satomi. Measuring organic matter retained by aquatic invertebrates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2467–2471, November 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-193.

Vallentyne:1967:SML

[997] J. R. Vallentyne. A simplified model of a lake for instructional use. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2473–2479, November 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-194.

Mercer:1967:WPW

[998] M. C. Mercer. Wintering of pilot whales, Globicephala melaena, in Newfoundland inshore waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2481–2484, November 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-195.

Hart:1967:FLW

[999] J. L. Hart. Fecundity and length–weight relationship in lingcod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2485–2489, November 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-196.

Keyte:1967:CVM

[1000] F. K. Keyte. Computer vs. manual calculations of temperatures and depths. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2491–2495, November 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-197.

Gilgan:1967:IMH

[1001] Michael W. Gilgan and M. Elizabeth Retallack. Insect molt-hormone activity of an extract of the Japanese yew (Taxus cuspidata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2497–2499, November 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-198.

Anonymous:1967:RAIk

[1002] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(11):2501–2502, November 1967. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f67-199.

Mitchell:1967:CBW

[1003] Edward Mitchell and Warren J. Houck. Cuvier’s beaked whale (Ziphius cavirostris) stranded in Northern California. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2503–2513, December 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-200.

Miller:1967:SCA

[1004] Bruce Stuart Miller. Stomach contents of adult starry flounder and sand sole in East Sound, Orcas Island, Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2515–2526, December 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-201.

Paulik:1967:EMS

[1005] G. J. Paulik, A. S. Hourston, and P. A. Larkin. Exploitation of multiple stocks by a common fishery. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2527–2537, December 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-202.

Berst:1967:LPR

[1006] A. H. Berst and A. A. Wainio. Lamprey parasitism of rainbow trout in Southern Georgian Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2539–2548, December 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-203.

Bishop:1967:UWS

[1007] C. M. Bishop and P. H. Odense. The ultrastructure of the white striated myotomal muscle of the cod, Gadus morhua. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2549–2553, December 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-204.

Jonas:1967:PAR

[1008] R. E. E. Jonas and E. Bilinski. Phospholipase a activity in rainbow trout muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2555–2562, December 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-205.

Ackman:1967:CMO

[1009] R. G. Ackman, C. A. Eaton, and P. J. Ke. Canadian marine oils of low iodine value: Fatty acid composition of oils from Newfoundland turbot (Greenland halibut), certain Atlantic herring, and a sablefish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2563–2572, December 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-206.

Patriquin:1967:BGM

[1010] David G. Patriquin. Biology of Gadus morhua in Ogac Lake, a landlocked fiord on Baffin Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12): 2573–2594, December 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-207.

Saunders:1967:ETF

[1011] Richard L. Saunders and K. Radway Allen. Effects of tagging and of finclipping on the survival and growth of Atlantic salmon between smolt and adult stages. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12): 2595–2611, December 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-208.

Raleigh:1967:GCL

[1012] Robert F. Raleigh. Genetic control in the lakeward migrations of sockeye salmon (Oncorhynchus nerka) fry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2613–2622, December 1967. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f67-209.

Stewart:1967:VSL

[1013] James E. Stewart and John W. Cornick. In vitro susceptibilities of the lobster pathogen Gaffkya homari to various disinfectants and antibiotics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2623–2626, December 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-210.

McMullen:1967:BKC

[1014] John C. McMullen. Breeding king crabs Paralithodes camtschatica located in ocean environment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2627–2628, December 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-211.

Bernard:1967:CCN

[1015] F. Bernard. Cuspidaria cowani, a new septibranch mollusc from the Northeastern Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12): 2629–2630, December 1967. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f67-212.

Mercer:1967:RAW

[1016] M. C. Mercer. Records of the Atlantic walrus, Odobenus rosmarus rosmarus, from Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2631–2635, December 1967. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f67-213.

Quayle:1967:PRC

[1017] D. B. Quayle, J. Flury, B. Ackerman, and F. Bernard. A portable recompression chamber for treatment of victims of diving accidents. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2637–2638, December 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-214.

Jensen:1967:ASC

[1018] Jens Møller Jensen. Atlantic salmon caught in the Irminger Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2639–2640, December 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f67-215.

Peden:1967:ECS

[1019] Alex E. Peden. Erratum: Correct spelling of Alectridium aurantiacum. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 24(12):2641, December 1967. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f67-216.

Fenderson:1968:CAF

[1020] Owen C. Fenderson, W. Harry Everhart, and Kenneth M. Muth. Comparative agonistic and feeding behavior of hatchery-reared and wild salmon in aquaria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1): 1–14, January 1968. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f68-001.

Caillouet:1968:LAC

[1021] Charles W. Caillouet, Jr. Lactic acidosis in channel catfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1):15–23, January 1968. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f68-002.

Hunn:1968:TUE

[1022] Joseph B. Hunn, Richard A. Schoettger, and Wayne A. Willford. Turnover and urinary excretion of free and acetylated M.S. 222 by rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1):25–31, January 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-003.

Hartman:1968:DJS

[1023] G. F. Hartman and C. A. Gill. Distributions of juvenile steelhead and cutthroat trout (Salmo gairdneri and S. clarki clarki) within streams in Southwestern British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1):33–48, January 1968. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f68-004.

Dahlberg:1968:IDO

[1024] Michael L. Dahlberg, Dean L. Shumway, and Peter Doudoroff. Influence of dissolved oxygen and carbon dioxide on swimming performance of largemouth bass and Coho salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1):49–70, January 1968. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f68-005.

Donaldson:1968:FMD

[1025] Edward M. Donaldson, Ulf H. Fagerlund, and Peter J. Schmidt. Fluorimetric method for the determination of cortisol in small quantities of salmonid plasma. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25 (1):71–79, January 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-006.

Halsey:1968:AWL

[1026] T. G. Halsey. Autumnal and over-winter limnology of three small eutrophic lakes with particular reference to experimental circulation and trout mortality. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1): 81–99, January 1968. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f68-007.

Nelson:1968:HIM

[1027] J. S. Nelson. Hybridization and isolating mechanisms between Catostomus commersonii and C. macrocheilus (Pisces: Catostomidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1):101–150, January 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-008.

Smith:1968:HCC

[1028] Charlie E. Smith. Hematological changes in Coho salmon fed a folic acid deficient diet. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25 (1):151–156, January 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-009.

Narver:1968:IME

[1029] David W. Narver. The isopod Mesidotea entomon in the Chignik Lakes, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1):157–167, January 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-010.

Dugal:1968:PRF

[1030] L. C. Dugal. Pesticide residue in freshwater fish oils and meals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1):169–172, January 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-011.

Houston:1968:HCR

[1031] A. H. Houston and Mary Anne DeWilde. Hematological correlations in the rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1):173–176, January 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-012.

Beckett:1968:HAT

[1032] J. S. Beckett. A harpoon adapter for tagging large free-swimming fish at the surface. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1): 177–179, January 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-013.

Peden:1968:TNS

[1033] Alex E. Peden. Two new specimens of the notacanthid fish Macdonaldia challengeri in the Eastern North Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1):181–188, January 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-014.

Duffy:1968:DRM

[1034] J. R. Duffy and D. O’Connell. DDT residues and metabolites in Canadian Atlantic Coast fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25 (1):189–195, January 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-015.

Almeida:1968:MST

[1035] L. J. Almeida, E. J. Da Silva, and Yvonne M. Freitas. Microorganisms from some tropical fish diseases. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1):197–201, January 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-016.

Phinney:1968:WRE

[1036] Duane E. Phinney and Michael L. Dahlberg. Western range extension of the surf smelt, Hypomesus pretiosus pretiosus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1):203–204, January 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-017.

Berkeley:1968:REV

[1037] C. Berkeley. Records of earthworms from Vancouver, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1):205, January 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-018.

Anonymous:1968:RAIa

[1038] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(1):207, January 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-019.

Smith:1968:EPF

[1039] M. W. Smith and J. W. Saunders. Effect of pond formation on catches of brook trout from a coastal stream system. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):209–238, February 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-020.

Fraser:1968:NDO

[1040] Doris I. Fraser, Dianne P. Pitts, and W. J. Dyer. Nucleotide degradation and organoleptic quality in fresh and thawed mackerel muscle held at and above ice temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):239–253, February 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-021.

Lawler:1968:MSF

[1041] G. H. Lawler and M. Fitz-Earle. Marking small fish with stains for estimating populations in Heming Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):255–266, February 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-022.

Ackman:1968:ORD

[1042] R. G. Ackman and H. J. Hingley. The occurrence and retention of dimethyl-β-propiothetin in some filter-feeding organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):267–284, February 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-023.

Keast:1968:FBB

[1043] Allen Keast. Feeding biology of the black crappie, Pomoxis nigromaculatus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):285–297, February 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-024.

Peters:1968:ESR

[1044] J. A. Peters, J. W. Slavin, J. P. Lane, W. A. MacCallum, E. J. Laishley, W. J. Dyer, D. R. Idler, and D. I. Fraser. Effect of stage of rigor and of freezing–thawing processes on storage quality of refrozen cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):299–320, February 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-025.

Kabata:1968:SCC

[1045] Z. Kabata. Some Chondracanthidae (Copepoda) from fishes of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):321–345, February 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-026.

Squires:1968:DCQ

[1046] H. J. Squires. Decapod crustacea from the Queen Elizabeth and nearby islands in 1962. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2): 347–362, February 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-027.

Truscott:1968:SZP

[1047] B. Truscott, D. R. Idler, R. J. Hoyle, and H. C. Freeman. Sub-zero preservation of Atlantic salmon sperm. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):363–372, February 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-028.

John:1968:RRB

[1048] Kenneth R. John and David M. Gring. Retinomotor rhythms in the bluegill, Lepomis macrochirus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):373–381, February 1968. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f68-029.

Wolf:1968:IPN

[1049] Ken Wolf, M. C. Quimby, C. P. Carlson, and G. L. Bullock. Infectious pancreatic necrosis: Selection of virus-free stock from a population of carrier trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2): 383–391, February 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-030.

Hunter:1968:ELS

[1050] John R. Hunter. Effects of light on schooling and feeding of Jack mackerel, Trachurus symmetricus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):393–407, February 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-031.

Nelson:1968:DNN

[1051] Joseph S. Nelson. Distribution and nomenclature of North American Kokanee, Oncorhynchus nerka. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):409–414, February 1968. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f68-032.

Nelson:1968:VGN

[1052] Joseph S. Nelson. Variation in gillraker number in North American Kokanee, Oncorhynchus nerka. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):415–420, February 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-033.

Halliday:1968:OPT

[1053] R. G. Halliday. Occurrence of Parasudis truculentus (Goode and Bean) 1895 (Iniomi: Chlorophthalmidae) off LaHave Bank, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):421–422, February 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-034.

Campbell:1968:IBT

[1054] Gordon Campbell and R. M. MacKelvie. Infection of brook trout (Salvelinus fontinalis) by Nocardiae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):423–425, February 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-035.

Scarratt:1968:DLL

[1055] D. J. Scarratt. Distribution of lobster larvae (Homarus americanus) off Pictou, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2): 427–430, February 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-036.

Truscott:1968:BAH

[1056] B. Truscott and D. R. Idler. Biosynthesis of aldosterone and 18hydroxycorticosterone from corticosterone by interrenal tissue of a teleost (Clupea harengus harengus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):431–435, February 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-037.

Anonymous:1968:RAIb

[1057] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(2):437, February 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-038.

Hoffman:1968:DLR

[1058] Ethelwyn G. Hoffman. Description of laboratory-reared larvae of Paralithodes platypus (Decapoda, Anomura, Lithodidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):439–455, March 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-039.

Taylor:1968:RMT

[1059] F. H. C. Taylor. The relationship of midwater trawl catches to sound scattering layers off the coast of Northern British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):457–472, March 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-040.

Gronlund:1968:BLC

[1060] William D. Gronlund, Harold O. Hodgins, Raymond C. Simon, and Douglas D. Weber. Blood lactate concentrations and mortality in sockeye and Chinook salmon (Oncorhynchus nerka and O. tshawytscha) after exercise. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):473– 484, March 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-041.

Kerns:1968:BDS

[1061] Orra E. Kerns, Jr. and John R. Donaldson. Behavior and distribution of spawning sockeye salmon on island beaches in Iliamna Lake, Alaska, 1965. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):485–494, March 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-042.

Alderdice:1968:SES

[1062] D. F. Alderdice and C. R. Forrester. Some effects of salinity and temperature on early development and survival of the English sole (Parophrys vetulus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):495– 521, March 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-043.

Brown:1968:ACC

[1063] Seward R. Brown. Absorption coefficients of chlorophyll derivatives. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):523–540, March 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-044.

Bernard:1968:ASP

[1064] F. R. Bernard. The aquiferous system of Polinices lewisi (Gastropoda, Prosobranchiata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25 (3):541–546, March 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-045.

Strasdine:1968:GSP

[1065] G. A. Strasdine and Joanne Melville. Growth and spore production of Clostridium botulinum type E in chemically defined media. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):547–553, March 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-046.

Paloheimo:1968:ASG

[1066] J. E. Paloheimo and A. C. Kohler. Analysis of the Southern Gulf of St. Lawrence cod population. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):555–578, March 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-047.

Whitworth:1968:EDF

[1067] Walter R. Whitworth. Effects of diurnal fluctuations of dissolved oxygen on the growth of brook trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):579–584, March 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-048.

Alderdice:1968:DCE

[1068] D. F. Alderdice and F. P. J. Velsen. Design of a controlled-environment incubator for small marine fish eggs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):585–588, March 1968. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f68-049.

Taylor:1968:BHS

[1069] F. H. C. Taylor. Behaviour of herring schools in response to a midwater trawl. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):589–590, March 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-050.

Eberhardt:1968:PEG

[1070] L. L. Eberhardt and R. E. Nakatani. A postulated effect of growth on retention time of metabolites. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):591–596, March 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-051.

Powles:1968:IIY

[1071] P. M. Powles, D. G. Garnett, G. D. Ruggieri S. J., and R. F. Nigrelli. Ichthyophonus infection in yellowtail flounder (Limanda ferruginea) off Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):597– 598, March 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-052.

Sparrow:1968:FRC

[1072] R. A. H. Sparrow. A first report of chum salmon fry feeding in fresh water of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):599–602, March 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-053.

Nakano:1968:ACC

[1073] T. Nakano and N. Tomlinson. Addendum: Catecholamine and carbohydrate concentrations in rainbow trout (Salmo gairdneri) in relation to physical disturbance. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3): 603, March 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-054.

Freese:1968:UBF

[1074] M. Freese and D. Makow. Ultrasound backscatter in fresh and thawed animal tissue. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3): 605–606, March 1968. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f68-055.

Stewart:1968:CHC

[1075] James E. Stewart and J. R. Dingle. Characteristics of hemolymphs of Cancer irroratus, C. borealis, and Hyas coarctatus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):607–610, March 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-056.

Todd:1968:ISP

[1076] W. R. Todd, L. E. Laastuen, and L. G. Fowler. Induction of skin pallor in juvenile salmonids by guanidine compounds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):611–612, March 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-057.

Anonymous:1968:RAIc

[1077] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(3):613, March 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-058.

Hogan:1968:SEP

[1078] James W. Hogan and Charles O. Knowles. Some enzymatic properties of brain acetylcholinesterase from bluegill and channel catfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):615–623, April 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-059.

Seki:1968:RBP

[1079] Humitake Seki. Relation between production and mineralization of organic matter in Aburatsubo Inlet, Japan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):625–637, April 1968. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f68-060.

Castell:1968:HMI

[1080] C. H. Castell and D. M. Spears. Heavy metal ions and the development of rancidity in blended fish muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):639–656, April 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-061.

Mackay:1968:AGR

[1081] Isobel Mackay and G. Power. Age and growth of round whitefish (Prosopium cylindraceum) from Ungava. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):657–666, April 1968. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f68-062.

Smith:1968:SSF

[1082] Stanford H. Smith. Species succession and fishery exploitation in the Great Lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4): 667–693, April 1968. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f68-063.

Cornick:1968:IPG

[1083] John W. Cornick and James E. Stewart. Interaction of the pathogen Gaffkya homari with natural defense mechanisms of Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):695– 709, April 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-064.

Hergenrader:1968:ICS

[1084] Gary L. Hergenrader and Arthur D. Hasler. Influence of changing seasons on schooling behavior of yellow perch. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):711–716, April 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-065.

Rafail:1968:SAR

[1085] Samir Zaky Rafail. A statistical analysis of ration and growth relationship of plaice (Pleuronectes platessa). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):717–732, April 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-066.

MacCallum:1968:CNT

[1086] W. A. MacCallum, June I. Jaffray, D. N. Churchill, and D. R. Idler. Condition of Newfoundland trap-caught cod and its influence on quality after single and double freezing. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):733–755, April 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-067.

Parker:1968:MMS

[1087] Robert R. Parker. Marine mortality schedules of pink salmon of the Bella Coola River, Central British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):757–794, April 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-068.

Cornick:1968:PGH

[1088] John W. Cornick and James E. Stewart. Pathogenicity of Gaffkya homari for the crab Cancer irroratus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):795–799, April 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-069.

Waldron:1968:ELC

[1089] Kenneth D. Waldron. Early larvae of the canary rockfish, Sebastodes pinniger. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):801– 803, April 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-070.

Ackman:1968:RIA

[1090] R. G. Ackman and P. H. Odense. Reinterpretation of isopropyl alcohol retention in fish protein concentrate. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):805–806, April 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-071.

Johnson:1968:FBS

[1091] Clifford Ray Johnson. Food of the buffalo sculpin, Enophrys bison. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):807–812, April 1968. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f68-072.

Gruchy:1968:SDA

[1092] C. G. Gruchy and V. D. Vladykov. Sexual dimorphism in anal fin of brown trout, Salmo trutta, and close relatives. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):813–815, April 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-073.

Fraser:1968:VRA

[1093] Doris I. Fraser, Sylvia G. Simpson, and W. J. Dyer. Very rapid accumulation of hypoxanthine in the muscle of redfish stored in ice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):817–821, April 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-074.

Caulfield:1968:MCC

[1094] W. Caulfield. Method for counting Coho fry in small streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):823, April 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-075.

Grinols:1968:FBT

[1095] Richard B. Grinols and Charles D. Gill. Feeding behavior of three oceanic fishes (Oncorhynchus kisutch, Trachurus symmetricus, and Anoplopoma fimbria) from the Northeastern Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):825–827, April 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-076.

Dyer:1968:TDF

[1096] W. J. Dyer, Doris I. Fraser, and M. Greenwell. Thaw drip in frozen American plaice fillets, its seasonal variation and development during storage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):829–833, April 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-077.

Anonymous:1968:RAId

[1097] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(4):835, April 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-078.

Beamish:1968:GLA

[1098] F. W. H. Beamish. Glycogen and lactic acid concentrations in Atlantic cod (Gadus morhua) in relation to exercise. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):837–851, May 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-079.

Liu:1968:GMC

[1099] D. Liu and P. M. Townsley. Glucose metabolism in the caecum of the marine borer Bankia setacea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):853–862, May 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-080.

Edelhauser:1968:UTC

[1100] Henry F. Edelhauser and Kenneth A. Siegesmund. Ultrastructure of trout cornea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5): 863–866, May 1968. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f68-081.

Groves:1968:ROV

[1101] Alan B. Groves, Gerald B. Collins, and Parker S. Trefethen. Roles of olfaction and vision in choice of spawning site by homing adult Chinook salmon (Oncorhynchus tshawytscha). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):867–876, May 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-082.

Templeman:1968:RMF

[1102] Wilfred Templeman. A review of the morid fish genus Halargyreus with first records from the Western North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):877–901, May 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-083.

Scott:1968:FFH

[1103] W. B. Scott and S. N. Tibbo. Food and feeding habits of swordfish, Xiphias gladius, in the Western North Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):903–919, May 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-084.

Castell:1968:PTF

[1104] C. H. Castell, D. M. Bishop, and Wanda E. Neal. Production of trimethylamine in frozen cod muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):921–933, May 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-085.

Dambergs:1968:CFA

[1105] N. Dambergs, P. Odense, and R. Guilbault. Changes in free amino acids in skeletal muscle of cod (Gadus morhua) under conditions simulating gillnet fishing. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5): 935–942, May 1968. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f68-086.

Steele:1968:AAA

[1106] Donald H. Steele and Pierre Brunel. Amphipoda of the Atlantic and Arctic coasts of North America: Anonyx (Lysianassidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):943–1060, May 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-087.

Ackman:1968:CRF

[1107] R. G. Ackman and P. J. Ke. Commercial redfish and flatfish (flounder) oils: Comparative features of fatty acid composition. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):1061–1065, May 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-088.

Bilton:1968:COS

[1108] H. T. Bilton and D. W. Jenkinson. Comparison of the otolith and scale methods for aging sockeye (Oncorhynchus nerka) and chum (O. keta) salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):1067– 1069, May 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-089.

Jones:1968:OLM

[1109] Leslie T. Jones. Occurrence of the larvae of Meganyctiphanes norvegica (Crustacea, Euphausiacea) off West Greenland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):1071–1073, May 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-090.

Scott:1968:OPS

[1110] W. B. Scott and S. N. Tibbo. An occurrence of the pelagic stingray Dasyatis violacea in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):1075–1076, May 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-091.

Yesaki:1968:ERR

[1111] Mitsuo Yesaki and Robert J. Wolotira, Jr. Extension of recorded range of butter sole, Isopsetta isolepis, into the Bering Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):1077–1078, May 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-092.

Lawrence:1968:OIO

[1112] D. J. Lawrence. Observation of inertial oscillation in the Northwest Atlantic at 42^{o}N. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5): 1079–1083, May 1968. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f68-093.

Manzer:1968:FPS

[1113] J. I. Manzer. Food of Pacific salmon and steelhead trout in the Northeast Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5): 1085–1089, May 1968. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f68-094.

Stewart:1968:MOP

[1114] Kenneth W. Stewart and Catherine B. Levin. A method of obtaining permanent dry mounted chromosome preparations from teleost fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):1091–1093, May 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-095.

Anonymous:1968:RAIe

[1115] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(5):1095, May 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-096.

Rao:1968:NOB

[1116] Desiraju B. Rao. Natural oscillations of the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1097–1114, June 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-097.

Harvey:1968:SRK

[1117] H. H. Harvey, W. S. Hoar, and C. R. Bothern. Sounding response of the Kokanee and sockeye salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1115–1131, June 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-098.

Keast:1968:DFP

[1118] Allen Keast and Linda Welsh. Daily feeding periodicities, food uptake rates, and dietary changes with hour of day in some lake fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1133–1144, June 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-099.

Spangler:1968:AHM

[1119] George R. Spangler. Angler harvest and mortality of Esox masquinongy in Pigeon and Sturgeon Lakes, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1145–1154, June 1968. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f68-100.

Buerkle:1968:RPT

[1120] Udo Buerkle. Relation of pure tone thresholds to background noise level in the Atlantic cod (Gadus morhua). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1155–1160, June 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-101.

Crisp:1968:DPI

[1121] D. J. Crisp. Distribution of the parasitic isopod Hemioniscus balani with special reference to the East Coast of North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1161–1167, June 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-102.

Graham:1968:SLH

[1122] Joseph J. Graham and Paul M. W. Venno. Sampling larval herring from tidewaters with buoyed and anchored nets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1169–1179, June 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-103.

Powers:1968:SBE

[1123] Charles F. Powers and Andrew Robertson. Subdivisions of the benthic environment of the Upper Great Lakes, with emphasis on Lake Michigan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1181–1197, June 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-104.

Keast:1968:FSG

[1124] Allen Keast. Feeding of some Great Lakes fishes at low temperatures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1199–1218, June 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-105.

Mathews:1968:EOM

[1125] Stephen B. Mathews. An estimate of ocean mortality of Bristol Bay sockeye salmon three years at sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1219–1227, June 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-106.

Duthie:1968:EPL

[1126] Hamish C. Duthie. Ecology of phytoplankton in Lake Belwood, a storage reservoir in Southern Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1229–1245, June 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-107.

Bell:1968:DTA

[1127] Gordon R. Bell. Distribution of transaminases (aminotransferases) in the tissues of Pacific salmon (Oncorhynchus), with emphasis on the properties and diagnostic use of glutamic-oxalacetic transaminase. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1247–1268, June 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-108.

Bursa:1968:SO

[1128] Adam S. Bursa. Starch in the oceans. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1269–1284, June 1968. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f68-109.

Chaston:1968:ILA

[1129] I. Chaston. Influence of light on activity of brown trout (Salmo trutta). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1285–1289, June 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-110.

Kutty:1968:EAE

[1130] M. Krishnan Kutty. Estimation of the age of exploitation at a given fishing mortality. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25 (6):1291–1294, June 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-111.

Hoyle:1968:PRF

[1131] R. J. Hoyle and D. R. Idler. Preliminary results in the fertilization of eggs with frozen sperm of Atlantic salmon (Saltno salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1295–1297, June 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-112.

Topp:1968:EFW

[1132] Robert W. Topp. An estimate of fecundity of the winter flounder, Pseudopleuronectes americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1299–1302, June 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-113.

Knight:1968:AGE

[1133] William Knight. Asymptotic growth: An example of nonsense disguised as mathematics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25 (6):1303–1307, June 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-114.

Anonymous:1968:RAIf

[1134] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(6):1309, June 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-115.

Pearcy:1968:EPR

[1135] William G. Pearcy and Lawrence F. Small. Effects of pressure on the respiration of vertically migrating crustaceans. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1311–1316, July 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-116.

Aspinwall:1968:IMP

[1136] Nevin Aspinwall and H. Tsuyuki. Inheritance of muscle proteins in hybrids between the redside shiner (Richardsonius balteatus) and the peamouth chub (Mylocheilus caurinum). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1317–1322, July 1968. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f68-117.

Bouck:1968:CEP

[1137] G. R. Bouck and R. C. Ball. Comparative electrophoretic patterns of lactate dehydrogenase in three species of trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1323–1331, July 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-118.

Ophel:1968:MFS

[1138] I. L. Ophel and J. M. Judd. Marking fish with stable strontium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1333–1337, July 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-119.

Morrison:1968:USM

[1139] C. M. Morrison and P. H. Odense. Ultrastructure of the striated muscle of the scallop (Placopecten magellanicus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1339–1345, July 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-120.

Ryder:1968:DEM

[1140] R. A. Ryder. Dynamics and exploitation of mature walleyes, Stizostedion vitreum vitreum, in the Nipigon Bay Region of Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1347–1376, July 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-121.

Dryer:1968:RLT

[1141] William R. Dryer and George R. King. Rehabilitation of lake trout in the Apostle Islands Region of Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1377–1403, July 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-122.

Fager:1968:RGA

[1142] E. W. Fager and A. R. Longhurst. Recurrent group analysis of species assemblages of demersal fish in the Gulf of Guinea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1405–1421, July 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-123.

Ricker:1968:NAS

[1143] W. E. Ricker and H. H. Ross. North American species of Taeniopteryx (Plecoptera, Insecta). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7): 1423–1439, July 1968. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f68-124.

Richard:1968:FAP

[1144] Joseph D. Richard. Fish attraction with pulsed low-frequency sound. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1441–1452, July 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-125.

Edmundson:1968:PSJ

[1145] Eldon Edmundson, F. E. Everest, and D. W. Chapman. Permanence of station in juvenile Chinook salmon and steelhead trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1453–1464, July 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-126.

Fagerlund:1968:EMP

[1146] Ulf H. M. Fagerlund, J. R. McBride, and Edward M. Donaldson. Effect of metopirone on pituitary–interrenal function in two teleosts, sockeye salmon (Oncorhynchus nerka) and rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1465–1474, July 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-127.

Sutton:1968:USP

[1147] A. H. Sutton and J. M. Ogilvie. Uptake of sodium and phosphorus, and weight changes in prerigor cod muscle dipped in sodium tripolyphosphate solutions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25 (7):1475–1484, July 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-128.

Gee:1968:ABL

[1148] J. H. Gee. Adjustment of buoyancy by longnose dace (Rhinichthys cataractae) in relation to velocity of water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1485–1496, July 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-129.

Donaldson:1968:EHP

[1149] Edward M. Donaldson, Fumio Yamazaki, and W. Craig Clarke. Effect of hypophysectomy on plasma osmolarity in goldfish and its reversal by ovine prolactin and a preparation of salmon pituitary “Prolactin”. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1497–1500, July 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-130.

Boulva:1968:PDS

[1150] Jean Boulva and André Simard. Présence du Salvelinus namaycush (Pisces: Salmonidae) dans les eaux marines de l’Arctique occidental canadien. (French) [Occurrence of the Salvelinus namaycush (Pisces: Salmonidae) in the marine waters of the western Canadian Arctic]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1501–1504, July 1968. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f68-131.

Glenn:1968:WWM

[1151] C. L. Glenn and F. J. Ward. “Wet” weight as a method for measuring stomach contents of walleyes, Stizostedion vitreum vitreum. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1505–1507, July 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-132.

Bernard:1968:CCN

[1152] F. R. Bernard. Cyclopecten carlottensis, a new species of pectinidae from the Northeastern Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1509–1510, July 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-133.

Royer:1968:AGL

[1153] L. M. Royer, F. M. Atton, and J.-P. Cuerrier. Age and growth of lake sturgeon in the Saskatchewan River Delta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1511–1516, July 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-134.

Knowles:1968:ESF

[1154] Charles O. Knowles, Suresh K. Arurkar, and James W. Hogan. Electrophoretic separation of fish brain esterases. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1517–1519, July 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-135.

McDermott:1968:ASI

[1155] L. A. McDermott. Aeromonas sp. infection in Great Lakes lampreys. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1521–1522, July 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-136.

Anonymous:1968:RAIg

[1156] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(7):1523, July 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-137.

Nowlan:1968:PIQ

[1157] Sandra S. Nowlan, W. J. Dyer, and Doris I. Fraser. Problems involved in the quantitative extraction of glycogen and high-energy phosphorus compounds in frozen cod muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1525–1538, August 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-138.

Tarr:1968:PDG

[1158] H. L. A. Tarr. Postmortem degradation of glycogen and starch in fish muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1539–1554, August 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-139.

Bilinski:1968:ESF

[1159] E. Bilinski and L. J. Gardner. Effect of starvation on free fatty acid level in blood plasma and muscular tissues of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1555–1560, August 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-140.

Sipos:1968:JCC

[1160] J. C. Sipos and R. G. Ackman. Jellyfish (Cyanea capillata) lipids: Fatty acid composition. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8): 1561–1569, August 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-141.

Hogan:1968:DOF

[1161] James W. Hogan and Charles O. Knowles. Degradation of organophosphates by fish liver phosphatases. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1571–1579, August 1968. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f68-142.

Gould:1968:MEE

[1162] Edith Gould. Malic enzyme: Evidence for two molecular forms in the sarcoplasm of fish skeletal muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1581–1589, August 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-143.

Miles:1968:REA

[1163] S. G. Miles. Rheotaxis of elvers of the American eel (Anguilla rostrata) in the laboratory to water from different streams in Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1591–1602, August 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-144.

Ackman:1968:MPF

[1164] R. G. Ackman, C. S. Tocher, and J. McLachlan. Marine phytoplankter fatty acids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1603–1620, August 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-145.

Butler:1968:EIS

[1165] Jerry A. Butler, Raymond E. Millemann, and Nelson E. Stewart. Effects of the insecticide Sevin on survival and growth of the cockle clam Clinocardium nuttalli. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1621–1635, August 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-146.

Tencati:1968:PAS

[1166] John R. Tencati and Stephen R. Geiger. Pelagic amphipods of the slope waters of Northeast Greenland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1637–1650, August 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-147.

Huang:1968:BII

[1167] Chau-Ting Huang and Cleveland P. Hickman, Jr. Binding of inorganic iodide to the plasma proteins of teleost fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1651–1666, August 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-148.

Booke:1968:CSC

[1168] Henry E. Booke. Cytotaxonomic studies of the Coregonine fishes of the Great Lakes, USA: DNA and karyotype analysis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1667–1687, August 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-149.

Kutty:1968:RQG

[1169] Methil Narayan Kutty. Respiratory quotients in goldfish and rainbow trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1689–1728, August 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-150.

McLeese:1968:OCS

[1170] D. W. McLeese and J. Watson. Oxygen consumption of the spider crab (Chionoecetes opilio) and the American lobster (Homarus americanus) at a low temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8): 1729–1732, August 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-151.

McLeese:1968:TRS

[1171] D. W. McLeese. Temperature resistance of the spider crab Chionoecetes opilio. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1733–1736, August 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-152.

Ellis:1968:SCP

[1172] Derek V. Ellis. A series of computer programs for summarising data from quantitative benthic investigations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1737–1738, August 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-153.

Weis:1968:TCP

[1173] Judith S. Weis. Taxonomic correlation of the position of spinal ganglia in teleost fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8): 1739–1740, August 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-154.

Applegate:1968:WCP

[1174] Richard L. Applegate, Alfred C. Fox, and Victor J. Starostka. A water core plankton sampler. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1741–1742, August 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-155.

Levings:1968:FEB

[1175] C. D. Levings. Fertilized eggs of the butter sole, Isopsetta isolepis, in Skidegate Inlet, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1743–1744, August 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-156.

Boyd:1968:EFH

[1176] J. W. Boyd and B. A. Southcott. Evaluation of fish homogenates for use as substrates in Clostridium botulinum type E investigations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1745–1747, August 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-157.

McAlice:1968:OCR

[1177] Bernard J. McAlice. An occurrence of ciliate red water in the Gulf of Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1749–1751, August 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-158.

Boyd:1968:CEE

[1178] J. W. Boyd and B. A. Southcott. Comparative effectiveness of ethylenediaminetetraacetic acid and chlortetracycline for fish preservation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1753–1758, August 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-159.

Anonymous:1968:CRP

[1179] Anonymous. Corrections for recent publications of The Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1759–1760, August 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-160.

Anonymous:1968:RAIh

[1180] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(8):1761, August 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-161.

Stewart:1968:ACI

[1181] James E. Stewart and H. J. Squires. Adverse conditions as inhibitors of ecdysis in the lobster Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1763–1774, September 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-162.

Scott:1968:MDG

[1182] J. S. Scott. Morphometrics, distribution, growth, and maturity of offshore sand launce (Ammodytes dubius) on the Nova Scotia Banks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1775–1785, September 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-163.

Macek:1968:RBT

[1183] Kenneth J. Macek. Reproduction in brook trout (Salvelinus fontinalis) fed sublethal concentrations of DDT. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1787–1796, September 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-164.

Gakstatter:1968:RAD

[1184] Jack H. Gakstatter. Rates of accumulation of ^{14}C-dieldrin residues in tissues of goldfish exposed to a single sublethal dose of ^{14}C-aldrin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1797–1801, September 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-165.

Brawn:1968:CCS

[1185] V. M. Brawn, D. L. Peer, and R. J. Bentley. Caloric content of the standing crop of benthic and epibenthic invertebrates of St. Margaret’s Bay, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9): 1803–1811, September 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-166.

Legault:1968:FPE

[1186] Roméo-O. Legault and Claude Delisle. La fraye d’une population d’éperlans géants, Osmerus eperlanus mordax, au lac Heney, Comté de Gatineau, Québec. (French) [The spawning of a population of giant perlans, Osmerus eperlanus mordax, at Lake Heney, Gatineau County, Québec]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1813–1830, September 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-167.

Simard:1968:ACD

[1187] André Simard and Étienne Magnin. Âge et croissance du maskinongé, Esox masquinongy, dans quatre lacs du Québec. (French) [Age and growth of the muskellunge, Esox masquinongy, in four lakes in Quebec]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1831–1842, September 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-168.

Mitchell:1968:MPP

[1188] Edward Mitchell. The mio-Pliocene pinniped Imagotaria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1843–1900, September 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-169.

Johnston:1968:ITP

[1189] C. E. Johnston and J. G. Eales. Influence of temperature and photoperiod on guanine and hypoxanthine levels in skin and scales of Atlantic salmon (Salmo salar) during parr–smolt transformation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1901–1909, September 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-170.

Sweers:1968:TMD

[1190] H. E. Sweers. Two methods of describing the “average” vertical temperature distribution of a lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1911–1922, September 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-171.

Parker:1968:RDC

[1191] R. R. Parker, Z. Kabata, L. Margolis, and M. D. Dean. A review and description of Caligus curtus Müller, 1785 (Caligidae: Copepoda), type species of its genus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9): 1923–1969, September 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-172.

Godfrey:1968:TAA

[1192] H. Godfrey, D. D. Worlund, and H. T. Bilton. Tests on the accuracy of ageing Chinook salmon (Oncorhynchus tshawytscha) from their scales. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1971–1982, September 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-173.

McCauley:1968:SPI

[1193] Robert W. McCauley. Suggested physiological interaction among rainbow trout fingerlings undergoing thermal stress. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1983–1986, September 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-174.

Lindsey:1968:TRW

[1194] C. C. Lindsey. Temperatures of red and white muscle in recently caught marlin and other large tropical fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1987–1992, September 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-175.

Turner:1968:APS

[1195] C. E. Turner and H. T. Bilton. Another pink salmon (Oncorhynchus gorbuscha) in its third year. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1993–1996, September 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-176.

Sinclair:1968:DLO

[1196] D. C. Sinclair. Diel limnetic occurrence of young Cottus asper in two British Columbia lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):1997–2000, September 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-177.

Ali:1968:CSR

[1197] M. A. Ali and M. Anctil. Corrélation entre la structure rétinienne et l’habitat chez Stizostedion vitreum vitreum et S. canadense. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):2001–2003, September 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-178.

Reimers:1968:SBA

[1198] Paul E. Reimers. Social behavior among juvenile fall Chinook salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):2005–2008, September 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-179.

Anonymous:1968:RAIi

[1199] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(9):2009, September 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-180.

Smith:1968:FPC

[1200] M. W. Smith. Fertilization and predator control to increase growth rate and yield of trout in a natural lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2011–2036, October 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-181.

Stockner:1968:AGP

[1201] John G. Stockner. Algal growth and primary productivity in a thermal stream. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2037–2058, October 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-182.

Anderson:1968:RBP

[1202] Margaret L. Anderson and Elinor M. Ravesi. Relation between protein extractability and free fatty acid production in cod muscle aged in ice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2059–2069, October 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-183.

Power:1968:UEC

[1203] H. E. Power, R. Sinclair, and K. Savagaon. Use of EDTA compounds for the preservation of haddock fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2071–2082, October 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-184.

Addison:1968:DFA

[1204] R. F. Addison, R. G. Ackman, and J. Hingley. Distribution of fatty acids in cod flesh lipids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10): 2083–2090, October 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-185.

Qadri:1968:GRL

[1205] S. U. Qadri. Growth and reproduction of the lake whitefish, Coregonus clupeaformis, in Lac la Ronge, Saskatchewan. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2091–2100, October 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-186.

Billaud:1968:NFU

[1206] Vera A. Billaud. Nitrogen fixation and the utilization of other inorganic nitrogen sources in a subarctic lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2101–2110, October 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-187.

Hogman:1968:AFS

[1207] Walter J. Hogman. Annulus formation on scales of four species of coregonids reared under artificial conditions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2111–2122, October 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-188.

Caddy:1968:UOS

[1208] J. F. Caddy. Underwater observations on scallop (Placopecten magellanicus) behaviour and drag efficiency. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2123–2141, October 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-189.

Miles:1968:LEO

[1209] S. G. Miles. Laboratory experiments on the orientation of the adult American eel, Anguilla rostrata. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2143–2155, October 1968. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f68-190.

Jonas:1968:DFM

[1210] R. E. E. Jonas. Dissolution of fish muscle homogenates by lysolecithin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2157–2164, October 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-191.

Murray:1968:SSA

[1211] A. R. Murray. Smolt survival and adult utilization of Little Codroy River, Newfoundland, Atlantic salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2165–2218, October 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-192.

Wallen:1968:LRD

[1212] D. G. Wallen and G. H. Geen. Loss of radioactivity during storage of 14C-labelled phytoplankton on membrane filters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2219–2224, October 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-193.

Sheri:1968:RWP

[1213] A. N. Sheri and G. Power. Reproduction of white perch, Roccus americanus, in the Bay of Quinte, Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2225–2231, October 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-194.

Duncan:1968:TMF

[1214] Richard N. Duncan and Ivan J. Donaldson. Tattoo-marking of fingerling salmonids with fluorescent pigments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2233–2236, October 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-195.

Pletcher:1968:SDT

[1215] F. T. Pletcher. A subcutaneous dart tag for fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2237–2240, October 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-196.

Taylor:1968:PTP

[1216] F. J. R. Taylor. Parasitism of the toxin-producing dinoflagellate Gonyaulax catenella by the endoparasitic dinoflagellate Amoebophrya ceratii. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2241–2245, October 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-197.

Kennedy:1968:EFT

[1217] W. A. Kennedy. Effect of the FP-66 tag on sablefish growth. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2247–2249, October 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f68-198.

Cobb:1968:DML

[1218] J. Stanley Cobb. Delay of moult by the larvae of Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2251–2253, October 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-199.

Anonymous:1968:RAIj

[1219] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(10):2255, October 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-200.

Budd:1968:SML

[1220] J. C. Budd, F. E. J. Fry, and J. B. Smith. Survival of marked lake trout in Lake Manitou, Manitoulin Island, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2257–2268, November 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-201.

Powell:1968:BPA

[1221] N. A. Powell. Bryozoa (Polyzoa) of Arctic Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2269–2320, November 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-202.

Greer:1968:DDA

[1222] G. L. Greer and U. Paim. Degradation of DDT in Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2321–2326, November 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-203.

Loring:1968:OSI

[1223] D. H. Loring and D. J. G. Nota. Occurrence and significance of iron, manganese, and titanium in glacial marine sediments from the estuary of the St. Lawrence River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2327–2347, November 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-204.

Cushing:1968:DEF

[1224] D. H. Cushing. Direct estimation of a fish population acoustically. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2349–2364, November 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-205.

vanderLand:1968:TNS

[1225] Jacob van der Land and Wilfred Templeman. Two new species of Gyrocotyle (Monogenea) from Hydrolagus affinis (Brito Capello) (Holocephali). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2365–2385, November 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-206.

Symons:1968:IAS

[1226] Philip E. K. Symons. Increase in aggression and in strength of the social hierarchy among juvenile Atlantic salmon deprived of food. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2387–2401, November 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-207.

Vanstone:1968:SMB

[1227] W. E. Vanstone and J. R. Markert. Some morphological and biochemical changes in Coho salmon, Oncorhynchus kisutch, during parr–smolt transformation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11): 2403–2418, November 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-208.

Jangaard:1968:PFA

[1228] P. M. Jangaard and P. J. Ke. Principal fatty acids of Depot fat and milk lipids from harp seal (Pagophilus groenlandica) and hooded seal (Cystophora cristata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2419–2426, November 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-209.

Colton:1968:RTS

[1229] John B. Colton, Jr. Recent trends in subsurface temperatures in the Gulf of Maine and contiguous waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2427–2437, November 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-210.

White:1968:ASS

[1230] H. C. White and J. C. Medcof. Atlantic salmon scales as records of spawning history. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11): 2439–2441, November 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-211.

Macek:1968:GRS

[1231] Kenneth J. Macek. Growth and resistance to stress in brook trout fed sublethal levels of DDT. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2443–2451, November 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-212.

Brooke:1968:SDS

[1232] Richard O. Brooke, Joseph M. Mendelsohn, and Frederick J. King. Significance of dimethyl sulfide to the odor of soft-shell clams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2453–2460, November 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-213.

Cross:1968:DTT

[1233] F. A. Cross, S. W. Fowler, J. M. Dean, L. F. Small, and C. L. Osterberg. Distribution of ^{65}Zn in tissues of two marine crustaceans determined by autoradiography. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11): 2461–2466, November 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-214.

Allen:1968:MAC

[1234] Richard L. Allen, Thomas K. Meekin, Gilbert B. Pauley, and M. Paul Fujihara. Mortality among Chinook salmon associated with the fungus Dermocystidium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11): 2467–2475, November 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-215.

Tsuyuki:1968:CPE

[1235] H. Tsuyuki, E. Roberts, R. H. Lowes, W. Hadaway, and S. J. Westrheim. Contribution of protein electrophoresis to rockfish (Scorpaenidae) systematics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2477–2501, November 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-216.

Hourston:1968:ACG

[1236] A. S. Hourston. Abnormal cessation of growth in a herring otolith. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2503–2504, November 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-217.

Bell:1968:ULS

[1237] R. K. Bell and F. T. Ward. Use of liquid scintillation counting for measuring self-absorption of carbon-14 in Daphnia pulex. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2505–2508, November 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-218.

Westrheim:1968:FRT

[1238] S. J. Westrheim. First records of three rockfish species (Sebastodes aurora, S. ciliatus, and Sebastolobus altivelis) from waters off British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2509–2513, November 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-219.

Wissing:1968:CVS

[1239] Thomas E. Wissing and Arthur D. Hasler. Calorific values of some invertebrates in Lake Mendota, Wisconsin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2515–2518, November 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-220.

Owens:1968:RMT

[1240] Eben L. Owens and Gary A. Gebhardt. A rapid marking technique for identification of individual fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2519–2520, November 1968. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f68-221.

Burt:1968:MFO

[1241] M. D. B. Burt and R. E. Drinnan. A microturbellarian found in oysters off the coast of Prince Edward Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2521–2522, November 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-222.

Lawler:1968:TNB

[1242] G. H. Lawler. Triaenophorus nodulosus in burbot, Lota lota, from Heming Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11): 2523–2524, November 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-223.

Anonymous:1968:RAIk

[1243] Anonymous. Recent articles on investigations of the Fisheries Research Board of Canada published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(11):2525, November 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-224.

MacCrimmon:1968:WDB

[1244] Hugh R. MacCrimmon and T. L. Marshall. World distribution of brown trout, Salmo trutta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12): 2527–2548, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-225.

Idler:1968:ISR

[1245] D. R. Idler and B. J. Szeplaki. Interrenalectomy and stress in relation to some blood components of an elasmobranch (Raja radiata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2549–2560, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-226.

Fitch:1968:FOC

[1246] John E. Fitch and Robert L. Brownell, Jr. Fish otoliths in cetacean stomachs and their importance in interpreting feeding habits. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2561–2574, December 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-227.

Kalff:1968:SPC

[1247] J. Kalff. Some physical and chemical characteristics of Arctic fresh waters in Alaska and Northwestern Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2575–2587, December 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-228.

Semakula:1968:AGF

[1248] S. N. Semakula and P. A. Larkin. Age, growth, food, and yield of the white sturgeon (Acipenser transmontanus) of the Fraser River, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2589–2602, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-229.

Hart:1968:RWG

[1249] J. S. Hart. Respiration of the Winnipeg goldeye (Hiodon alosoides). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2603–2608, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-230.

Hulburt:1968:SMC

[1250] Edward M. Hulburt. Stratification and mixing in coastal waters of the Western Gulf of Maine during summer. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2609–2621, December 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-231.

Mounib:1968:CPA

[1251] M. S. Mounib, P. C. Hwang, and D. R. Idler. Cryogenic preservation of Atlantic cod (Gadus morhua) sperm. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2623–2632, December 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-232.

Crisp:1968:DBN

[1252] D. J. Crisp. Differences between North American and European populations of Balanus balanoides revealed by transplantation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2633–2641, December 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-233.

McDermott:1968:EPB

[1253] L. A. McDermott and A. H. Berst. Experimental plantings of brook trout (Salvelinus fontinalis) from furunculosis-infected stock. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2643–2649, December 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-234.

Wilkins:1968:MHA

[1254] N. P. Wilkins. Multiple haemoglobins of the Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2651–2663, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-235.

Day:1968:SAB

[1255] Donald S. Day and William G. Pearcy. Species associations of benthic fishes on the Continental Shelf and Slope off Oregon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2665–2675, December 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-236.

Anderson:1968:ESD

[1256] John M. Anderson. Effect of sublethal DDT on the lateral line of brook trout, Salvelinus fontinalis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2677–2682, December 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-237.

Scarratt:1968:ARL

[1257] D. J. Scarratt. An artificial reef for lobsters (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2683–2690, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-238.

Pauley:1968:MRF

[1258] Gilbert B. Pauley and Roy E. Nakatani. Metabolism of the radioisotope 65Zn in the freshwater mussel Anodonta californiensis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2691–2694, December 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-239.

Withler:1968:ECS

[1259] F. C. Withler and R. B. Morley. Effects of chilled storage on viability of stored ova and sperm of sockeye and pink salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2695–2699, December 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-240.

Allen:1968:SMC

[1260] K. Radway Allen. Simplification of a method of computing recruitment rates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2701–2702, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-241.

Reisman:1968:RIM

[1261] Howard M. Reisman. Reproductive isolating mechanisms of the blackspotted stickleback, Gasterosteus wheatlandi. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2703–2706, December 1968. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f68-242.

Beckett:1968:NRB

[1262] J. S. Beckett. New records of the barnacle Conchoderma virgatum in the Northwest Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12): 2707–2710, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-243.

Gosselin-Rey:1968:LCK

[1263] C. Gosselin-Rey, G. Hamoir, and R. K. Scopes. Localization of creatine kinase in the starch–gel and moving-boundary electrophoretic patterns of fish muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12): 2711–2714, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-244.

Scopes:1968:PCM

[1264] R. K. Scopes and C. Gosselin-Rey. Polymorphism in carp muscle creatine kinase. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2715–2716, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-245.

Garside:1968:ULT

[1265] E. T. Garside and C. M. Jordan. Upper lethal temperatures at various levels of salinity in the euryhaline cyprinodontids Fundulus heteroclitus and F. diaphanus after isosmotic acclimation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2717–2720, December 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-246.

Barrett:1968:FOG

[1266] B. E. Barrett. First occurrence of Greenland halibut (Reinhardtius hippoglossoides) in the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2721–2722, December 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-247.

Squires:1968:LFD

[1267] H. J. Squires and G. P. Ennis. Lobsters from 95 fath deep off St. John’s, Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12): 2723–2724, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-248.

Thomas:1968:OAL

[1268] M. L. H. Thomas. Overwintering of American lobsters, Homarus americanus, in burrows in Bideford River, Prince Edward Islandq. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2725–2727, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-249.

Ennis:1968:OSS

[1269] G. P. Ennis. Occurrences of the staghorn sculpin (Gymnocanthus tricuspis) in Newfoundland waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2729–2731, December 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-250.

Delisle:1968:FRS

[1270] Claude Delisle and W. Van Vliet. First records of the sculpins Myoxocephalus thompsonii and Cottus ricei from the Ottawa Valley, Southwestern Quebec. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12): 2733–2737, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f68-251.

Park:1968:SPS

[1271] P. Kilho Park, Stephen W. Hager, Jacques E. Pirson, and David S. Ball. Surface phosphate and silicate distributions in the Northeastern Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2739–2741, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-252.

Pienaar:1968:EMW

[1272] J. V. Pienaar and W. K. Ricker. Estimating mean weight from length statistics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2743–2747, December 1968. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f68-253.

Wiles:1968:PEH

[1273] Michael Wiles. Possible effects of the harbour seal bounty on codworm infestations of Atlantic cod in the Gulf of St. Lawrence, the Strait of Belle Isle, and the Labrador Sea. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2749–2753, December 1968. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f68-254.

Anonymous:1968:IP

[1274] Anonymous. Index to publications for 1968. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 25(12):2755–2849, December 1968. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f68-255.

Stewart:1969:EIG

[1275] James E. Stewart, April Dockrill, and John W. Cornick. Effectiveness of the integument and gastric fluid as barriers against transmission of Gaffkya homari to the lobster Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):1–14, January 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-001.

Hodgins:1969:VLD

[1276] H. O. Hodgins, W. E. Ames, and F. M. Utter. Variants of lactate dehydrogenase isozymes in sera of sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):15–19, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-002.

Haines:1969:RYS

[1277] Terry A. Haines and Robert L. Butler. Responses of yearling smallmouth bass (Micropterus dolomieui) to artificial shelter in a stream aquarium. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):21–31, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-003.

Northcote:1969:LMY

[1278] T. G. Northcote. Lakeward migration of young rainbow trout (Salmo gairdneri) in the Upper Lardeau River, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):33–45, January 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-004.

Cherrington:1969:VDD

[1279] A. D. Cherrington, U. Paim, and O. T. Page. In vitro degradation of DDT by intestinal contents of Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):47–54, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-005.

Lie:1969:SCB

[1280] Ulf Lie. Standing crop of benthic infauna in Puget Sound and off the coast of Washington. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26 (1):55–62, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-006.

Mason:1969:HSP

[1281] J. C. Mason. Hypoxial stress prior to emergence and competition among Coho salmon fry. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26 (1):63–91, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-007.

Peterson:1969:ITC

[1282] R. H. Peterson and J. M. Anderson. Influence of temperature change on spontaneous locomotor activity and oxygen consumption of Atlantic salmon, Salmo salar, acclimated to two temperatures. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):93–109, January 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-008.

Smith:1969:FAA

[1283] Charlie E. Smith and John E. Halver. Folic acid anemia in Coho salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):111–114, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-009.

Wedemeyer:1969:SME

[1284] Gary Wedemeyer, A. J. Ross, and Lynwood Smith. Some metabolic effects of bacterial endotoxins in salmonid fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):115–122, January 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-010.

Pienaar:1969:AWL

[1285] L. V. Pienaar and J. A. Thomson. Allometric weight–length regression model. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):123–131, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-011.

Phillips:1969:FTM

[1286] Robert W. Phillips and K. V. Koski. A fry trap method for estimating salmonid survival from egg deposition to fry emergence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):133–141, January 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-012.

Beck:1969:OGS

[1287] B. Beck and A. W. Mansfield. Observations on the Greenland shark, Somniosus microcephalus, in Northern Baffin Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):143–145, January 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-013.

Pequegnat:1969:EZR

[1288] John E. Pequegnat, Scott W. Fowler, and Lawrence F. Small. Estimates of the zinc requirements of marine organisms. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):145–150, January 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-014.

Heinle:1969:CCC

[1289] Donald R. Heinle. Culture of calanoid copepods in synthetic sea water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):150–153, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-015.

Scott:1969:LTF

[1290] K. R. Scott. Low-temperature fish storage facility with precise temperature control. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):

154–161, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-016.

Silliman:1969:CBG

[1291] R. P. Silliman. Comparison between Gompertz and von Bertalanffy curves for expressing growth in weight of fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):161–165, January 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-017.

Muir:1969:GDF

[1292] B. S. Muir. Gill dimensions as a function of fish size. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):165–170, January 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-018.

Sheldon:1969:PCM

[1293] R. W. Sheldon, D. H. Loring, and S. Deleu. Physiographic changes in Malpeque Bay, Prince Edward Island, between 1845 and 1955, and the possible effects on oyster production. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):171–175, January 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-019.

Crossman:1969:INP

[1294] E. J. Crossman and John M. Casselman. Identification of northern pike and muskellunge from axial skeletons, scales, and epipleurals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):175–178, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-020.

Allen:1969:ACE

[1295] K. Radway Allen. An application of computers to the estimation of exploited populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26 (1):179–189, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-021.

Bernard:1969:PCM

[1296] F. R. Bernard. The parasitic copepod Mytilicola orientalis in British Columbia bivalves. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26 (1):190–191, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-022.

Forrester:1969:SPS

[1297] C. R. Forrester. Sinistrality in Platichthys stellatus off British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):191–196, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-023.

Anonymous:1969:RAFa

[1298] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(1):197, January 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-024.

McAllister:1969:AEZ

[1299] C. D. McAllister. Aspects of estimating zooplankton production from phytoplankton production. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):199–220, February 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-025.

Ball:1969:DLT

[1300] Ian R. Ball. Dugesia lugubris (Tricladida: Paludicola), a European immigrant into North American fresh waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):221–228, February 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-026.

McDonald:1969:DGS

[1301] J. G. McDonald. Distribution, growth, and survival of sockeye fry (Oncorhynchus nerka) produced in natural and artificial stream environments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):229–267, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-027.

Saunders:1969:CSN

[1302] Richard L. Saunders. Contributions of salmon from the Northwest Miramichi River, New Brunswick, to various fisheries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):269–278, February 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-028.

Bishop:1969:UMB

[1303] John E. Bishop and H. B. N. Hynes. Upstream movements of the benthic invertebrates in the Speed River, Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):279–298, February 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-029.

Lighthart:1969:PBB

[1304] Bruce Lighthart. Planktonic and benthic bacteriovorous protozoa at eleven stations in Puget Sound and adjacent Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):299–304, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-030.

Hopkins:1969:ZBR

[1305] Thomas L. Hopkins. Zooplankton biomass related to hydrography along the drift track of Arlis II in the Arctic Basin and the East Greenland Current. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):305–310, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-031.

Kabata:1969:FLC

[1306] Z. Kabata. Four Lernaeopodidae (Copepoda) parasitic on fishes from Newfoundland and West Greenland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):311–324, February 1969. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f69-032.

Henderson:1969:DFS

[1307] Nancy E. Henderson and Richard E. Peter. Distribution of fishes of Southern Alberta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2): 325–338, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-033.

Braaten:1969:RDP

[1308] Duane O. Braaten. Robustness of the DeLury population estimator. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):339–355, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-034.

Schmidt:1969:IPS

[1309] P. J. Schmidt and E. G. Baker. Indirect pigmentation of salmon and trout flesh with canthaxanthin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):357–360, February 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-035.

Sameoto:1969:CEL

[1310] D. D. Sameoto. Comparative ecology, life histories, and behaviour of intertidal sand-burrowing amphipods (Crustacea: Haustoriidae) at Cape Cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):361–388, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-036.

Reed:1969:CME

[1311] Paul H. Reed. Culture methods and effects of temperature and salinity on survival and growth of Dungeness crab (Cancer magister) larvae in the laboratory. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2): 389–397, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-037.

Gray:1969:EWC

[1312] J. R. A. Gray and J. M. Edington. Effect of woodland clearance on stream temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2): 399–403, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-038.

Narver:1969:PVT

[1313] David W. Narver. Phenotypic variation in threespine sticklebacks (Gasterosteus aculeatus) of the Chignik River System, Alaska. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):405–412, February 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-039.

Hwang:1969:SMC

[1314] P. C. Hwang and D. R. Idler. A study of major cations, osmotic pressure, and pH in seminal components of Atlantic salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):413–419, February 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-040.

Mujib:1969:COB

[1315] Khawaja Abdul Mujib. Cranial osteology of Brosme brosme with notes on its position within the subfamily Lotinae (Gadidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):421–432, February 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-041.

Gardner:1969:SBM

[1316] George R. Gardner and Paul P. Yevich. Studies on the blood morphology of three estuarine cyprinodontiform fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):433–447, February 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-042.

Bigg:1969:CPS

[1317] Michael A. Bigg. Clines in the pupping season of the harbour seal, Phoca vitulina. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):449–455, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-043.

Bidgood:1969:LTG

[1318] B. F. Bidgood and A. H. Berst. Lethal temperatures for Great Lakes rainbow trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2): 456–459, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-044.

Garside:1969:EER

[1319] E. T. Garside and J. J. Kerekes. Eastern extension of the range of the blackspotted stickleback, Gasterosteus wheatlandi, in Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):460–461, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-045.

Li:1969:FAR

[1320] M. F. Li and Carol Jordan. Factors affecting rainbow trout ovary cells cultivated in vitro. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2): 461–463, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-046.

Kalff:1969:DPO

[1321] J. Kalff. A diel periodicity in the optimum light intensity for maximum photosynthesis in natural phytoplankton populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):463–468, February 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-047.

Halliday:1969:OMC

[1322] R. G. Halliday. Occurrence and morphometries of Chlorophthalmus agassizi (Iniomi: Chlorophthalmidae) off Southwestern Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):469–472, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-048.

Musick:1969:OSS

[1323] J. A. Musick and K. W. Able. Occurrence and spawning of the sculpin Triglops murrayi (Pisces, Cottidae) in the Gulf of Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):473–475, February 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-049.

Anonymous:1969:RAFb

[1324] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(2):477, February 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-050.

Ricker:1969:ESS

[1325] W. E. Ricker. Effects of size-selective mortality and sampling bias on estimates of growth, mortality, production, and yield. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):479–541, March 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-051.

Woods:1969:PSV

[1326] Susan M. Woods. Polyteny and size variation in the copepod Pseudocalanus from two semi-landlocked fiords on Baffin Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):543–556, March 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-052.

Leong:1969:LSP

[1327] Roderick J. H. Leong and Charles P. O’Connell. A laboratory study of particulate and filter feeding of the northern anchovy (Engraulis mordax). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):557–582, March 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-053.

Brawn:1969:FBC

[1328] V. M. Brawn. Feeding behaviour of cod (Gadus morhua). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):583–596, March 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-054.

Anctil:1969:SRC

[1329] Michel Anctil. Structure de la rétine chez quelques téléostéens marins du plateau continental. (French) [Structure of the retina in some marine teleosts of the Continental Shelf]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):597–628, March 1969. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f69-055.

Paterson:1969:IPC

[1330] W. D. Paterson, R. J. Douglas, I. Grinyer, and L. A. McDermott. Isolation and preliminary characterization of some Aeromonas salmonicida bacteriophages. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3): 629–632, March 1969. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f69-056.

Beamish:1969:ADW

[1331] R. J. Beamish and H. H. Harvey. Age determination in the white sucker. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):633–638, March 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-057.

Butler:1969:CSR

[1332] D. G. Butler. Corpuscles of Stannius and renal physiology in the eel (Anguilla rostrata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26 (3):639–654, March 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-058.

Mathisen:1969:VCE

[1333] Ole A. Mathisen and Tor Gunnerød. Variance components in the estimation of potential egg deposition of sockeye salmon escapements. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):655–670, March 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-059.

LaLanne:1969:ADS

[1334] John J. LaLanne and Gunnar Safsten. Age determination from scales of chum salmon (Oncorhynchus keta). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):671–681, March 1969. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f69-060.

Steele:1969:ACG

[1335] D. H. Steele. Additional comments on the genus Anonyx (Crustacea: Amphipoda) of the Atlantic and Arctic coasts of North America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):683–686, March 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-061.

Kwain:1969:AVF

[1336] Wen hwa Kwain and Hugh R. MacCrimmon. Age and vision as factors in bottom colour selection by rainbow trout, Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):687–693, March 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-062.

Saunders:1969:MMB

[1337] J. W. Saunders. Mass mortalities and behaviour of brook trout and juvenile Atlantic salmon in a stream polluted by agricultural pesticides. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):695–699, March 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-063.

Thomas:1969:MME

[1338] M. L. H. Thomas and G. N. White. Mass mortality of estuarine fauna at Bideford, P.E.I., associated with abnormally low salinities. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):701–704, March 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-064.

Rao:1969:NRC

[1339] S. V. Suryanarayana Rao, J. R. Rangaswamy, and N. L. Lahiry. Nucleotides and related compounds in canned shrimp. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):704–706, March 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-065.

Heaton:1969:TMS

[1340] LeRoy H. Heaton and Gilbert B. Pauley. Two modified stains for use in oyster pathology. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26 (3):707–709, March 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-066.

Woronecki:1969:FRO

[1341] David E. Woronecki. First record of occurrence of the northern redbelly dace, Chrosomus eos, within Prince Edward Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):709–710, March 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-067.

Anonymous:1969:RAFc

[1342] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(3):711, March 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-068.

Margolis:1969:F

[1343] L. Margolis. FOREWORD. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4), April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-069.

Anonymous:1969:WMC

[1344] Anonymous. T. W. M. Cameron — an appreciation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):713–716, April 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-070.

Baer:1969:DPT

[1345] Jean G. Baer. Diphyllobothrium pacificum, a Tapeworm from sea lions endemic in man along the coastal area of Peru. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):717–723, April 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-071.

Stunkard:1969:SPR

[1346] Horace W. Stunkard. The sporozoa: With particular reference to infections in fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4): 725–739, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f69-072.

Orr:1969:MSS

[1347] T. S. C. Orr and C. A. Hopkins. Maintenance of Schistocephalus solidus in the laboratory with observations on rate of growth of, and proglottid formation in, the plerocercoid. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):741–752, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-073.

Szidat:1969:SDB

[1348] Lothar Szidat. Structure, development, and behaviour of new strigeatoid metacercariae from subtropical fishes of South America. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):753–786, April 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-074.

Manter:1969:TNG

[1349] Harold W. Manter. Three new genera of hemiurid trematodes from Australian fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4): 787–792, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f69-075.

Myers:1969:NFA

[1350] Betty June Myers and Robert E. Kuntz. Nematodes of fishes, amphibians, and reptiles taken by U.S. Naval Medical Research Unit No. 2 Expedition to North Borneo (Malaysia). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):793–797, April 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-076.

Arai:1969:NTG

[1351] Hisao P. Arai. A new trematode of the genus Lepidophyllum (Digenea: Steganodermatidae) from a cottid fish, Hemilepidotus hemilepidotus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):799–803, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-077.

Threlfall:1969:SPE

[1352] William Threlfall. Some parasites from elasmobranchs in Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):805–811, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-078.

Boyce:1969:PFP

[1353] Norbert P. J. Boyce. Parasite fauna of pink salmon (Oncorhynchus gorbuscha) of the Bella Coola River, Central British Columbia, during their early sea life. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4): 813–820, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f69-079.

Lawler:1969:ABT

[1354] G. H. Lawler. Aspects of the biology of Triaenophorus nodulosus in yellow perch, Perca flavescens, in Heming Lake, Manitoba. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):821–831, April 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-080.

Tedla:1969:OSC

[1355] Shibru Tedla and C. H. Fernando. Observations on the seasonal changes of the parasite fauna of yellow perch (Perca flavescens) from the Bay of Quinte, Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4): 833–843, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f69-081.

Yamaguti:1969:SMN

[1356] Satyu Yamaguti. Special modes of nutrition in some digenetic trematodes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):845–848, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-082.

Ko:1969:RGC

[1357] Ronald C. Ko and Roy C. Anderson. A revision of the genus Cystidicola Fischer, 1798 (Nematoda: Spiruroidea) of the swim bladder of fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):849–864, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-083.

Dechtiar:1969:TNS

[1358] Alex O. Dechtiar. Two new species of monogenetic trematodes (Trematoda: Monogenea) from nasal cavities of catostomid fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):865–869, April 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-084.

Freeman:1969:OTD

[1359] Reino S. Freeman and Bruce H. Thompson. Observations on transmission of Diphyllobothrium sp. (Cestoda) to lake trout in Algonquin Park, Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):871– 878, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-085.

Scott:1969:TPA

[1360] J. S. Scott. Trematode populations in the Atlantic Argentine, Argentina silus, and their use as biological indicators. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):879–891, April 1969. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f69-086.

Margolis:1969:LSM

[1361] L. Margolis and N. P. Boyce. Life span, maturation, and growth of two hemiurid trematodes, Tubulovesicula lindbergi and Lecithaster gibbosus, in Pacific salmon (genus Oncorhynchus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):893–907, April 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-087.

Pippy:1969:PLZ

[1362] John H. C. Pippy. Pomphorhynchus laevis (Zoega) Müller, 1776 (Acanthocephala) in Atlantic salmon (Salmo salar) and its use as a biological tag. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):909– 919, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-088.

Kabata:1969:PCW

[1363] Z. Kabata. Phrixocephalus cincinnatus Wilson, 1908 (Copepoda: Lernaeoceridae): morphology, metamorphosis, and host–parasite relationship. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):921–934, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-089.

Davey:1969:MPN

[1364] K. G. Davey. Moulting in a parasitic nematode, Phocanema decipiens. V. Timing of feeding during the moulting cycle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):935–939, April 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-090.

Adams:1969:MRI

[1365] James R. Adams. Migration route of invasive juvenile Philonema oncorhynchi (Nematoda: Philometridae) in young salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):941–946, April 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-091.

Rausch:1969:DCH

[1366] Robert L. Rausch. Diphyllobothriid cestodes from the Hawaiian monk seal, Monachus schauinslandi Matschie, from Midway Atoll. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):947–956, April 1969. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f69-092.

CaballeroyC:1969:ETR

[1367] Eduardo Caballero y C. and Guillermina Caballero R. Études des trématodes récoltés par Howard A. Winter chez les poissons marins de l’Océan Pacifique du Mexique et des États Unis. II. (French) [Study of trematodes collected by Howard A. Winter in marine fishes of the Pacific Ocean of Mexico and the United States. II]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):957–963, April 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-093.

Gupta:1969:NAA

[1368] S. P. Gupta and Vijay Laxmi Kajaji. A new acanthocephalan, Acanthosentis cameroni n. sp., from the intestine of a freshwater fish, Rohtee cotio Skyes, 1839, from Lucknow. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):965–974, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-094.

Burt:1969:BBD

[1369] M. D. B. Burt and I. M. Sandeman. Biology of Bothrimonus (= Diplocotyle) (Pseudophyllidea: Cestoda) Part I. History, description, synonymy, and systematics. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):975–996, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-095.

Roberts:1969:EAS

[1370] Larry S. Roberts. Ergasilus arthrosis n. sp. (Copepoda: Cyclopoida) and the taxonomic status of Ergasilus versicolor Wilson, 1911, Ergasilus elegans Wilson, 1916, and Ergasilus celestis Mueller, 1936, from North American fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26 (4):997–1011, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-096.

Parker:1969:VBC

[1371] Robert R. Parker. Validity of the binomen Caligus elongatus for a common parasitic copepod formerly misidentified with Caligus rapax. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):1013–1035, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-097.

Dollfus:1969:QEC

[1372] Robert Ph. Dollfus. Quelques espèces de cestodes Tétrarhynques de la côte atlantique des États Unis, dont l’une n’était pas connue à l’état adulte. (French) [Some species of Tetrarhynchus cestodes from the Atlantic Coast of the United States, one of which was not known as an adult]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):1037– 1061, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-098.

Llewellyn:1969:CSD

[1373] J. Llewellyn and Christine M. Tully. A comparison of speciation in diclidophorinean monogenean gill parasites and in their fish hosts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):1063–1074, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-099.

Laird:1969:MFH

[1374] Marshall Laird and Wilbur L. Bullock. Marine fish haematozoa from New Brunswick and New England. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):1075–1102, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-100.

Mawson:1969:SNA

[1375] Patricia M. Mawson. Some nematodes from Australian gulls and terns. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(4):1103–1111, April 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-101.

Buerkle:1969:AMC

[1376] Udo Buerkle. Auditory masking and the critical band in Atlantic cod (Gadus morhua). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26 (5):1113–1119, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-102.

Medcof:1969:CAO

[1377] J. C. Medcof and M. L. H. Thomas. Canadian Atlantic oyster drills (Urosalpinx) — distribution and industrial importance. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1121–1131, May 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-103.

Thomas:1969:PNE

[1378] William H. Thomas. Phytoplankton nutrient enrichment experiments off Baja California and in the Eastern Equatorial Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1133–1145, May 1969.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-104.

McBride:1969:CCP

[1379] J. R. McBride and A. P. van Overbeeke. Cytological changes in the pituitary gland of the adult sockeye salmon (Oncorhynchus nerka) after gonadectomy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5): 1147–1156, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f69-105.

Mackay:1969:AIC

[1380] Rosemary J. Mackay. Aquatic insect communities of a small stream on Mont St. Hilaire, Quebec. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1157–1183, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-106.

Lee:1969:UAU

[1381] Agnes S. K. Lee, William E. Vanstone, John R. Markert, and Naval J. Antia. UV-absorbing and UV-fluorescing substances in the belly skin of fry of Coho salmon (Oncorhynchus kisutch). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1185–1198, May 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-107.

Bilton:1969:ADS

[1382] H. T. Bilton and D. W. Jenkinson. Age determination of sockeye (Oncorhynchus nerka) and chum (O. keta) salmon from examination of pectoral fin rays. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5): 1199–1203, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f69-108.

Powles:1969:SCM

[1383] P. M. Powles. Size changes, mortality, and equilibrium yields in an exploited stock of American plaice (Hippoglossoides platessoides). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1205–1235, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-109.

Grinols:1969:NSA

[1384] Richard B. Grinols. A new species of Acantholiparis (Pisces: Liparidae) from the Eastern subarctic Pacific region, with distribution notes for the genus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5): 1237–1242, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f69-110.

Jahn:1969:MHC

[1385] Lawrence A. Jahn. Movements and homing of cutthroat trout (Salmo clarki) from open-water areas of Yellowstone Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1243–1261, May 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-111.

Weber:1969:EFS

[1386] Douglas Weber and Roy J. Wahle. Effect of finclipping on survival of sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1263–1271, May 1969. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f69-112.

Kohler:1969:CFC

[1387] A. C. Kohler and D. N. Fitzgerald. Comparisons of food of cod and haddock in the Gulf of St. Lawrence and on the Nova Scotia Banks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1273–1287, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-113.

Bailey:1969:AGM

[1388] Merryll M. Bailey. Age, growth, and maturity of the longnose sucker Catostomus catostomus, of Western Lake Superior. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1288–1298, May 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-114.

Pitt:1969:MAP

[1389] T. K. Pitt. Migrations of American plaice on the Grand Bank and in St. Mary’s Bay, 1954, 1959, and 1961. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1301–1319, May 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-115.

Sameoto:1969:SAE

[1390] D. D. Sameoto. Some aspects of the ecology and life cycle of three species of subtidal sand-burrowing amphipods (Crustacea: Haustoriidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1321–1345, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-116.

Lewis:1969:SCF

[1391] A. G. Lewis and A. Ramnarine. Some chemical factors affecting the early developmental stages of Euchaeta japonica (Crustacea: Copepoda: Calanoida) in the laboratory. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1347–1362, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-117.

Narver:1969:POL

[1392] David W. Narver. Productivity of Owikeno Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1363–1368, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-118.

Manohar:1969:SPF

[1393] S. V. Manohar. Some properties of the fluorescence of fish muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1368–1371, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-119.

Pattie:1969:EKN

[1394] B. H. Pattie and C. S. Baker. Extensions of the known northern range limits of ocean whitefish, Caulolatilus princeps, and California halibut, Paralichthys californicus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1371–1372, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-120.

Larkin:1969:FSS

[1395] P. A. Larkin and A. Walton. Fish school size and migration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1372–1374, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-121.

Collette:1969:FMB

[1396] Bruce B. Collette and John A. MacPhee, Jr. First Massachusetts Bay record of the Arctic shanny Stichaeus punctatus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1375–1377, May 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-122.

Li:1969:USC

[1397] M. F. Li and Carol Jordan. Use of spinner culture cells to detect water pollution. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1378– 1382, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-123.

Lander:1969:STS

[1398] R. H. Lander. Swimming thrust of sockeye salmon (Oncorhynchus nerka) in relation to selectivity of gillnets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1383–1385, May 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-124.

Stewart:1969:CGH

[1399] James E. Stewart, Diane M. Foley, and R. G. Ackman. Characteristics of Gaffkya homari, the causative agent of the lobster disease gaffkemia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1385–1389, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-125.

Garside:1969:DIF

[1400] E. T. Garside. Distribution of insular fishes of Sable Island, Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1390–1392, May 1969.

CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-126.

Stewart:1969:PLR

[1401] James E. Stewart and Diane M. Foley. A precipitin-like reaction of the hemolymph of the lobster Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1392–1397, May 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-127.

Naevdal:1969:BPP

[1402] Gunnar Nævdal. Blood protein polymorphism in harp seals off Eastern Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1397– 1399, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-128.

Graybill:1969:LFS

[1403] J. R. Graybill and H. F. Horton. Limited fertilization of steelhead trout eggs with cryo-preserved sperm. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1400–1404, May 1969. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f69-129.

Anonymous:1969:RAFd

[1404] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(5):1405, May 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-130.

Kabata:1969:TFN

[1405] Z. Kabata. Tanypleuridae fam.nov. (Copepoda: Caligoida), parasitic on fishes in the Canadian Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1407–1414, June 1969. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f69-131.

Mills:1969:CCM

[1406] Eric L. Mills. The community concept in marine zoology, with comments on continua and instability in some marine communities: A review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1415–1428, June 1969. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f69-132.

Allen:1969:DAE

[1407] K. Radway Allen. Distinctive aspects of the ecology of stream fishes: A review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1429– 1438, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-133.

Willock:1969:DLF

[1408] Thomas A. Willock. Distributional list of fishes in the Missouri drainage of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1439– 1449, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-134.

Bernard:1969:HFS

[1409] F. R. Bernard and J. W. Bagshaw. Histology and fine structure of the accessory boring organ of Polinices lewisi (Gastropoda, Prosobranchiata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1451–1457, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-135.

Billi:1969:QCP

[1410] James L. Billi and Ken Wolf. Quantitative comparison of peritoneal washes and feces for detecting infectious pancreatic necrosis (IPN) virus in carrier brook trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26 (6):1459–1465, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-136.

Gruchy:1969:CRW

[1411] C. G. Gruchy. Canadian records of the warty poacher Occa verrucosa, with notes on the standardization of plate terminology in agonidae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1467–1472, June 1969. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f69-137.

Hunt:1969:OSW

[1412] Robert L. Hunt. Overwinter survival of wild fingerling brook trout in Lawrence Creek, Wisconsin. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1473–1483, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-138.

McLaren:1969:PPE

[1413] Ian A. McLaren. Population and production ecology of zooplankton in Ogac Lake, a landlocked fiord on Baffin Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1485–1559, June 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-139.

McLaren:1969:PPN

[1414] Ian A. McLaren. Primary production and nutrients in Ogac Lake, a landlocked fiord on Baffin Island. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1561–1576, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-140.

Addison:1969:FFA

[1415] R. F. Addison, R. G. Ackman, and J. Hingley. Free fatty acids of herring oils: Possible derivation from both phospholipids and triglycerides in fresh herring. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26 (6):1577–1583, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-141.

Leggett:1969:DBT

[1416] William C. Leggett and G. Power. Differences between two populations of landlocked Atlantic salmon (Salmo salar) in Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1585–1596, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-142.

Dyer:1969:NDF

[1417] W. J. Dyer and Doris I. Hiltz. Nucleotide degradation in frozen swordfish muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1597– 1603, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-143.

McQueen:1969:RZS

[1418] Donald J. McQueen. Reduction of zooplankton standing stocks by predaceous Cyclops bicuspidatus thomasi in Marion Lake, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1605–1618, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-144.

Phinney:1969:FTF

[1419] Duane E. Phinney and Stephen B. Mathews. Field test of fluorescent pigment marking and finclipping of Coho salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1619–1624, June 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-145.

Rao:1969:OEC

[1420] D. V. Subba Rao and T. Platt. Optimal extraction conditions of chlorophylls from cultures of five species of marine phytoplankton. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1625–1630, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-146.

LeBrasseur:1969:GJC

[1421] R. J. LeBrasseur. Growth of juvenile chum salmon (Oncorhynchus keta) under different feeding regimes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1631–1645, June 1969. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f69-147.

Houde:1969:SSA

[1422] Edward D. Houde. Sustained swimming ability of larvae of walleye (Stizostedion vitreum vitreum) and yellow perch (Perca flavescens). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1647–1659, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-148.

Ducharme:1969:ASR

[1423] L. J. A. Ducharme. Atlantic salmon returning for their fifth and sixth consecutive spawning trips. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1661–1664, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-149.

Westenhouse:1969:DRD

[1424] Ray G. Westenhouse. Dichromate reagent decomposition and carbon estimation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1664– 1667, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-150.

Hamilton:1969:MSI

[1425] Andrew L. Hamilton. A method of separating invertebrates from sediments using longwave ultraviolet light and fluorescent dyes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1667–1672, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-151.

Nursall:1969:ASS

[1426] J. R. Nursall and Morley E. Pinsent. Aggregations of spottail shiners and yellow perch. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26 (6):1672–1676, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-152.

Dragovich:1969:RBL

[1427] Alexander Dragovich and John P. Wise. Relations between length of the pen and other measurements in the squid Lolliguncula brevis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1676–1679, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-153.

Eagle:1969:FRN

[1428] Rodney J. Eagle. First records from the Northeastern Pacific of the deep-sea ophidioid fish Parabassogigas grandis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1680–1685, June 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-154.

Hamilton:1969:NTE

[1429] Andrew L. Hamilton. A new type of emergence trap for collecting stream insects. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1685– 1689, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-155.

Ennis:1969:OLS

[1430] G. P. Ennis. Occurrences of the little sculpin, Myoxocephalus aeneus, in Newfoundland waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1689–1694, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-156.

Murray:1969:MDA

[1431] A. R. Murray, D. R. L. White, and R. S. Whitaker. Muscle damage in adult Atlantic salmon caught in nylon gillnets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1694–1696, June 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-157.

Anonymous:1969:RAFe

[1432] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(6):1697, June 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-158.

MacCrimmon:1969:WDB

[1433] Hugh R. MacCrimmon and J. Scott Campbell. World distribution of brook trout, Salvelinus fontinalis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1699–1725, July 1969. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f69-159.

Kyte:1969:SKR

[1434] Michael A. Kyte. A synopsis and key to the recent Ophiuroidea of Washington State and Southern British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1727–1741, July 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-160.

Fraser:1969:EFS

[1435] J. H. Fraser. Experimental feeding of some Medusae and Chaetognatha. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1743–1762, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-161.

Northcote:1969:SCL

[1436] T. G. Northcote and T. G. Halsey. Seasonal changes in the limnology of some meromictic lakes in Southern British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1763–1787, July 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-162.

Donaldson:1969:CSR

[1437] Edward M. Donaldson and Ulf H. M. Fagerlund. Cortisol secretion rate in gonadectomized female sockeye salmon (Oncorhynchus nerka): Effects of estrogen and cortisol treatment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1789–1799, July 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-163.

Windell:1969:DRR

[1438] John T. Windell, David O. Norris, James F. Kitchell, and James S. Norris. Digestive response of rainbow trout, Salmo gairdneri, to pellet diets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1801–1812, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-164.

Savitz:1969:ETB

[1439] Jan Savitz. Effects of temperature and body weight on endogenous nitrogen excretion in the bluegill sunfish (Lepomis macrochirus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1813–1821, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-165.

Grimm:1969:SST

[1440] A. S. Grimm, M. J. O’Halloran, and D. R. Idler. Stimulation of sodium transport across the isolated toad bladder by 1α-hydroxycorticosterone from an elasmobranch. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1823–1835, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-166.

McKeown:1969:ILA

[1441] B. A. McKeown and A. P. van Overbeeke. Immunohistochemical localization of ACTH and prolactin in the pituitary gland of adult migratory sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1837–1846, July 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-167.

Houston:1969:SPC

[1442] A. H. Houston, Mary Anne DeWilde, and Jane A. Madden. Some physiological consequences of aortic catheterization in the brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7): 1847–1856, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f69-168.

Bilinski:1969:LAT

[1443] E. Bilinski and Y. C. Lau. Lipolytic activity toward long-chain triglycerides in lateral line muscle of rainbow trout (Salmo gairdneri). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1857–1866, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-169.

Symons:1969:GDW

[1444] Philip E. K. Symons. Greater dispersal of wild compared with hatcheryreared juvenile Atlantic salmon released in streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1867–1876, July 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-170.

Wolfert:1969:MFW

[1445] David R. Wolfert. Maturity and fecundity of walleyes from the Eastern and Western Basins of Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1877–1888, July 1969. CODEN JFRBAK.

ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/10.
1139/f69-171.

Messieh:1969:SON

[1446] Shoukry N. Messieh. Similarity of otolith nuclei in spring- and autumnspawning Atlantic herring in the Southern Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1889–1898, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-172.

Squires:1969:DCB

[1447] H. J. Squires. Decapod crustacea of the Beaufort Sea and Arctic waters eastward to Cambridge Bay, 1960–65. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1899–1918, July 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-173.

Dambergs:1969:IWM

[1448] N. Dambergs. Isopropanol–water mixtures for the production of fish protein concentrate from Atlantic herring (Clupea harengus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1919–1923, July 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-174.

Dambergs:1969:IWA

[1449] N. Dambergs. Isopropanol–water azeotrope as solvent in the production of fish protein concentrate from herring (Clupea harengus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1923–1926, July 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-175.

Eales:1969:CSP

[1450] J. G. Eales. A comparative study of purines responsible for silvering in several freshwater fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1927–1931, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-176.

Scarratt:1969:LLP

[1451] D. J. Scarratt. Lobster larvae off Pictou, Nova Scotia, not affected by bleached kraft mill effluent. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1931–1934, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-177.

Paterson:1969:OCA

[1452] Colin G. Paterson. Occurrence of Coregonus artedii and C. zenithicus in Barrow Lake, Alberta. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1934–1939, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-178.

Fromm:1969:UDI

[1453] Paul O. Fromm and Robert C. Hunter. Uptake of dieldrin by isolated perfused gills of rainbow trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1939–1942, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-179.

Dunn:1969:AAW

[1454] R. S. Dunn and A. V. Tyler. Aspects of the anatomy of the winter flounder ovary with hypotheses on oocyte maturation time. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1943–1947, July 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-180.

Schindler:1969:TUD

[1455] D. W. Schindler. Two useful devices for vertical plankton and water sampling. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1948– 1955, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-181.

Best:1969:REF

[1456] E. A. Best and Peter J. Eldridge. Range extension of flag rockfish (Sebastodes rubrivinctus) to Aleutian Islands. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1955–1956, July 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-182.

Pugh:1969:ETS

[1457] John R. Pugh, Alan B. Groves, and Wesley J. Ebel. Experimental tank to simulate certain reservoir conditions. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1956–1959, July 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-183.

Hughes:1969:AIS

[1458] Roger N. Hughes. Appraisal of the iodate–sulphuric acid wet-oxidation procedure for the estimation of the caloric content of marine sediments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7):1959–1964, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-184.

Herold:1969:HSS

[1459] Richard C. Herold. Hermaphrodite specimen of the sand dollar, Echinarachnius parma. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26 (7):1965–1966, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-185.

Anonymous:1969:RAFf

[1460] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(7): 1967, July 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-186.

Apollonio:1969:BFG

[1461] S. Apollonio. Breeding and fecundity of the glass shrimp, Pasiphaea multidentata (Decapoda, Caridea), in the Gulf of Maine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):1969–1983, August 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-187.

Alverson:1969:DFE

[1462] Dayton L. Alverson and Walter T. Pereyra. Demersal fish explorations in the Northeastern Pacific Ocean — an evaluation of exploratory fishing methods and analytical approaches to stock size and yield forecasts. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):1985–2001, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-188.

Hargrave:1969:EAP

[1463] Barry T. Hargrave. Epibenthic algal production and community respiration in the sediments of Marion Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2003–2026, August 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-189.

MacCallum:1969:NCP

[1464] W. A. MacCallum, D. R. Adams, R. G. Ackman, P. J. Ke, W. J. Dyer, D. I. Fraser, and S. Punjamapirom. Newfoundland capelin: Proximate composition. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2027– 2035, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f69-190.

Ackman:1969:NCL

[1465] R. G. Ackman, P. J. Ke, W. A. MacCallum, and D. R. Adams. Newfoundland capelin lipids: Fatty acid composition and alterations during frozen storage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8): 2037–2060, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-191.

McInerney:1969:RBB

[1466] John E. McInerney. Reproductive behaviour of the blackspotted stickleback, Gasterosteus wheatlandi. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2061–2077, August 1969. CODEN JFRBAK. ISSN

0015-296X. URL https://cdnsciencepub.com/doi/full/10.1139/
f69-192.

Brawn:1969:BAP

[1467] Vivien M. Brawn. Buoyancy of Atlantic and Pacific herring. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2077–2091, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-193.

Pinder:1969:SBC

[1468] Lesley J. Pinder and J. G. Eales. Seasonal buoyancy changes in Atlantic salmon (Salmo salar) parr and smolt. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2093–2100, August 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-194.

Barlow:1969:CSP

[1469] John Barlow and George J. Ridgway. Changes in serum protein during the molt and reproductive cycles of the American lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8): 2101–2109, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-195.

Oshima:1969:ODN

[1470] Kiyoshi Oshima, William E. Hahn, and Aubrey Gorbman. Olfactory discrimination of natural waters by salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2111–2121, August 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-196.

Oshima:1969:EOR

[1471] Kiyoshi Oshima, William E. Hahn, and Aubrey Gorbman. Electroencephalographic olfactory responses in adult salmon to waters traversed in the homing migration. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2123–2133, August 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-197.

Patalas:1969:CHD

[1472] K. Patalas. Composition and horizontal distribution of crustacean plankton in Lake Ontario. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8): 2135–2164, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-198.

Chaston:1969:SAF

[1473] I. Chaston. Seasonal activity and feeding pattern of brown trout (Salmo trutta) in a Dartmoor stream in relation to availability of food. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2165–2171, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-199.

Hasler:1969:OWO

[1474] Arthur D. Hasler, Edward S. Gardella, Ross M. Horrall, and H. Francis Henderson. Open-water orientation of white bass, Roccus chrysops, as determined by ultrasonic tracking methods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2173–2192, August 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-200.

Matton:1969:EOD

[1475] Pierre Matton and Quentin N. LaHam. Effect of the organophosphate dylox on rainbow trout larvae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2193–2200, August 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-201.

Sergeant:1969:TWW

[1476] D. E. Sergeant and P. F. Brodie. Tagging white whales in the Canadian Arctic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2201–2205, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-202.

Jonkel:1969:WWW

[1477] C. J. Jonkel. White whales wintering in James Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2205–2207, August 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-203.

Rayner:1969:MTP

[1478] Martin D. Rayner, Morris H. Baslow, and Thomas I. Kosaki. Marine toxins from the Pacific — ciguatoxin: not an in vivo anticholinesterase. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2208–2210, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-204.

Pereyra:1969:SFS

[1479] Walter T. Pereyra, Forrest E. Carvey, Jr., and William G. Pearcy. Sebastodes flavidus, a shelf rockfish feeding on mesopelagic fauna, with consideration of the ecological implications. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2211–2215, August 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-205.

Wellings:1969:MIE

[1480] S. R. Wellings, L. E. Ashley, and G. E. McArn. Microsporidial infection of English sole, Parophrys vetulus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2215–2218, August 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-206.

Manzer:1969:SCJ

[1481] J. I. Manzer. Stomach contents of juvenile Pacific salmon in Chatham Sound and adjacent waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2219–2223, August 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-207.

Holland:1969:ERS

[1482] Geoffrey L. Holland. Effect of the rate of storm growth on subsequent surge elevations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8): 2223–2227, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-208.

Gee:1969:SST

[1483] J. H. Gee and V. G. Bartnik. Simple stream tank simulating a rapids environment. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2227–2230, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-209.

Bernard:1969:PDN

[1484] F. R. Bernard. Preliminary diagnoses of new septibranch species from the Eastern Pacific (Bivalvia, Anomalodesmata). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2230–2234, August 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-210.

Saito:1969:CRC

[1485] A. Saito. Color in raw and cooked Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2234–2236, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-211.

Jonas:1969:EMS

[1486] R. E. E. Jonas. Effect of magnesium sulphate and cryoprotective agents on free fatty acid formation in cold-stored fish muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2237–2240, August 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-212.

Sameoto:1969:OSS

[1487] D. D. Sameoto and L. O. Jaroszynski. Otter surface sampler: a new neuston net. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8): 2240–2244, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-213.

Johnson:1969:PRF

[1488] Eugene A. Johnson and Kenneth K. Chew. Preliminary report on the fecundity of Mytilicola orientalis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2245–2246, August 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-214.

Lightner:1969:MCI

[1489] Donald Lightner and George Post. Morphological characteristics of infectious pancreatic necrosis yirus in trout pancreatic tissue. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2247–2250, August 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-215.

Orr:1969:MLC

[1490] T. S. C. Orr and C. A. Hopkins. Maintenance of the life cycle of Ligula intestinalis in the laboratory. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2250–2251, August 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-216.

Muth:1969:AGB

[1491] K. M. Muth. Age and growth of broad whitefish, Coregonus nasus, in the Mackenzie and Coppermine Rivers, N.W.T. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2252–2256, August 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-217.

Sepic:1969:HGC

[1492] K. Sepic and W. A. MacCallum. Heading and gutting of capelin with special hand shears. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8): 2257–2259, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-218.

Buchwald:1969:TCA

[1493] D. G. Buchwald and J. R. Nursall. Triaenophorus crassus in Arctic lampreys of the Northwest Territories, Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2260–2261, August 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-219.

Anonymous:1969:CRP

[1494] Anonymous. Corrections for recent publications of The Fisheries Research Board of Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2263–2264, August 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-220.

Anonymous:1969:RAFg

[1495] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(8):2265, August 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-221.

Allen:1969:ABG

[1496] K. Radway Allen. Application of the von Bertalanffy growth equation to problems of fisheries management: A review. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2267–2281, September 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-222.

Sameoto:1969:PTB

[1497] D. D. Sameoto. Physiological tolerances and behaviour responses of five species of Haustoriidae (Amphipoda: Crustacea) to five environmental factors. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2283–2298, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-223.

Castell:1969:EHC

[1498] C. H. Castell and D. M. Bishop. Effect of hematin compounds on the development of rancidity in muscle of cod, flounder, scallops, and lobster. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2299–2309, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-224.

McFadden:1969:EDT

[1499] T. W. McFadden. Effective disinfection of trout eggs to prevent egg transmission of Aeromonas liquefaciens. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2311–2318, September 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-225.

Arai:1969:PRP

[1500] Hisao P. Arai. Preliminary report on the parasites of certain marine fishes of British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9): 2319–2337, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-226.

Wilson:1969:RFA

[1501] Dennis C. Wilson and Raymond E. Millemann. Relationships of female age and size to embryo number and size in the shiner perch, Cymatogaster aggregata. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2339–2344, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-227.

Platt:1969:CCE

[1502] Trevor Platt, Vivien M. Brawn, and Brian Irwin. Caloric and carbon equivalents of zooplankton biomass. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2345–2349, September 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-228.

Tsuyuki:1969:STS

[1503] H. Tsuyuki, E. Roberts, and E. A. Best. Serum transferrin systems and the hemoglobins of the Pacific halibut (Hippoglossus stenolepis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2351–2362, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-229.

Brett:1969:GRB

[1504] J. R. Brett, J. E. Shelbourn, and C. T. Shoop. Growth rate and body composition of fingerling sockeye salmon, Oncorhynchus nerka, in relation to temperature and ration size. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2363–2394, September 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-230.

Culley:1969:PIR

[1505] Dudley D. Culley, Jr. and Denzel E. Ferguson. Patterns of insecticide resistance in the mosquitofish, Gambusia affinis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2395–2401, September 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-231.

Levings:1969:ZLP

[1506] C. D. Levings. The zoarcid Lycodopsis pacifica in Outer Burrard Inlet, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9): 2403–2412, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-232.

Budd:1969:FOS

[1507] J. C. Budd, F. E. J. Fry, and P. S. M. Pearlstone. Final observations on the survival of planted lake trout in South Bay, Lake Huron. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2413–2424, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-233.

Eastman:1969:PCV

[1508] Joseph T. Eastman. Progressive changes in the ventral aorta of the carp, Cyprinus carpio. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9): 2425–2430, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-234.

Nelson:1969:GVB

[1509] Joseph S. Nelson. Geographic variation in the brook stickleback, Culaea inconstans, and notes on nomenclature and distribution. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2431–2447, September 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-235.

Mount:1969:CTC

[1510] Donald I. Mount and Charles E. Stephan. Chronic toxicity of copper to the fathead minnow (Pimephales promelas) in soft water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2449–2457, September 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-236.

Davis:1969:SDC

[1511] Charles C. Davis. Seasonal distribution, constitution, and abundance of zooplankton in Lake Erie. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2459–2476, September 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-237.

Yuen:1969:ETB

[1512] Kenneth B. Yuen. Effect of tidal barriers upon the M_{2} tide in the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2477–2492, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-238.

Dorfman:1969:EFL

[1513] D. Dorfman and W. R. Whitworth. Effects of fluctuations of lead, temperature, and dissolved oxygen on the growth of brook trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2493–2501, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-239.

Stewart:1969:ITG

[1514] James E. Stewart, John W. Cornick, and B. M. Zwicker. Influence of temperature on gaffkemia, a bacterial disease of the lobster Homarus americanus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2503–2510, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-240.

Wolf:1969:IPN

[1515] Ken Wolf and M. C. Quimby. Infectious pancreatic necrosis: Clinical and immune response of adult trouts to inoculation with live virus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2511–2516, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-241.

Farrington:1969:ETA

[1516] John W. Farrington and Dwight L. Baker. Effect of temperature on the activity of glyceraldehyde-3-phosphate dehydrogenase; a comparison of halibut and rabbit enzymes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2517–2521, September 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-242.

Chen:1969:RHL

[1517] Maynard Y. Chen. A record of hermaphroditism in lake whitefish, Coregonus clupeaformis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9): 2521–2523, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-243.

Varga:1969:ESC

[1518] S. Varga and C. M. Blackwood. Effect of seawater chilling on landed quality of scallop meat. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9): 2523–2526, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-244.

Bourne:1969:DSL

[1519] Neil Bourne and M. A. Pope. Deep-sea line fishing off British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2527–2531, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-245.

Nyman:1969:PSE

[1520] O. L. Nyman. Polymorphic serum esterases in two species of freshwater fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2532–2534, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-246.

Pippy:1969:KDJ

[1521] J. H. C. Pippy. Kidney disease in juvenile Atlantic salmon (Salmo salar) in the Margaree River. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2535–2537, September 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-247.

MacKay:1969:MAA

[1522] K. T. MacKay and E. T. Garside. Meristic analyses of Atlantic mackerel, Scomber scombrus, from the North American coastal populations. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2537–2540, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-248.

Anonymous:1969:RAFh

[1523] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(9):2541, September 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-249.

Carter:1969:LCL

[1524] John C. H. Carter. Life cycles of Limnocalanus macrurus and Senecella calanoides, and seasonal abundance and vertical distributions of various planktonic copepods, in Parry Sound, Georgian Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2543–2560, October 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-250.

Sergeant:1969:BSW

[1525] D. E. Sergeant and P. F. Brodie. Body size in white whales, Delphinapterus leucas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10): 2561–2580, October 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-251.

Brinkhurst:1969:MSP

[1526] R. O. Brinkhurst, K. E. Chua, and E. Batoosingh. Modifications in sampling procedures as applied to studies on the bacteria and tubificid oligochaetes inhabiting aquatic sediments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2581–2593, October 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-252.

Banse:1969:AFP

[1527] Karl Banse. Acrocirridae n. fam. (Polychaeta sedentaria). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2595–2620, October 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-253.

Nowlan:1969:GNC

[1528] Sandra S. Nowlan and W. J. Dyer. Glycolytic and nucleotide changes in the critical freezing zone, −0.8 to −5 C, in prerigor cod muscle frozen at various rates. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10): 2621–2632, October 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-254.

Tsuyuki:1969:MPP

[1529] H. Tsuyuki and E. Roberts. Muscle protein polymorphism of sablefish from the Eastern Pacific Ocean. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2633–2641, October 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-255.

Glass:1969:DCP

[1530] Norman R. Glass. Discussion of calculation of power function with special reference to respiratory metabolism in fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2643–2650, October 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-256.

Pelroy:1969:EET

[1531] Gretchen A. Pelroy and John P. Seman, Jr. Effect of EDTA treatment on spoilage characteristics of petrale sole and ocean perch fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2651–2657, October 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-257.

Brinkhurst:1969:PIE

[1532] Ralph O. Brinkhurst and Kian E. Chua. Preliminary investigation of the exploitation of some potential nutritional resources by three sympatric tubificid oligochaetes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10): 2659–2668, October 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-258.

Mettrick:1969:NCC

[1533] D. F. Mettrick and J. B. Jennings. Nutrition and chemical composition of the rhabdocoel turbellarian Syndesmis franciscana, with notes on the taxonomy of S. antillarum. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2669–2679, October 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-259.

McCombie:1969:SES

[1534] A. M. McCombie and A. H. Berst. Some effects of shape and structure of fish on selectivity of gillnets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2681–2689, October 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-260.

Halliday:1969:RMO

[1535] R. G. Halliday and W. B. Scott. Records of mesopelagic and other fishes from the Canadian Atlantic with notes on their distribution. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2691–2702, October 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-261.

Edelstein:1969:PSS

[1536] T. Edelstein, J. S. Craigie, and J. McLachlan. Preliminary survey of the sublittoral flora of Halifax County. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2703–2713, October 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-262.

Tautz:1969:SES

[1537] Arthur Tautz, P. A. Larkin, and W. E. Ricker. Some effects of simulated long-term environmental fluctuations on maximum sustained yield. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2715–2726, October 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-263.

Anderson:1969:RFF

[1538] Margaret L. Anderson and Elinor M. Ravesi. Reaction of free fatty acids with protein in cod muscle frozen and stored at −29 C after aging in ice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2727–2736, October 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-264.

McMullen:1969:EDM

[1539] John C. McMullen. Effects of delayed mating on the reproduction of king crab, Paralithodes camtschatica. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2737–2740, October 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-265.

Mathisen:1969:AWL

[1540] Ole A. Mathisen. Allocation of weight and length measurements for estimation of the average weight of juvenile sockeye salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2741–2746, October 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-266.

Colton:1969:TCG

[1541] John B. Colton, Jr. Temperature conditions in the Gulf of Maine and adjacent waters during 1968. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2746–2751, October 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-267.

Butler:1969:TBS

[1542] T. H. Butler and R. W. Sheldon. Trawl-board sediment sampler. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2751–2753, October 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f69-268.

Narver:1969:ASS

[1543] David W. Narver. Age and size of steelhead trout in the Babine River, British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10): 2754–2760, October 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-269.

Murchelano:1969:BFS

[1544] Robert A. Murchelano and Carolyn Brown. Bacterial flora of some algal foods used for rearing bivalve larvae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2760–2764, October 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-270.

Mighell:1969:RCB

[1545] James L. Mighell. Rapid cold-branding of salmon and trout with liquid nitrogen. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2765–2769, October 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-271.

MacKay:1969:FRA

[1546] K. T. MacKay and Gethin Thomas. First records of Ariomma bondi, Caranx crysos, and Selar crumenopthalmus (Pisces) in the Gulf of St. Lawrence. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2769–2771, October 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-272.

Anonymous:1969:RAFi

[1547] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(10):2773, October 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-273.

Greenberg:1969:MTD

[1548] David A. Greenberg. Modification of the M_{2} tide due to barriers in the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11): 2775–2783, November 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-274.

Baker:1969:HEM

[1549] Jeremy T. P. Baker. Histological and electron microscopical observations on copper poisoning in the winter flounder (Pseudopleuronectes americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):2785–2793, November 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-275.

Mackay:1969:FTC

[1550] I. Mackay and K. H. Mann. Fecundity of two cyprinid fishes in the River Thames, Reading, England. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):2795–2805, November 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-276.

Farmer:1969:OCT

[1551] G. J. Farmer and F. W. H. Beamish. Oxygen consumption of Tilapia nilotica in relation to swimming speed and salinity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):2807–2821, November 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-277.

Fessler:1969:SMB

[1552] James L. Fessler and Harry H. Wagner. Some morphological and biochemical changes in steelhead trout during the parr–smolt transformation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):2823–2841, November 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-278.

Riffenburgh:1969:SMI

[1553] Robert H. Riffenburgh. A stochastic model of interpopulation dynamics in marine ecology. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11): 2843–2880, November 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-279.

Chlebeck:1969:HST

[1554] Anne Chlebeck and Gary L. Phillips. Hematological study of two buffalofishes, Ictiobus cyprinellus and I. bubalus (Catostomidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):2881–2886, November 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-280.

Godin:1969:TET

[1555] G. Godin. Theory of the exploitation of tidal energy and its application to the Bay of Fundy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):

2887–2957, November 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-281.

Craigie:1969:SSI

[1556] J. S. Craigie. Some salinity-induced changes in growth, pigments, and cyclohexanetetrol content of Monochrysis lutheri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):2959–2967, November 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-282.

Dabrowski:1969:CCS

[1557] Teofil Dabrowski, Edward Kolakowski, and Barbara Karnicka. Chemical composition of shrimp flesh (Parapenaeus spp.) and its nutritive value. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):2969–2974, November 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-283.

McBride:1969:HIT

[1558] J. R. McBride and A. P. van Overbeeke. Hypertrophy of the interrenal tissue in sexually maturing sockeye salmon (Oncorhynchus nerka) and the effect of gonadectomy. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):2975–2985, November 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-284.

Kabata:1969:RGS

[1559] Z. Kabata. Revision of the genus Salmincola Wilson, 1915 (Copepoda: Lernaeopodidae). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11): 2987–3041, November 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-285.

Kabata:1969:CNS

[1560] Z. Kabata. Chondracanthus narium sp. n. (Copepoda: Chondracanthidae), a parasite of nasal cavities of Ophiodon elongatus (Pisces: Teleostei) in British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11): 3043–3047, November 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-286.

Clayton:1969:LDI

[1561] J. W. Clayton and J. H. Gee. Lactate dehydrogenase isozymes in longnose and blacknose dace (Rhinichthys cataractae and R. atratulus) and their hybrid. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):3049–3053, November 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-287.

Das:1969:TBC

[1562] B. S. Das, S. G. Reid, J. L. Betts, and K. Patrick. Tetrachloro-o-benzoquinone as a component in bleached kraft chlorination effluent toxic to young salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11): 3055–3067, November 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-288.

Knight:1969:FBG

[1563] William Knight. A formulation of the von Bertalanffy growth curve when the growth rate is roughly constant. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):3069–3072, November 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-289.

Jhingran:1969:DAL

[1564] V. G. Jhingran and A. V. Natarajan. Derivation of average lengths of different age-groups in fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):3073–3076, November 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-290.

Manion:1969:ELL

[1565] Patrick J. Manion. Evaluation of lamprey larvicides in the Big Garlic River and Saux Head Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):3077–3082, November 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-291.

Durkin:1969:DDM

[1566] Joseph T. Durkin, Wesley J. Ebel, and Jim Ross Smith. A device to detect magnetized wire tags in migrating adult Coho salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):3083–3088, November 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-292.

Light:1969:ERM

[1567] William J. Light. Extension of range for Manayunkia aestuarina (Polychaeta: Sabellidae) to British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):3088–3091, November 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-293.

Leggett:1969:PLA

[1568] William C. Leggett. Pugheadedness in landlocked Atlantic salmon (Salmo salar). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11): 3091–3093, November 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-294.

Dunn:1969:OOB

[1569] J. R. Dunn and C. R. Hitz. Oceanic occurrence of black rockfish (Sebastodes melanops) in the Central North Pacific. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):3094–3097, November 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-295.

Anonymous:1969:RAFj

[1570] Anonymous. Recent articles on FRB investigations published in other periodicals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(11):3099, November 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-296.

Smith:1969:CEB

[1571] M. W. Smith. Changes in environment and biota of a natural lake after fertilization. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12): 3101–3132, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-297.

Pinhorn:1969:FBA

[1572] Allenby T. Pinhorn. Fishery and biology of Atlantic cod (Gadus morhua) off the southwest coast of Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3133–3164, December 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-298.

Seki:1969:MBO

[1573] H. Seki and Owen D. Kennedy. Marine bacteria and other heterotrophs as food for zooplankton in the Strait of Georgia during the winter. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3165–3173, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-299.

Gould:1969:AGD

[1574] Edith Gould. Alpha-glycerophosphate dehydrogenase as an index of icedstorage age of fresh, gutted haddock (Melanogrammus aeglefinus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3175–3181, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-300.

McPhail:1969:PES

[1575] J. D. McPhail. Predation and the evolution of a stickleback (Gasterosteus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3183–3208, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-301.

Buhler:1969:EFD

[1576] Donald R. Buhler and P. Benville. Effect of feeding and of DDT on the activity of hepatic glucose 6-phosphate dehydrogenase in two salmonids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3209–3216, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-302.

Hayward:1969:BCC

[1577] M. J. Hayward and W. A. MacCallum. Bacteria counts on cod and flounder fillets produced commercially from fish frozen at sea and thawed in water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3217–3231, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-303.

Kwain:1969:FOR

[1578] Wen hwa Kwain and Hugh R. MacCrimmon. Further observations on the response of rainbow trout, Salmo gairdneri, to overhead light. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3233–3237, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-304.

Rafail:1969:FAR

[1579] Samir Zaky Rafail. Further analysis of ration and growth relationship of plaice (Pleuronectes platessa). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3237–3241, December 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-305.

Wiles:1969:FCL

[1580] Michael Wiles. Fibrous and cystic lesions in the ovaries of aged Atlantic cod (Gadus morhua): a preliminary report. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3242–3246, December 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-306.

Bourne:1969:BHE

[1581] Neil Bourne and D. E. McAllister. The black hagfish, Eptatretus deani, from British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3246–3248, December 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-307.

Tibbo:1969:LSX

[1582] S. N. Tibbo and L. M. Lauzier. Larval swordfish (Xiphias gladius) from three localities in the Western Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3248–3251, December 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-308.

McIntyre:1969:SBB

[1583] John D. McIntyre. Spawning behavior of the brook lamprey, Lampetra planeri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3252–3254, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-309.

Truscott:1969:IEF

[1584] B. Truscott and D. R. Idler. An improved extender for freezing Atlantic salmon spermatozoa. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12): 3254–3258, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-310.

MacKelvie:1969:IPN

[1585] R. M. MacKelvie and H. Artsob. Infectious pancreatic necrosis virus in young salmonids of the Canadian Maritime Provinces. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3259–3262, December 1969. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f69-311.

Arita:1969:SDC

[1586] George S. Arita. Sexual dimorphism in the cyclopterid fish Eumicrotremus orbis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12): 3262–3265, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f69-312.

Lawler:1969:APS

[1587] G. H. Lawler. Activity periods of some fishes in Heming Lake, Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3266–3267, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-313.

Utter:1969:TVP

[1588] Fred M. Utter. Transferrin variants in Pacific hake (Merluccius productus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3268–3271, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-314.

Wood:1969:QEF

[1589] A. L. Wood and R. Legendre. Quality evaluations of freeze-dried cod steaks. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3271–3274, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-315.

Jenkins:1969:NFB

[1590] Thomas M. Jenkins, Jr. Night feeding of brown and rainbow trout in an experimental stream channel. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3275–3278, December 1969. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f69-316.

Curtis:1969:SPS

[1591] Mark A. Curtis. Synonymy of the polychaete Scoloplos acutus with S. arrniger. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3279–3282, December 1969. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f69-317.

Anonymous:1969:IP

[1592] Anonymous. Index to publications for 1969. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 26(12):3283–3379, December 1969. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f69-318.